A Complete Bibliography of Publications in the Journal of Mathematical Physics: 2005–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.20

Title word cross-reference

(1 + 1) [1242, 753]. (2 + 1) [1905, 822, 1547, 1101]. (2, 0) [575]. (3 + 1) [1203]. (G′/G) [1741, 1724]. (iz)^{m} [312]. (λ/4!)φ^{4} [611]. (n + 1) [888, 1222]. (p, q) [941, 941]. (q, γ) [1304]. (x, z) [247]. −d^{2}/dr^{2} − 1/(4r^{2}) [595]. 1 [1634]. 1/2 [631, 576, 67]. 10 [1036, 1608]. 11 [1036, 1608]. 120 [165]. 150 [1511, 1920]. 2 [100, 1011, 1634, 22, 2070, 1269, 1082, 1682]. 2 + 1 [2170, 242, 396, 933, 467, 64, 1348, 2171, 1012, 1172]. 2, 3, 5 [640]. 27 [36]. 2 ⊃ ⊗2 ⊃ ⊗d [1645]. 2 × 2 [1638, 2079]. 3 [1357, 2187, 1875, 2133, 1182, 603, 2160, 403, 450]. 3/2 [463]. 3 × 3 [732]. 4 [897]. 5 [897]. 50 [1863]. ^{∗} [1404]. ^{1}S_{3} [504]. _{1}F_{1}(a, b, z) [1526]. _{i} [250]. A [1261, 1526]. A_{2} [1900]. A_{2}^{(1)} [2182]. A_{2n}^{(2)} [102]. AdS(3)/CFT(2) [1831]. AdS/CFT [1624]. AdS_{2} [1254]. AdS_{d+1} → AdS_{d} [372]. α [1069]. α^{2} [253]. Aut(F_{4}) [596]. AX − XA = C [2002]. b [1526]. B(m, n) [1783]. bc [2105]. β [658]. BF [1387]. BF CG [1387]. δ̆^{o}(A): \underline{Σ} → \underline{R}^{⪰} [1489]. C [1474, 1509, 369]. c = 26 [1275]. C^{∗} [1213, 1413, 1789, 1761, 1641, 795, 1525, 565]. C^{0} [1427]. C^{n} [1580]. D [1130, 968, 412, 749, 608, 1204, 1565, 787, 2148, 549, 1815, 899].

D(2, 1; α) [1217]. D < 2 [917]. d = 1 [188]. D = 2 + 1 [472]. ẍ + 3xẋ + x^{3} = 0 [1969]. ẍ + f(x)˙(x) + g(x) = 0 [2044]. ẍ + f(x)ẋ + g(x) = 0 [1983]. δ [135]. Δ(3n^{2}) [1094]. Δ(6n^{2}) [1746]. E_{6} [36, 1312, 294, 1220]. E_{7} [392, 1220]. E_{8} [803, 1098]. E_{τ,η}(A_{2}^{(2)}) [1299]. ϵ [250]. F [462]. F_{L}(η, ρ) [572]. G(2) [431]. G_{2} [337, 345, 389]. γ [1747]. g ≥ 2 [1786]. GL(2, F_{q}) [1297]. gl(2|2) [35]. GL(M, C) [1520]. gl(m|n)_{k} [1038]. GL_{q,j}(1|1) [1364]. H [689, 387]. H^{∗}(T^{2}, R) [1238]. H^{1/2} [1718]. H^{2} [180, 441, 440, 2070]. H_{n} [280]. R̂(θ) [257]. ŝl_{n} [1026]. ŝu_{n} [1026]. ô_{N} [842]. ISL(n, R) [252]. j [1258, 1521]. K [2040, 1718, 219, 269, 455, 181]. K(m, n) [2179]. κ [1006, 1191, 635, 2041, 1925, 1128]. KP [1966]. L [258]. L^{2} [68, 947]. L^{p} [338, 1630, 1246]. L_{2} [463]. L_{p} [728, 109, 212]. Λ [1649]. λϕ^{4} [19]. ln tan [1603]. M + λD [1664]. M/W [1135]. C^{n}/Z_{m} [1765]. CP^{1} [2029]. CP^{n} [2163]. R [1537]. R^{1+n} [1153]. R^{2+n} [668, 488]. R^{2D} [474]. R^{3} [1279, 1376, 1652, 382, 912]. R^{4} [72]. s_{m} [955]. Z [2071]. Z_{2} [875]. ⌉(∈) [625]. }↕(∞|\) [1545]. ∫↕_{∈} [380]. ∫≀(∋) [536]. ∫≀(▽) [536]. ∫⊓(∋) [1549]. μ [571, 538]. N [854, 1349, 1875, 764, 1910, 788, 69, 1590, 1540, 376, 873, 413, 1144, 1033, 1896, 1995, 216, 23, 1665, 1064, 1696, 1290, 365, 609, 1408, 2086, 2054]. N + 1 [376]. N = 1 [404, 1029]. N = 2 [983, 1968, 1708, 806]. N = 3 [471]. N = 6 [1976, 391, 501]. NK [1149, 1857]. ν [1580]. O(1) [1621]. o(4, C) [1178]. O(5) [204]. O(p, q) [1893]. O(ε^{2}) [676]. O^{⋆} [1494]. O_{2} [1439, 1709, 295, 1636, 1805, 1902]. O_{∞} [1185, 1805]. OSp_{q}(1/2) [397, 855]. p [1321, 1977, 1850, 109, 212, 479]. p, q [1476]. Φ^{4} [1447, 755, 871]. ϕ^{6} [1517]. P T [367, 63, 221, 2147, 253]. Q [855, 1150, 985, 258, 1198, 1926, 493, 269, 923, 809, 715, 466, 231]. q > 1 [1150]. Q^{n} [1705]. R [1328, 1363, 1220, 134, 1956, 1514, 564, 1267, 924, 1808]. R^{2} [2070]. R^{4} [275]. R^{N^{2}−1} [1266]. S [842, 627, 602, 1971, 1144]. S^{1|2} [1217]. S^{2} [180, 441, 1254, 440]. S^{2} → CP^{N−1} [1266]. S^{3} × R [27]. S_{3} [471]. S_{n} [2022]. σ [278, 1852]. σ(M) [1664]. sl(2) [986, 1182]. SL(2, F_{q}) [1297]. sl(2, R) [926]. sl(3) [986]. sl(4, C) [1178]. sl(p^{2}, C) [497]. SO(10) [92]. so(2, 1) [798, 6, 104]. SO(3) [988]. SO_{0}(1, d + 1) [566]. Sp(1) [1948]. sp(4, C) [1178]. \sqrt{−Δ + m^{2}} − m [559]. SU(1, 1) [408, 1667, 1867]. SU(2) [371, 1958, 292, 1148, 1811, 1198]. SU(2) × SU(2) [204]. SU(2^{N}) [310]. SU(3) [1884, 814, 266, 988, 431]. SU(3)/SU(2) [814]. SU(N) [599, 1854, 101, 374, 399]. su_{q}(1, 1) [759]. SU_{q}(3) [698]. T [855, 1380, 1848, 1896]. T^{∗}T^{n} [681]. T^{2} [1133]. t_{1x} = t + x + d(t, t_{1}) [2053]. τ [734, 1751, 1026, 1966]. U(1) [917]. U(2Ω) ⊃ U(Ω) ⊃ ⊗SU(2) [1022]. U(3) [2046, 600]. U(8) ⊃ O(8) ⊃ SU(3) [663]. U(N) [1557, 141]. U_{q}(D_{4}^{(3)}) [984]. U_{q}(gl(1|1)) [485]. U_{q}(so_{3}) [466]. U_{q}(so_{5}) [466]. U_{q,p}( ˆ∫↕_{2})_{k} [1451]. U_{q}[osp(m|n)] [459]. U_{q}su_{(n,n)} [231]. ϑ [665]. W [771, 602, 1555, 1659, 1856, 2032]. W (2, 2) [1659]. W (2, 2p − 1) [1096]. W/M [812, 1790]. W_{1+∞} [1304]. XY [1678]. Y (}↕(\)) [1740]. y^{2} = x^{5} − x [2082]. Z [175]. Z^{3} [1364]. |x| + |z| [247].

-1 [1850]. -2-Hahn [1926]. -algebra [1476, 258, 1659, 565]. -algebraic [1213]. -algebras [1665, 2133, 1404, 1474, 1494, 771, 1761]. -analog [985]. -analysis [923]. -ary [1995]. -Barut [1867]. -body [2054, 1875, 1910, 413]. -brane [749]. -branes [479, 1130]. -compact [947]. -component [269]. -cosympletic [455]. -cycle [2160]. -deformed [571, 1198, 269, 1925, 466, 538]. -dependent [1191]. -designs [1082]. -differential [231]. -dimensional [1896, 2148, 888, 23, 1850, 1222, 1064, 1242, 753, 640, 822, 1547, 1101, 608, 764, 1565, 787]. -dynamo [253]. -electron [219]. -entropies [1150]. -equivalence [1408]. -expansion [1741, 1971, 1724]. -extended [1856]. -flat [369]. -functional [1026]. -functions [734, 665]. -graded [1364, 2071]. -group [1977]. -Hermite [658]. -integrable [1144]. -invariant [1893]. -Jacobi [855]. -Kauffman [1149, 1857]. -KdV [899, 549]. -Kepler [1621, 1948]. -Laguerre [658]. -level [2086, 1682]. [?]LiebRybkin:2005:SCL. -manifolds [376]. -matrices [1956, 564, 1267, 924, 1808, 627, 1363, 1220, 134]. -matrix [1514, 855]. -metrics [1204]. -Minkowski [1006, 635, 2041]. -model [1852]. -models [759]. -modules [1413, 1641]. -norm [1269]. -norms [1321, 109, 212]. -orthogonal [466, 1815]. -oscillator [493, 1476]. -plurality [1380, 1848]. -point [897, 1290, 609]. -property [250]. -quaternions [1198]. -qubit [216, 365]. -separable [1328]. -shape [1135]. -shape-peaks [1790, 812]. -soliton [854, 788]. -space [1357, 2070]. -spectral [68]. -sphere [1357]. -spheres [403]. -stars [809]. -state [1590]. -structures [1811]. -subalgebras [1789]. -surfaces [1011]. -symbols [1521]. -symmetric [253, 917, 367, 221, 2147]. -symplectic [181]. -system [2105]. -ternary [795, 1525]. -theorem [689, 387]. -topology [1896]. -type [135]. -uniform [1630]. -vector [941]. -vortex [1349]. -wave [1144].

16th [1690].

43 [211, 2208, 37]. 44 [1330, 934]. 45 [1609, 263, 38, 104]. 46 [2206, 441, 302, 2212, 212, 568, 2095, 301, 467, 347, 1117, 501, 440]. 47 [628, 630, 860, 2141, 1236, 900, 668, 1790, 1116, 1227, 2213, 1118, 629, 1157, 899]. 48 [2216, 1498, 2196, 1499, 1415, 987, 2215, 1365, 1608, 1329, 1414, 961]. 49 [1588, 2195, 1566, 1456, 2214, 1758, 1834, 1940, 2202].

50 [2201, 2197, 2203, 2210, 2207, 2096, 2209, 2205, 2217, 2211, 2204, 2198, 2097, 1907, 2200, 1941].

80 [2199].

ABACUS [2028]. Abelian [513, 272, 173, 1871, 1392, 2182, 748, 856, 1463, 1303, 545, 477, 450, 667, 2151]. Ablowitz [2171, 1523]. Absence [632, 403]. Absolute [1761, 1031]. Absolutely [1662]. absorption [1371, 1544]. Abstract [1890, 126, 1985]. accelerated [379]. acceleration [1325, 525]. acceleration-dependent [525]. accuracy [2035]. acoustic [1901, 1728]. acoustics [653, 587]. acting [1300, 1378]. action [2039, 1426, 662, 948, 2038, 1658, 143, 553, 2217, 72, 365, 1834, 450, 621]. action-angle [948]. actionlike [1493]. actions [1424, 686, 1623]. active [989]. activity [1511]. actualization [2089]. added [1759]. Addendum [467, 1414]. addition [1123]. Additivity [1617, 1073, 563]. ADHMN [923]. Adiabatic [2035, 186, 1196, 265, 743]. Adiabatically [1616]. adjacent [583]. adjoint [1599, 852, 897, 1252]. adjointness [770, 1234, 1553]. Adler [905, 1441]. admitting [2085]. AdS [326]. advection [1067]. aeroengine [1057]. affgebroids [1156]. affine [736, 2158, 1527, 910]. against [1857]. Aharonov [186, 1015, 2150, 1421, 573, 1335, 999, 680]. AKNS [94]. al [2198]. al. [2150]. algebra [1821, 1439, 1432, 798, 389, 1895, 276, 1364, 1424, 258, 697, 1451, 197, 1481, 625, 1977, 1177, 1667, 2029, 1606, 495, 1998, 698, 932, 1900, 1971, 1773, 535, 295, 1185, 1636, 1902, 101, 1022, 344, 1659, 1316, 609, 834, 1304, 392, 294, 534, 391, 501, 1865, 1476, 1624, 565, 1815, 6, 104, 638]. Algebraic [46, 162, 1299, 1074, 1213, 238, 1136, 1604, 283, 1949, 1782, 1003, 1737, 1264, 1389, 651, 252, 924, 1808, 1924]. Algebras [1329, 736, 1096, 955, 1474, 1995, 2131, 1494, 624, 771, 2099, 848, 797, 1132, 1524, 1725, 2193, 1572, 1261, 289, 1527, 706, 794, 194, 1816, 1025, 1550, 1665, 2133, 498, 1297, 983, 1761, 1182, 1931, 2104, 439, 1900, 1845, 1404, 1555, 856, 339, 1250, 1363, 228, 1484, 620, 1220, 881, 1693, 2016, 2076, 1027, 122, 341, 859, 256, 1904, 1754, 432, 795, 2198, 2136, 1237, 1336, 1525, 813, 466, 1830, 667, 1267, 88, 979, 1477, 2189, 77, 1401, 288, 741, 638]. algebro [1747, 80]. algebro-geometric [1747, 80]. algebroids [761]. algorithm [1817, 2028, 310, 398, 1827]. Algorithmic [464]. aligned [807]. alloy [1164]. alloy-type [1164]. almost [1173, 152, 274, 1396, 1950, 931, 1293]. almost-commutative [1396, 1950, 931]. almost-symplectic [1173]. alphabetic [2189]. Alternative [108]. alternatives [940]. ambiguity [1463]. among [737]. amplitude [2109, 959, 1285, 1561, 154, 747]. amplitudes [1009, 1506, 2146, 223, 1723]. analog [985, 1507, 1521, 824, 1304, 671, 1575]. analogical [1687]. analogs [1607]. Analysis [2148, 2059, 9, 197, 1365, 1578, 2122, 517, 571, 1109, 196, 786, 2109, 1516, 1676, 508, 245, 649, 68, 164, 579, 914, 1100, 29, 1154, 1058, 2139, 1412, 1319, 1187, 923, 716, 26, 644, 1915, 8, 38, 538, 774, 746, 1157, 1101, 1044]. Analytic [705, 445, 158, 378, 408, 639, 1626, 996, 220, 1226, 980, 568, 120, 960, 83]. Analytical [1860, 343, 1054, 549, 2160, 1069, 1653, 1758, 899, 1988, 1393, 520, 1447, 711, 155]. Analyticity [1087, 747, 2159, 1290]. analyzing [1821]. Anatomy [1171]. and/or [448]. Anderson [973, 963]. angle [948]. angles [599, 337]. Angular [2116, 24, 1190, 606]. anharmonic [1661, 965, 42]. Anisotropic [1649, 1914, 282, 1594, 2000, 1785, 1581, 1840]. annealing [1680]. annihilation [895, 765]. annihilation-creation [765]. Announcement [1691, 1759, 1692, 1305, 1610, 1643, 1690, 2142]. annular [1147]. Anomalous [1990, 2184, 1541]. Anomaly [710, 642]. Anomaly-free [710]. Anosov [1087]. ansatz [2026, 1262, 626, 1299, 924, 1808, 140]. Anti [1179, 1552, 1454, 752, 1504]. anti-commuting [1504]. anti-de [752]. anti-Poisson [1454]. antibracket [1532]. anticommutation [1237]. antifield [400, 1409]. antimonopole [313]. antiparticle [517]. antisymmetric [1698]. antisymplectic [1964, 857]. antithesis [7, 1609]. any [1526, 1232, 480, 1128, 915]. Anyons [561]. aperture [685]. appearances [208]. appearing [2091]. Appell [1296]. Application [719, 367, 1040, 980, 1675, 1025, 1658, 80, 684, 839, 388, 2126, 91, 850, 1331, 67]. Applications [1934, 1460, 1324, 1226, 1194, 468, 87, 1817, 703, 1992, 1196, 2132, 1319, 1405, 2167, 1890, 1224, 1826, 2189, 1401, 774, 2118, 1579, 2194]. applied [167, 559]. Approach [1667, 1213, 1194, 791, 233, 2092, 1518, 464, 1349, 537, 1438, 28, 1949, 2026, 1881, 811, 1823, 120, 1544, 707, 228, 1887, 910, 336, 187, 1231, 1433, 181, 370, 98, 446, 746, 1157]. approaches [1845, 121, 1441]. Approximate [1600, 2198, 1525, 919, 1830, 58, 849, 1128]. approximately [306]. Approximating [1372, 2027]. Approximation [1628, 490, 1619, 1607, 941, 1105, 1578, 443, 1196, 2035, 393, 963, 415, 2143, 1803, 1495, 1205]. approximations [356]. arbitrary [1847, 1146, 1748, 482, 629, 2161, 699, 150]. Areal [1133]. arguments [1748, 882]. arising [1375, 2008, 1579, 1603, 1788, 1636, 1882, 2200]. Arnol’d [2057]. array [954]. arrival [829, 1331, 1333]. arrows [1489]. ary [1995]. Askey [262]. Aspects [258, 1669, 1593, 1079, 159]. associated [1032, 841, 1375, 955, 186, 1400, 1942, 790, 333, 1641, 1900, 2071, 510, 923, 147, 534, 870, 88, 429, 979, 1864, 2077, 984, 1928]. association [1542]. asymmetric [189, 1634, 138, 1582, 974, 2018]. Asymptotic [2109, 1137, 641, 1214, 526, 385, 68, 1611, 51, 1058, 1573, 1533, 1175, 2137, 1581, 1954, 2050, 2186, 1837, 131, 237, 896, 1787, 2019, 1546, 1483, 1940]. asymptotically [515, 1133, 515]. Asymptotics [230, 816, 1195, 156, 1938, 57, 168]. Atlas [1139]. atmospheric [283, 730]. atom [125, 508, 543, 271, 1039, 763, 507]. atomic [1501, 1367, 1613, 817, 6, 104, 91]. Atomistic [1477]. atoms [1314, 264, 540, 1167, 175, 6, 104]. attachment [1245]. attractive [2031]. attractor [1049, 1982, 1630, 1800]. Attractors [2049, 682, 730, 618, 947]. attribution [1681]. Auger [484]. August [1690]. author [1691]. authorship [1681]. automata [1934, 1647]. automaton [1511, 1920, 517, 984]. automorphic [1580]. Automorphisms [773, 1815, 564]. autonomous [1638]. auxiliary [1792]. available [1691]. average [1854, 1818, 1941]. averaged [1980]. averages [896]. averaging [125, 1360]. axial [656, 1227]. axial-to-planar [656, 1227]. axioms [1906]. axisymmetric [1343, 1053].

B [1252]. B-quadrilateral [1252]. background [417, 1703, 1166, 1498]. backgrounds [2084]. Bäcklund [1222, 1729]. backscattering [1160]. Baker [1817]. balance [884, 584]. Balanced [1811]. Balian [820]. ball [52, 626, 120]. Banach [2198, 1830]. band [308, 633]. bandwidth [1717].

Bardeen [905, 1808]. bare [124]. Bargmann [571, 538, 640]. barotropic [2165, 587, 1072]. Barrett [516]. barrier [352, 2030]. barriers [1961, 1235, 699]. bars [1121]. Barut [1415, 1233, 1867]. Baryon [907]. baryons [698]. base [1023, 984]. based [1496, 470, 1773, 408, 2107, 1288, 1682, 1266]. bases [44, 993, 1794, 820, 439, 1082, 1699]. Basic [389, 801, 2058, 429, 1169, 2118]. basis [2020, 904, 817, 536]. Bavand [2198]. Baxter [794, 2003, 1852, 724]. Baxterisation [929]. Bayesian [824]. be [7, 1609, 1332]. beam [1131, 1933]. beams [1281]. bearing [1339]. Becchi [1132, 1555]. behavior [19, 1214, 1133, 1054, 1286, 1984, 234, 2137, 175, 1807, 873, 507, 879]. Beilinson [2151]. being [864]. Bell [1973, 796, 336, 224, 1306, 1083]. Bell-type [796]. Beltrami [323]. belts [2124]. Bénard [84]. bending [1785]. Benenti [552, 1001]. Benjamin [1718, 788]. Benney [1114]. Berezin [1028, 1528]. Bernoulli [268]. Berry [2003, 265, 1864]. Bertotti [1466]. Bessel [1109, 2181, 1326, 149]. Best [215]. Bethe [2026, 626, 1299, 140, 924, 1808]. between [1321, 468, 196, 1853, 1792, 1671, 253, 1793, 626, 347, 335, 121, 2152, 795, 437, 1525, 1441, 2103, 1278, 1023, 1928, 1903]. beyond [1995, 597, 1801, 1306]. BGK [1246]. Bi [1593, 1209, 1962, 1828, 533, 1747]. Bi-Hamiltonian [1593, 1962, 1828, 533, 1747]. Bi-invariant [1209]. bialgebra [1998]. Bianchi [1386, 773, 581, 284, 520, 1904]. bias [1616]. bidimensional [695]. bidimensionally [854]. Bifurcation [29, 359, 877, 1071, 358]. Bifurcations [210]. Bilinear [1411, 1633]. binary [1875, 1149]. binding [444]. Biorthogonal [1163, 1165]. Biorthonormal [1654]. bipartite [1417, 442, 708, 826]. bipartitite [215]. Birational [1720]. Birkhoff [681, 659, 281]. Birman [559]. birth [2072]. bisection [1689]. bistable [1656, 871]. bistochastic [1683]. bisymplectic [320]. bivaluations [509]. bivectors [630, 551]. BKP [777, 1265]. black [1134, 1703, 1426, 1871, 1202, 1768, 768, 1429, 1170, 479, 579, 1574, 2217, 1979, 1834, 1169, 808]. blackhole [378]. Blasius [568, 83]. blob [1111]. Bloch [145, 1412]. Block [660, 1794, 839]. Block-circulant [1794]. Block-diagonalizability [660]. blocks [1794]. Blow [1778, 296, 1410, 1043, 1836, 489, 2070, 1780, 1715, 1652, 1502, 1053, 1660]. Blow-up [1778, 489, 2070, 1780, 1715, 1652, 1502, 1053, 1660]. Blowup [1997]. bodies [885, 930, 1224]. body [351, 384, 959, 776, 519, 402, 1651, 1501, 838, 878, 1875, 157, 1984, 1910, 2035, 1855, 930, 1224, 1723, 763, 413, 1034, 2054]. Bogoliubov [901, 121]. Bogolubov [2031]. Bogomolnyi [2096, 1937]. Bohm [186, 1015, 2150, 1421, 573, 1335, 999, 680]. Bohr [508, 1367]. Boltzmann [1633, 783, 247, 190, 1243, 2062, 721, 1093, 619]. bombardments [1066]. Bondi [51]. Boolean [2075, 1857]. Borcherds [920]. Borel [240, 682, 1447]. Born [1574, 127, 386, 379, 229, 942, 393]. Bose [1213, 87, 866, 1959, 1908, 965, 1188, 1563, 1773, 1356, 2033, 1187, 227, 544, 324, 844]. boson [1733, 1615, 2195, 1185, 336, 1437, 1772]. Bosonic [1781, 1362, 277, 722, 895, 16, 166, 226, 1923, 106, 1795]. bosons [1825, 1884, 1009, 1636, 1902, 72, 782]. bottom [1568]. Bound [1039, 798, 460, 916, 1283, 469, 407, 45, 1306, 2097]. boundaries [478, 353, 1879, 2011, 5, 263]. Boundary [2131, 329, 1802, 1993, 1380, 226, 986, 886, 1515, 1206, 568, 891, 2129, 674, 900, 1285, 132, 378, 33, 1545, 1445, 123, 1067, 83, 1483, 1940, 1728, 666, 57, 312, 85, 2062, 690, 721, 1093, 2012, 922, 2050]. boundary-boundary [85]. Boundary-layer [1802]. Bounded [895, 351, 1126, 797, 1337, 2120, 2149, 323, 1298, 1564, 2007, 1806]. Boundedness [283]. Bounds [1196, 309, 107, 363, 843, 1049, 1181, 510, 904, 1472, 747, 2097, 1907, 1796]. Boussinesq [1586, 1071, 1600, 1783]. Boussinesq-like [1783]. Bowen [1087]. box [768, 1125, 52, 626]. box-ball [52, 626]. BPS [1029, 1573]. bra [506]. bra-ket [506]. brackets [1352, 1492, 523, 790, 133, 491, 858]. braid [1218, 1032, 258, 257]. braided [2124, 1262]. braids [2124]. branch [1875]. branched [1161]. branches [1615, 2195, 1664]. Branching [1439]. brane [749, 1648, 1538]. brane-induced [1648]. branes [1130, 1086, 1591, 479, 1574, 1084, 1765, 317]. Brans [1574]. break [1520]. Breakdown [381, 349]. breaking [971, 227, 544, 1075]. Breather [1727, 2015]. breathers [2114]. Bregman [2199, 2119]. Bright [1563]. Bright-matter [1563]. broken [729]. brought [267]. Brownian [1918, 1656, 90, 1175, 1654, 1104]. BRST [207, 371, 251, 273, 597]. Bryant [1548]. Bryuno [682]. building [36, 92]. built [623]. Bulk [1938]. bumps [1714]. bundle [417, 1257, 265]. bundles [1732, 1041]. Bures [862, 1996]. Burgers [155, 146, 1905]. Burnett [550, 1486]. Burnside [2082]. Büttiker [110, 1616, 952].

caged [1594]. Calabi [1130, 1381, 1763]. Calculating [1542]. calculation [1633, 62, 1773, 727, 639, 572, 91]. calculations [1368]. calculi [723, 1107]. Calculus [1745, 791, 530, 950, 198, 1481, 2156, 823, 1247]. Caldirola [1711]. Calogero [1118, 1956, 93, 502, 1855, 392, 294, 480, 762, 241, 471]. calorons [1142]. Camassa [136, 61, 1747, 1715, 1410, 1708, 489, 735, 462]. Campbell [1817]. can [1332]. Canonical [862, 1914, 770, 1180, 1749, 266, 216, 2061, 994, 511, 420, 1074, 2136, 1237, 146]. Cantor [521]. capability [1759]. capacity [1159, 1945]. Caputo [1584]. CAR [295, 95]. Carlo [398]. carrying [1505]. Cartan [1732, 2193, 723, 59, 310, 251, 244, 1888, 288]. Cartesian [111]. cartography [773]. case [902, 816, 964, 1522, 1281, 773, 1794, 1960, 1170, 21, 467, 64, 1066, 1441, 2151, 1034, 107, 570]. cases [987, 887, 215, 1983]. Casimir [48, 47, 459, 547, 1127, 1199, 364, 252]. Categories [1041, 986, 1490]. category [404]. Cauchy [2008, 897, 171, 1212, 577, 1531, 616, 434, 86, 1114, 1663, 1246]. Causal [453, 233, 270, 454, 679, 945, 1951]. causality [747]. caustic [1799]. caustics [1978]. cavity [48, 47, 1676]. Cayley [882, 297]. cells [583]. cellular [1511, 1920, 1647, 517, 984]. center [1628]. centers [1841]. Central [1143, 180, 441, 1469, 1683, 440, 2089, 76, 1532, 787, 2054]. central-extended [2089]. centralizer [1220]. centralizers [929]. Certain [562, 1283, 1069, 2093, 1865, 1411, 1476, 1278, 1467, 1862]. chain [842, 1596, 1719, 1283, 1377, 2023, 1779]. chains [790, 1020, 1534, 1627, 2053, 2163, 1545, 858, 564, 1267]. Chandra [1316]. Chandrasekhar [24]. Changes [267, 1857]. changing [547]. channel [1159]. channeling [656, 1227]. channels [864, 218, 1614, 109, 212, 469, 40, 1269, 1945, 305, 1617]. chaos [716]. Chaotic [2164, 1984, 1348, 381, 684, 839]. Character [643, 1528, 872, 1562, 1221]. characteristic [1862]. characteristics [994, 1928]. Characterization [1461, 914, 2128, 634, 779, 1791, 807, 1844]. characters [1691, 392]. Charge [74, 1432, 1132, 269, 1505]. Charged [1131, 353, 573, 1984]. charges [749, 160, 2115]. Charging [178]. Charney [342]. Chern [770, 124, 803, 117, 1952, 20, 1214, 925, 451, 2011, 450, 1889]. Chernoff [1283]. Chetaev [950]. Chinese [1691]. chiral [174, 1327, 771, 242, 1029, 1720, 1075]. chord [217, 1023]. Christoffel [1773]. chromodynamics [1535, 1425]. Chua [1626]. circuits [1008]. circulant [1794, 839]. circular [1393, 1922, 2116, 604]. circulation [283]. CKP [1265, 1605]. Clarification [1957]. Class [1608, 1584, 842, 1568, 1087, 1143, 1474, 2186, 272, 1454, 1137, 552, 1241, 1468, 130, 258, 1917, 139, 1036, 1100, 1776, 120, 676, 2048, 1115, 2093, 1053, 26, 762, 1118, 1872, 1707, 1862, 1482, 1397, 1800]. classes [955]. Classical [1620, 351, 1132, 1992, 1705, 1329, 430, 1612, 1061, 1015, 971, 1317, 589, 103, 2164, 28, 1512, 56, 1320, 1631, 1880, 1208, 199, 200, 394, 603, 896, 881, 655, 1945, 181, 480, 1278, 446, 564, 1808, 88, 429, 763, 1369, 2069, 1401, 2178, 992, 1174]. classical-quantum [1945]. Classification [893, 1499, 1847, 790, 983, 1659, 1590, 655, 979, 1370, 1328, 695, 585, 1627, 1776, 152, 274, 1396, 1950, 603, 1979, 184, 1904, 858, 88, 429]. Clausen [1448]. Clean [311]. Clebsch [36, 663, 392, 2118]. Clifford [163, 1307, 848, 1895, 803, 1647, 2189, 761, 853]. clique [591]. clockwork [1367, 1613]. cloning [304, 2036]. Closed [2091, 982, 664, 236, 535, 71, 319]. Closed-form [2091, 982, 71]. closest [1030]. closures [1067]. cluster [1655, 2116, 1890, 817, 185, 413]. clusters [55, 286, 1855]. co [1732]. coarse [369]. coat [1291]. cocycles [1641]. codes [1786, 992, 1827]. Codeword [1827]. codimension [1978, 813]. codimensions [704]. coefficient [683, 1100, 1536, 2126]. coefficients [36, 550, 1486, 115, 566, 663, 1026, 1123]. coframe [21]. coherence [222, 953]. Coherency [1867]. Coherent [700, 1849, 1, 211, 1298, 2208, 208, 722, 1375, 980, 938, 1189, 1233, 1334, 1415, 1456, 1180, 941, 847, 443, 512, 408, 1064, 269, 1348, 1066, 361, 1230, 114, 509, 870, 1160, 447, 536, 828, 1274]. coherent-state [938, 1189]. cohomology [90, 2151]. coin [2148]. Coleman [698]. collapse [1429, 1702, 1315]. Collective [1875, 40]. Colliding [479]. Collinear [2054]. collineations [1201]. collision [846]. collisions [1875, 701]. Colombeau [1631]. color [1692, 277]. color-flavor [277]. colored [2061]. coloring [1158]. column [18]. Combinatorial [336, 87, 166, 2006, 446, 2024]. Combinatorics [609]. Comment [2201, 2216, 1498, 441, 2196, 2197, 1415, 1236, 1831, 2209, 568, 2095, 2217, 1790, 1116, 2198, 2208, 440, 2202, 1157, 899, 2097, 1907]. Comments [1456, 2217, 1834]. communication [446]. Commutation [259, 2136]. commutative [864, 634, 2029, 152, 274, 1396, 1950, 931, 822]. commutators [230]. commute [265]. Commuting [179, 341, 1345, 1504, 1928]. Compact [2040, 1786, 1522, 289, 1463, 1046, 592, 947]. compactification [131, 1099]. compactifications [151]. compactified [19, 1993]. comparative [1964]. Comparing [299, 1457]. Compatible [491]. compendious [1697]. competing [882]. Complementary [867, 157]. Complete [245, 2053, 2147, 163, 1947, 306, 79, 1076, 35, 1980]. complete-positive [163]. completely [948, 1761, 1378]. Completeness [2020, 2079, 1787, 1339]. completion [679]. Complex [1112, 633, 1861, 400, 630, 551, 893, 1499, 1939, 1248, 68, 1041, 1264, 1982, 1976, 1700, 1319, 347, 335, 1623, 1082, 970, 1249]. complexified [2001]. complexity [1895, 2194]. component [649, 648, 2139, 1455, 1042, 269, 2033, 390, 735, 1899, 2180]. composition [852]. compound [119]. compressible [2059, 1771, 1455, 1353, 2174]. compression [1449]. Computation [2061, 847, 1196, 2035, 159, 39, 1318, 2142]. computational [1895]. computations [1052]. Computing [1709, 1817, 1642]. concentrated [1059]. concentration [1684]. concentric [604]. concomitants [630, 551]. Concurrence [1193]. concurring [1797]. condensate [1214, 1563, 1356, 227, 544]. condensates [965, 1188, 2033]. condensation [225, 1187, 324]. condensed [1449]. condition [682, 674, 900, 33, 2012]. Conditional [1222, 147, 1822, 58]. Conditioning [1842]. Conditions [109, 212, 1380, 1516, 226, 1127, 23, 160, 891, 1545, 123, 1510, 2152, 1067, 32, 1728, 666, 2, 37, 312, 690, 922]. conduction [951, 1775, 1446]. conductivity [1935]. conductors [2000]. cone [2021]. Cones [1911]. configuration [797, 1390, 1228, 165]. Configurational [182]. configurations [1469, 2054]. Confined [1760, 2100, 1284]. confinement [355, 1535]. confluent [1526, 1585, 2177, 2207, 428]. Conformal [128, 2092, 602, 646, 986, 643, 23, 1251, 2215, 1277, 1099, 2107, 2106, 1005, 1027, 227, 544, 1754]. conformally [49, 130, 199, 200, 394, 603, 758, 1300]. confusion [805]. Congress [1690]. conical [48]. conjecture [1068, 456, 1310, 2093, 1169, 463]. conjugacy [955]. conjugate [2183]. Conley [850]. connected [1655]. Connection [1278, 1515, 699, 1853]. connections [2157, 2188, 369, 461, 455, 1857]. connectivity [1245]. Connes [276]. consequence [316, 913, 1480]. consequences [2089]. Conservation [943, 1047, 404, 877, 299, 1989, 1044, 1365, 179, 1105, 1640, 142]. conservative [981, 1139]. conserved [52, 626, 2162, 1877]. considerations [1457]. Consistent [78, 28, 1891, 1601, 1026, 1368]. constant [145, 180, 441, 1390, 1713, 1133, 1935, 632, 1286, 504, 1993, 1587, 1336, 440, 2077, 966, 2068]. constants [195, 751, 1037, 91]. Constitutive [2000]. Constrained [418, 950, 586, 28, 1987, 1265]. constraint [1686]. Constraints [442, 734, 272, 315, 1454, 1018, 670, 1044, 1365, 1742, 1138, 1065, 422, 720, 1433]. Constructing [148, 1241, 323]. Construction [1032, 327, 114, 817, 462, 29, 1076, 923, 35]. constructions [1073]. constructive [310]. contact [1903]. Contextual [10, 223]. Continuity [363, 1460, 1761]. Continuous [238, 205, 301, 1685, 1552, 518, 628, 1078, 1926, 1577, 448, 1694, 1662, 403, 390, 204, 431, 601, 1120, 754]. continuous-time [1078, 1120]. continuum [16, 815, 945, 369, 187, 367, 2056, 1016]. contour [408, 882]. contours [188]. Contraction [729, 205, 301]. Contractions [859, 1401, 278, 1527]. Contractivity [728]. contrast [1486]. contributions [1448]. Control [159, 67, 1338, 684, 839, 158, 2025, 716, 39, 447, 1259, 743]. controllability [158]. controlled [1232, 657]. controlled-NOT [1232]. convection [1058, 84, 142]. convection-diffusion [1058]. conventional [2073]. Convergence [1161, 1558, 2174, 1321, 1655, 1652, 1755, 904, 248]. converging [1516]. conversion [272, 1454]. convex [1280, 2093]. convexity [1556]. convolution [1291]. convolutions [1646]. Cooper [1808]. coordinates [1866, 1191, 860, 647, 948, 994]. core [701, 1019]. correction [1913, 992, 40]. corrections [385, 1152, 1288]. Correlated [1416, 1283, 1377]. Correlation [2028, 1471, 1991, 2090, 801, 1644, 1290, 560, 85, 485, 1803]. correlations [1244, 951, 917, 782, 185]. correspondence [1367, 1613, 626, 1074, 1441]. corresponding [257, 1179, 288]. coset [862]. cosine [204, 431]. cosmic [1239]. cosmological [1386, 1133, 1917, 649, 648, 1344, 1587, 1649, 8, 38, 1980]. cosmologies [773, 581, 318]. cost [1159]. cosympletic [455]. Couette [1802]. Coulomb [445, 1515, 261, 504, 2116, 1339, 904, 1231, 220, 1336, 1309, 572]. Coulomb-plus-linear [220]. counting [768, 1573]. Coupled [287, 1203, 816, 1778, 581, 1715, 1044, 1365, 1029, 1539, 755, 974, 1517, 320, 785, 818, 1072, 72, 1502, 1535, 1116, 1874, 67, 1666, 1846, 1769, 1495, 2126, 1423, 248, 1425]. Coupling [937, 1684, 1286, 861, 863, 632, 1993, 883, 1249]. couplings [94, 620, 150]. covariance [1960, 712]. Covariant [528, 511, 1016, 231, 482, 629, 864, 198, 1522, 745, 1458, 289, 12, 448, 1228, 1925, 655, 1459, 584]. Coxeter [1608, 1036, 1260]. CPT [328, 1167]. CR [726]. Crane [1183, 516]. Creating [2191]. creation [895, 765]. criteria [2088, 1780]. criterion [990, 1106, 422, 2180]. Critical [19, 1706, 76, 2098, 1993, 234]. criticality [1289]. cross [885, 249]. cross-section [249]. crossed [1131, 697]. Crum [827]. crystal [354, 439, 656, 1227, 984]. crystalline [409]. Crystallographic [932]. crystals [562, 185]. cubic [1939, 2166, 2074, 427, 1557, 1693, 2016, 1249]. cubic-quartic [2166]. Cummings [941, 443, 2086]. Cumulants [1821]. Cuntz [1439, 295, 339, 1185, 1636, 1902]. Curie [245]. curl [436, 1147, 1812]. Curlander [2097, 1907, 1796]. current [1860, 1518, 876, 801, 1008, 2018, 1038]. current-current [801]. currents [748, 583]. curvature [1011, 1409, 130, 180, 441, 1390, 1847, 1747, 56, 1737, 504, 440, 2077, 966, 2068]. curve [2082, 1726, 1376, 1592, 491]. Curved [2162, 2039, 2001, 2038, 1933, 876, 2105, 376]. curves [427, 2064]. cuspons [1790, 812, 1135]. cutoff [1744, 721, 1093]. cutoffs [1915]. cycle [2160]. cycles [1959, 1420, 844]. cyclic [2019, 1275]. cylinder [2017, 675, 1063, 15, 1023, 604]. cylinders [123]. cylindrical [436, 1147, 1168]. Cylindrically [1766]. Czech [1690].

D [1875, 603]. damped [1706, 2216, 946, 1880, 2055, 1877, 1482]. damping [1483, 1940, 1999, 1728, 1874]. Darboux [2043, 1627, 2053, 914, 340, 1773, 1389, 2048, 243, 1523, 102]. Daseinisation [1488]. data [732, 1292, 1880, 128, 176, 2134, 1663, 329]. Davey [495, 1323, 1637]. day [2217]. death [2072]. decay [2091, 1244, 22, 1738, 917, 2125, 1290, 702, 1999, 1728, 312, 185, 2069, 489, 1119, 1333]. decaying [1844]. Decoherence [1458, 249, 112]. Decomposition [448, 494, 310, 1257, 1250, 1027, 252, 368, 849, 1083]. Decompositions [708, 216, 1549, 1318]. Decoupling [581, 1420]. decreasing [1529]. decuplet [698]. defects [748, 1168]. defined [156]. definition [1268, 2060]. Deformation [1293, 2100, 1361, 2123, 903, 409, 2064, 161, 1587, 2105]. Deformations [2133, 1532, 1177, 1606, 941, 2121, 427, 2045, 2205, 694, 1904, 134, 1223]. Deformed [1091, 571, 1588, 1503, 1667, 1198, 1873, 93, 269, 1925, 2060, 466, 538]. Degasperis [1259, 1807]. Degenerate [870, 208, 2029, 1548, 969, 1457, 2190, 2211]. Degenerations [2099]. DeGennes [900, 674]. degradable [1614]. degree [617, 1310, 1146, 1818, 1941]. delay [1584, 1268, 1955, 1985, 2052]. Deligne [90, 2151]. Delone [1861]. Delta [1893, 1961, 1235, 709]. demonstration [1339]. Dense [1688]. densities [219, 1844]. density [541, 219, 779, 1820, 1771, 2037, 456, 526, 385, 916, 2204, 1590, 1975, 560, 1436, 548, 1988]. density-dependent [1771, 1988]. dependence [1181]. dependent [1991, 186, 1015, 351, 353, 1518, 1421, 1191, 617, 2196, 794, 1771, 23, 2030, 1335, 764, 158, 1146, 2125, 1993, 1531, 1108, 422, 720, 1336, 1002, 1366, 525, 1013, 1034, 1988]. depolarized [1269]. Dereziński [2130]. Derivation [1057, 229, 386, 328, 516, 1123, 1210]. derivations [2198, 1525, 1830]. derivative [2201, 2080, 2203, 1460, 1435]. Derivatives [1526, 1584, 786, 719, 1031, 2006, 1622, 977]. derived [1726, 1604, 1923, 1420, 1846]. Derrick [1801]. descent [1898]. describe [1864]. describing [201, 1174, 1822]. Description [333, 1300, 2216, 2028, 946, 355, 52, 383, 1016, 65, 517, 1549]. descriptions [108, 2024]. design [1394]. designs [991, 1082]. desired [2191]. detailed [884]. detection [1883]. Determinant [485]. determinantal [605]. determinants [2201, 1372, 2080, 2203, 156, 1247, 123, 2009].

Determination [2043, 1285, 1146, 617, 1082]. determine [975]. deterministic [669]. Development [219]. deviations [1283]. devices [1394]. diagonal [433, 1797, 1960]. diagonalizability [660]. diagonalization [162]. diagram [29]. Diagrammatic [141]. diagrams [279, 806, 1859]. Diamagnetic [733]. diatomic [936]. Dicke [1574, 2086]. dictionary [1130]. dielectric [287, 1394, 778]. dielectrics [483]. diffeomorphisms [513, 771, 486]. difference [1272, 1411, 1271, 91, 1446]. different [1853, 1838, 1122, 1149]. Differentiable [1951]. Differential [671, 1347, 1107, 1810, 1584, 530, 2186, 819, 737, 59, 1579, 1837, 957, 486, 34, 1736, 1782, 1986, 240, 975, 115, 1969, 2064, 1300, 144, 1985, 1892, 231, 1495, 2161, 1205, 1570, 1397, 1724]. differentials [251]. Differentiation [1399, 1044, 1365]. difficulty [26]. Diffraction [685, 97, 1235]. Diffusion [2143, 189, 529, 458, 780, 53, 2184, 767, 1613, 1990, 1222, 1058, 246, 331, 1541, 437, 142, 919, 2137, 1836, 338, 1630, 248]. diffusion-convection [142]. diffusions [1143, 1438, 57]. diffusive [2062]. dilogarithm [1448]. dilogarithmic [1788]. Dimension [1706, 1104, 1847, 1337, 1419, 1982, 2211, 167, 365, 1447, 1231, 1657]. Dimensional [1352, 692, 806, 1218, 1032, 1470, 1896, 841, 2148, 1375, 36, 2017, 178, 902, 145, 1913, 1063, 1342, 880, 1302, 1932, 1868, 180, 239, 441, 888, 1501, 1052, 1871, 1566, 1428, 1448, 1054, 1625, 1284, 706, 547, 585, 1080, 608, 1346, 1771, 643, 2088, 625, 1780, 748, 23, 1270, 1897, 1550, 362, 1055, 1248, 1491, 1850, 291, 445, 1427, 1986, 546, 1050, 2156, 2049, 764, 933, 830, 1937, 2096, 479, 157, 418, 522, 51, 1354, 1888, 1222, 2129, 1873, 1041, 199, 200, 758, 1264, 969, 896, 1099, 358, 1565, 1557, 1064, 235, 576, 1601, 72, 954, 1587]. dimensional [511, 2167, 859, 1322, 772, 1904, 11, 89, 271, 1242, 1775, 67, 813, 906, 2171, 376, 787, 165, 1846, 2025, 440, 1012, 85, 642, 134, 753, 1378, 664, 873, 1707, 640, 822, 1779, 966, 1547, 1101, 1139, 1520, 35, 879, 102, 473, 268, 653, 239, 394, 1356]. dimensionality [267]. dimensionally [23]. dimensions [1218, 172, 1762, 1307, 624, 1634, 1203, 257, 2170, 1819, 515, 1566, 1428, 986, 2123, 704, 396, 1190, 756, 330, 2095, 543, 467, 64, 1946, 917, 2190, 255, 1348, 939, 1045, 996, 477, 715, 1436, 1954, 2069]. Diophantine [2159]. dipole [1327, 2177, 2207]. Dirac [1498, 1116, 172, 1466, 976, 709, 174, 1452, 911, 1166, 799, 1263, 28, 1234, 1897, 2173, 1047, 41, 1936, 785, 1602, 123, 616, 592, 904, 1322, 2152, 692, 1289, 437, 906, 1254, 1403, 1128, 1012, 714, 761, 1963, 774, 1301, 268]. Dirac-delta [709]. Dirac-harmonic [1263]. Dirac-type [28]. Direct [1339, 1073, 1600, 1482, 744]. direction [1051]. Directional [1972]. Dirichlet [329, 886, 1982, 1250, 144]. disassortative [1818, 1941]. disc [1117]. discontinuity [1460]. Discrete [1947, 521, 346, 1619, 1060, 598, 204, 431, 601, 1874, 2161, 1388, 233, 1552, 238, 1537, 247, 2114, 1596, 1704, 897, 1589, 1894, 2079, 448, 1412, 1756, 620, 1066, 1529, 765, 1657, 1729, 665, 1846, 870, 1769, 1120, 565, 754, 949, 1905, 1727, 2147, 268, 818]. discrete- [1120]. discretization [877, 241, 1350]. discretizations [78, 28]. disentanglement [826]. disk [1206, 408, 999, 1602, 426]. dislocations [1833]. disordered [2061, 782, 1673]. dispersion [1110, 1398, 1783]. Dispersionful [1575]. dispersionless [1537, 777, 418, 753]. dispersive [80, 1657]. displacement [1860]. dissipation [1056, 918, 2125]. dissipative [2067, 240, 682, 981, 1410, 1997, 1067, 32, 1877, 1835, 1929, 1278, 489, 438]. dissipative/damped [1877]. distance [218, 169, 1773]. distance-regular [1773]. distances [1679]. Distant [816]. distillation [1416]. distinguishability [1533, 1617, 2097, 1907, 1796]. Distinguished [1917]. Distinguishing [215]. Distributed [493, 991, 290]. Distribution [1719, 1386, 1017, 1279, 1402, 1734, 688, 1245, 2132, 1023, 1743, 947]. distributional [96, 463, 1810]. distributions [1822, 2138, 155, 1519, 1893, 1565, 505]. disturbing [701]. divergence [1843]. Divergences [414, 2119, 2199]. divergent [240]. divine [1367, 1613]. DNA [2008]. Dobiński [2006]. does [304, 265]. Domain [755, 974, 1517, 809, 611, 852, 1814, 1029, 1937, 2096, 323, 1812, 96, 2012, 922]. domains [1126, 738, 651, 2093, 2110, 5, 263, 1630, 1446]. dominated [1133]. dot [343]. Double [888, 1395, 178, 343, 732, 1585, 428, 1248, 1583, 1938, 974, 1793, 772, 778, 344]. double-complex [1248]. double-confluent [1585, 428]. double-sphere [778]. doubles [794]. Dowker [982]. dressing [1887, 2126, 621, 1040]. drift [919]. drift-diffusion [919]. Drinfeld [209]. driven [1015, 975, 357]. Dual [2057, 868, 1971, 575, 20, 944, 1926, 2047, 124, 1047, 114, 1520]. Duality [1828, 951, 1258, 1911]. Dualization [1313]. duals [1756]. due [680, 1239]. Duffin [2196, 1342, 641, 1108]. Dullin [1660]. dyadic [1068]. Dym [1593]. Dynamic [1063, 1838, 1543, 1478, 424, 2206, 1285, 1071]. Dynamical [1391, 326, 30, 2112, 2209, 286, 1904, 1420, 879, 1798, 1956, 521, 866, 1468, 1089, 652, 623, 1044, 1365, 1024, 1278, 1140, 265, 1205, 1425]. Dynamics [1015, 1136, 1876, 1061, 1494, 145, 373, 1062, 1684, 1088, 1579, 1458, 708, 683, 283, 2026, 2047, 244, 1138, 418, 1071, 261, 815, 2004, 286, 1535, 82, 2056, 1366, 731, 1442, 265, 743, 1832, 1951, 1048]. dynamo [253]. dyon [1505]. Dyson [1507].

edge [2011]. Edgeworth [1402, 1734]. Editorial [1863]. EDS [791]. effect [1612, 1860, 2150, 1595, 1421, 496, 547, 1184, 1127, 1335, 1084, 674, 900, 1199, 364, 999, 680, 484, 2062, 1477, 413]. Effective [540, 1648, 1744, 1249, 2039, 62, 2038, 1542, 1773, 1356, 989]. effects [349, 355, 419, 214, 1236, 1461, 1167, 409, 1186, 1993, 800, 1120, 883, 766, 871, 1239]. efficient [1817]. Efimov [413]. Ehrenfest [1974]. eigenbasis [1654]. Eigenfunctions [436, 1147, 1552, 802, 1444, 1578, 493, 294, 1289, 1777]. eigensolutions [1342]. eigenspinoren [1301]. eigenstates [842, 1339]. Eigenvalue [1804, 684, 1664, 1402, 1734, 385, 1362, 592, 510, 32, 781, 168, 904]. Eigenvalues [459, 650, 24, 1738, 1960, 1936, 1286, 789, 312]. eigenvector [510]. eight [1470, 165, 2012]. eight-dimensional [165]. eight-vertex [2012]. Eightfold [1425, 1535]. Einstein [1732, 178, 280, 965, 1188, 1563, 23, 230, 1229, 1356, 1457, 235, 2033, 127, 1187, 227, 544, 671, 810, 128, 1872, 1295, 1467, 129, 761, 2134, 1980]. either [247]. elastic [718, 846]. elasticity [584, 1350]. electric [145, 2177, 2207, 686, 615, 1285, 967, 168]. electrical [1616, 892, 89]. Electrodynamics [1357, 468, 2039, 1695, 1919, 2038, 116, 414, 375, 1915, 73, 1172]. Electromagnetic [1933, 750, 1914, 1327, 1131, 840, 1918, 1713, 878, 1335, 2172, 1710, 1008, 644, 482, 629, 1423, 1810]. Electromagnetism [1828, 229, 1648, 686]. electron [219, 343, 959, 1281, 1631, 1544, 267, 952, 6, 104]. electron-spin [1544]. Electronic [956, 727]. electrons [145, 71, 484]. electroweak [118]. elementary [2043, 1519, 1736]. elements [445, 416]. ellipse [1770]. ellipsoid [1349]. ellipsoidal [1730, 757]. Ellipsoids [2046]. Elliptic [1669, 1451, 888, 1367, 1613, 1206, 93, 502, 2064, 1299, 491, 2111]. embedded [521, 988, 101, 1022, 1152]. Embedding [1217, 957, 1089, 1625, 1805, 2191]. embeddings [1576]. Emden [2078, 1711]. Emden-type [1711]. Emergence [350, 1066, 1149]. emission [1371]. encapsulated [376]. encoding [1216]. encryption [1824, 740]. Endomorphisms [835, 1439, 1709]. endpoints [210]. energies [654, 639, 1408, 1785, 507]. Energy [325, 1106, 1056, 1577, 1105, 232, 1072, 782, 48, 47, 877, 148, 2039, 219, 1634, 972, 1672, 1279, 515, 197, 1908, 131, 1049, 1181, 1923, 712, 916, 164, 51, 1443, 448, 674, 900, 330, 2095, 467, 64, 1946, 357, 702, 1673, 1053, 140, 375, 870, 817, 747, 650, 619]. energy-momentum [1908]. Engle [2109]. enhanced [444]. ensemble [1402, 1734, 245, 101, 1022, 836]. ensembles [689, 658, 526, 1922]. enstrophy [1049]. entangled [343]. Entanglement [470, 1909, 1417, 1416, 605, 708, 1867, 868, 2146, 826, 2003, 1696, 2128, 1332]. entanglements [1032]. entangler [1308]. entries [1922, 86]. entropic [1504]. entropies [1150, 527, 1945, 538]. Entropy [1216, 363, 1426, 1159, 456, 1195, 1721, 531, 469, 2217, 1024, 1834, 1379, 95, 563, 952, 657, 1288, 2, 37, 1243, 457]. entry [352]. envelope [1792]. Environment [1417, 1518, 1656]. environments [1683]. enzyme [884]. epistemic [1569]. Epstein [2066, 228, 1556]. equal [711]. equality [1645, 1020, 2, 37]. equation [172, 1961, 1466, 458, 780, 1586, 24, 968, 137, 1703, 2165, 567, 1166, 1498, 1057, 22, 60, 819, 854, 1421, 641, 280, 1726, 1465, 1241, 2067, 1757, 1399, 2184, 1930, 135, 1328, 865, 460, 888, 1106, 1413, 2043, 2138, 1939, 61, 1668, 794, 1722, 321, 1049, 1184, 1112, 2002, 897, 1077, 190, 2177, 2207, 2030, 1990, 1897, 987, 887, 1741, 342, 1343, 875, 975, 253, 296, 703, 1410, 1718, 1190, 2074, 1935, 80, 2173, 1323, 1997, 1891, 41, 891, 2129, 725, 1600, 654, 1982, 1559, 549, 909, 1113, 1356, 668, 488, 577, 789, 331, 1373, 2190, 2211, 195, 1445, 2125]. equation [123, 1790, 788, 1551, 1531, 696, 154, 1339, 79, 671, 1176, 775, 898, 1983, 2044, 1483, 1940, 1728, 1657, 220, 666, 812, 1315, 906, 1985, 2171, 293, 1886, 1581, 1103, 833, 1004, 2213, 388, 970, 58, 661, 1267, 1836, 1012, 2161, 1654, 1259, 1807, 1967, 2062, 4, 690, 934, 1642, 307, 721, 1114, 1093, 977, 2069, 2134, 177, 489, 1663, 1547, 1450, 1318, 899, 619, 1905, 1777, 1882, 2200, 711, 1040, 1144, 1520, 1246, 1660, 1446, 1798, 102, 894, 2050, 329]. Equations [1153, 1975, 1584, 275, 734, 2186, 1706, 351, 1676, 353, 1134, 424, 2206, 518, 628, 299, 913, 852, 1814, 737, 2083, 2078, 59, 777, 1778, 2008, 2170, 1052, 1579, 1837, 193, 942, 247, 289, 1648, 1284, 155, 587, 784, 2141, 1771, 1508, 486, 1780, 34, 1736, 1585, 1335, 428, 1782, 1715, 1652, 2123, 1986, 240, 2166, 1050, 2049, 764, 495, 300, 730, 1876, 435, 1629, 1880, 845, 1635, 522, 914, 1091, 1354, 1071, 1100, 1776, 340, 568, 1047, 1838, 1222, 1225, 1558, 1755, 1058, 1969, 2139, 1582, 1885, 1561, 659, 1051, 2045, 2205, 184, 320, 1455, 1708, 785, 235, 1541]. equations [94, 1043, 1348, 626, 2064, 1653, 1758, 616, 1710, 1701, 557, 499, 2058, 1272, 1053, 1116, 1045, 83, 2044, 724, 1686, 1999, 2110, 996, 82, 2143, 144, 437, 119, 142, 1347, 1353, 2085, 2052, 1411, 1846, 806, 1271, 1016, 1155, 128, 1769, 1495, 667, 1872, 1295, 2126, 146, 73, 483, 1467, 1442, 822, 1205, 2179, 1034, 1714, 374, 1630, 1570, 1101, 2174, 1988, 1783, 462, 1899, 1397, 91, 1724, 1928, 879, 1727, 1903, 1033, 2168, 2180, 43, 1980, 1983, 2044]. Equilibria [1785, 2159, 1443]. Equilibrium [802, 262, 1317, 688, 912, 528, 1211, 411]. Equivalence [972, 848, 245, 689, 486, 1850, 1344, 1345, 1408]. equivalences [1792]. Equivariant [1406]. erased [1842]. Ergodic [432, 1687]. Ernst [177]. Erratum [2195, 2206, 1330, 1499, 1566, 2141, 2203, 2210, 2207, 987, 2215, 1365, 1609, 2212, 211, 2214, 1608, 900, 668, 2211, 2204, 1758, 1329, 1117, 1940, 37, 263, 1227, 2213, 1118, 961, 934, 2097, 2199, 2200, 104]. Error [1377, 1913, 992, 1121, 40, 2097, 1907, 1796]. Escort [1844]. Essential [1553, 210, 751, 1341]. estimate [1164, 2125]. Estimates [1050, 115, 1436, 1452, 444, 1599, 802, 188, 2114, 1384, 84, 1657, 2097, 1907, 1796, 338]. estimation [503, 1611]. estimations [287]. estimators [1212]. ETM [590]. Euclidean [2217, 1834, 1426, 1328, 1847, 625, 555, 1264, 2066, 1565, 66, 1355, 473]. Euler [599, 337, 1652, 1755, 949, 2174]. Eulerian [1016]. Eutactic [1030]. Evaluation [736, 1603, 2091, 6, 104]. even [617, 247, 140, 1218]. Evenly [991]. event [556]. evolution [343, 1726, 2083, 2008, 2170, 419, 1511, 649, 1741, 1776, 1920, 69, 1590, 1711, 657, 437, 2085, 2052, 181, 1536, 1278, 368, 849, 1903, 1442]. evolutions [1842, 708, 981]. Exact [201, 936, 70, 1741, 157, 568, 555, 921, 543, 1113, 1124, 83, 1001, 1012, 39, 129, 711, 458, 2165, 1281, 888, 1579, 951, 1100, 1539, 1637, 235, 626, 2171, 1411, 1467, 2179, 1783, 1899, 102, 780, 331]. Exactly [1943, 2072, 1496, 1823]. example [1681, 15]. examples [819]. exceeds [885]. Exceptional [391, 389, 1220, 392, 501, 979]. exchange [1232]. exchangeable [1694]. excitation [1327, 2114, 1419, 140]. excitations [590, 89]. exclusion [529, 2018]. Existence [1584, 840, 117, 1308, 1850, 300, 730, 847, 330, 2095, 322, 1483, 1940, 413, 1093, 1494, 1106, 2008, 190, 1410, 1880, 1502, 2110, 1353, 656, 1227, 176, 808, 2069, 1889]. exit [352]. exotic [1762, 590]. Expanding [856, 2193, 918]. expansion [1595, 589, 412, 1725, 1402, 1734, 738, 1741, 381, 1562, 1971, 1753, 896, 676, 2019, 1528, 235, 279, 1540, 1175, 1890, 607, 817, 185, 1724, 1033]. expansion/Fokker [589]. expansions [236, 316, 733, 1655, 149]. expectation [1912, 700]. expectations [147]. experimental [54, 732, 360]. experiments [2150]. Explicit [2002, 2070, 2037, 686, 764, 1356, 2204, 1097, 480, 1757, 292, 1295, 1806]. exploring [148]. exponent [1706]. Exponential [1244, 248, 2145, 2035, 1290, 2060, 1243]. exponents [518, 628, 973, 1377]. expression [1806]. expulsive [1563]. Extended [990, 863, 1240, 1248, 553, 940, 163, 1307, 2089, 913, 777, 1579, 1716, 1528, 1601, 1887, 426, 1117, 1856]. Extending [1568, 314]. Extension [819, 850, 189, 1174, 1256, 396, 1968, 2065, 2107, 2106, 1480, 1957]. Extensions [295, 589, 402, 594, 944, 498, 1532, 122, 1369]. Extensive [457, 1079]. exterior [5, 263]. external [780, 1672, 190, 1113, 755, 1423, 1966, 618, 947, 1678]. extra [715]. extracted [1332]. extraction [636]. extraspecial [1977]. Extremal [1522, 745, 668, 488, 1979, 1445]. extremals [1072].

facile [1697]. factorizable [842]. factorization [659]. factorizations [600]. factors [2023]. failure [95]. false [1119]. families [1220, 509]. family [186, 1583, 914, 114, 1480]. far [1901, 2172]. far-field [1901, 2172]. Faraday [496]. Fateev [140]. Fay [760]. Features [869]. Fedosov [1732, 1752, 81, 606]. feeded [1563]. fees [1692]. Fermi [1314, 1014, 1033]. fermion [73, 485]. Fermionic [1374, 722, 2020, 1195, 1383, 1533]. fermionization [1902]. Fermions [1764, 1439, 935, 710, 1820, 1636, 2101]. ferromagnetic [1596, 1478]. Feynman [1194, 1829, 1273, 1919, 1589, 1923, 691, 398, 1597, 279, 1074, 1296, 844, 254]. fiber [486]. Fibonacci [1511]. fidelity [218, 539]. field [2040, 386, 1825, 125, 1914, 472, 2073, 2031, 1409, 87, 126, 950, 1509, 1675, 631, 169, 1672, 1464, 1241, 1792, 865, 1713, 581, 942, 1603, 986, 1788, 1648, 203, 302, 1512, 1901, 643, 1127, 1240, 686, 1607, 421, 328, 230, 1385, 1881, 1712, 2104, 21, 1838, 1384, 71, 1582, 1885, 755, 2041, 512, 1685, 228, 1326, 167, 2172, 1573, 1005, 161, 696, 609, 834, 154, 1710, 13, 1008, 671, 455, 1026, 545, 800, 677, 181, 2101, 1368, 806, 98, 128, 168, 667, 1872, 2185, 731, 1423, 1682, 1772, 883, 1744, 747, 1467, 949, 1038, 507]. field [1924, 1239, 1980]. field-antifield [1409]. Fields [379, 841, 1131, 726, 1828, 911, 18, 145, 1203, 770, 540, 1424, 1361, 807, 686, 421, 2001, 811, 1429, 1933, 956, 2214, 967, 1019, 1507, 1455, 2048, 2005, 680, 1345, 1276, 426, 1117, 50, 1649, 383, 482, 629, 915, 1772, 8, 38, 1407, 2202, 746, 1157, 1870, 1928, 1301, 1678]. Fierz [444, 373]. fight [2217]. filament [1141]. film [203, 302]. filter [327]. Final [845, 2074, 1471]. Finding [1736, 1579, 1724]. Fine [2019, 1178, 497]. Finetti [1330, 3, 1822, 449, 1697]. Finite [303, 1822, 61, 625, 1127, 1152, 963, 542, 1905, 871, 472, 902, 866, 602, 832, 964, 570, 1819, 794, 2084, 1607, 1719, 15, 830, 439, 308, 633, 449, 1717, 651, 123, 1303, 1497, 67, 509, 35, 1446]. finite-dimensional [830, 67]. finite-gap [2084]. Finite-time [303]. Finitely [709]. Finsler [1732, 761, 1293]. Finslerian [669]. First [1688, 2043, 1736, 1208, 960, 592, 1425]. Fisher [1311, 608, 1721, 156, 1081, 787, 505, 1570]. Fitzhugh [58]. five [1978, 841, 178, 1566, 1428, 838, 1625, 1427, 2156, 479, 235, 906, 473]. five-body [838]. five-dimensional [841, 178, 1566, 1428, 1625, 1427, 2156, 479, 235, 906, 473]. fivebrane [225]. fixed [1068, 109, 212, 1471, 1841, 1753, 1545, 2094, 1010]. fixing [1406, 1129]. Flat [1937, 2096, 350, 49, 130, 695, 515, 199, 200, 394, 603, 758, 1750, 369, 557, 499, 376, 1864]. flatness [51, 1954]. Flavor [1094, 1746, 1303, 277]. flip [1595]. flip-over [1595]. Floquet [390]. Flow [587, 557, 1011, 1057, 1056, 1054, 1548, 546, 688, 1221, 1072, 1802, 323, 1536, 757]. flows [1087, 2059, 1141, 1060, 587, 2088, 139, 1658, 179, 1065, 420, 1067, 1442, 1963, 1223]. Fluctuation [918, 779, 1024]. Fluctuation-dissipation [918]. Fluctuations [1672, 783, 1918, 658, 2018]. Fluid [424, 2206, 1048, 1386, 478, 1281, 581, 587, 649, 1044, 1365, 300, 553, 1435, 1071, 756, 568, 1065, 1449, 1059, 322, 83, 1766, 1649, 618, 947]. Fluid-dynamic [424, 2206]. fluid-dynamical [1044, 1365]. fluids [2059, 2164, 1878, 648, 1702]. flux [186, 1015, 352, 1989]. fluxes [419]. focal [2183]. Fock [2031, 295, 1636, 366, 1034]. Foias [65]. Fokker [1092, 589, 1091]. Foldy [2152]. foliation [1133]. force [780, 913, 190, 644]. forced [1056]. forces [1008, 618, 947]. Forchheimer [2059]. forcing [1961, 240]. Form [1492, 1762, 2091, 2042, 1499, 1638, 982, 526, 2053, 71, 1079, 2023, 664, 1810]. Form-invariant [1492]. Formal [520, 1487]. formalism [458, 1513, 1409, 506, 239, 1191, 318, 729, 1084, 1992, 347, 335, 511, 181, 804]. formalisms [1464]. formation [201, 1400, 959, 1429, 1170, 1169, 1723]. Forms [1178, 183, 2193, 893, 1398, 421, 769, 1277, 727, 1250, 1238, 2004, 432, 1411, 1810, 1148, 966]. formula [799, 570, 110, 1616, 203, 302, 1359, 592, 952, 2153, 1230, 558, 699, 850]. formulas [44, 166, 777, 2002, 643, 2192, 298]. formulation [1530, 591, 270, 1971, 1169, 1016]. Fornberg [1967]. found [1980]. foundation [1487, 1488, 1489, 1490, 1680]. Foundations [804, 1840]. Four [291, 2109, 713, 522, 1879, 235, 72, 511, 1699]. Four-dimensional [291, 522, 235, 72, 511]. four-punctured [713]. four-simplex [2109]. four-state [1879]. Fourier [1552, 1179, 665, 1699]. fourth [486, 693, 1776, 2190, 2211, 1411, 1446, 2050]. fourth-order [1776, 1446]. Fowler [2078]. fractal [622, 1056, 1348]. fractal-forced [1056]. Fractional [786, 478, 957, 831, 1493, 481, 132, 837, 1430, 1584, 2186, 1399, 2184, 1858, 1930, 1077, 1460, 719, 34, 1127, 703, 2087, 1435, 1245, 1576, 1765, 2132, 331, 1326, 1541, 2167, 437, 1622, 2161, 977]. fractions [149]. frame [810]. frameness [2034]. frames [1853, 1370, 2001, 1756, 1207, 1122]. Framework [464, 819, 238, 1878, 1156, 320, 2185, 1295]. Francoise [1956]. Francoise-type [1956]. Frank [357]. Fredholm [1642]. Free [1038, 935, 710, 1174, 326, 1735, 1408, 1673, 2127, 560, 140, 1254, 383, 642, 1916, 1818, 1941, 485]. frequency [364, 583]. Freund [499]. frictionless [1092]. Friedel [576]. Friedmann [319, 548, 1538]. Friedrichs [863]. Frobenius [339, 1763]. Fuchs [314, 909, 898]. Full [1129, 1637]. fully [1783, 462]. Function [370, 1961, 1991, 1526, 1947, 1687, 201, 506, 1751, 202, 1672, 1402, 1734, 738, 768, 1831, 958, 1471, 1469, 1154, 1544, 290, 1247, 1851, 632, 1561, 595, 1556, 1693, 2075, 1975, 1175, 664, 2032, 107, 2012, 572, 462, 473, 298]. Functional [2201, 2080, 2203, 1953, 123, 1194, 219, 325, 1661, 2138, 2166, 1026, 1985]. functionals [219, 1044, 1365, 1963]. functions [841, 1109, 1991, 734, 2017, 2181, 798, 433, 9, 1858, 1930, 1191, 2196, 2028, 1233, 2090, 1448, 2135, 643, 1496, 1460, 1415, 1580, 1631, 115, 593, 1245, 1750, 1179, 639, 1290, 1108, 609, 2152, 1300, 2060, 560, 1955, 220, 607, 1803, 149, 1149, 665, 474, 1698, 1243, 1296, 817, 85, 259, 411, 2097, 1907, 2119, 2199, 1966, 150, 1570, 1857, 153, 485]. fundamental [1465, 1582, 2117, 2185]. Further [1055]. fusion [1977, 1856]. future [2127]. fuzziness [1958].

G [344]. G-double [344]. Gabor [851, 820, 1756]. Galilean [384, 1458]. Galois [1347, 742]. galore [1294]. games [1079]. Gap [1927, 1613, 2084, 2061, 917, 1664, 1295]. gapped [2027]. gaps [308, 633]. Gardner [404]. gas [1633, 87, 1088, 1579, 1908, 321, 2061, 1175, 844]. gases [733, 1284, 267]. gates [1791]. Gaudin [380, 70, 564, 1267]. Gauge [1265, 406, 714, 841, 1406, 1432, 1828, 126, 1085, 675, 717, 2083, 803, 289, 686, 1712, 1387, 1582, 1463, 1115, 2106, 2004, 2154, 451, 545, 292, 477, 1002, 1892, 1865, 474, 1129, 129, 1864, 1148, 1924]. gauge-invariant [1582]. Gauged [475, 171, 642]. Gauss [423, 1357, 1430, 966]. Gaussian [1402, 1734, 526, 493, 101, 1022, 2005, 1795, 2136, 1441]. Gaussians [290]. Gazeau [1456, 1334, 114]. Gelfand [1404]. Gell [698, 1000]. Gell-Mann [698, 1000]. General [1739, 2214, 1156, 1245, 2125, 79, 1999, 1407, 2202, 1011, 581, 1508, 1248, 1583, 1491, 1850, 2166, 1953, 452, 266, 71, 1885, 543, 1910, 409, 910, 387, 2185, 2054, 496, 78, 1992]. Generalization [1995, 1317, 1562, 44, 276, 1559, 1243, 1843, 1101, 1144]. generalizations [1470, 2158, 1969, 1877]. Generalized [219, 687, 523, 987, 887, 886, 1020, 1268, 531, 1229, 1885, 1700, 1412, 1186, 1945, 1754, 1289, 1002, 564, 1267, 447, 2119, 2199, 977, 1037, 966, 1352, 1375, 2059, 202, 1672, 2184, 2008, 2043, 2138, 738, 1324, 1791, 1990, 1923, 546, 1180, 1362, 1737, 1631, 495, 725, 1582, 543, 1212, 1557, 1637, 1887, 2194, 52, 336, 121, 11, 204, 431, 601, 560, 497, 484, 1892, 1666, 31, 1155, 2086, 88, 429, 979, 2126, 664, 2097, 1907, 1796, 828, 1114, 150, 1318, 462, 879, 1523, 638, 696]. Generalizing [1005]. Generally [1228, 270]. generated [841, 1789, 156, 497, 1149]. generates [312]. Generating [1991, 2193, 834, 197]. generation [1323, 407]. generator [1613]. generators [625, 297]. generic [1712, 1378, 536, 1978]. Genericity [1429]. genetic [1149]. Gennes [121]. genus [100, 1786, 1905]. geodesic [1086, 814, 1767]. Geodesics [524, 2107]. Geometric [2140, 848, 878, 784, 2141, 1737, 1571, 872, 2098, 1747, 2146, 932, 470, 80, 1221, 1162, 1433, 1145, 1414, 98, 1549, 1350]. Geometrical [695, 634, 532, 1704, 347, 335]. geometries [152, 274, 1396, 1950, 1979, 96, 751]. geometrodynamics [377, 2212]. Geometrothermodynamics [890]. Geometry [188, 369, 1378, 989, 307, 1981, 46, 1406, 1391, 1386, 613, 874, 1964, 1017, 151, 857, 1454, 1312, 1110, 1554, 605, 453, 646, 2176, 635, 2183, 556, 445, 2047, 68, 56, 1888, 1873, 1035, 577, 1500, 821, 1152, 1592, 436, 1147, 1766, 1892, 644, 931, 757]. geophysical [1071]. geophysics [730]. Gepner [749]. Gérard [2130]. ghost [1449, 2127]. ghost-free [2127]. ghosts [1418]. GHZ [636]. Gibbs [1243, 1674, 55, 1019, 1685]. Gibbsianness [1678]. Ginzburg [349, 1130, 351, 353, 354, 355, 1939, 356, 1632, 1982, 358, 1993, 348, 317]. Girardello [1415, 1233, 1867]. given [1809, 1030]. Glaser [2066, 228]. Glashow [698]. glass [1671, 532]. glasses [1674, 1675, 1682]. Glauber [1991, 815]. Glauber-type [815]. Gleason [824]. Global [2008, 1771, 190, 658, 948, 514, 1292, 1880, 1190, 1841, 358, 960, 2110, 1353, 1807, 2069, 1800, 1106, 283, 342, 1410, 914, 1502, 1467, 1547]. Globally [712]. Going [1306]. golden [1033]. goldfish [519, 402]. Goldfishing [776, 717]. Goldstone [121]. good [1586]. Gordan [36, 663, 392]. Gordon [1703, 888, 1668, 1127, 2015, 1880, 553, 1635, 2074, 789, 616, 1728, 383, 970, 711, 1798]. Gott [132]. Gottwald [1660]. governing [1016]. grade [1878]. graded [1364, 2071, 344, 2076, 563]. gradient [2180]. grading [497]. Gradings [1527, 1178]. graining [369]. grand [1110]. granular [424, 2206]. graph [591, 1655, 1553, 414, 800, 1689]. graph-Laplacian [1553]. graphical [1910]. graphs [1688, 1934, 670, 1164, 1677, 228, 834, 724, 1679]. Grassmann [1750, 1257]. Grassmannian [333]. gravitating [756, 1984]. Gravitation [622]. Gravitational [1427, 1702, 615, 580, 680, 1434, 1426, 1770, 648, 1429, 579, 1582, 1885, 2217, 1834, 415, 50, 677, 715, 750, 1010, 1239]. gravitons [2140]. gravity [1732, 233, 1828, 1203, 803, 1704, 1248, 1583, 1491, 1427, 1702, 752, 2157, 553, 51, 21, 1587, 319, 772, 1872, 1012, 129, 281, 1833, 1980]. Green [841, 201, 798, 433, 1858, 115, 1262, 664, 473, 298]. Greitzer [1057]. Gribov [1463]. Gross [1190, 2139, 1113, 1356, 1455]. Grothendieck [1306]. Ground [916, 2130, 375, 366, 861, 2027, 674, 900, 727, 457]. Group [1100, 707, 184, 1178, 513, 864, 163, 1307, 2089, 1312, 1522, 570, 326, 1628, 1871, 794, 12, 1977, 1746, 944, 1607, 92, 1215, 2063, 1776, 205, 301, 1561, 1507, 405, 192, 1910, 2188, 1207, 1094, 1299, 1200, 451, 561, 1322, 1026, 1983, 2044, 252, 2171, 165, 370, 565, 277, 1832]. group-covariant [12]. Group-theoretical [707]. Groups [1608, 430, 901, 1392, 1491, 1562, 1209, 1320, 2163, 1036, 932, 1260, 1909, 1637, 651, 1303, 1097, 14, 420, 204, 431, 601, 1151]. growing [1245, 254]. growth [1735, 651]. growths [193]. GUE [385]. guides [1131]. Günther [181]. Gutzwiller [570].

Hadamard [993, 1917, 1976]. Hadamards [1699]. Hahn [1926, 1309]. Half [1042, 2017, 1174, 1629, 157, 1255, 557, 13, 1459, 835]. half-flat [557]. half-infinite [2017]. half-line [1629, 157, 1255]. half-plane [1174]. half-sided [835]. Half-Skyrmions [1042]. half-space [13]. Hall [46, 1015, 423]. Hamilton [1732, 137, 1241, 1112, 1354, 1888, 1207, 79, 82, 1420]. Hamiltonian [2216, 1011, 186, 786, 1828, 681, 533, 582, 1174, 946, 586, 594, 1716, 237, 1816, 1240, 1747, 1173, 790, 1593, 421, 814, 1639, 1987, 418, 1514, 80, 1900, 243, 69, 367, 1076, 825, 1877, 182, 25, 370, 1430, 753, 1962, 1707, 1140, 1351, 2111, 1523, 2147]. Hamiltonians [842, 627, 962, 208, 186, 1126, 540, 901, 1868, 863, 1749, 68, 409, 1571, 11, 989]. handed [907]. Hankel [156]. hard [885, 701, 1019, 1021, 721]. hard-core [701]. hard-rod-type [1021]. Hardy [1234]. Harish [1316]. Harish-Chandra [1316]. Harmonic [1154, 818, 1545, 426, 1117, 8, 38, 468, 1388, 2216, 2184, 1191, 1390, 861, 946, 1106, 1263, 539, 1419, 1563, 1658, 253, 401, 1162, 1904, 1877, 67, 1434, 542, 661, 1772, 2077, 374, 1450, 1266]. harmonics [589, 978, 928, 1667, 1867, 1748, 1743]. Harper [1793]. Harry [1593]. Hartree [2031, 1620, 1034]. Hartwig [156]. Hasegawa [342, 879]. Hausdorff [1817]. having [1136, 193, 625, 1983, 2044]. Heat [738, 688, 1384, 155, 951, 1935, 595, 2000, 119, 1775]. Heat-kernel [738]. Hecke [202, 1261]. Hedin [996]. Heisenberg [1465, 706, 1481, 1596, 2121, 701, 2023, 659, 994, 1316, 1124, 565, 1815]. helically [4, 690, 934]. helically-reduced [4, 934]. helices [2092]. helicity [139]. helicoid [1166, 1498]. helium [1543, 763]. Hellmann [1923]. Helmholtz [664]. helping [152]. Hermite [658]. Hermitian [1927, 468, 1197, 1811, 1571, 367, 1475, 781]. Hermiticity [671, 739]. Herring [1722]. heteropolymer [29]. Heun [2207, 968, 1585, 2177, 428]. Hiatus [1376]. Hidden [872, 942, 1422, 1248, 562]. hierarchies [734, 552, 1970, 136, 2182, 1747, 693, 179, 1265, 2045, 2205]. Hierarchy [142, 465, 2159, 777, 831, 2084, 1593, 396, 933, 418, 1900, 1601, 1887, 788, 1729, 1135, 1575, 1605, 150, 1523, 2168]. Higgs [770, 1952, 173, 1214, 124]. high [1671, 885, 654, 1436, 1378, 747]. high-dimensional [1378]. high-energy [747]. Higher [1218, 1032, 748, 1987, 1208, 1846, 1504, 1799, 1470, 1762, 36, 2201, 1474, 624, 786, 1828, 126, 1085, 600, 1725, 2193, 1018, 515, 1871, 419, 2080, 2203, 1346, 643, 1658, 1702, 1387, 704, 2121, 51, 1041, 2190, 2211, 2144, 319, 477, 1954, 1864]. Higher-dimensional [748]. higher-order [1799, 1474, 1725, 2193, 319]. highest [231]. Hilbert [1236, 1006, 1853, 1913, 1934, 623, 362, 1641, 1216, 766]. Hill [1961]. Hirota [777, 2126]. Hiscock [132]. histories [1078, 112]. hole [1134, 1703, 1426, 1429, 579, 2217, 1834, 808]. holes [1871, 1202, 1768, 1170, 479, 1979, 1169, 484]. Holevo [2097, 1907, 1269, 1796]. Holm [136, 61, 1747, 1715, 1410, 1708, 489, 735, 462, 1660]. Holm- [1747]. holomorphic [1041]. holonomic [670, 159, 39]. Holonomy [1892, 1515, 1571]. homoclinic [2113, 242]. Homogeneous [318, 1576, 2020, 1302, 1908, 1188, 1638, 454, 1986, 811, 1429, 1702, 1485, 243, 1347, 1467, 1642, 2054, 153]. Homological [992]. homomorphisms [1177, 1525]. Homotopy [1783, 2157, 535]. honeycomb [1252]. Hooft [669]. Hopf [955, 276, 1572, 1364, 237, 1931, 1845, 228, 344, 609, 834, 1095, 288, 638]. Hopfield [1687]. Hopfions [27]. hopping [1959]. horizon [556, 1979]. horizons [227, 544]. Horn [1565]. Horseshoes [1769]. host [424, 2206]. hot [2093]. Hove [260]. Hubbard [1213, 1959, 1479, 916]. Hudson [830]. Hulthén [1128]. Hunter [1708]. Husimi [505]. Hybrid [264, 1539]. Hydrodynamic [1629, 858, 660, 533, 1492, 790, 1440, 1968, 1577, 1994]. hydrodynamical [1288]. hydrodynamics [138]. hydrogen [543, 271, 1039, 507]. hydrogenic [608, 1912]. hydrogenlike [1167]. hydromagnetic [1989]. Hyers [2138]. hyperbola [271]. Hyperbolic [100, 532, 191, 1479, 280, 180, 441, 1191, 1357, 2070, 712, 68, 223, 1221, 1602, 1540, 440, 107, 532]. hyperelliptic [258, 1591]. hyperfunctions [2138]. hypergeometric [1526, 201, 1191]. Hyperincursive [1388]. Hyperspherical [1748, 978]. hypersurfaces [752]. Hypothesis [1795, 386, 1377, 2063]. hysteresis [285].

I. [1915]. Ice [1632]. Ice-water [1632]. ICMP09 [1690]. Ideal [904, 587, 1780, 139]. Ideality [422]. identification [1983, 2044]. identify [537]. identities [400, 1964, 1400, 23, 502, 1220, 1801, 1200, 2006, 149]. identity [624, 1261, 1655, 1773, 1405, 760, 1826]. II [1978, 627, 637, 1331, 151, 567, 22, 1465, 617, 1695, 1934, 515, 1794, 155, 1252, 1488, 1081, 1170, 769, 200, 921, 1705, 2016, 544, 1026, 2044, 2153, 1004, 970, 1723, 1442, 1951]. III [1225, 1333, 36, 773, 1489, 645, 274, 394, 1146]. images [2153]. imaginary [2197, 1921, 1505]. imaginary-charge-carrying [1505]. imaginary-real [2197, 1921]. immersion [743]. impenetrable [5, 263]. Implementation [12]. implicit [1561]. implies [1090]. imprisonment [1510]. Improved [1716]. Improvement [113]. impulsive [586, 422, 720, 1433, 2052]. impurity [135]. Inadequacies [2073]. Ince [428]. incidence [2155]. including [2045, 2205, 121]. inclusions [349, 354, 835]. Incomplete [1158]. Incompressible [1455, 1054, 756, 1558, 1755, 1059, 618, 947]. Incursive [877]. indecomposable [2169]. indefinite [291, 500, 1423]. Independent [952, 86]. index [1709, 850]. indices [2181]. indirect [462]. Induced [376, 1595, 1417, 53, 1648, 583]. inefficient [469]. Inequalities [360, 2, 37, 571, 1452, 1973, 796, 1684, 1234, 217, 712, 95, 224, 1306]. inequality [55, 1181, 868, 463]. Inequivalent [471]. inertia [1280]. Infeld [229, 942, 1574, 127]. infimum [1908, 214, 1236, 2120, 1186, 1564, 766]. Infinite [2181, 1322, 1568, 1674, 2017, 1913, 866, 1959, 203, 302, 190, 362, 1248, 1491, 1410, 295, 896, 286, 882, 1053, 1985, 134]. Infinite-dimensional [1322, 1913, 1248, 1491, 896, 134]. infinite-range-hopping [1959]. infinite-volume [1674]. Infinitely [838, 1023]. infinitesimal [655, 297]. infinity [1954]. inflation [578]. Influence [1138, 1194]. Influences [1818, 1941]. Information [978, 527, 2142, 1824, 1809, 1311, 608, 1374, 1721, 56, 1081, 469, 787, 505, 309, 1570]. Information-theoretic [978]. informationally [163, 1947, 306]. Infraparticle [1695]. infrared [843, 1200]. inhomogeneities [2191]. inhomogeneity [1818, 1941]. inhomogeneous [1235, 1020, 1749, 409, 451, 1401, 338, 329]. Initial [1445, 112, 886, 1920, 32, 57, 128, 176, 2069, 2050]. Initial-boundary [1445, 886, 57]. Inner [1095, 1413, 304, 2147]. inspired [1431, 1924]. Instabilities [320, 648, 883]. Instability [512, 1899]. instantons [1198, 1700, 255, 1869, 477]. instruments [1819]. insulators [2011]. integer [1109, 109, 212]. Integrability [513, 1726, 2159, 284, 2086, 275, 1432, 1828, 16, 239, 2043, 1508, 206, 549, 93, 1852, 520, 1626, 1389, 2048, 243, 2081, 144, 1347, 1666, 899]. Integrable [1537, 897, 1252, 1173, 2121, 435, 1206, 427, 94, 1610, 1643, 25, 792, 1223, 1762, 552, 1473, 1898, 2028, 537, 2170, 1177, 790, 948, 1629, 1627, 2053, 1900, 1539, 1969, 1559, 2045, 2205, 620, 1348, 1097, 1790, 434, 256, 1983, 2044, 724, 812, 1135, 1411, 1846, 924, 1808, 134, 1962, 241, 822, 853, 150, 1318, 75, 1144, 1928, 618, 947, 2014]. Integral [1329, 119, 936, 569, 1661, 1603, 1788, 1357, 1589, 1398, 691, 41, 881, 1231, 1355, 1296, 1495, 699, 1139, 572]. integrals [2091, 2181, 1829, 2043, 585, 1080, 1736, 1562, 1320, 1208, 676, 2019, 398, 1528, 960, 2144, 1693, 2016, 1597, 79, 107, 6, 104, 91, 254]. Integration [278, 1750, 2186, 1273, 878, 2175, 2065, 2126]. intense [185]. Interacting [1284, 575, 1923, 1429, 71, 1994, 1175, 1772, 1714]. Interaction [1505, 18, 1126, 959, 1063, 861, 1671, 1376, 2098, 1232, 1180, 800, 85, 118, 1547, 1040]. interactions [468, 709, 2031, 1085, 653, 188, 882, 1803, 1797, 750]. interesting [1397]. interface [126, 418]. Interference [2115]. interferometer [677]. interior [1327, 1802]. intermediate [983, 388]. International [1690]. interpolation [253, 437]. interpretation [1949, 1571]. Intersecting [752]. interspersion [221]. intertwiner [516]. Intertwining [1853, 1096, 1868, 981]. interval [1809]. Intrinsic [556, 807]. introduces [2142]. Introduction [1670, 1048, 2014, 348]. invariance [662, 1823, 262, 827, 11]. Invariant [980, 1328, 461, 1835, 1226, 611, 1244, 1492, 1370, 1136, 1098, 831, 237, 283, 2187, 1204, 948, 1893, 1209, 470, 492, 1582, 339, 308, 1750, 1507, 1389, 1533, 420, 1300, 1095, 1368, 1278, 450, 817, 1567, 2192]. Invariants [1512, 617, 258, 459, 1813, 2210, 1440, 1968, 914, 769, 1146, 451, 182, 2151, 1401, 1981]. Inverse [1809, 1912, 2084, 1335, 1583, 891, 666, 146, 1034, 962, 791, 1935, 654, 1019, 1597, 1775, 813, 31, 388]. Inversible [256]. Inversion [105]. inverted [2177, 2207]. Investigation [610, 1711, 954]. Inviscid [1068, 1060, 2058]. invisibility [1394]. involutive [564]. involving [1400, 1343, 1045, 2180]. ionized [2058]. Irreducible [1884, 630, 551, 1343, 152, 274, 1396, 1950, 405, 192, 1659, 392, 482, 629, 742, 35]. Irreversibility [1090]. irreversible [1139]. Ishimori [1891]. Ishimori-I [1891]. Ising [1991, 2017, 188, 1448, 1719, 85]. Isochronous [582, 1174]. isochrony [2159]. isoholonomic [39]. isolated [1211]. isometries [1789]. isometry [1871]. isomonodromic [1751]. isomorphism [1607, 1636, 14]. Isomorphisms [795]. isoperimetric [1726, 217]. isospectral [1868]. Israel [837]. Issue [1670, 1048, 2014, 1305, 1610, 1643]. iterated [1154]. Iteration [1965, 641, 1511, 719]. iterative [499]. IV [1490, 1396, 603]. Ivanov [1932]. Iwahori [1261]. Iwatsuka [168]. IX [284, 520].

J [1588, 2201, 2195, 2216, 1498, 2206, 628, 441, 630, 2196, 1330, 1499, 2197, 1566, 302, 860, 2141, 2203, 2210, 2207, 1415, 1456, 987, 2215, 1365, 1609, 2212, 1236, 212, 2096, 2209, 211, 2214, 1608, 568, 900, 2095, 301, 467, 668, 2205, 2211, 2204, 347, 1758, 1790, 1329, 1834, 1116, 1117, 1940, 2198, 1414, 501, 2208, 37, 263, 1227, 2213, 440, 1118, 629, 38, 961, 934, 2097, 2199, 2202, 1157, 899, 2200, 104, 1941, 1907]. J. [2217]. Jacobi [855, 624, 137, 1241, 1112, 382, 522, 914, 1354, 1888, 79, 898, 1560, 82, 665, 1360]. Japanese [1691]. Jarlskog [2187, 2192]. Jarzynski [1020]. Jaynes [941, 443, 2086]. jets [1652]. Jimbo [693]. Johnson [543]. Joint [1419, 1121]. joint-measurement [1121]. Jones [1709]. Jordan [2198, 1830, 77]. jordanian [380]. Journal [1759, 2142, 1863]. Jump [1598, 1805, 1516]. jumps [2143, 1985].

Kac [729, 1527, 191, 495, 398, 1754, 2153, 638]. Kadomtsev [567, 725, 1318]. Kähler [2096, 173, 291, 1937, 1238, 1293]. Kalb [178]. Kaluza [2047, 614, 910, 473]. Kanai [1711]. Kantor [534]. Karhunen [1216]. Karman [2110]. Karpelevich [1528]. Kato [1060]. Kato-type [1060]. Kauffman [1149, 1857]. Kaup [2171, 293, 150]. Kazhdan [944]. KdV [899, 404, 549]. Kemmer [1342, 641, 2196, 1108]. Kepler [441, 939, 440, 180, 282, 1621, 1948, 828, 640, 1013]. kernel [1706, 490, 777, 738, 1641, 1384, 595, 696, 213, 119, 1642]. kernels [1641, 1685, 378]. Kerr [911, 178, 556, 579, 616, 774]. Kertész [1472]. ket [506]. Khokhlov [2067]. kicked [1793, 390, 69]. Killing [524, 695, 515, 646, 206, 546, 1345, 227, 544]. Kinematical [1524]. kinematics [106, 487, 542]. Kinetic [1722, 1991, 219, 347, 335, 720, 484, 883]. kinetics [1656, 884]. Kinks [2163]. Kirchhoff [1800]. Kirkpatrick [1671]. Klauder [1415, 1456, 1233, 1334, 114]. Klein [1880, 616, 1703, 1668, 1127, 2047, 553, 1635, 2074, 614, 789, 1728, 383, 711, 1798, 473]. Knizhnik [875]. knot [908, 451, 450]. knotlike [748]. known [44]. Kochen [1158, 360, 165]. Kolmogorov [2057]. Kondo [1084]. Korean [1691]. Korteweg [961, 2159, 1970, 1878, 683, 845, 1559, 1536, 1886, 970, 889, 894]. Korteweg-type [1878]. Koszul [400]. KP [1887]. Krajewski [1859]. Kramers [1317]. Kravchuk [1415, 1233]. Kreimer [276]. Krein [770]. Krieger [339]. Kullback [1843]. Kupershmidt [790, 293].

Labeling [1934]. Ladder [443, 928]. Ladik [1523]. ladungskonjugationsoperators [1301]. Lagrange [1732, 81, 1293]. Lagrangian [2201, 2216, 1499, 2203, 2042, 239, 1018, 946, 893, 957, 2080, 1712, 1639, 273, 1064, 525, 1139]. Lagrangians [1294, 1560]. Laguerre [1400, 658, 526, 2006]. Laguerre-type [2006]. lamb [1787, 588]. Lambert [2032]. Lamé [1103]. Lanczos [1508]. Landau [423, 349, 1375, 186, 1130, 351, 353, 354, 355, 1939, 356, 1632, 68, 1580, 1982, 358, 1993, 2081, 348, 317]. Landau-like [68]. Landauer [110, 1616, 952]. landscapes [148, 1338]. Lane [1711]. Langevin [2184, 1930, 1990, 696]. languages [1487]. Laplace [1328, 593, 2009]. Laplace-type [2009]. Laplacian [280, 1485, 1553, 651]. Laplacians [1964]. Large [771, 412, 1142, 1283, 57, 873, 507, 283, 156, 105, 1066, 1540, 32, 175, 1707]. Large- [412, 175]. Large-order [507]. large-scale [283]. Large-time [57]. largest [1402, 1734, 1960, 86]. laser [1595]. laser-induced [1595]. last [898, 1560]. lattice [350, 990, 770, 1252, 1014, 933, 956, 1206, 2061, 769, 1879, 1539, 74, 917, 1601, 334, 1533, 1535, 1425, 457]. lattices [2148, 1898, 2114, 897, 1252, 1705, 1795]. Laurent [676, 1026]. law [975, 1047, 1024, 1844, 1679]. laws [1439, 943, 404, 299, 1738, 179, 1047, 1640, 781, 142, 584]. Lax [2131, 1956, 179, 2079, 788, 65, 753]. layer [704, 1070, 1802, 2062]. layered [1327]. layers [837, 1046, 721, 1093]. leaky [217]. Least [903]. leaves [1598]. Lee [1270]. Leffler [1930]. Lefschetz [1010]. Left [907, 1584]. Left- [907]. Leibler [1843]. Leibniz [1606]. lemma [297]. Lenard [1779]. length [1781, 1231, 1039, 995, 2185, 1567, 1023, 1818, 1941]. lengths [825, 1254]. lens [1717]. lensing [579, 1955, 1010]. Lenz [2162]. Lepton [907]. Leutwyler [1528]. Level [99, 1612, 862, 125, 264, 1699, 2086, 1682]. levels [1375, 325, 221]. Lévy [975]. Lewis [645]. liberation [1488]. Lichnerowicz [1964]. Lie [1380, 275, 2099, 2165, 798, 389, 103, 602, 1358, 1524, 2078, 1725, 2193, 59, 1837, 289, 1816, 1177, 1606, 1550, 1665, 498, 1003, 1209, 1998, 1848, 1900, 1156, 2071, 856, 1971, 1222, 620, 1220, 1097, 2076, 1587, 1276, 14, 434, 420, 859, 294, 491, 1904, 534, 1983, 2044, 204, 431, 601, 1217, 813, 667, 88, 429, 979, 1151, 1401]. Lieb [1599, 1363, 463]. Liénard [1983, 2044]. Liénard-type [1983, 2044]. Lifshitz [2081]. Lifting [224]. lifts [1732]. Light [2021, 685, 1953, 953, 185]. Light-cone [2021]. light-mass [185]. like [913, 68, 668, 488, 235, 437, 177, 1783]. likelihood [1212, 2119, 2199]. Limit [87, 1159, 1592, 1961, 1620, 937, 1825, 1143, 2089, 238, 373, 76, 1279, 861, 1952, 1060, 260, 863, 1613, 2177, 2207, 793, 2023, 1683, 1059, 369, 560, 1395, 2022, 1807]. Limiting [781, 155, 987, 887, 1685]. limits [1760, 428, 1142, 1455, 1187, 1120, 2178]. Lindstedt [1014, 2057, 481, 381]. line [1651, 1989, 156, 1629, 157, 1255, 1041, 1544, 892, 89, 1472, 1155, 1459]. Linear [1164, 2216, 946, 719, 1248, 693, 1958, 2045, 2205, 2064, 415, 781, 220, 1581, 8, 38]. linearization [280, 1225]. linearized [1828, 1465, 210, 587, 650]. linearly [969]. lines [165]. link [258, 2151]. link-invariants [258]. linking [1357]. Liouville [1943, 1263, 713, 2129]. Liouvillian [1202, 1768].

Lippmann [543]. liquid [354, 1632, 1543]. liquid-vapor [1632]. liquids [29]. list [2053]. little [855]. Liu [1256]. lives [2217]. Livine [2109]. Lobatto [1586]. Local [60, 273, 1753, 616, 176, 2131, 136, 1716, 748, 1398, 712, 314, 357, 1685, 2004, 365, 1345, 1339, 1076, 1763, 446, 22, 1607]. localizable [309]. Localization [843, 800, 1698, 268]. localized [590, 847]. locally [1688, 183, 49, 618, 947]. LOCC [215]. Loève [1216]. Loewner [1842]. log [1175, 2119, 2199]. log-likelihood [2119, 2199]. log-partition [1175]. logarithm [2060, 1243]. Logarithmic [1096, 944, 912, 1856, 2137, 2105]. logarithmically [702]. logic [10, 1232, 2149]. logics [2036]. logistic [2160]. Long [1959, 683, 1546, 1757, 188, 80, 388, 1023, 1797, 1547, 1040]. long-range [1797]. Long-time [683, 1546]. long-wave [1757]. longitudinal [1131, 883]. look [1262]. loop [611, 1842, 1704, 1025, 2038, 2163, 1900, 676, 476, 834, 772, 1892, 399]. loops [217, 106, 1507, 1175]. Lorentz [468, 170, 913, 1607, 2001, 1167, 405, 192, 2154, 2101, 1480]. Lorentz-force-like [913]. Lorentz-violating [468]. Lorentzian [1499, 874, 880, 1302, 1390, 893, 1102, 1346, 2183, 1665, 1592]. Lorenz [1136]. Loss [1678]. Lotka [1473, 960]. Lovelock [1427, 752, 281]. Low [2039, 1946, 583, 1967, 843, 1486, 885, 1550, 916, 467, 64, 859, 560, 820, 1000]. low-dimensional [859]. Low-energy [2039, 1946, 467, 64]. Low-frequency [583]. Low-regularity [1967]. Lower [1615, 2195, 2114, 1520, 510]. Lower-dimensional [1520]. Lüders [1324, 1457, 1340, 2094]. LUE [385]. Luke [1114]. Lusztig [944]. Lusztig-dual [944]. Lyapunov [628, 518, 973, 788]. Lyra [1386, 1766].

Macdonald [2029]. macroscopic [1546, 483, 1980]. macrostates [689]. made [1479, 290]. magnet [659]. Magnetic [1141, 138, 476, 1008, 743, 631, 540, 865, 193, 686, 1850, 956, 615, 1838, 1384, 674, 900, 71, 1851, 580, 167, 161, 154, 13, 1045, 800, 817, 883, 507, 473]. magnetizable [1914]. magnetization [1719]. magnetoencephalography [1730]. magnetohydrodynamic [1577, 1069]. magnetohydrodynamic- [1069]. magnetohydrodynamics [1780, 322, 2058]. magnification [1799, 1978, 1981]. magnitude [1580]. Majumdar [580]. malicious [1171]. Mallows [2022]. Mandelstam [164]. manifold [841, 1942, 326, 585, 1080, 244, 1026, 743]. manifolds [709, 726, 151, 1932, 695, 1628, 1112, 1346, 1270, 81, 1897, 291, 546, 1515, 1463, 1257, 1412, 597, 1043, 557, 2162, 1145, 1414, 1835, 376, 1763, 2105]. Manin [1430]. Manipulation [1774]. Mann [698, 1000]. Many [1949, 930, 1224, 709, 938, 1189, 776, 519, 402, 1281, 838, 157, 2035, 2086]. Many-body [930, 1224, 776, 519, 402, 157, 2035]. many-mode [2086]. Map [754, 1453, 886, 2160, 1763, 1378, 2161]. Mapping [1312]. mappings [1992, 1720, 1336, 1846]. Maps [1379, 238, 1370, 997, 1263, 1374, 2070, 1658, 381, 1761, 1359, 574, 1298, 724, 728, 1955, 1911, 1378, 374, 1266]. marginally [893, 1499]. margins [111]. Markov [779, 1020, 1534, 563, 2136, 884].

Markovian [1317, 432, 657, 1237, 1654]. Mass [1181, 116, 1011, 1346, 764, 1993, 1336, 383, 185, 307, 1695]. masses [907]. Massive [931, 278, 911, 16, 676, 1296, 374]. Massless [1509, 880, 1385, 611, 1452, 1828, 18, 1127, 2023, 2130, 366]. master [2143, 1886, 1654, 43]. material [584]. materials [409]. Math [1588, 2201, 2195, 2216, 1498, 2206, 628, 441, 630, 2196, 1330, 1499, 2197, 1566, 302, 860, 2141, 2203, 2210, 2207, 1415, 1456, 987, 2215, 1365, 1609, 2212, 1236, 212, 2096, 2209, 211, 2214, 1608, 568, 900, 2095, 301, 467, 668, 2205, 2217, 2211, 2204, 347, 1758, 1790, 1329, 1834, 1116, 1117, 1940, 2198, 1414, 501, 2208, 37, 263, 1227, 2213, 440, 1118, 629, 38, 961, 934, 2097, 1907, 2199, 2202, 1157, 899, 2200, 104, 1941]. Mathematical [1759, 508, 164, 1680, 1690, 2142, 1863, 716, 1681, 942, 1170, 958, 1084, 535, 1169, 1955, 2153, 2185, 483, 1724, 1048]. mathematics [1061]. matrices [627, 1218, 1032, 1956, 993, 1973, 258, 257, 1794, 732, 2187, 1960, 1562, 2087, 395, 492, 839, 1976, 101, 1022, 1363, 1220, 1590, 1475, 497, 564, 1267, 924, 1808, 1287, 86, 2192, 134, 1966, 1862, 288, 1360]. Matrix [2155, 1028, 77, 842, 855, 1927, 1804, 1751, 2002, 732, 925, 1025, 658, 1593, 526, 445, 1938, 1514, 2021, 1854, 891, 162, 101, 1507, 782, 332, 781, 1720, 1441, 1395, 2097, 1907]. Matrix-valued [1028, 332]. matter [578, 1203, 1563, 1449, 376, 808, 1767]. Maurer [2193, 59, 251]. Max [386, 256]. Max-Plus [256]. Maxima [1444]. Maximal [1973, 109, 212, 1269, 2044]. Maximally [1007]. maximization [330, 2095]. maximizer [1431]. Maximizing [2117, 1883]. Maximum [1212, 864, 591, 1721, 1288]. Maxwell [178, 770, 1860, 1134, 852, 1814, 318, 124, 1047, 1529, 1664, 1872, 483, 2174]. Mean [1825, 1011, 1923, 1772, 2031, 87, 1675, 55, 1847, 2187, 217, 1881, 1685, 1682, 1844]. mean-chord [217]. Mean-curvature [1011]. Mean-field [1772, 1685]. means [222, 898]. measure [1087, 1947, 1661, 2175, 912, 2146, 801, 1219, 1696, 2194, 741]. measurement [1121, 1219, 2172, 1459]. measurements [1821, 829, 1331, 1333, 745, 12, 469, 701, 1285, 1457, 1082]. measures [1825, 163, 1674, 634, 311, 1522, 12, 978, 2061, 339, 261, 1646, 306, 2128, 1533]. mechanical [1513, 1017, 681, 1881, 1138, 1500, 671, 1211, 1890, 1874, 5, 263, 1278, 368, 65, 1139]. Mechanics [1670, 2014, 1305, 1610, 1643, 637, 872, 1866, 196, 1530, 346, 524, 506, 384, 108, 76, 1604, 523, 1501, 1197, 270, 1949, 1460, 1669, 1367, 1974, 371, 1849, 2102, 452, 1208, 612, 568, 1156, 1228, 1906, 1618, 528, 1207, 1122, 1693, 2016, 1701, 1975, 655, 367, 14, 744, 765, 1124, 83, 422, 720, 1433, 66, 1016, 804, 1459, 1840, 873, 2178, 1870]. mechanism [1797]. media [2059, 840, 550, 1486, 1054, 1749, 119, 1581]. medium [222, 1914, 1327, 424, 2206, 1994]. Melnikov [2113, 1650]. membrane [229, 1103]. membranes [2010, 77]. memory [696, 2110, 2161]. Menger [2166]. Meromorphic [753]. Mesner [1773]. Mesonic [1535]. mesoscopic [352, 355]. metallic [685]. metaplectic [850]. metastable [588]. Method [1898, 91, 1860, 641, 1792, 2113, 888, 1628, 2043, 197, 1077, 719, 1741, 1583, 1894, 891, 1600, 684, 839, 1449, 1528, 347, 335, 2144, 499, 2167, 323, 1026, 2127, 45, 1775, 716, 2126, 85, 619, 1783, 604, 462, 1040, 1724, 1928]. methods [1586, 972, 1970, 2164, 162, 1447]. metric [862, 911, 1509, 1166, 1498, 22, 645, 1164, 1277, 910, 1423, 774]. metric-affine [910]. metrics [968, 1302, 1381, 1811, 1204, 1427, 1209, 1344, 1221, 2068]. metriplectic [1137]. MHD [253]. MIC [828]. Michelson [677]. microlensing [1955, 2153]. micropolar [2088]. Microscopic [1994, 1518]. MICZ [939]. MICZ-Kepler [939]. Mills [785, 1116, 1406, 726, 803, 371, 1850, 1557, 72, 370, 873, 374, 1520]. Mima [342, 879]. mimetic [78]. minima [1556]. Minimal [1673, 836, 1567, 631, 1566, 1428, 1781, 944, 1280, 434, 825, 1231, 1856, 1254, 1039, 995, 2069]. Minimally [701, 72]. minimization [1721]. minimizers [357]. Minimizing [1812]. Minimum [365, 1024, 2097, 1907, 1676, 500, 1796]. Minimum-error [2097, 1907, 1796]. Minkowski [2215, 668, 1006, 1896, 530, 475, 635, 1251, 333, 1354, 1888, 1153, 2041, 488, 1445]. mirror [314, 471]. mirror- [471]. Mittag [1930]. Mittag-Leffler [1930]. Miura [737]. Miwa [693]. Mixed [2168, 123, 113, 2097, 1907, 1796, 1663, 617, 1146]. mixing [442, 1152]. mixtures [1633, 247]. mKP [1887]. modality [1949]. mode [125, 581, 1453, 121, 2086, 883]. model [1470, 610, 1213, 349, 611, 1991, 36, 1615, 2195, 380, 174, 2017, 1386, 613, 874, 1314, 354, 1634, 373, 1751, 16, 53, 1959, 2020, 1672, 70, 1465, 519, 402, 1281, 2046, 508, 1068, 475, 245, 242, 706, 27, 2182, 283, 1270, 1632, 2026, 1591, 92, 1044, 1365, 1583, 1491, 941, 1938, 951, 916, 1170, 958, 2121, 124, 1435, 171, 443, 1070, 266, 988, 1517, 223, 1852, 358, 1507, 502, 2051, 782, 1543, 407, 1993, 821, 1152, 1447, 1551, 2081, 882, 963, 952, 392, 294, 2130, 1169, 1720, 140, 919, 1288, 348, 1865, 656, 1227, 366, 1075, 168]. model [2022, 1859, 85, 1423, 1682, 8, 38, 241, 2012, 1313, 471, 485, 221]. modeling [1548]. Models [1610, 1643, 842, 627, 943, 908, 1762, 1733, 278, 759, 1183, 1804, 589, 1619, 749, 303, 280, 653, 188, 2028, 925, 1496, 719, 356, 658, 1989, 944, 30, 1248, 1180, 1881, 333, 973, 2163, 1823, 1854, 1065, 755, 974, 1099, 1685, 1546, 1069, 1408, 319, 1540, 781, 1437, 1856, 1021, 1929, 1649, 1476, 2103, 1395, 63, 2086, 1567, 642, 1538, 185, 669, 1293, 922, 1800, 871, 457, 1678, 2014]. modes [1452, 1768, 2116]. Modified [667, 961, 889, 1970, 1110, 242, 683, 1775, 388, 970, 1318, 894]. modular [769, 1238, 835, 1148]. Modulational [1899]. module [986]. modules [1413, 983, 1641, 2071, 2169, 1659, 1316]. Moduli [1550, 784, 2141]. Moerbeke [1441]. moist [730]. molecular [1501]. molecule [936, 884]. molecules [1314, 1443, 213, 393]. Moment [111, 1821, 1281, 1739, 1280, 672, 1646]. momenta [585, 994]. Moments [1862, 101, 1687]. momentum [1908, 608, 1912, 1077, 1190, 66, 1459, 606]. monodromies [909]. monodromy [1735, 2064, 1430]. monodromy-free [1735]. Monogamy [1645, 868]. Monomial [1320]. Monopole [313, 928, 476, 127, 1322, 806, 473]. Monopole-antimonopole [313]. monopoles [1850, 580, 743]. monotone [2097, 1907]. monotones [470]. monotonicity [779, 56, 1462]. Monte [398]. Moody [495, 729, 1527, 191, 1754, 638]. Moore [1057]. Morse [1456, 1334, 579, 1336]. Morse-theoretical [579]. Moser [2057, 93, 1855]. most [1697]. Motion [573, 614, 1329, 1918, 1726, 1174, 2164, 1713, 586, 1086, 1989, 814, 2010, 1153, 2122, 881, 2144, 1693, 2016, 1654, 1767]. motions [229]. Mott [1419]. Moutard [897]. Movable [1782]. move [2127]. moving [841, 2073, 886, 1933]. Moyal [869, 927, 2123, 143, 72, 383]. Mullins [1722, 2050]. Multi [1514, 255, 1869, 425, 470, 753, 1523]. Multi-Hamiltonian [1514, 753, 1523]. Multi-instantons [255, 1869]. multi-overlap [425]. multi-qubit [470]. multiboson [926]. multibranching [1598]. multicomponent [1025, 1605, 150]. multiconnected [1812]. multicuts [1869]. Multidimensional [1177, 962, 1195, 654, 1541, 1560, 1249, 1144]. multifractional [1104]. multilayered [1901]. multilevel [702]. multilinear [144, 2085]. multimaterial [1449]. Multimedia [1759]. multimonopoles [292]. multioverlaps [1671]. Multiparameter [2100, 1218, 1470, 627]. Multiparticle [1272, 1124]. multipartite [636, 1310, 446]. multiphase [320]. Multiple [460, 285, 2178, 1248, 676]. multiples [1275]. multiplet [1029]. multiplets [575]. Multiplicativity [1269, 109, 212, 574]. multiplier [898, 1560]. Multipolar [236]. Multipole [876, 1972]. multiqubit [541, 1083]. Multiresolution [1109, 196]. Multispecies [189]. multistate [627]. Multisymplectic [1814, 1586, 852, 1757, 1240, 421, 725]. multitime [560]. multivariable [1091]. multivariate [1552, 1179, 398]. multivector [421]. multivortex [20]. Muoniumlike [73]. mutual [469]. Mutually [993, 44, 1794].

Nagumo [58]. Nagy [65]. Nahm [2123]. Naimark [1404]. naked [514]. Nambu [1995, 121]. names [1691]. nanostructure [343]. nature [2092]. Navier [1758, 1060, 1052, 1049, 1771, 1652, 1055, 1050, 2049, 1558, 1755, 1051, 1653, 1053, 1353, 1988, 2180]. Near [279, 350, 1738, 1443, 1979, 1289]. near-horizon [1979]. Necessary [23, 160]. necessity [1906]. neck [1548]. need [1864]. negative [1747, 2146, 504]. negativity [793]. neighborhood [1086]. Nelson [1367, 1613, 366]. nerve [1446]. Nested [627, 257]. networks [1687, 1688, 1245, 1542, 1773, 1149, 1857, 1818, 1941]. Neumann [2170, 1444, 1127, 1195, 886, 1457, 1250, 2117, 334, 13, 563, 432]. Neumark [902]. neural [1687]. neutral [1985, 2052]. neutrino [1152]. neutrinos [907, 931]. neutron [2031, 846]. Neveu [225]. Newell [2171, 150]. Newman [556, 579, 774]. Newton [654, 1347, 1714]. Newtonian [1965, 618, 947]. Nichols [1297]. Nicole [610]. nilgeometric [8, 38]. Nilpotent [430, 1391, 1097]. nilradical [813]. Nine [1618]. NLS [193, 650]. No [18, 1125, 304, 2036, 44, 1008]. No-cloning [304, 2036]. Nodal [1045, 460, 1743]. Noether [637, 400, 207]. noise [1612, 975, 2075, 560, 185]. Non [468, 272, 651, 1447, 1510, 545, 450, 1808, 1927, 2131, 1317, 1599, 1392, 2182, 748, 2102, 2146, 1303, 477, 366, 667, 924, 618, 947, 1571]. Non-Abelian [272, 545, 450, 1392, 2182, 748, 1303, 477, 667]. Non-Borel [1447]. non-Fock [366]. Non-Hermitian [468, 1927, 1571]. Non-imprisonment [1510]. Non-Laplacian [651]. non-Markovian [1317]. non-negative [2146]. non-Newtonian [618, 947]. non-plane-wave [2102]. non-self-adjoint [1599]. Non-skew-symmetric [1808, 924]. non-ultra-local [2131]. nonanalytic [1255, 1395]. nonassociative [1110, 406]. Nonassociativity [1422]. nonautonomous [1706, 1513, 2049, 2173, 2048, 1630, 618, 947]. nonbacktracking [1287]. noncanonical [988]. noncentral [711]. nonclassical [269, 1836]. Noncommutative [1861, 613, 1970, 1424, 1361, 452, 407, 821, 1764, 1866, 530, 874, 1530, 1110, 1554, 258, 475, 2015, 328, 1849, 179, 1041, 1228, 403, 1152, 461, 477, 806, 474, 747, 129, 1963, 822]. noncommuting [1003]. noncompact [738, 948]. noncylindrical [2110]. nondecreasing [1668]. Nondegenerate [1264, 246, 1457]. nondistinct [1739]. Nondistributive [1604]. Nonequilibrium [815, 1656, 419]. Nonexponential [1738]. nonextensive [1604]. Nonholonomic [1742, 244, 1442, 1017, 523, 1839, 81, 1138, 1156, 2045, 2205, 82, 761, 1963, 1156]. nonhomogeneous [1638, 324]. Nonideal [1433, 1989]. nonincreasing [997]. noninertial [1207]. Nonintegrability [548, 763]. noninvariant [1209]. nonisentropic [2174]. Nonisolated [2183]. nonisospectral [396, 933]. Nonlinear [458, 780, 329, 1635, 331, 32, 2085, 661, 2062, 721, 278, 299, 60, 1726, 1132, 737, 239, 135, 537, 1106, 2008, 2170, 210, 2114, 590, 719, 1741, 1180, 1631, 2074, 1100, 568, 891, 1222, 892, 1787, 659, 184, 320, 1072, 1042, 2125, 1043, 1502, 1705, 267, 616, 83, 1999, 666, 1353, 114, 1846, 57, 1155, 1769, 63, 1450, 1101, 1882, 2200, 1783, 462, 1899, 1724, 1928, 1727, 894, 1033, 1877]. nonlinear-Schrödinger-type [1101]. Nonlinearizations [894]. Nonlinearizing [2045, 2205]. Nonlocal [1969, 293, 404, 325, 464, 60, 819, 796, 136, 1256, 321, 1125, 1373, 1339, 1886, 970]. nonlocality [419, 2214, 1407, 2202]. nonlocally [819]. Nonmalleable [1824]. nonmaximal [1983]. nonmerging [221]. nonmonotonic [1090]. nonorthogonal [304]. Nonperturbative [2038, 905, 2039, 43]. Nonrelativistic [1270, 2089, 1952, 1695, 1820, 553, 116, 1309, 2178]. Nonresonant [883]. nonscalar [272]. nonseparable [1332]. nonskew [1267]. nonsmooth [254]. Nonstandard [1516, 455, 1877, 498, 1162, 746, 1157]. nonstationary [1518, 854]. nonsymmetric [467, 64]. nonthreshold [816]. Nontopological [124, 1889]. nontranslational [611]. Nontrivial [1906]. Nonunitary [981, 1218]. nonvanishing [891, 666]. Nordstrom [1768]. norm [1269]. Normal [226, 2004, 349, 2116, 336]. normed [2166]. norms [1321, 109, 212, 728]. Note [1588, 628, 630, 302, 860, 212, 875, 2096, 2055, 2071, 301, 2205, 347, 1066, 501, 629, 38, 1941, 857, 1703, 131, 371, 1236, 2149, 2060, 1985, 558, 766, 2068, 2111]. Notes [1160]. notion [1712]. Novel [519, 402, 590, 1394]. nuclear [1723]. Null [1027, 137, 1170, 2107, 1169]. Number [1149, 55, 1050, 308, 633, 1983, 1355]. numbers [885, 1105, 336]. Numerical [1381, 356, 1052, 1447]. NUT [177]. NUT-like [177]. Nutku [1498, 1166].

objective [1435]. oblate [678]. oblique [1282]. Observability [249]. observable [2007, 1338]. Observables [1183, 606, 503, 2120, 111, 2149, 805, 1865, 1504, 1564, 1806]. observation [54]. Observers [805, 1341]. obstacle [1055]. obstacles [1482]. obstructions [1347]. obtain [1783]. obtained [1260]. obtaining [197, 2144]. Obukhov [342]. occupied [1368]. occurrence [1090]. octagons [100]. octapolic [1730]. octet [698]. octic [1098]. Octonic [1710, 1701]. octonionic [1110]. Odd [1409, 1307, 247, 257]. off [1464, 1797]. off-diagonal [1797]. off-shell [1464]. Ohm [1047]. Okubo [698]. On-diagonal [433]. on-resonance [1793]. on-shell [1464, 1925]. One [1986, 125, 554, 145, 506, 1342, 901, 706, 1419, 1771, 686, 2038, 546, 157, 612, 676, 358, 1051, 1231, 89, 1657, 271, 1135, 2025, 2180, 268, 653, 239, 1356]. One-dimensional [1986, 145, 1342, 706, 1771, 157, 358, 89, 271, 2025, 268, 653, 239, 1356]. one-loop [2038, 676]. one-mode [125]. one-parameter [901]. one-peak [1135]. ones [732]. only [686, 312]. Ono [1718, 788]. Onsager [1068]. Open [535, 236, 1317, 226, 952, 1338, 43]. Open-closed [535]. operation [2094]. Operational [218]. operations [448, 1340, 446]. Operator [316, 1633, 163, 174, 1947, 444, 170, 634, 506, 311, 1522, 12, 99, 1901, 1376, 2037, 1761, 981, 2087, 1384, 1936, 2065, 674, 900, 1255, 543, 1851, 595, 672, 306, 1662, 1602, 2204, 390, 1529, 1664, 592, 692, 1124, 1045, 181, 1666, 664, 268, 298]. operator-valued [306]. operators [1096, 976, 541, 1809, 1307, 895, 1853, 1452, 902, 166, 799, 103, 1358, 1599, 433, 1451, 1934, 957, 1704, 141, 1234, 2098, 928, 1897, 1362, 1737, 115, 1164, 801, 1578, 443, 1555, 2122, 339, 111, 1646, 1507, 994, 1793, 1573, 332, 500, 1300, 765, 1298, 1289, 1021, 846, 970, 231, 1403, 995, 259, 1083, 1963, 2009, 774, 850]. Oppenheimer [393]. Optical [644, 2102, 1957]. optics [532, 347, 335, 426, 1117, 1882, 2200]. Optimal [1111, 503, 1459, 1259, 1279, 1611, 398, 1775, 1436, 39, 1082]. optimality [740]. optimally [847]. Optimization [718, 425]. orbifolds [1275]. orbit [1180, 1362, 365, 800]. orbital [678]. orbits [681, 838, 1097]. Order [1329, 1799, 1526, 1474, 786, 993, 959, 1524, 239, 1725, 2193, 1018, 419, 99, 486, 1736, 2187, 1782, 1343, 1702, 1894, 1182, 1987, 2121, 115, 1208, 1776, 199, 200, 394, 555, 603, 758, 921, 2190, 2211, 1326, 881, 2144, 1693, 2016, 319, 1701, 2052, 1411, 1886, 1723, 1797, 915, 2161, 1570, 507, 1446, 1033, 2050]. ordering [226, 336]. Orders [986, 233, 1399, 1650]. ordinary [1837, 1736, 1782, 1969, 11, 1495]. Ornstein [687, 387]. orthoalgebras [509]. orthoatomistic [1477]. orthochronous [405, 192]. Orthogonal [1354, 1475, 494, 1734, 2135, 1481, 12, 526, 1938, 1922, 215, 1082, 466, 1287, 2077, 1815]. orthogonally [1345]. orthonormalization [1226, 980]. oscillating [1211]. oscillations [2159]. Oscillator [1336, 1388, 759, 2216, 1661, 1342, 2184, 1191, 1390, 861, 946, 539, 1594, 1781, 253, 2055, 493, 969, 401, 1545, 1162, 1597, 1711, 1904, 1877, 67, 1254, 1476, 1434, 542, 2077]. oscillators [468, 1572, 537, 1419, 818, 639, 1874]. Oseen [357]. OSRC [604]. Osserman [291]. Ostrogradsky [2127]. OSV [836]. other [2031]. output [1269]. Overlap [783, 425, 1154].

packet [2005]. packets [998, 75]. packings [1436]. Page [24]. Painlevé [645, 1343, 693, 1225, 1753, 2139, 1602, 2016, 898, 1729, 1101]. pair [343]. pairs [2131, 1868, 2079]. pants [90]. Papapetrou [807, 580]. parabolic [71, 2190, 2211]. parabosons [1262]. parafermionic [2024]. parallel [1847, 1199, 1892]. parameter [1526, 901, 794, 1976, 1587, 1243, 1210]. parameterization [2118]. parameters [1961]. Parametric [828, 1274, 1135]. Parametric-time [828, 1274]. Parametrization [645, 862, 345, 395, 492, 482, 629]. parametrized [950, 270]. paraparticles [915]. parity [170]. Parr [302, 203]. Parseval [327]. part [615]. Partial [149, 1155, 1809, 233, 819, 737, 1579, 1789, 34, 975, 1285, 2128, 1129, 1205, 1397, 1724]. partially [189, 582, 1879, 2058, 953, 1144]. Particle [2124, 874, 506, 631, 229, 861, 2196, 1656, 326, 573, 286, 161, 1108, 67, 906, 787, 1254, 517, 750, 1889]. particles [236, 1131, 424, 2206, 938, 880, 701, 2108, 1994, 45, 731, 1505, 525, 772]. particular [275, 1170]. Partition [474, 2012, 2017, 1687, 1751, 1672, 643, 958, 1471, 1175]. partitions [87, 1739, 2024]. partly [846]. passage [1688]. Passive [1341]. past [1055]. Pasta [1014]. Path [41, 52, 936, 569, 691, 679, 398, 1326, 1597, 1231, 1002, 804, 1818, 1941]. Paths [2024, 1538]. patterns [97, 1783]. Pauli [172, 444, 373, 1973, 1421, 825, 497]. PDE’s [552]. peak [1135]. peakons [1715]. peaks [1790, 812]. peculiar [1689]. Peel [1291]. Peierls [188]. penetrable [604]. perceptron [1682]. percolation [1688]. Pereira [2109]. Perelomov [1415, 1233]. perfect [1386, 733, 581, 1702, 1644, 1717, 1394]. performance [1611, 604]. period [1142]. Periodic [242, 2129, 1551, 681, 539, 550, 1486, 838, 573, 240, 2049, 956, 1071, 1997, 1578, 1113, 308, 633, 755, 974, 1517, 1756, 52, 626, 1529, 800, 5, 1967, 1779, 774, 1660, 1545]. periodically [390, 69]. Periods [1829]. permittivity [1285]. permutation [1709, 1275]. permutative [1902]. perpendicular [1841]. Perron [339]. perspective [474]. Perturbation [1413, 1033, 816, 412, 1078, 1349, 1837, 99, 1738, 1376, 1600, 598, 1783, 507, 1832]. perturbations [554, 1509, 1202, 282, 2098, 649, 1285, 378, 548, 1013]. Perturbative [2104, 43, 1464, 228]. Perturbed [839, 1062, 854, 1600, 1058, 1529, 1664, 58]. Petiau [2196, 1342, 641, 1108]. Petrova [1932]. Petviashvili [567, 725, 1318]. Phase [169, 1078, 672, 1557, 2033, 744, 1698, 1495, 572, 1307, 316, 2176, 245, 1671, 1197, 270, 1632, 249, 2003, 29, 1841, 111, 1543, 1478, 1571, 187, 154, 1975, 1689, 271, 699, 265, 1864]. Phase-integral [572, 699]. phases [2140, 492, 2192]. phenomena [1864, 1660]. Phillips [65]. photon [1371, 1816, 612, 1355, 1382, 1774]. photons [1008]. Phys [1588, 2201, 2195, 2216, 1498, 2206, 628, 441, 630, 2196, 1330, 1499, 2197, 1566, 302, 860, 2141, 2203, 2210, 2207, 1415, 1456, 987, 2215, 1365, 1609, 2212, 1236, 212, 2096, 2209, 211, 2214, 1608, 568, 900, 2095, 301, 467, 668, 2205, 2211, 2204, 347, 1758, 1790, 1329, 1834, 1116, 1117, 1940, 2198, 1414, 501, 2208, 37, 263, 1227, 2213, 440, 1118, 629, 38, 961, 934, 2097, 1907, 2199, 2202, 1157, 899, 2200, 104, 1941]. Phys. [2217]. Physical [1423, 208, 1822, 1489, 686, 703, 1275, 549, 899]. Physics [1690, 874, 1487, 1488, 1489, 1490, 535, 561, 1724, 1759, 2142, 1863]. Picard [314, 909]. Pickands [1037]. piecewise [1935]. piecewise-constant [1935]. pinch [1548]. Pinning [349]. pipe [1933]. Pitaevskii [1190, 2139, 1113, 1356, 1455]. Planar [247, 685, 1841, 960, 2093, 561, 656, 1227]. Planck [589, 1092, 1091]. Planckian [407]. Plane [1582, 1861, 798, 1726, 1174, 180, 441, 1191, 2090, 1049, 705, 594, 2102, 1885, 72, 677, 440]. planes [143]. plasma [1860, 1595, 300, 883]. plasmas [2058]. plate [718]. plates [1199]. playground [1061]. plurality [1380, 1848]. plus [220, 256]. Podolsky [1229, 1457]. Pohozaev [1826]. Poincaré [2089, 681, 1174, 1424, 1182, 999, 949, 1832, 1951]. Point [1061, 1126, 653, 841, 1307, 1063, 1068, 897, 1831, 1631, 1577, 1058, 1290, 69, 609, 1983, 2044, 292, 915, 1772, 1010]. pointed [955]. pointlike [2115]. points [1325, 1279, 2183, 769, 243, 2094]. Pointwise [746, 1157]. Poiseuille [1802]. Poisson [1092, 2209, 1329, 438, 1414, 1352, 1380, 2131, 1454, 1492, 410, 289, 1184, 1771, 1591, 421, 2112, 522, 1848, 2055, 1558, 1755, 1958, 133, 2132, 881, 1587, 858, 560, 1145, 1985, 2174]. polar [1191]. Polarizability [778]. polarizable [1914]. polarized [953]. polaron [375]. polyatomic [213, 393]. polygons [2090]. Polyhedra [1260]. Polyhedral [439, 2172]. polylogarithms [676]. polymers [1673]. Polynomial [881, 1329, 1856, 617, 193, 2002, 1813, 2210, 1986, 1146, 243]. polynomials [855, 1400, 170, 1481, 925, 2029, 1926, 1938, 493, 262, 1475, 466, 1287, 1309, 2077, 1815, 742, 1862]. Polytropic [1088]. polyvector [803]. polyvector-valued [803]. population [1960]. porous [2059, 1054]. portraits [1841]. posedness [1184, 1715, 1292, 1190, 1728, 1807, 2134]. Position [995, 608, 764, 1336]. position-dependent [764, 1336]. positions [262]. Positive [518, 628, 1346, 163, 1947, 634, 311, 1522, 515, 1133, 131, 12, 1747, 1761, 306, 2128, 510, 1298, 728, 1911, 1378, 2161, 4, 934]. positivity [1462, 952]. positron [959]. positron-electron [959]. Positronium [959, 1723, 2091]. possibility [54]. Possible [487, 1797, 2189]. Postbuckling [1393]. posteriori [1052]. posts [233]. postulates [1457]. Potential [1640, 1568, 1784, 936, 464, 854, 62, 2196, 173, 1106, 1334, 1579, 705, 767, 1563, 573, 2177, 2207, 2030, 1456, 1938, 1070, 1469, 71, 1113, 261, 1946, 1019, 504, 243, 267, 1108, 1529, 1944, 1231, 220, 312, 661, 1160, 699, 2054, 1450, 1249, 1903]. potentials [1568, 1809, 76, 832, 964, 180, 441, 193, 1668, 831, 1335, 1735, 707, 308, 633, 1455, 1693, 2016, 1339, 500, 25, 2191, 787, 463, 42, 5, 263, 1366, 1434, 1155, 440, 1395, 1274, 721, 1093, 1249, 711, 1727, 254]. Potts [245, 882, 1720]. power [1090, 2197, 1794, 1245, 1921, 1844, 1679]. power-law [1679]. power-law-decaying [1844]. practical [2113]. Prague [1690]. Prasad [1937, 2096]. pre [2099]. pre-Lie [2099]. precession [678]. preferential [1245]. prehomogeneous [1107]. Preissman [1814]. preliminaries [555]. Preliminary [1776]. preparation [1338]. prepotential [1823]. prequantization [784, 2141]. prescribed [1302]. prescriptions [950]. presence [315, 1421, 1781, 1879, 1113, 825, 720, 45, 1649]. presentations [2189]. preserving [486, 304, 2064, 1379, 728]. pressure [1290, 1988]. prime [1307, 1794]. Principal [1740, 265, 1313]. principle [1311, 851, 1919, 1367, 1721, 903, 1081, 164, 1344, 559, 1024, 420, 1288]. principles [183, 1839, 789, 1425]. print [1692]. probabilistic [2166, 1079]. probabilities [1927, 722, 829, 940, 1331, 1333, 1879]. Probability [637, 1822, 1419, 2132, 1883, 1975, 1369, 1844]. problem [54, 1860, 791, 423, 840, 660, 776, 591, 180, 441, 1349, 2008, 838, 210, 878, 1419, 282, 486, 1875, 217, 1055, 1236, 296, 845, 2074, 1935, 2129, 1841, 330, 2095, 1982, 793, 93, 672, 1019, 659, 577, 1910, 33, 1445, 1345, 1531, 616, 336, 187, 319, 13, 904, 434, 1322, 1053, 426, 1117, 1775, 813, 57, 440, 1128, 146, 39, 640, 650, 1013, 1114, 1034, 2054, 1663, 766, 1246, 2050]. problems [1943, 1501, 1837, 718, 425, 1493, 1335, 886, 1073, 1362, 693, 15, 1206, 171, 568, 684, 939, 1621, 1948, 2167, 83, 32, 1686, 763, 1309, 1482, 894]. Proca [1870]. procedure [1226, 980, 1579, 1256, 738, 1392, 1971, 1249]. Procesi [1259, 1807]. process [1371, 1471, 387, 2018, 338]. processes [1598, 1733, 189, 529, 687, 53, 779, 1215, 1212, 1994, 2072, 884, 248]. produced [1933]. product [494, 316, 697, 1413, 1086, 927, 1461, 2021, 304, 1186, 497, 558, 1549, 850]. production [1024, 952, 657]. products [36, 880, 1519, 1097, 149, 2147]. profile [1778, 526]. profiles [1520]. programming [309]. projection [141]. projective [2155, 1942, 1700, 426, 1117, 1082, 165]. projector [1383]. projectors [627, 257]. Prolongation [2179]. Proof [656, 1227, 110, 1000, 1170]. proofs [904]. Propagating [810]. propagation [841, 1825, 61, 1410, 1686]. Propagator [832, 964, 938, 1189, 361, 1074, 1230, 1007]. Propagators [186, 1412]. proper [405, 192]. Properties [1901, 174, 935, 1784, 170, 316, 169, 1371, 1695, 2135, 155, 1009, 527, 1055, 1953, 801, 469, 1384, 595, 1685, 1326, 269, 332, 483, 1467, 1843, 2118, 268]. property [673, 1917, 1596, 563, 432, 250, 2136, 743]. pseudo [1847, 1112, 367, 671, 739, 966]. pseudo-Euclidean [1847]. pseudo-Hermitian [367]. pseudo-Hermiticity [671, 739]. pseudo-Riemannian [1112, 966]. Pseudodifferential [1255, 103, 1358, 123, 1217]. Pseudoduality [2103]. pseudogroups [59]. pseudorelativistic [1384, 175]. pseudospin [1128]. pseudosupersymmetry [63]. Pseudosymmetry [1201]. PT [971, 308]. PT-invariant [308]. Publisher [1588, 628, 630, 302, 860, 212, 2096, 301, 2205, 347, 501, 629, 38, 1941].

Puiseux [1650]. Pullback [1762]. pulsars [201, 1400]. pulse [465, 1551]. punctured [713]. Pure [69, 415, 1148, 2037, 2146, 826, 639, 1883, 2204, 365]. pure-state [2037, 826, 2204]. Purely [1588, 1503]. pyramidal [47].

Q [120]. Q-ball [120]. QCD [74]. QED4 [1200]. QMA [1076]. quad [724]. quad-graphs [724]. Quadratic [2145, 1816, 630, 551, 2138, 1709, 585, 1080, 206, 1893, 727, 133, 341, 1139]. quadratures [1473]. quadrilateral [1252]. quadruple [1491]. Qualitative [1649]. Quantitative [444]. quantities [1489, 52, 626]. Quantization [1942, 1028, 1914, 1732, 272, 389, 770, 1572, 355, 371, 7, 1609, 903, 273, 704, 266, 672, 1162, 1238, 597, 161, 696, 1623, 2127, 50, 45, 1145, 1414, 1622, 1293]. quantizations [471]. quantized [125, 1423, 606]. Quantum [397, 637, 125, 940, 1913, 16, 866, 1370, 605, 2176, 929, 1080, 689, 670, 206, 712, 106, 823, 1534, 2175, 2063, 769, 2014, 988, 93, 1363, 2188, 1046, 1122, 2108, 1610, 1643, 392, 1168, 271, 305, 1288, 5, 263, 924, 505, 2185, 1431, 1870, 430, 736, 46, 1612, 862, 2148, 937, 1786, 1588, 1503, 2100, 722, 872, 468, 1824, 829, 1331, 1333, 1311, 2089, 233, 759, 363, 530, 196, 1494, 2140, 1530, 218, 1760, 506, 569, 343, 108, 992, 169, 589, 1619, 103, 1358, 76, 733, 1464, 1943, 1174, 600, 1191, 135, 3, 1330]. quantum [723, 1364, 697, 1451, 1501, 1695, 326, 1645, 1261, 1819, 453, 1827, 1603, 539, 1159, 1614, 1163, 1197, 1165, 1788, 708, 503, 1284, 794, 1917, 885, 270, 1704, 10, 1949, 1460, 1488, 1791, 1181, 1161, 2037, 944, 1974, 1849, 214, 56, 1081, 2102, 487, 830, 1461, 157, 2027, 2214, 1310, 452, 698, 116, 2104, 439, 531, 612, 1992, 469, 1196, 448, 162, 1873, 2120, 40, 555, 758, 921, 414, 896, 306, 449, 1697, 1228, 1909, 1906, 994, 1618, 158, 228, 502, 344, 1186, 1546, 2149, 2035, 818, 1545, 2194, 159, 1500, 2204, 1299, 1693, 2016, 390, 2081, 609]. quantum [834, 1701, 2036, 1975, 1680, 563, 1945, 367, 671, 14, 740, 341, 744, 952, 1535, 147, 772, 294, 765, 1124, 800, 66, 67, 2101, 919, 1865, 2, 37, 375, 1647, 1278, 804, 1459, 542, 1107, 1403, 1569, 564, 88, 429, 979, 1120, 368, 65, 1369, 1378, 1915, 39, 1654, 447, 73, 185, 669, 873, 747, 2097, 1907, 1796, 2032, 265, 743, 989, 2077, 742, 1407, 2202, 1564, 2007, 1338, 1806, 920, 1996, 2178, 741, 43, 1425, 1239, 2142]. Quantum-inspired [1431]. Quantum-mechanical [5, 263]. Quantumlike [1079]. quark [174]. quartic [2166, 1938, 1557, 639, 1597]. quartic-cubic [1557]. Quasi [2078, 1197, 1496, 935, 318, 1845, 1916]. Quasi-exactly [1496]. quasi-free [935, 1916]. Quasi-Hermitian [1197]. quasi-Hopf [1845]. Quasi-Lie [2078]. quasi-Maxwell [318]. quasiclassical [427]. quasicoherent [895]. quasideterminants [1362]. quasiexact [1816, 555, 921]. Quasiexactly [1271, 1272]. Quasifinite [1358]. quasifree [1244, 1374, 1533, 1237]. quasigeostrophic [1070]. quasigraded [667]. quasihomogeneous [1469]. Quasilattice [1803]. quasilinear [460, 1638, 296, 1999]. quasilocal [1011]. quasinormal [1768]. Quasiperiodic [1939, 682, 518, 628, 240, 33]. Quasiseparation [865]. quasispherical [176]. quasistable [1832, 1951]. quasisymplectic [597]. quaternion [1986, 162]. Quaternionic [767, 998, 596, 1587, 108, 705, 162, 833, 1004, 1745, 2213]. quaternions [1198, 2001, 1260]. qubit [216, 1611, 470, 574, 365]. qubits [1192, 1813, 2210, 2003, 1696, 867]. qudits [1838]. quench [2026]. Quintics [1497]. Quiver [1845, 477].

Rabi [1816]. Racah [101, 1022, 566]. radial [976, 41, 774]. radiance [1733]. Radiating [1767]. Radiation [1327, 1172, 2073, 373, 615, 378, 415, 2115]. radiative [1315]. Rakhmanov [2095]. Ramond [178, 983]. Random [233, 1670, 425, 1305, 1287, 915, 1966, 1961, 1927, 783, 1804, 1842, 354, 1688, 1672, 526, 801, 1879, 1677, 2122, 101, 1022, 1683, 782, 1066, 500, 1686, 1689, 1617, 1441, 86, 1160, 1772, 1862, 1743, 1818, 1941, 1679, 1955, 2153]. randomness [1210]. range [1959, 832, 964, 188, 5, 263, 1797]. Rank [1608, 554, 600, 1337, 1036, 69, 601]. rank- [69]. rapid [312]. Rapidly [965, 1188, 254]. rare [582]. Rate [1755, 1317, 1997, 2125, 489, 1119]. rates [1321, 1991, 1458, 1728]. ratio [2197, 1921]. Rational [1220, 1736, 1735, 1693]. ratios [1862]. ray [201, 1400]. Raychaudhuri [2043]. Rayleigh [84]. rays [1953, 415]. reaction [189, 529, 959, 2051, 1723, 1630, 1172]. reaction-diffusion [189, 529, 1630]. reactive [1449]. reading [1861]. Real [630, 551, 361, 1178, 2197, 1112, 1550, 156, 512, 960, 1921, 312, 894]. real-valued [894]. Realization [1819, 46, 1740, 671, 534, 1038]. realizations [1737, 439, 1754]. really [1704]. reciprocity [1957]. recollapse [319]. Reconstruction [976, 209]. reconstructions [902]. recovering [945]. Recovery [2172, 1155, 1678, 732]. rectangular [1147, 1446]. recurrence [1683]. recursion [1970, 1666, 970]. Recursive [62, 1773, 492, 1185, 395]. redefinition [836]. redshift [2134]. Reduced [570, 1065, 1566, 1428, 4, 690, 934, 1692]. reducible [400]. reducing [267]. Reduction [1638, 399, 1352, 990, 303, 2170, 2176, 814, 1639, 1894, 21, 469, 1600, 692, 1929, 1666, 1886, 806, 1296, 1916, 949]. reductions [418, 1225, 867, 1520]. reductive [598]. redundant [327]. reference [1370, 1122]. Refining [1731]. reflecting [1879]. Reflection [541, 1462, 1176, 840, 964, 194, 632, 1282, 651, 1536, 1267]. reflectionless [707]. reflections [494]. Regge [22, 504]. regime [1687, 802, 1486, 2051]. regimes [588]. regular [1773, 127, 1962, 2009, 572]. regular-singular [2009]. Regularity [2088, 1052, 1953, 330, 2095, 1051, 1967, 2180, 1578]. regularization [738, 1875, 640]. regularized [1757, 1383, 1247]. regulated [230]. Reissner [1768]. related [208, 895, 1537, 819, 523, 982, 2182, 217, 306, 1363, 502, 344, 1602, 294, 491, 1298, 505, 638, 298]. Relating [1464]. Relation [121, 1023, 1121, 1159, 495, 1408, 2136, 1237, 787, 1903, 921]. Relations [1321, 196, 1327, 552, 2159, 1868, 1110, 1398, 1258, 698, 113, 2006, 1296, 1911, 1039, 1567, 259, 1504]. relationship [347, 335, 1928]. relationships [1793]. relative [363, 1443, 1379, 1945]. Relativistic [1438, 1713, 401, 833, 1004, 2213, 1120, 1832, 1951, 236, 872, 1143, 17, 641, 76, 1281, 2164, 1292, 124, 2010, 1153, 654, 543, 576, 1701, 2108, 1855, 387, 375, 804, 1840, 1309, 883, 1767, 711, 2178]. relativity [1011, 1508, 1953, 452, 1885, 1618, 910, 79, 545, 78, 1207]. relaxation [1994, 323, 1211]. relic [2140]. remark [1674, 1982]. Remarkable [502]. Remarks [410, 1780, 1092, 1250, 13, 1450, 1101, 49, 1404, 2118]. Removable [785, 1116]. Removal [969]. Renormalization [1200, 1628, 623, 116, 2066, 1561]. reparametrization [1003]. repeatedly [1291]. Representation [673, 1474, 1398, 223, 3, 1330, 1077, 1489, 1667, 1359, 2120, 266, 1636, 1079, 449, 476, 1710, 1067, 66, 305, 366, 1564, 741, 485, 850]. Representations [103, 602, 794, 1484, 466, 430, 736, 1006, 855, 36, 209, 1273, 1358, 1078, 625, 1747, 1404, 295, 339, 1902, 205, 301, 566, 405, 192, 408, 1097, 1322, 2152, 1002, 1476, 231, 753, 536, 853, 1951, 966, 920, 35]. Reproducing [1641]. Republic [1690]. repulsive [410]. require [304]. resistances [1542, 1773]. resolvent [595]. Resonance [768, 1544, 1793, 702, 849, 1864, 1547, 1040]. Resonances [1851, 393, 125, 1784, 1161, 969, 368, 65, 413]. resonant [1225, 1013]. resources [796]. respect [864, 1526, 2181, 1122]. Response [441]. rest [1435]. Restricted [2017, 1223]. result [985]. results [1809, 1494, 2150, 706, 1253, 574, 331, 294, 2058, 1104, 2159]. resummation [62, 1249]. retarded [1944]. reversal [216, 2034]. Reverse [571]. reversibility [884]. reversible [779, 2051]. review [2, 37]. revised [53]. revisited [1088, 1402, 496, 1193, 1199, 1276, 970, 297, 1944]. revolution [1279, 15, 912, 1434]. Reynolds [1050]. rhomboidal [838]. Riccati [239, 1315]. Ricci [2096, 1302, 617, 1548, 546, 1937, 1146, 1442, 1963]. Rice [2153]. Richardson [2026]. Riemann [2141, 771, 617, 784, 1009, 1506, 1508, 897, 1591, 691, 1440, 1968, 244, 1146, 376, 742]. Riemann-flat [376]. Riemannian [709, 1391, 1752, 423, 1964, 1390, 1554, 1112, 1270, 1412, 516, 751, 966]. Riesz [1279, 1326]. right [1584, 907]. right-handed [907]. Rigid [384, 515, 2000]. Rigorous [506, 373, 1194, 1676, 110, 1479, 1187, 1472, 1120, 2150]. Rikitake [1389]. ring [1874]. rings [1393, 1856, 313]. Robertson [319, 548, 1538, 1918, 280, 1102, 811, 1277]. Robinson [1466, 775]. robustness [1149, 1857]. rod [1997, 1021]. Role [1187, 1457]. root [596, 480]. roots [430, 736, 258]. Rosen [1456, 1334, 230, 1229, 1457]. Rosochatius [1223]. Rotating [1574, 1466, 1703, 965, 1188, 941, 1071, 443, 1455, 1874, 1072]. rotating-wave [443]. rotation [40, 969, 1322]. Rotational [1443, 1190]. Rotational-vibrational [1443]. rotationally [1370, 1328, 817]. Rotations [1368]. rotor [1793]. Rouet [1132, 1555]. rough [2011]. roughening [1722]. Routh [1392]. Rovelli [2109]. Ruelle [1087]. rule [2197, 1511, 1920, 1921, 984, 1033]. ruled [1046]. rules [1261, 1820]. Runge [2162]. Rusk [181].

S [2095, 1725]. Safko [2108]. Sahi [985]. Sample [1326, 1960]. Sasa [970, 1777]. Sasakian [2068]. satisfiability [1686]. Satsuma [970, 1777].

Saturated [165]. Savadkouhi [2198]. sawtooth [238]. Saxton [1708]. Scalar [472, 1203, 675, 1519, 2041, 1409, 581, 1361, 811, 1429, 56, 676, 1906, 512, 2005, 1300, 1067, 715, 1649, 98, 1872, 746, 1157, 711, 2068, 473]. scalar-tensor [715, 1872]. Scale [662, 1109, 2164, 283, 1639, 1142, 1066, 545, 817, 471, 1818, 1941]. scale-free [1818, 1941]. scales [435, 1962]. Scaling [1801, 1501, 1187, 2167, 1395]. scatterer [1901]. scatterers [2172]. Scattering [911, 854, 798, 1668, 213, 154, 984, 962, 1784, 1235, 1695, 885, 652, 1901, 768, 2084, 30, 249, 1583, 1268, 2102, 1635, 891, 654, 467, 64, 1946, 1787, 267, 367, 2108, 666, 930, 1224, 5, 263, 388, 747, 1034, 1663]. SCD [97]. Schatten [1321]. Schauder [820]. scheme [852, 1814, 1757, 897, 725, 1779]. schemes [2078, 1819, 1252, 362, 1542, 740, 1420]. Schmidt [1337]. Schouten [533]. Schramm [1842]. Schrieffer [1808]. Schrieffer-type [1808]. Schrödinger [628, 2207, 900, 438, 2200, 1809, 1964, 518, 854, 1599, 433, 329, 135, 865, 460, 1106, 1778, 210, 2114, 1184, 2070, 1077, 2177, 2030, 1335, 1376, 2098, 764, 15, 296, 703, 1998, 2087, 1164, 801, 340, 1578, 891, 701, 674, 1851, 659, 1662, 320, 1373, 1484, 1042, 1043, 1502, 1531, 332, 154, 1339, 500, 1045, 1657, 220, 666, 1846, 1624, 1155, 1769, 661, 1495, 307, 977, 1714, 1450, 1101, 1882, 1899, 1727, 894, 1033, 298]. Schur [1471]. Schwarz [225]. Schwarzschild [1134, 22, 2010, 577, 2217]. Schwinger [1884, 1507, 559]. screen [685]. search [1716]. Second [693, 199, 200, 394, 603, 758, 45, 2052, 207, 272, 1454, 959, 239, 1572, 1878, 99, 1736, 1782, 115, 152, 555, 921, 101, 1701, 1729, 26, 1723, 1570, 1139]. second-class [272, 1454]. second-grade [1878]. Second-order [199, 758, 2052, 959, 239, 1782, 555, 1701, 1723, 1570]. section [885, 249, 2142]. sections [1706]. sectors [74]. Segal [571, 1404, 538]. segment [1629]. segregation [53]. Segur [2171]. Seiberg [275, 499, 806, 1148]. Selberg [799]. selected [1784]. Selection [1261]. Self [1234, 1026, 575, 1595, 18, 770, 1599, 852, 20, 897, 1252, 1453, 1429, 124, 1891, 756, 1553, 1984, 1546, 1601, 234, 1368, 644, 1520]. self-adjoint [852, 897, 1252]. Self-adjointness [1234, 770, 1553]. self-consistent [1891, 1601, 1368]. Self-consistent-field [1026]. self-dual [575, 20, 124, 1520]. self-force [644]. self-gravitating [756, 1984]. self-interacting [1429]. self-interaction [18]. self-similar [1595, 1453, 1546]. self-similarity [234]. Semi [1952, 203, 302, 882, 884]. semi-infinite [203, 302, 882]. semi-Markov [884]. Semi-nonrelativistic [1952]. semibounded [797]. Semiclassical [938, 1189, 193, 607, 802, 700, 361, 1288, 1774]. semiconductors [919, 1288]. semidefinite [309]. Semidensities [1454, 857]. Semidirect [620]. semigroup [368, 849, 1951]. semigroups [866, 856, 432, 846, 65]. Semilattice [344]. semilinear [1483, 1940]. semilocal [117]. semirelativistic [1620]. semirings [1521]. semisimple [601]. Semispectral [1646]. sensitivity [2075]. separability [1310]. separable [1328, 594, 2085, 1996]. Separation [530, 732, 1804, 137, 552, 590, 1112, 1354, 921, 2033, 1242]. separations [1802]. sequence [627, 257]. sequences [1861]. sequential [1461, 1186]. series [1817, 1650, 1014, 2057, 983, 481, 240, 676, 405, 334, 392, 144]. servomechanisms [523]. sesquilinear [630, 551]. sesquilinear-quadratic [630, 551]. set [233, 1416, 2187, 1926, 1046, 167, 945, 2147]. sets [710, 521, 1279, 454, 912, 1756, 904, 1045, 1378, 1743]. seven [2123, 255]. several [1492, 1813, 2210]. sewing [1009]. sextic [1735]. Shabat [1286]. Shannon [1243, 538]. Shape [1823, 827, 1111, 831, 1544, 1790, 262, 11, 812, 1135, 1567]. Sharp [1502, 902]. sharpness [456]. Sheaf [741]. Sheaf-theoretic [741]. shear [1067, 2153]. sheets [1104]. shell [2040, 1464, 1695, 1925, 1103]. Sherrington [1671]. shift [99, 1851, 1533]. shift-invariant [1533]. shifted [1731]. Shiota [1441]. Shock [54, 1449, 1282, 1516]. Short [649, 169, 465, 1551, 5, 263, 1547, 1040]. short-range [5, 263]. Shouldn’t [7, 1609]. sided [835, 2097, 1907, 1796]. sigma [1591, 333, 2163, 171, 1540, 2103, 1313]. sign [500]. signals [892]. signature [547]. Silver [1565]. similar [1595, 1453, 1546]. Similarity [1666, 1886, 623, 981, 1600, 234]. Simons [1952, 1214, 124, 770, 803, 117, 20, 925, 451, 450, 1889]. Simple [1303, 1005, 906, 1123, 1087, 602, 345, 537, 1497, 1920, 1965, 656, 1227, 450, 2018]. simplex [2109]. simplification [819]. Simplified [1170, 1169, 300]. Simplifying [1073]. simultaneous [1875, 720]. Sinai [1087]. sine [601, 2015, 970]. sine-Gordon [2015, 970]. Single [1232, 2172, 787, 884]. single-molecule [884]. single-particle [787]. Single-step [1232]. singlet [1306]. Singular [2157, 499, 937, 97, 539, 1376, 115, 195, 42, 86, 2009]. singularities [1799, 433, 1426, 155, 1782, 1548, 514, 1170, 1171, 2217, 785, 2106, 1834, 1169, 1116, 1168]. singularity [1753, 2139]. singularly [1600, 1058]. sinh [888]. sinh-Gordon [888]. sinusoidal [1568]. sites [453]. Sitter [417, 17, 1134, 379, 1385, 752, 235, 2005, 2154]. Sitter-like [235]. six [993, 1408, 813, 1699, 642]. six-dimensional [813, 642]. six-level [1699]. six-vertex [1408]. sixth [1886]. size [871]. skew [1938, 924, 1808]. skew-orthogonal [1938]. skin [1860]. Skinner [181]. Skyrme [943, 266, 988]. Skyrme-type [943]. Skyrmions [1042]. Slater [1372]. slow [702, 2069]. small [2159, 1292, 1880, 1285, 1066, 930, 1224, 2191]. Smilga [1528]. smoothing [1184]. Smoothness [411]. Smorodinsky [1274]. Sobolev [1452]. soft [1797, 1093]. Sojourn [554]. Soldered [417]. solid [1072]. Solitary [892, 1366, 683, 300, 1318, 1783]. solitary-wave [1318]. Soliton [908, 567, 46, 174, 854, 590, 1583, 1937, 2096, 988, 94, 788, 1026, 517, 984, 1223, 2168]. Solitonlike [89]. solitons [2040, 1563, 1548, 1427, 1718, 1801, 1790, 1855, 812, 1135]. Solution [1466, 1465, 1930, 1531, 1393, 1860, 780, 178, 201, 343, 70, 20, 773, 581, 1715, 1583, 975, 914, 80, 71, 1910, 235, 898, 234, 996, 220, 1411, 1128, 1012, 2062, 489, 1547, 1905, 711, 2050, 1980]. Solutions [1585, 1373, 1541, 374, 275, 458, 351, 1134, 2165, 1203, 567, 641, 1241, 2067, 1868, 888, 1778, 2008, 1579, 1837, 1214, 1939, 1650, 289, 242, 590, 794, 705, 2002, 2070, 1771, 1077, 2088, 1014, 2030, 1747, 807, 1055, 1741, 342, 645, 1343, 764, 1440, 730, 1880, 479, 157, 124, 2010, 522, 1071, 1997, 1100, 568, 120, 1222, 2129, 1539, 549, 1113, 1356, 755, 974, 1517, 785, 1637, 818, 1545, 1042, 127, 1043, 1348, 626, 1653, 1758, 788, 1551, 499, 1026, 1124, 1116, 1045, 83, 1135, 1353, 2085, 2171, 919, 1649, 480, 58, 1872, 1295, 39, 1807, 1967, 690, 1467, 129, 2069, 2179, 774]. solutions [177, 735, 1605, 1318, 899, 1882, 2200, 1988, 1783, 462, 1899, 75, 1724, 1246, 1727, 102, 1520]. solvability [1816, 555, 921, 1001]. Solvable [529, 2014, 1610, 1643, 221, 1568, 189, 776, 519, 1943, 402, 1468, 1496, 926, 1823, 765, 1272, 1437, 1271, 2072]. solver [1446]. Solving [135, 1887]. Some [208, 166, 49, 2067, 1792, 1077, 1253, 703, 1404, 2132, 294, 1826, 2192, 483, 1239, 908, 2089, 1804, 1134, 1675, 1858, 1817, 1501, 2135, 1923, 1741, 156, 1562, 1637, 1920, 95, 1296, 1107, 185, 1864, 298, 1494, 1472]. Sommerfeld [2096, 1937, 690]. Source [1323, 841, 1860, 1113, 2125, 1483, 1940, 1966]. sources [2073, 379, 581, 1933, 2157, 1891, 1601, 1810]. Space [578, 622, 1896, 100, 2039, 530, 316, 169, 797, 49, 280, 913, 1078, 130, 1110, 1328, 893, 1102, 1499, 1934, 773, 581, 515, 1133, 2176, 1197, 635, 1625, 1284, 547, 27, 1357, 784, 2141, 270, 2070, 927, 1077, 34, 2001, 2038, 1652, 1236, 2049, 975, 2157, 333, 876, 932, 1354, 1888, 1153, 1216, 555, 1958, 896, 1099, 111, 672, 1035, 668, 488, 2217, 504, 2190, 2211, 1445, 1005, 1345, 616, 1925, 1834, 13, 499, 2167, 744, 250, 1168, 271, 1985, 376, 165, 1698, 2103, 176, 482, 629, 747, 977, 2069, 966, 766, 947, 473]. space-fractional [34]. space-time [2217]. spacelike [2183]. spaces [2155, 1006, 1732, 1853, 1913, 2083, 180, 441, 1390, 803, 258, 1847, 1086, 1625, 131, 608, 623, 23, 1550, 362, 2166, 1485, 199, 200, 394, 603, 758, 1636, 1700, 1228, 1565, 406, 592, 1892, 440, 1107, 1293, 2077]. spacetime [1466, 1426, 2010, 2041, 1510, 2108, 804]. spacetimes [1918, 384, 1917, 712, 811, 2156]. spanning [1676]. Spatial [1102, 1244, 1492, 1847, 1915, 1954, 584]. Special [1670, 1048, 1608, 2014, 1305, 1610, 1643, 1496, 139, 1036, 277, 1570, 1783]. species [247]. specific [1356]. Specker [1158, 360, 165]. Specker-type [165]. spectators [1848]. Spectra [1307, 1090, 2197, 197, 1704, 1577, 1443, 1457, 1340, 1921, 870]. Spectral [174, 1400, 1371, 635, 2120, 917, 1793, 846, 1083, 1963, 774, 268, 1360, 1809, 1615, 2195, 1784, 201, 662, 779, 706, 794, 283, 1791, 1613, 556, 143, 68, 2061, 1384, 1851, 390, 1067, 1657, 463, 1915, 894]. spectrum [1820, 197, 1908, 210, 658, 448, 1662, 364, 69, 1529, 13, 500, 2011, 140]. speed [1325, 61, 1410]. sphere [180, 441, 1191, 1357, 1009, 1506, 713, 1557, 1072, 778, 1436, 440, 664, 606]. spheres [397, 1474, 478, 1669, 1, 211, 403, 2208, 1291, 1403]. Spherical [2010, 153, 589, 1667, 1867, 1972, 583, 1168, 1743]. spherically [1203, 127, 1353]. spheroidal [987, 887]. spheroids [678]. spike [1042]. spike-vortex [1042]. spiked [1960]. Spin [1854, 1733, 1991, 1615, 2195, 1674, 1244, 1828, 126, 1085, 343, 22, 1973, 1675, 631, 1671, 1596, 1897, 328, 1180, 1362, 1881, 2027, 1283, 1377, 1992, 2071, 1544, 205, 301, 1022, 1363, 1685, 576, 592, 800, 67, 140, 1230, 817, 564, 1267, 1128, 1682, 792, 151, 557]. spin- [22, 576, 67]. spin-boson [1733, 1615, 2195]. spin-cluster [817]. spin-orbit [1180, 1362, 800]. spin-spin [1244]. spin-statistics [328]. spinning [1509, 750]. Spinor [913, 2001, 2156, 130, 617, 1257, 1146, 1649, 1301]. spinorial [591, 913]. spinors [805]. spins [1189, 105]. splittings [326]. Spontaneous [971, 407, 227, 544]. spots [2093]. springs [818]. Square [75, 2030, 253, 891, 1097, 2147]. Square-integrable [75]. square-potential [2030]. square-well [2147]. Squared [1777]. squeezed [1588, 1503, 2140]. Squire [253]. Stability [1314, 160, 1718, 317, 908, 1137, 972, 280, 2138, 1453, 2166, 1092, 1435, 1072, 359, 788, 1483, 1940, 2110, 1965, 2062, 721, 2050]. stabilized [1827]. stabilizer [636, 1992]. stable [468, 114]. Stäckel [552, 200, 603]. staggered [2169]. Standard [498, 613, 874, 1250, 821, 1152, 432, 1075, 1859, 134]. star [1030, 1097]. Stargenfunctions [270]. Stark [962]. stars [2031, 967, 580, 809, 715]. state [673, 798, 938, 1189, 861, 1284, 1367, 1613, 2037, 1055, 916, 2027, 845, 1611, 2074, 2146, 1435, 826, 1879, 674, 900, 727, 2204, 1590, 882, 361, 1298, 1306, 1230, 1082, 375, 366, 1128, 536, 1882, 2200, 457]. states [1588, 1503, 208, 722, 1375, 1333, 278, 1244, 895, 980, 2140, 990, 264, 1416, 636, 997, 3, 1330, 442, 460, 1695, 1233, 1334, 193, 1917, 700, 1374, 285, 1415, 1456, 1180, 941, 1849, 956, 1, 211, 1310, 470, 1283, 1377, 847, 1229, 1644, 756, 443, 306, 449, 512, 408, 304, 269, 2005, 1883, 365, 1920, 2128, 1332, 2024, 1379, 1533, 1795, 367, 215, 147, 2130, 113, 45, 2136, 1237, 2208, 114, 870, 112, 446, 1569, 1039, 849, 1341, 447, 1916, 2097, 1907, 1796, 828, 1274, 1774, 1832, 1951, 1996]. static [49, 954, 808, 1203]. Stationary [260, 1343, 1215, 1325, 2159, 2050]. Statistical [722, 2014, 1305, 1610, 1643, 1211, 842, 627, 1604, 1669, 223, 528, 2194, 1890, 1803, 1670]. statistics [87, 328, 1721, 1925, 781, 560, 88, 429, 979, 2032]. Steady [756, 1882, 2200, 1055]. steady-state [1055]. stellar [1965, 1315]. step [705, 767, 1232]. steplike [2084]. Stewartson [495, 1323, 1637]. Stieltjes [925, 31]. stiff [1203]. Stirling [336]. Stochastic [1273, 1089, 155, 927, 1989, 696, 1623, 1619, 303, 1684, 732, 1367, 588, 975, 1876, 1215, 2143, 1955, 2153, 1985, 2052, 185, 307, 1205]. Stokes [1771, 1558, 1755, 1758, 1060, 1052, 1049, 1652, 1055, 1050, 2049, 1558, 1051, 195, 1653, 1053, 1353, 757, 1988, 2180]. Stora [1132, 1555]. Stratified [1071, 840]. Stratonovich [1479, 823]. strength [76]. stress [138]. Strict [161, 952]. String [734, 1924, 229, 1275, 132, 626, 1239]. strings [575, 226, 1152]. Strong [861, 168, 540, 342, 1065, 154, 563, 1874, 1803, 1249]. Strong-coupling [861]. Strong-electric-field [168]. Strongly [1714, 965, 240, 682, 1535, 1425]. Structure [1085, 783, 1942, 169, 1416, 913, 1364, 1827, 942, 1614, 2037, 251, 991, 1029, 1138, 418, 876, 826, 2055, 199, 394, 2139, 1958, 358, 2019, 405, 378, 1557, 2204, 1066, 2076, 1074, 1689, 1095, 1647, 1075, 1369, 146, 265, 1864, 822, 1523]. Structures [1305, 1032, 2155, 1675, 130, 1604, 425, 1747, 1003, 2112, 382, 2209, 1998, 1514, 522, 1900, 727, 1064, 1587, 1877, 1962, 761, 2179, 1482, 1924, 1670, 1811].

Studies [1062, 2064]. Study [42, 85, 699, 1964, 2079, 2160, 1069, 140, 2077, 6, 104]. studying [446]. Sturm [1943, 2129]. sub [1391]. sub-Riemannian [1391]. Subadditivity [1150]. subalgebra [1900, 1720]. subalgebras [103, 1789]. subcritical [1677]. subdiffusive [2051]. subductions [1261]. subgrid [1067]. subgroup [513]. subgroups [596]. subharmonic [1650]. sublattices [897, 1252]. sublinear [2186]. submanifolds [948]. Subnormalized [997]. subquantum [222]. Subrepresentation [1521]. subspace [1423]. subspaces [1337, 1368, 989]. subtrees [1673]. sufficient [23]. Sum [1820, 1372, 1073, 1718]. summability [240, 682]. summable [1447]. Summation [240, 982, 298]. sums [1794, 982, 1926, 620, 334, 954]. Super [909, 1624, 1733, 803, 856, 979, 241, 404]. super-integrable [241]. super-radiance [1733]. superalgebra [1358, 1532, 1217, 1038]. superalgebras [602, 1524, 191, 2071, 2013, 429, 920]. supercharge [543]. superconducting [351, 352]. Superconductivity [222, 356, 167, 359, 1975, 348]. superconductors [355, 285]. Superconformal [575, 983, 1573, 122, 391, 501]. Supercoset [1035]. supercritical [1106]. superdense [967]. Superdiffusivity [1634]. superfield [273]. supergravity [1566, 1428, 1029, 642]. supergroup [1364]. supergroups [1528]. superheating [203, 302]. Superintegrability [1594, 1693, 2016, 239, 2086]. Superintegrable [600, 1651, 1329, 585, 1080, 199, 200, 394, 555, 603, 758, 921, 1264, 1405, 881, 2144, 1001, 762, 1118, 1351, 792]. superlinear [2111]. supermagnetic [2121]. supermanifolds [1752, 1359, 909]. superoscillatory [9]. superplane [723]. Superselection [315, 2034, 74]. superspace [475]. superspaces [1251, 2215, 1750, 153]. supersphere [98]. superstable [2160, 1803]. superstring [151]. Supersymmetric [1375, 1165, 321, 1440, 922, 70, 1968, 1708, 1500, 1693, 2016, 806, 474, 873, 2105, 457]. supersymmetrical [1868]. Supersymmetry [2144, 1418, 1151, 1375, 831, 1422, 543]. Supertetrahedra [2013]. supertrace [1405]. superversions [1528]. supplementary [405]. Support [912, 1279]. supported [1376, 1893]. supporting [1673]. supremum [2007, 1806]. Surface [1432, 352, 1136, 1722, 784, 2141, 2098, 809, 2069]. Surfaces [1266, 47, 1786, 1011, 771, 423, 1554, 893, 1102, 1499, 1847, 283, 1591, 1893, 1280, 15, 333, 668, 488, 1046, 1445, 1389, 434, 1434, 961, 889, 742, 966]. survey [1686]. survey-propagation [1686]. Susceptibility [1677]. SUSY [131]. Sutherland [502, 392, 294, 480]. switched [1616]. switched-on [1616]. Symbol [1247, 1255]. symbolic [1318]. symbols [1258, 156, 994, 1521, 1217]. Symmetric [163, 1552, 1947, 1203, 1349, 1328, 1133, 1448, 253, 1142, 2146, 2010, 701, 306, 1179, 917, 449, 127, 592, 367, 1766, 1353, 2101, 2103, 312, 63, 1267, 924, 1808, 4, 934, 1007, 221, 2147]. symmetric-positive [4, 934]. Symmetries [637, 775, 96, 953, 1929, 565, 172, 275, 541, 207, 2165, 299, 464, 60, 819, 1579, 942, 1566, 1428, 729, 1716, 1204, 1248, 1712, 875, 1222, 1969, 403, 1497, 1005, 655, 562, 1983, 2044, 724, 497, 293, 1666, 1886, 970, 128, 58, 129, 1798, 1903]. Symmetry [697, 929, 1897, 1637, 2171, 1916, 575, 971, 1522, 570, 59, 326, 2158, 1392, 314, 526, 1003, 1491, 926, 495, 2063, 1600, 1873, 1582, 1412, 1363, 2004, 227, 544, 11, 165, 1128, 1480, 471]. Symplectic [1586, 1406, 1942, 1361, 1173, 526, 1248, 1583, 1491, 1138, 181, 2101, 26]. system [1612, 877, 1015, 1136, 410, 1952, 2008, 314, 588, 1440, 1968, 300, 1292, 1092, 1611, 1635, 1215, 80, 2003, 1225, 1558, 1755, 1185, 1787, 1984, 232, 1637, 818, 1072, 284, 520, 960, 1626, 1389, 1502, 1348, 626, 390, 1590, 434, 534, 1211, 1666, 480, 1378, 754, 4, 934, 828, 2105, 735, 150, 1889, 1899]. Systematic [954, 516]. Systems [2014, 1329, 1610, 1643, 673, 862, 1620, 236, 1588, 1503, 2100, 1391, 2031, 327, 786, 1513, 1828, 1784, 1017, 1760, 521, 352, 681, 1137, 866, 552, 1537, 464, 819, 660, 533, 1473, 1317, 797, 582, 1468, 600, 239, 1018, 537, 1651, 1645, 957, 1089, 1839, 1638, 1163, 1165, 851, 708, 590, 585, 1080, 594, 608, 1912, 237, 270, 652, 10, 1490, 1177, 1240, 1923, 1195, 1173, 948, 1248, 682, 926, 1894, 1987, 830, 1138, 2112, 2209, 157, 1275, 2027, 1514, 1992, 1156, 448, 1154, 199, 200, 394, 555, 603, 758, 921, 1264, 969]. systems [306, 1286, 596, 917, 1694, 1906, 1363, 2045, 2205, 158, 1064, 2035, 1545, 576, 1042, 2194, 1500, 1478, 1405, 243, 52, 694, 881, 2144, 702, 95, 563, 1533, 952, 1994, 1560, 256, 1890, 1877, 1965, 182, 25, 1835, 67, 1803, 1230, 787, 2025, 480, 1278, 1001, 26, 817, 1699, 924, 1808, 762, 1118, 134, 753, 1962, 447, 73, 1707, 1083, 1140, 742, 853, 792, 1338, 1139, 2111, 43, 1239, 438]. syzygies [617, 1146]. Sz. [65].

’t [669]. tachyon [225]. tail [1339]. Tamm [54, 164]. tangent [1732]. Tate [400]. Taylor [1802]. technique [598]. techniques [1910, 910, 1929, 1360]. Teichmüller [100]. telegraph [299, 60, 1399, 1100]. teleparallel [2042, 21]. Teleportation [362]. temperature [472, 843, 1671, 1127, 1993, 1152]. Temperley [1363]. temporally [114]. Tensor [1866, 2158, 36, 575, 138, 533, 198, 1302, 617, 630, 551, 748, 615, 1582, 1146, 1220, 497, 715, 1872, 1549]. tensors [524, 695, 646, 206, 399]. term [1133, 1730, 1190, 1373, 1483, 1940, 1728, 1649]. terms [240, 1050, 826, 676, 2125]. ternary [1931, 1319, 795, 2198, 1525, 1830]. Test [525, 15]. testing [1377, 2063, 1447, 1795]. tests [360]. Tetrads [377, 2212]. tetrahedron [1448]. Teukolsky [968]. textures [1762]. th [1033]. their [54, 349, 1956, 895, 2181, 1760, 1917, 155, 1508, 897, 1252, 684, 839, 1099, 1646, 269, 858, 474, 1911, 2189, 920, 1857]. Theorem [637, 1799, 1978, 1584, 1143, 186, 207, 1060, 1822, 515, 131, 689, 860, 647, 1324, 1923, 2057, 1974, 2102, 823, 830, 820, 1561, 1697, 1683, 576, 2107, 2106, 1801, 824, 905, 2036, 1000, 360, 387, 1564, 1010, 1957, 281]. theorems [87, 1263, 918, 1739, 1346, 328, 1850, 1618, 1592, 322]. theoretic [386, 978, 741]. theoretical [527, 707, 1910, 1983, 2044, 579]. theories [2040, 1764, 2201, 219, 1432, 950, 803, 1822, 986, 28, 2080, 2203, 1487, 1488, 1489, 1490, 1240, 328, 1702, 614, 1463, 1115, 127, 1573, 319, 715, 1002, 806, 474, 98, 1872, 714, 949, 1924]. theory [19, 622, 1406, 1927, 2100, 472, 17, 1474, 575, 2031, 178, 137, 126, 373, 675, 972, 169, 412, 1464, 1241, 1792, 1078, 717, 2083, 20, 1214, 1603, 1448, 496, 1650, 586, 1668, 1788, 99, 1512, 1374, 1901, 1738, 643, 1488, 421, 92, 1850, 1268, 2047, 1712, 1387, 1631, 2175, 713, 1574, 470, 2104, 1888, 162, 2065, 199, 394, 603, 2041, 228, 1072, 2000, 2048, 1005, 406, 267, 1447, 910, 696, 609, 834, 511, 455, 451, 561, 1994, 1026, 545, 765, 1657, 1211, 1955, 2153, 292, 477, 953, 181, 305, 2101, 484, 1441, 370, 1581, 846, 1103, 1039, 995]. theory [2185, 65, 1369, 1744, 747, 536, 77, 1148, 1832, 1924, 1295]. theory-inspired [1924]. There [44, 7, 1609, 1008]. thermal [1935, 411]. Thermodynamic [2022, 1432, 1878, 2023]. thermodynamics [419, 1256]. theta [202]. thickness [699]. Thin [704, 937, 1161]. Third [1329, 1343, 881, 1693, 2016]. third-order [881]. Thirring [16, 1599, 463]. Thomas [1314]. Three [1302, 1932, 1868, 1873, 772, 172, 47, 862, 264, 1062, 959, 880, 1651, 1501, 1052, 1566, 1428, 1448, 1284, 2088, 1780, 1270, 1055, 445, 546, 1831, 160, 1190, 1354, 1888, 2003, 1838, 1225, 758, 1264, 2139, 1946, 1984, 167, 1587, 882, 1076, 1855, 1904, 1045, 11, 1723, 763, 1707, 1034, 966, 1899, 879, 394]. three-body [959, 1651, 1501, 1984, 1855, 1723, 763, 1034]. three-component [1899]. Three-dimensional [1302, 1932, 1868, 772, 880, 1052, 1448, 1284, 2088, 1780, 1055, 445, 546, 1354, 1888, 758, 1264, 1587, 1904, 1707, 966, 879, 394]. three-level [862]. three-local [1076]. three-point [1831]. three-wave [1225]. Threshold [332, 2114, 1738, 279, 413]. thresholds [1502]. Tikekar [967]. Time [829, 1331, 1333, 353, 1518, 631, 1421, 216, 1511, 34, 2034, 622, 2148, 1991, 578, 940, 186, 554, 1015, 2039, 351, 343, 303, 913, 1078, 852, 1814, 1110, 2196, 539, 1625, 547, 27, 683, 2030, 1335, 2001, 2038, 1268, 981, 975, 164, 435, 876, 448, 105, 1099, 668, 488, 2217, 158, 1546, 2125, 1005, 407, 1031, 69, 1531, 616, 1590, 1108, 1925, 1834, 1711, 422, 720, 1955, 1366, 57, 1278, 176, 1120, 368, 482, 629, 1962, 747, 828, 1274, 1013, 977, 1034, 774, 338, 473]. Time- [34]. time-covariant [448]. Time-dependent [353, 1421, 1991, 186, 1015, 351, 2196, 2030, 1335, 158, 2125, 1531, 1108, 422, 720, 1366, 1013]. time-domain [852, 1814]. time-energy [164]. time-extended [940]. time-inhomogeneous [338]. time-like [668, 488]. Time-of-arrival [829, 1331, 1333]. time-periodic [539, 774]. Time-reversal [2034]. time-varying [2148]. timelike [1445]. times [1331, 49, 130, 1102, 515, 1133, 1345, 250, 1168]. Toda [2182, 2084, 933, 1206, 418, 1601, 667, 102]. Toeplitz [1028, 706, 2087]. tomography [2037, 2204]. tones [2117]. tongues [805]. Tonomura [2150]. Tools [1821]. top [417, 17]. Topological [1786, 1387, 50, 1382, 20, 1731, 748, 1404, 696, 1168, 1095, 450, 85]. Topologically [357]. topology [1896, 2124, 1253, 679, 945, 821, 118]. topos [1487, 1488, 1489, 1490]. tori [1474, 1041]. toric [151]. torsion [416, 810, 2151]. torus [697, 986, 1153]. Total [885, 2018]. toy [1447]. Trace [593, 1230, 799, 997, 728, 1642]. trace-nonincreasing [997]. trace-preserving [728]. traces [341]. Tracy [1441]. trains [320]. trajectories [631, 700, 869, 652, 504, 361]. trajectory [730]. transcendent [645, 2016]. transcendents [1343, 1753, 1602]. Transfer [1021, 842]. transform [854, 851, 2084, 200, 603, 204, 431, 601, 666, 31, 388, 640]. transformation [494, 170, 569, 202, 1479, 981, 2152, 827, 621, 277, 1523]. transformations [901, 737, 1265, 340, 1805, 1074, 1729, 1892, 446, 102]. Transforming [2154]. transforms [1552, 593, 1216, 1179, 1685, 665]. Transition [1009, 1991, 1506, 656, 1227]. Transitional [1879]. transitions [245, 1632, 1543, 1478, 187]. transitive [1345]. translation [1244]. translational [515]. transmission [840, 1596, 892, 89]. transmutation [490]. Transparent [1067]. Transport [935, 1656, 2025, 885, 232, 1892, 1581, 846]. transportation [307]. transposition [2128]. transverse [883]. trap [1455]. trapped [893, 1499]. traps [965, 1188, 1356, 1772]. Trautman [775]. traveling [2067, 1741, 1724]. treatment [2073, 936, 585, 409]. treatments [1174]. tree [1676, 1662, 882]. treelike [1688]. trees [1598, 1673, 399]. Tri [594]. Tri-Hamiltonian [594]. triangles [2117]. triangular [350, 1252]. triangularity [258]. triatomic [1443]. trigonometric [1334, 1456, 1179, 392, 294]. trilinear [1411]. triple [534]. Tsallis [458, 527, 1721]. Tube [1833]. tunable [221]. Tunneling [1235, 1331, 2030]. Turaev [1183]. turbulence [1068]. turbulent [688]. turning [1058]. turns [1863]. twist [380, 389, 2041]. Twisted [1403, 1572, 1424, 929, 194, 1936, 1316, 1176, 1145, 1414, 564, 667]. twisting [415]. twistorlike [2155]. twists [209]. Two [130, 1025, 648, 1070, 71, 1976, 1099, 11, 1243, 2097, 1907, 1796, 1318, 172, 1612, 611, 1375, 2017, 1192, 18, 1751, 264, 1063, 1371, 180, 239, 441, 1778, 1054, 547, 585, 1080, 878, 1419, 1816, 625, 1897, 686, 649, 1639, 1050, 2049, 1937, 2096, 300, 157, 418, 1900, 340, 756, 2129, 1841, 200, 330, 2095, 969, 1946, 1557, 1455, 576, 1042, 1601, 2033, 1345, 954, 322, 601, 867, 1775, 813, 440, 85, 664, 73, 873, 413, 2069, 1779, 177, 735, 150, 1889, 2115, 1139, 6, 104, 102, 125, 1284, 199, 758]. Two- [1284, 758]. two-body [878]. two-cluster [413]. two-column [18]. Two-component [648, 649, 1455, 1042, 2033, 735]. two-coupled [1778]. Two-dimensional [1099, 1375, 2017, 1063, 180, 239, 441, 1054, 547, 585, 1897, 1050, 2049, 1937, 2096, 418, 2129, 200, 969, 1601, 954, 1775, 440, 85, 664, 1779, 1139, 199]. two-electron [6, 104]. two-fermion [73]. two-fluid [322]. Two-layer [1070]. two-level [1612, 125]. two-loop [611]. two-matrix [1751]. Two-parameter [1976, 1243]. two-particle [1889]. two-photon [1371, 1816]. two-scale [1639]. Two-sided [2097, 1907, 1796]. Two-spinor [130]. Type [773, 807, 1075, 943, 1956, 1386, 968, 151, 660, 533, 1492, 602, 796, 135, 2170, 1060, 1261, 581, 1878, 28, 285, 1867, 691, 1206, 1629, 1275, 845, 1164, 614, 2139, 815, 1711, 1983, 2044, 2006, 1021, 165, 1495, 1808, 762, 1118, 65, 714, 2009, 1101, 566]. type- [968]. type-I [1386]. types [1025, 439, 1900, 1318]. Tyutin [1132, 1555]. Tzitzeica [1855, 621].

Uhlenbeck [687, 387]. Ulam [2138, 1014]. ultra [2131]. ultradiscrete [793]. ultrametric [246]. ultrarelativistic [1314]. ultraviolet [373, 1744]. umbral [1481]. Unbiased [1699, 44, 993, 1794]. unbounded [2122, 1630, 1727]. Uncertainty [1311, 851, 1081, 787, 1121, 903, 164, 113, 1039, 1567, 1504]. unconditional [1661]. unconstrained [1372, 370]. underlying [1542]. undistillable [1416]. unification [36, 1110, 452, 1618]. Unified [1513, 585, 765, 537, 92, 474, 35]. Uniform [84, 618, 71, 1728, 883, 1630, 947]. uniformization [2082, 713]. uniformly [841, 1563, 947]. unifying [1438]. unilateral [422, 720, 1433]. Uniqueness [20, 438, 690, 1584, 840, 289, 1850]. unistochastic [732, 2187]. Unit [1325, 1809, 1159, 160, 408, 664]. unitaries [864, 991]. Unitary [1636, 1218, 1032, 1402, 708, 732, 1562, 991, 973, 1105, 395, 492, 101, 1022, 405, 192, 365, 1617, 231, 2192, 277]. unity [430, 736, 258]. Universal [855, 1675, 1121, 1049, 454, 40, 1902, 1799, 1978, 209, 1308, 1125, 1981]. Universality [120, 467, 64, 1679, 1938]. universe [918, 1385, 2005, 1766, 548]. Unknown [3, 1330]. unmixed [955]. unoriented [77]. unrestricted [1739]. unstable [1333, 702]. unusual [595]. uplifting [557]. Upper [76, 510, 2114, 916]. use [2113, 2065, 727, 1910, 6, 104]. Useful [1332, 166]. using [219, 1077, 2001, 1073, 701, 1449, 215, 85].

V [581, 1275, 1950, 758]. vacuum [613, 351, 1918, 773, 190, 807, 1383, 1427, 1029, 1979, 821, 1119, 21]. Vaidya [232]. vakonomic [346, 950, 244]. Validity [95, 699]. value [886, 1206, 568, 2129, 1445, 83, 57, 2069, 2050]. valued [1028, 163, 1947, 634, 311, 2083, 1522, 803, 12, 1761, 306, 332, 1810, 894]. values [247, 1912, 700, 86, 1844]. Vanishing [1059, 139]. vapor [1632]. Variable [137, 590, 683, 1100, 1694, 1326, 2085, 383, 1745, 2126, 1744, 1905]. variable-coefficient [2126]. Variable-separation [137]. variables [1492, 865, 1112, 1422, 1044, 1365, 1354, 340, 921, 1750, 1242]. Variant [923]. variants [519]. variation [91]. Variational [183, 1919, 811, 789, 910, 73, 2092, 950, 412, 62, 1839, 719, 1493, 1072, 287, 420, 146, 1482]. variations [791, 1138]. varieties [605, 1264]. varying [2148]. Vector [536, 208, 1375, 2083, 1847, 1385, 546, 1180, 941, 1507, 2048, 94, 666, 25, 1107, 1007, 1123, 711, 1899, 1928]. vector-valued [2083]. vectors [2145, 515, 2158, 876, 2162]. velocity [1133, 84, 2180]. velocity-term [1133]. Vernon [1194]. version [895, 874, 1440, 1355, 669]. versus [950, 244, 1855, 1418]. vertex [1470, 624, 1451, 1731, 2104, 1408, 2012, 922]. vertical [1017]. very [540, 595]. VI [1225, 1753, 1602, 898]. via [798, 1970, 729, 590, 658, 23, 1234, 2104, 1888, 1600, 609, 834, 79, 114, 1278, 309, 2097, 1907, 1796]. vibration [1482]. vibrational [1443]. view [1569]. viewpoint [1742]. violating [468]. violation [1973, 1167]. Virasoro [1998, 734, 944, 1658, 495, 2169, 1484, 1316]. Viro [1183]. Virtual [1733]. virtue [1923]. viscoelastic [719, 1435, 2125, 1999]. viscosity [1771, 1059, 1988]. viscous [1057, 1878, 1766, 1649, 1259, 1807]. visibility [514]. Visualization [1501]. Visualizing [1192]. Vlasov [1292, 1092, 1880]. void [2217]. voids [1426, 1834]. Volkov [75]. Volterra [1473, 396, 960, 1666, 1779]. Volume [2187, 416, 1762, 1674, 963, 1996]. vortex [1061, 1141, 1062, 1349, 784, 2141, 106, 1042, 313]. Vortices [167, 1063, 117, 173, 477, 1889]. vorticity [2165, 615, 1070, 1059]. Vries [970, 961, 2159, 1970, 683, 845, 1559, 1536, 1886, 889, 894]. VSI [250].

Waals [1944]. walk [1879]. Walker [860, 1918, 280, 1102, 647, 811, 1277, 319, 548, 1538]. walks [2148, 1842, 1683, 1287, 1120, 1818, 1941]. wall [350, 755, 974, 1517, 1545, 96, 2012, 922]. walls [1029, 1937, 2096, 809]. Wannier [290]. Ward [1200, 1663]. Warped [1086, 880]. water [1632]. Wave [2196, 1221, 1108, 841, 1706, 506, 22, 2067, 1757, 9, 705, 998, 885, 1460, 1453, 987, 887, 941, 2102, 479, 80, 1354, 443, 2129, 1225, 577, 184, 320, 195, 639, 2005, 1975, 2152, 1483, 1940, 437, 220, 677, 930, 1224, 906, 1366, 833, 1004, 2213, 388, 73, 4, 690, 934, 411, 1547, 1318, 572, 1040, 75, 1724, 1144, 1520]. wave-short [1547, 1040]. wavefunction [1372]. waveguides [937, 816, 1161, 287]. wavelength [649]. wavelet [1109, 684, 839, 716]. wavelets [327, 1972]. Waves [731, 54, 1516, 1820, 683, 230, 1749, 300, 1577, 1582, 1885, 1282, 1705, 1874]. way [1535, 1425]. ways [1699]. Weak [21, 1140, 288, 920, 1821, 863, 2088, 30, 730, 344, 680, 739, 1999, 1772, 1249, 638]. Weak-Hamiltonian [1140]. weakened [1885]. weaker [739]. weakly [186, 1256, 1410, 1997, 2125, 489]. webs [1328]. wedge [48, 47, 295]. Wegner [1164]. Weierstrass [966]. weight [1659, 231]. Weighted [1082]. Weil [1794]. Weingarten [2135]. Weiss [245]. Weitzenböck [1964]. Weizsäcker [1314]. Well [1184, 1715, 1728, 2134, 1568, 1938, 253, 1292, 1190, 1900, 974, 1807, 2147]. Well-posedness [1728, 2134]. wells [148]. Werner [1269]. Wess [475, 1084]. Weyl [347, 864, 1530, 569, 198, 797, 2042, 1625, 615, 701, 994, 335, 14, 1622, 1815, 850]. Weyl-covariant [198]. Wheeler [22, 1919]. which [265]. white [560]. Whitham [734, 694, 1575, 1967]. Whittaker [2181]. Widom [1441]. Wiener [1661, 927, 2065]. Wigert [925]. Wigner [347, 1825, 1947, 1530, 569, 1572, 410, 1413, 261, 101, 1022, 818, 1545, 335, 14, 607, 383, 107]. Wilson [170, 262]. window [816]. window-coupled [816]. Winternitz [1274]. without [529, 860, 647, 2001, 981, 509, 1482]. Witten [275, 1084, 499, 2108, 806, 1148]. Wong [823]. worlds [1949]. worst [215]. Wouthuysen [2152]. Wyman [234].

X [201, 1400]. X-ray [201, 1400]. XXZ [2020, 2023].

Yang [1116, 370, 1406, 726, 803, 794, 371, 1850, 2003, 1852, 1557, 785, 72, 724, 873, 374, 1520]. Yangian [1740]. Yangians [929, 194, 1176]. Yau [1130, 1381, 1763]. Yetter [1183]. yield [636]. Young [141, 806]. Yukawa [1292].

Zabolotskaya [2067]. Zak [851]. Zakai [823]. Zakharov [1876, 1286]. Zamolodchikov [875, 140]. Zassenhaus [558]. Zeeman [1253]. Zehnder [850]. Zero [702, 1452, 1747, 456, 1447, 996, 650]. zero-entropy-density [456]. Zeta [2009, 738, 1247, 595, 1556]. zeta-function [738, 1247]. Zumino [1084, 475].

References

Hall:2002:CSS

[1] Brian C. Hall and Jeffrey J. Mitchell. Coherent states on spheres. Journal of Mathematical Physics, 43(3):1211–1236, March 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [211] and comment [2208].

Ruskai:2002:IQE

[2] Mary Beth Ruskai. Inequalities for quantum entropy: A review with conditions for equality. Journal of Mathematical Physics, 43(9):4358– 4375, September 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [37].

Caves:2002:UQS

[3] Carlton M. Caves, Christopher A. Fuchs, and Rüdiger Schack. Unknown quantum states: The quantum de Finetti representation. Journal of Mathematical Physics, 43(9):4537–4559, September 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1330].

Torre:2003:HRW

[4] C. G. Torre. The helically-reduced wave equation as a symmetric-positive system. Journal of Mathematical Physics, 44(12):6223–6232, December 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [934].

Saenz:2004:QMS

[5] A. W. Sáenz. Quantum-mechanical scattering in exterior domains with impenetrable periodic boundaries and short-range potentials. Journal of Mathematical Physics, 45(4):1426–1446, April 2004. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v45/i4/p1426_s1. See er-

ratum [263].

Zamastil:2004:UAE

[6] J. Zamastil, J. Cízek, M. Kalhous, L. Skála, and M. Simánek. The use of so(2, 1) algebra for the evaluation of atomic integrals: The study of two-electron atoms. Journal of Mathematical Physics, 45(7):2674– 2693, July 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v45/i7/p2674_s1. See erratum [104].

Galapon:2004:STA

[7] Eric A. Galapon. Shouldn’t there be an antithesis to quantization? Journal of Mathematical Physics, 45(8):3180–3215, August 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i8/p3180_

s1. See erratum [1609].

Tanimoto:2004:HAL

[8] Masayuki Tanimoto. Harmonic analysis of linear fields on the nilgeometric cosmological model. Journal of Mathematical Physics, 45(12):4896– 4919, December 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/

jmapaq/v45/i12/p4896_s1. See publisher’s note [38].

Calder:2005:ASW

[9] Matt S. Calder and Achim Kempf. Analysis of superoscillatory wave functions. Journal of Mathematical Physics, 46(1):012101, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/
p012101_s1.

Domenech:2005:CLQ

[10] Graciela Domenech and Hector Freytes. Contextual logic for quantum systems. Journal of Mathematical Physics, 46(1):012102, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/
p012102_s1.

Panahi:2005:TTD

[11] H. Panahi and M. A. Jafarizadeh. Two and three dimensional Hamiltonians with generalized and ordinary shape invariance symmetry. Journal of Mathematical Physics, 46(1):012103, January 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i1/p012103_s1.

Decker:2005:IGC

[12] Thomas Decker, Dominik Janzing, and Martin Rötteler. Implementation of group-covariant positive operator valued measures by orthogonal measurements. Journal of Mathematical Physics, 46(1):012104, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i1/p012104_s1.

Morame:2005:RSN

[13] Abderemane Morame and Françoise Truc. Remarks on the spectrum of the Neumann problem with magnetic field in the half-space. Journal of Mathematical Physics, 46(1):012105, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i1/p012105_s1.

Mukunda:2005:WWI

[14] N. Mukunda, G. Marmo, A. Zampini, S. Chaturvedi, and R. Simon. Wigner–Weyl isomorphism for quantum mechanics on Lie groups. Journal of Mathematical Physics, 46(1):012106, January 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i1/p012106_s1.

Gravesen:2005:SPS

[15] Jens Gravesen, Morten Willatzen, and L. C. Lew Yan Voon. Schrödinger problems for surfaces of revolution — the finite cylinder as a test example. Journal of Mathematical Physics, 46(1):012107, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/p012107_s1.

Bhattacharyya:2005:QIB

[16] Tanaya Bhattacharyya. Quantum integrability of bosonic massive Thirring model in continuum. Journal of Mathematical Physics, 46(1): 012301, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i1/p012301_s1.

Armenta:2005:SRT

[17] J. Armenta and J. A. Nieto. The de Sitter relativistic top theory. Journal of Mathematical Physics, 46(1):012302, January 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i1/p012302_s1.

Bekaert:2005:NSI

[18] Xavier Bekaert, Nicolas Boulanger, and Sandrine Cnockaert. No selfinteraction for two-column massless fields. Journal of Mathematical Physics, 46(1):012303, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i1/p012303_s1.

Abreu:2005:CBC

[19] L. M. Abreu, C. de Calan, A. P. C. Malbouisson, J. M. C. Malbouisson, and A. E. Santana. Critical behavior of the compactified λϕ^{4} theory. Journal of Mathematical Physics, 46(1):012304, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/p012304_s1.

Choe:2005:UTM

[20] Kwangseok Choe. Uniqueness of the topological multivortex solution in the self-dual Chern–Simons theory. Journal of Mathematical Physics, 46(1):012305, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i1/p012305_s1.

Itin:2005:WFR

[21] Yakov Itin. Weak field reduction in teleparallel coframe gravity: Vacuum case. Journal of Mathematical Physics, 46(1):012501, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i1/p012501_s1.

Blue:2005:WES

[22] P. Blue and A. Soffer. The wave equation on the Schwarzschild metric II. Local decay for the spin-2 Regge–Wheeler equation. Journal of Mathematical Physics, 46(1):012502, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i1/p012502_s1.

Edgar:2005:NSC

[23] S. Brian Edgar. Necessary and sufficient conditions for n-dimensional conformal Einstein spaces via dimensionally dependent identities. Journal of Mathematical Physics, 46(1):012503, January 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i1/p012503_s1.

Batic:2005:ECP

[24] Davide Batic, Harald Schmid, and Monika Winklmeier. On the eigenvalues of the Chandrasekhar–Page angular equation. Journal of Mathematical Physics, 46(1):012504, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i1/p012504_s1.

Pucacco:2005:IHS

[25] Giuseppe Pucacco and Kjell Rosquist. Integrable Hamiltonian systems with vector potentials. Journal of Mathematical Physics, 46(1):012701, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i1/p012701_s1.

Shirzad:2005:DSA

[26] A. Shirzad and M. Mojiri. The difficulty of symplectic analysis with second class systems. Journal of Mathematical Physics, 46(1):012702, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i1/p012702_s1.

DeCarli:2005:MHS

[27] E. De Carli and L. A. Ferreira. A model for Hopfions on the space– time S^{3} × R. Journal of Mathematical Physics, 46(1):012703, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i1/p012703_s1.

DiBartolo:2005:DTA

[28] Cayetano Di Bartolo, Rodolfo Gambini, Rafael A. Porto, and Jorge Pullin. Dirac-type approach for consistent discretizations of classical constrained theories. Journal of Mathematical Physics, 46(1):012901, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i1/p012901_s1.

Ivanova:2005:BAC

[29] A. N. Ivanova, S. I. Kuchanov, and L. I. Manevitch. Bifurcation analysis for the construction of a phase diagram of heteropolymer liquids. Journal of Mathematical Physics, 46(1):013301, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/p013301_s1.

Field:2005:DMW

[30] Timothy R. Field and Robert J. A. Tough. Dynamical models of weak scattering. Journal of Mathematical Physics, 46(1):013302, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i1/p013302_s1.

Schwarz:2005:GST

[31] John H. Schwarz. The generalized Stieltjes transform and its inverse. Journal of Mathematical Physics, 46(1):013501, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/p013501_s1.

Papanicolaou:2005:NDE

[32] Vassilis G. Papanicolaou and Paul K. Newton. Nonlinear dissipative eigenvalue problems with large initial conditions. Journal of Mathematical Physics, 46(1):013502, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i1/p013502_s1.

Li:2005:QBC

[33] Y. Charles Li. On quasiperiodic boundary condition problem. Journal of Mathematical Physics, 46(1):013503, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i1/p013503_s1.

Duan:2005:TSF

[34] Jun-Sheng Duan. Time- and space-fractional partial differential equations. Journal of Mathematical Physics, 46(1):013504, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/
p013504_s1.

Zhang:2005:UCC

[35] Yao-Zhong Zhang and Mark D. Gould. A unified and complete construction of all finite dimensional irreducible representations of gl(2|2). Journal of Mathematical Physics, 46(1):013505, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/p013505_s1.

Anderson:2005:UMB

[36] Gregory W. Anderson and Tomás Blazek. E_{6} unification model building. III. Clebsch–Gordan coefficients in E_{6} tensor products of the 27 with higher dimensional representations. Journal of Mathematical Physics, 46(1):013506, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i1/p013506_s1.

Ruskai:2005:EIQ

[37] Mary Beth Ruskai. Erratum: “Inequalities for quantum entropy: A review with conditions for equality”, [J. Math. Phys. 43, 4358 (2002)]. Journal of Mathematical Physics, 46(1):019901, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/
p019901_s1. See [2].

Tanimoto:2005:PNH

[38] Masayuki Tanimoto. Publisher’s note: “Harmonic analysis of linear fields on the nilgeometric cosmological model”, [J. Math. Phys. 45, 4896 (2004)]. Journal of Mathematical Physics, 46(1):019902, January 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i1/
p019902_s1. See [8].

Tanimura:2005:ESI

[39] Shogo Tanimura, Mikio Nakahara, and Daisuke Hayashi. Exact solutions of the isoholonomic problem and the optimal control problem in holonomic quantum computation. Journal of Mathematical Physics, 46(2):022101, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i2/p022101_s1.

Junge:2005:UCR

[40] Marius Junge, Peter T. Kim, and David W. Kribs. Universal collective rotation channels and quantum error correction. Journal of Mathematical Physics, 46(2):022102, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i2/p022102_s1.

Ichinose:2005:PIR

[41] Takashi Ichinose. Path integral for the radial Dirac equation. Journal of Mathematical Physics, 46(2):022103, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i2/p022103_s1.

Saad:2005:SAS

[42] Nasser Saad, Richard L. Hall, and Qutaibeh D. Katatbeh. Study of anharmonic singular potentials. Journal of Mathematical Physics, 46(2):022104, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i2/p022104_s1.

Zhu:2005:PNM

[43] Wusheng Zhu and Herschel Rabitz. Perturbative and nonperturbative master equations for open quantum systems. Journal of Mathematical Physics, 46(2):022105, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i2/p022105_s1.

Archer:2005:TNG

[44] Claude Archer. There is no generalization of known formulas for mutually unbiased bases. Journal of Mathematical Physics, 46(2):022106, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i2/p022106_s1.

Peletminskii:2005:SQM

[45] Sergey V. Peletminskii and Yuriy V. Slyusarenko. Second quantization method in the presence of bound states of particles. Journal of Mathematical Physics, 46(2):022301, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i2/p022301_s1.

Abounasr:2005:AGR

[46] R. Abounasr, M. Ait Ben Haddou, A. El Rhalami, and E. H. Saidi. Algebraic geometry realization of quantum Hall soliton. Journal of Mathematical Physics, 46(2):022302, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i2/p022302_s1.

Ahmedov:2005:CEW

[47] H. Ahmedov and I. H. Duru. Casimir energy for a wedge with three surfaces and for a pyramidal cavity. Journal of Mathematical Physics, 46(2):022303, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i2/p022303_s1.

Ahmedov:2005:CEC

[48] H. Ahmedov and I. H. Duru. Casimir energy in a conical wedge and a conical cavity. Journal of Mathematical Physics, 46(2):022304, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i2/p022304_s1.

Brozos-Vazquez:2005:SRL

[49] Miguel Brozos-Vázquez, Eduardo García-Río, and Ramón Vázquez-Lorenzo. Some remarks on locally conformally flat static space–times. Journal of Mathematical Physics, 46(2):022501, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i2/p022501_s1.

Patino:2005:TQG

[50] Leonardo Patiño and Hernando Quevedo. Topological quantization of gravitational fields. Journal of Mathematical Physics, 46(2):022502, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i2/p022502_s1.

Hollands:2005:AFB

[51] Stefan Hollands and Akihiro Ishibashi. Asymptotic flatness and Bondi energy in higher dimensional gravity. Journal of Mathematical Physics, 46(2):022503, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i2/p022503_s1.

Mada:2005:PDC

[52] Jun Mada, Makoto Idzumi, and Tetsuji Tokihiro. Path description of conserved quantities of generalized periodic box-ball systems. Journal of Mathematical Physics, 46(2):022701, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i2/p022701_s1.

Blesgen:2005:RMD

[53] T. Blesgen. A revised model for diffusion induced segregation processes. Journal of Mathematical Physics, 46(2):022702, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i2/p022702_s1.

Afanasiev:2005:SWT

[54] G. N. Afanasiev, V. M. Shilov, and Yu. P. Stepanovsky. Shock waves in the Tamm problem and the possibility of their experimental observation. Journal of Mathematical Physics, 46(2):022901, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i2/p022901_s1.

Cammarota:2005:IGM

[55] Camillo Cammarota. An inequality for the Gibbs mean number of clusters. Journal of Mathematical Physics, 46(2):023301, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i2/
p023301_s1.

Gibilisco:2005:MSC

[56] Paolo Gibilisco and Tommaso Isola. On the monotonicity of scalar curvature in classical and quantum information geometry. Journal of Mathematical Physics, 46(2):023501, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i2/p023501_s1.

Salvarani:2005:LTA

[57] F. Salvarani and G. Toscani. Large-time asymptotics for nonlinear diffusions: the initial-boundary value problem. Journal of Mathematical Physics, 46(2):023502, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i2/p023502_s1.

Shih:2005:ACS

[58] M. Shih, E. Momoniat, and F. M. Mahomed. Approximate conditional symmetries and approximate solutions of the perturbed Fitzhugh– Nagumo equation. Journal of Mathematical Physics, 46(2):023503, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i2/p023503_s1.

Cheh:2005:MCE

[59] Jeongoo Cheh, Peter J. Olver, and Juha Pohjanpelto. Maurer–Cartan equations for Lie symmetry pseudogroups of differential equations. Journal of Mathematical Physics, 46(2):023504, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i2/p023504_s1.

Bluman:2005:LNS

[60] G. W. Bluman, Temuerchaolu, and R. Sahadevan. Local and nonlocal symmetries for nonlinear telegraph equation. Journal of Mathematical Physics, 46(2):023505, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i2/p023505_s1.

Constantin:2005:FPS

[61] Adrian Constantin. Finite propagation speed for the Camassa–Holm equation. Journal of Mathematical Physics, 46(2):023506, February 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i2/
p023506_s1.

Brandt:2005:RCE

[62] Sebastian F. Brandt, Hagen Kleinert, and Axel Pelster. Recursive calculation of effective potential and variational resummation. Journal of Mathematical Physics, 46(3):032101, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p032101_s1.

Sinha:2005:SMN

[63] A. Sinha and P. Roy. P T symmetric models with nonlinear pseudosupersymmetry. Journal of Mathematical Physics, 46(3):032102, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i3/p032102_s1.

Khuri:2005:ULE

[64] N. N. Khuri, André Martin, Pierre C. Sabatier, and Tai Tsun Wu. Universality of low-energy scattering in 2 + 1 dimensions: The nonsymmetric case. Journal of Mathematical Physics, 46(3):032103, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/

p032103_s1. See addendum [467].

Strauss:2005:SNF

[65] Y. Strauss. Sz.–Nagy–Foias theory and Lax–Phillips type semigroups in the description of quantum mechanical resonances. Journal of Mathematical Physics, 46(3):032104, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i3/p032104_s1.

Privault:2005:EQM

[66] Nicolas Privault and Jean-Claude Zambrini. Euclidean quantum mechanics in the momentum representation. Journal of Mathematical Physics, 46(3):032105, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i3/p032105_s1.

Rangan:2005:CFD

[67] C. Rangan and A. M. Bloch. Control of finite-dimensional quantum systems: Application to a spin-1/2 particle coupled with a finite quantum harmonic oscillator. Journal of Mathematical Physics, 46(3): 032106, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i3/p032106_s1.

Ghanmi:2005:ACH

[68] Allal Ghanmi and Ahmed Intissar. Asymptotic of complex hyperbolic geometry and L^{2}-spectral analysis of Landau-like Hamiltonians. Journal of Mathematical Physics, 46(3):032107, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p032107_s1.

McCaw:2005:PPS

[69] James McCaw and B. H. J. McKellar. Pure point spectrum for the time evolution of a periodically rank-N kicked Hamiltonian. Journal of Mathematical Physics, 46(3):032108, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p032108_s1.

Breiderhoff:2005:ESS

[70] Julia Breiderhoff, Fabio Musso, and Orlando Ragnisco. Exact solution to a supersymmetric Gaudin model. Journal of Mathematical Physics, 46(3):032109, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i3/p032109_s1.

Kandemir:2005:TIE

[71] B. S. Kandemir. Two interacting electrons in a uniform magnetic field and a parabolic potential: The general closed-form solution. Journal of Mathematical Physics, 46(3):032110, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p032110_s1.

Loikkanen:2005:YMA

[72] Juha Loikkanen and Cornelius Paufler. Yang–Mills action from minimally coupled bosons on R^{4} and on the four-dimensional Moyal plane. Journal of Mathematical Physics, 46(3):032301, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p032301_s1.

Terekidi:2005:VTF

[73] Andrei G. Terekidi and Jurij W. Darewych. Variational two-fermion wave equations in quantum electrodynamics: Muoniumlike systems. Journal of Mathematical Physics, 46(3):032302, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p032302_s1.

Kijowski:2005:CSS

[74] J. Kijowski and G. Rudolph. Charge superselection sectors for QCD on the lattice. Journal of Mathematical Physics, 46(3):032303, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i3/p032303_s1.

Zakowicz:2005:SIW

[75] Stephan Zakowicz. Square-integrable wave packets from the Volkov solutions. Journal of Mathematical Physics, 46(3):032304, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/
p032304_s1.

Brau:2005:ULC

[76] Fabian Brau. Upper limit on the critical strength of central potentials in relativistic quantum mechanics. Journal of Mathematical Physics, 46(3): 032305, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i3/p032305_s1.

Varga:2005:MTU

[77] Peter Varga. Matrix theory of unoriented membranes and Jordan algebras. Journal of Mathematical Physics, 46(3):032306, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/
p032306_s1.

DiBartolo:2005:CMD

[78] Cayetano Di Bartolo, Rodolfo Gambini, and Jorge Pullin. Consistent and mimetic discretizations in General Relativity. Journal of Mathematical Physics, 46(3):032501, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i3/p032501_s1.

Montiel-Pina:2005:GRC

[79] Enrique Montiel-Piña, Ezra Ted Newman, and Gilberto Silva-Ortigoza. General relativity via complete integrals of the Hamilton–Jacobi equation. Journal of Mathematical Physics, 46(3):032502, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/
p032502_s1.

Hon:2005:AGS

[80] Y. C. Hon and E. G. Fan. An algebro-geometric solution for a Hamiltonian system with application to dispersive long wave equation. Journal of Mathematical Physics, 46(3):032701, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p032701_s1.

Etayo:2005:LFN

[81] Fernando Etayo, Rafael Santamaría, and Sergiu I. Vacaru. Lagrange– Fedosov nonholonomic manifolds. Journal of Mathematical Physics, 46(3):032901, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i3/p032901_s1.

Pavon:2005:HJE

[82] Michele Pavon. Hamilton–Jacobi equations for nonholonomic dynamics. Journal of Mathematical Physics, 46(3):032902, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/p032902_s1.

Panayotounakos:2005:EAS

[83] D. E. Panayotounakos, N. B. Sotiropoulos, A. B. Sotiropoulou, and N. D. Panayotounakou. Exact analytic solutions of nonlinear boundary value problems in fluid mechanics (Blasius equations). Journal of Mathematical Physics, 46(3):033101, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i3/p033101_s1. See comment [568].

Nunez:2005:UEV

[84] Manuel Núñez. Uniform estimates on the velocity in Rayleigh–Bénard convection. Journal of Mathematical Physics, 46(3):033102, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/
p033102_s1.

Suzuki:2005:SBB

[85] Masuo Suzuki, Hidenori Suzuki, and Shu-Chiuan Chang. Study on boundary-boundary correlation functions of the two-dimensional Ising model using topological interaction method. Journal of Mathematical Physics, 46(3):033301, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i3/p033301_s1.

Soshnikov:2005:LSV

[86] Alexander Soshnikov and Yan V. Fyodorov. On the largest singular values of random matrices with independent Cauchy entries. Journal of Mathematical Physics, 46(3):033302, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p033302_s1.

Benfatto:2005:LTS

[87] Giuseppe Benfatto, Marzio Cassandro, I. Merola, and E. Presutti. Limit theorems for statistics of combinatorial partitions with applications to mean field Bose gas. Journal of Mathematical Physics, 46(3): 033303, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i3/p033303_s1.

Stoilova:2005:CGQa

[88] N. I. Stoilova and J. Van der Jeugt. A classification of generalized quantum statistics associated with classical Lie algebras. Journal of Mathematical Physics, 46(3):033501, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i3/p033501_s1.

Pelap:2005:SEO

[89] François B. Pelap and Mansour M. Faye. Solitonlike excitations in a one-dimensional electrical transmission line. Journal of Mathematical Physics, 46(3):033502, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i3/p033502_s1.

Leandre:2005:BPD

[90] Rémi Léandre. Brownian pants and Deligne cohomology. Journal of Mathematical Physics, 46(3):033503, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p033503_s1.

Zamastil:2005:MVC

[91] J. Zamastil, M. Simánek, J. Cízek, and L. Skála. Method of variation of constants for difference equations and its application to the calculation of atomic integrals. Journal of Mathematical Physics, 46(3): 033504, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i3/p033504_s1.

Fukuyama:2005:GTU

[92] Takeshi Fukuyama, Amon Ilakovac, Tatsuru Kikuchi, Stjepan Meljanac, and Nobuchika Okada. SO(10) group theory for the unified model building. Journal of Mathematical Physics, 46(3):033505, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/
p033505_s1.

Khodarinova:2005:QID

[93] L. A. Khodarinova. Quantum integrability of the deformed elliptic Calogero–Moser problem. Journal of Mathematical Physics, 46(3): 033506, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i3/p033506_s1.

Ma:2005:ICV

[94] Wen-Xiu Ma. Integrable couplings of vector AKNS soliton equations. Journal of Mathematical Physics, 46(3):033507, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/p033507_s1.

Moriya:2005:VFS

[95] Hajime Moriya. Validity and failure of some entropy inequalities for CAR systems. Journal of Mathematical Physics, 46(3):033508, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i3/p033508_s1.

Pantoja:2005:SDD

[96] Nelson Pantoja and Alberto Sanoja. Symmetries of distributional domain wall geometries. Journal of Mathematical Physics, 46(3):033509, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i3/p033509_s1.

Baake:2005:SPS

[97] M. Baake and D. Frettlöh. SCD patterns have singular diffraction. Journal of Mathematical Physics, 46(3):033510, March 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i3/p033510_s1.

Schunck:2005:GAS

[98] A. F. Schunck and Chris Wainwright. A geometric approach to scalar field theories on the supersphere. Journal of Mathematical Physics, 46(3): 033511, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i3/p033511_s1.

Derezinski:2005:LSO

[99] Jan Dereziński and Rafal Früboes. Level shift operator and second order perturbation theory. Journal of Mathematical Physics, 46(3): 033512, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i3/p033512_s1.

Aigon-Dupuy:2005:HOT

[100] Aline Aigon-Dupuy, Peter Buser, Michel Cibils, Alfred F. Künzle, and Frank Steiner. Hyperbolic octagons and Teichmüller space in genus 2. Journal of Mathematical Physics, 46(3):033513, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/p033513_s1.

Kota:2005:WRA

[101] V. K. B. Kota. SU(N) Wigner–Racah algebra for the matrix of second moments of embedded Gaussian unitary ensemble of random matrices. Journal of Mathematical Physics, 46(3):033514, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i3/p033514_s1.

Zhou:2005:DTE

[102] Zi-Xiang Zhou. Darboux transformations and exact solutions of two dimensional A_{2n}^{(2)} Toda equation. Journal of Mathematical Physics, 46(3): 033515, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i3/p033515_s1.

Boyallian:2005:RCL

[103] Carina Boyallian and José I. Liberati. Representations of classical Lie subalgebras of quantum pseudodifferential operators. Journal of Mathematical Physics, 46(3):033516, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i3/p033516_s1.

Zamastil:2005:EUA

[104] J. Zamastil, J. Cízek, M. Kalhous, L. Skála, and M. Simánek. Erratum: “The use of so(2, 1) algebra for the evaluation of atomic integrals: The study of two-electron atoms”, [J. Math. Phys. 45, 2674 (2004)]. Journal of Mathematical Physics, 46(3):039901, March 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i3/p039901_s1. See [6].

Karatsuba:2005:ITL

[105] Ekatherina A. Karatsuba and Paolo Moretti. Inversion time of large spins. Journal of Mathematical Physics, 46(4):042101, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/
p042101_s1.

Goldin:2005:QKB

[106] Gerald A. Goldin, Robert Owczarek, and David H. Sharp. Quantum kinematics of bosonic vortex loops. Journal of Mathematical Physics, 46(4): 042102, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p042102_s1.

Wood:2005:BIW

[107] J. G. Wood and A. J. Bracken. Bounds on integrals of the Wigner function: The hyperbolic case. Journal of Mathematical Physics, 46(4): 042103, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p042103_s1.

Blasi:2005:ADQ

[108] A. Blasi, G. Scolarici, and L. Solombrino. Alternative descriptions in quaternionic quantum mechanics. Journal of Mathematical Physics, 46(4):042104, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p042104_s1.

Giovannetti:2005:CMM

[109] Vittorio Giovannetti, Seth Lloyd, and Mary Beth Ruskai. Conditions for multiplicativity of maximal l_{p}-norms of channels for fixed integer p. Journal of Mathematical Physics, 46(4):042105, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/

p042105_s1. See publisher’s note [212].

Cornean:2005:RPL

[110] Horia D. Cornean, Arne Jensen, and Valeriu Moldoveanu. A rigorous proof of the Landauer–Büttiker formula. Journal of Mathematical Physics, 46(4):042106, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p042106_s1.

Kiukas:2005:MOC

[111] Jukka Kiukas, Pekka Lahti, and Kari Ylinen. Moment operators of the Cartesian margins of the phase space observables. Journal of Mathematical Physics, 46(4):042107, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i4/p042107_s1.

Scherer:2005:ISD

[112] Artur Scherer and Andrei N. Soklakov. Initial states and decoherence of histories. Journal of Mathematical Physics, 46(4):042108, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/
p042108_s1.

Park:2005:IUR

[113] Yong Moon Park. Improvement of uncertainty relations for mixed states. Journal of Mathematical Physics, 46(4):042109, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/p042109_s1.

Roknizadeh:2005:CDF

[114] R. Roknizadeh and M. K. Tavassoly. Construction of the dual family of Gazeau–Klauder coherent states via temporally stable nonlinear coherent states. Journal of Mathematical Physics, 46(4):042110, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i4/p042110_s1.

Haba:2005:EGF

[115] Z. Haba. Estimates on Green functions of second order differential operators with singular coefficients. Journal of Mathematical Physics, 46(4): 042301, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p042301_s1.

Hiroshima:2005:MRN

[116] Fumio Hiroshima and Herbert Spohn. Mass renormalization in nonrelativistic quantum electrodynamics. Journal of Mathematical Physics, 46(4):042302, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p042302_s1.

Chae:2005:ESC

[117] Dongho Chae. Existence of the semilocal Chern–Simons vortices. Journal of Mathematical Physics, 46(4):042303, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i4/p042303_s1.

Tolksdorf:2005:TEI

[118] Jürgen Tolksdorf. The topology of the electroweak interaction. Journal of Mathematical Physics, 46(4):042304, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i4/p042304_s1.

Pirozhenko:2005:IEH

[119] I. G. Pirozhenko, V. V. Nesterenko, and M. Bordag. Integral equations for heat kernel in compound media. Journal of Mathematical Physics, 46(4):042305, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p042305_s1.

Ioannidou:2005:UCQ

[120] T. A. Ioannidou, A. Kouiroukidis, and N. D. Vlachos. Universality in a class of Q-ball solutions: An analytic approach. Journal of Mathematical Physics, 46(4):042306, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p042306_s1.

Mine:2005:RBG

[121] M. Mine, M. Okumura, and Y. Yamanaka. Relation between generalized Bogoliubov and Bogoliubov–de Gennes approaches including Nambu– Goldstone mode. Journal of Mathematical Physics, 46(4):042307, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i4/p042307_s1.

Nagi:2005:ESA

[122] Jasbir Nagi. On extensions of superconformal algebras. Journal of Mathematical Physics, 46(4):042308, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i4/p042308_s1.

Loya:2005:FDD

[123] Paul Loya and Jinsung Park. Functional determinants for the Dirac equation with mixed pseudodifferential boundary conditions over finite cylinders. Journal of Mathematical Physics, 46(4):042309, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/
p042309_s1.

Han:2005:NBS

[124] Jongmin Han and Jaeduk Jang. Nontopological bare solutions in the relativistic self-dual Maxwell–Chern–Simons–Higgs model. Journal of Mathematical Physics, 46(4):042310, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i4/p042310_s1.

Amniat-Talab:2005:QAR

[125] M. Amniat-Talab, S. Guérin, and H. R. Jauslin. Quantum averaging and resonances: Two-level atom in a one-mode quantized field. Journal of Mathematical Physics, 46(4):042311, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i4/p042311_s1.

Bengtsson:2005:AIH

[126] A. K. H. Bengtsson. An abstract interface to higher spin gauge field theory. Journal of Mathematical Physics, 46(4):042312, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/p042312_s1.

Lombardo:2005:NSS

[127] D. J. Cirilo Lombardo. New spherically symmetric monopole and regular solutions in Einstein–Born–Infeld theories. Journal of Mathematical Physics, 46(4):042501, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p042501_s1.

Sharma:2005:CSE

[128] Ramesh Sharma. Conformal symmetries of Einstein’s field equations and initial data. Journal of Mathematical Physics, 46(4):042502, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i4/p042502_s1.

Vacaru:2005:ESN

[129] Sergiu I. Vacaru. Exact solutions with noncommutative symmetries in Einstein and gauge gravity. Journal of Mathematical Physics, 46(4): 042503, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p042503_s1.

Cardoso:2005:TSC

[130] J. G. Cardoso. Two-spinor curvature structures for a class of conformally flat space–times. Journal of Mathematical Physics, 46(4):042504, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i4/p042504_s1.

Dai:2005:NPE

[131] Xianzhe Dai. A note on positive energy theorem for spaces with asymptotic SUSY compactification. Journal of Mathematical Physics, 46(4): 042505, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p042505_s1.

Krisch:2005:FBG

[132] J. P. Krisch. Fractional boundary for the Gott–Hiscock string. Journal of Mathematical Physics, 46(4):042506, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i4/p042506_s1.

Kisisel:2005:QPB

[133] Ali Ulas Özgür Kisisel. On quadratic Poisson brackets. Journal of Mathematical Physics, 46(4):042701, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i4/p042701_s1.

Szablikowski:2005:DSM

[134] Blazej M. Szablikowski and Maciej Blaszak. On deformations of standard R-matrices for integrable infinite-dimensional systems. Journal of Mathematical Physics, 46(4):042702, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i4/p042702_s1.

Caudrelier:2005:SQN

[135] V. Caudrelier, M. Mintchev, and E. Ragoucy. Solving the quantum nonlinear Schrödinger equation with δ-type impurity. Journal of Mathematical Physics, 46(4):042703, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i4/p042703_s1.

Casati:2005:LNC

[136] Paolo Casati, Paolo Lorenzoni, Giovanni Ortenzi, and Marco Pedroni. On the local and nonlocal Camassa–Holm hierarchies. Journal of Mathematical Physics, 46(4):042704, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i4/p042704_s1.

Benenti:2005:VST

[137] S. Benenti, C. Chanu, and G. Rastelli. Variable-separation theory for the null Hamilton–Jacobi equation. Journal of Mathematical Physics, 46(4): 042901, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p042901_s1.

Billig:2005:MHA

[138] Yuly Billig. Magnetic hydrodynamics with asymmetric stress tensor. Journal of Mathematical Physics, 46(4):043101, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/p043101_s1.

Eshraghi:2005:SCV

[139] Homayoon Eshraghi and Yousefali Abedini. A special class in vanishing helicity ideal flows. Journal of Mathematical Physics, 46(4):043102, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i4/p043102_s1.

Ray:2005:BAS

[140] Subhankar Ray and J. Shamanna. A Bethe Ansatz study of free energy and excitation spectrum for even spin Fateev–Zamolodchikov model. Journal of Mathematical Physics, 46(4):043301, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/p043301_s1.

Elvang:2005:DYP

[141] Henriette Elvang, Predrag Cvitanović, and Anthony D. Kennedy. Diagrammatic Young projection operators for U(n). Journal of Mathematical Physics, 46(4):043501, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p043501_s1.

Popovych:2005:HCL

[142] Roman O. Popovych and Nataliya M. Ivanova. Hierarchy of conservation laws of diffusion-convection equations. Journal of Mathematical Physics, 46(4):043502, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p043502_s1.

Gayral:2005:SAM

[143] Victor Gayral and Bruno Iochum. The spectral action for Moyal planes. Journal of Mathematical Physics, 46(4):043503, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/p043503_s1.

Pickering:2005:DSI

[144] A. Pickering and J. Prada. Dirichlet series and the integrability of multilinear differential equations. Journal of Mathematical Physics, 46(4):

043504, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p043504_s1.

Bentosela:2005:DOD

[145] F. Bentosela, P. Duclos, V. Moldoveanu, and G. Nenciu. The dynamics of one-dimensional Bloch electrons in constant electric fields. Journal of Mathematical Physics, 46(4):043505, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i4/p043505_s1.

Talukdar:2005:IVP

[146] B. Talukdar, S. Ghosh, and U. Das. Inverse variational problem and canonical structure of Burgers equations. Journal of Mathematical Physics, 46(4):043506, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p043506_s1.

Niestegge:2005:CEA

[147] Gerd Niestegge. Conditional expectations associated with quantum states. Journal of Mathematical Physics, 46(4):043507, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/
p043507_s1.

Aubin:2005:CEW

[148] Jean-Pierre Aubin and Annick Lesne. Constructing and exploring wells of energy landscapes. Journal of Mathematical Physics, 46(4):043508, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i4/p043508_s1.

Rogers:2005:PFE

[149] Mathew D. Rogers. Partial fractions expansions and identities for products of Bessel functions. Journal of Mathematical Physics, 46(4):043509, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i4/p043509_s1.

Xia:2005:MGK

[150] Tiecheng Xia and Engui Fan. The multicomponent generalized Kaup– Newell hierarchy and its multicomponent integrable couplings system with two arbitrary functions. Journal of Mathematical Physics, 46(4):

043510, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p043510_s1.

Belhaj:2005:TGS

[151] Adil Belhaj and Jørgen Rasmussen. On toric geometry, Spin(7) manifolds, and type II superstring compactifications. Journal of Mathematical Physics, 46(4):043511, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i4/p043511_s1.

Jureit:2005:CIAa

[152] Jan-Hendrik Jureit and Christoph A. Stephan. On a classification of irreducible almost commutative geometries, a second helping. Journal of Mathematical Physics, 46(4):043512, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i4/p043512_s1.

Zhang:2005:SFH

[153] R. B. Zhang and Y. M. Zou. Spherical functions on homogeneous superspaces. Journal of Mathematical Physics, 46(4):043513, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i4/
p043513_s1.

Michel:2005:SAS

[154] Laurent Michel. Scattering amplitude and scattering phase for the Schrödinger equation with strong magnetic field. Journal of Mathematical Physics, 46(4):043514, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i4/p043514_s1.

Davies:2005:SHB

[155] Ian M. Davies, Aubrey Truman, and Huaizhong Zhao. Stochastic heat and Burgers equations and their singularities. II. Analytical properties and limiting distributions. Journal of Mathematical Physics, 46(4): 043515, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p043515_s1.

Garoni:2005:ASL

[156] T. M. Garoni. On the asymptotics of some large Hankel determinants generated by Fisher–Hartwig symbols defined on the real line. Journal of Mathematical Physics, 46(4):043516, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i4/p043516_s1.

Hallnas:2005:EST

[157] Martin Hallnäs and Edwin Langmann. Exact solutions of two complementary one-dimensional quantum many-body systems on the half-line. Journal of Mathematical Physics, 46(5):052101, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p052101_s1.

Lan:2005:ACT

[158] Chunhua Lan, Tzyh-Jong Tarn, Quo-Shin Chi, and John W. Clark. Analytic controllability of time-dependent quantum control systems. Journal of Mathematical Physics, 46(5):052102, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i5/p052102_s1.

Lucarelli:2005:CAH

[159] Dennis Lucarelli. Control aspects of holonomic quantum computation. Journal of Mathematical Physics, 46(5):052103, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p052103_s1.

Gridnev:2005:STU

[160] D. K. Gridnev, C. Greiner, and W. Greiner. Stability of three unit charges: Necessary conditions. Journal of Mathematical Physics, 46(5): 052104, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p052104_s1.

Mantoiu:2005:SDQ

[161] Marius Mantoiu and Radu Purice. Strict deformation quantization for a particle in a magnetic field. Journal of Mathematical Physics, 46(5): 052105, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p052105_s1.

Jiang:2005:AMD

[162] Tongsong Jiang. Algebraic methods for diagonalization of a quaternion matrix in quaternionic quantum theory. Journal of Mathematical Physics, 46(5):052106, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p052106_s1.

Appleby:2005:SIC

[163] D. M. Appleby. Symmetric informationally complete-positive operator valued measures and the extended Clifford group. Journal of Mathematical Physics, 46(5):052107, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i5/p052107_s1.

Gray:2005:MAM

[164] John E. Gray and Andrew Vogt. Mathematical analysis of the Mandelstam–Tamm time-energy uncertainty principle. Journal of Mathematical Physics, 46(5):052108, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i5/p052108_s1.

Ruuge:2005:SKS

[165] Artur E. Ruuge and Freddy van Oystaeyen. Saturated Kochen–Speckertype configuration of 120 projective lines in eight-dimensional space and its group of symmetry. Journal of Mathematical Physics, 46(5):052109, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i5/p052109_s1.

Blasiak:2005:SUC

[166] P. Blasiak, K. A. Penson, A. I. Solomon, A. Horzela, and G. H. E. Duchamp. Some useful combinatorial formulas for bosonic operators. Journal of Mathematical Physics, 46(5):052110, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p052110_s1.

Liu:2005:VSA

[167] Zuhan Liu. Vortices set and the applied magnetic field for superconductivity in dimension three. Journal of Mathematical Physics, 46(5): 052111, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p052111_s1.

Shirai:2005:SEF

[168] Shin ichi Shirai. Strong-electric-field eigenvalue asymptotics for the Iwatsuka model. Journal of Mathematical Physics, 46(5):052112, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/
p052112_s1.

Bostelmann:2005:PSP

[169] Henning Bostelmann. Phase space properties and the short distance structure in quantum field theory. Journal of Mathematical Physics, 46 (5):052301, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p052301_s1.

Bender:2005:WPL

[170] Carl M. Bender, Peter N. Meisinger, and Qinghai Wang. Wilson polynomials and the Lorentz transformation properties of the parity operator. Journal of Mathematical Physics, 46(5):052302, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p052302_s1.

Huh:2005:CPG

[171] Hyungjin Huh. Cauchy problems of the gauged sigma model. Journal of Mathematical Physics, 46(5):052303, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i5/p052303_s1.

Adam:2005:SDP

[172] C. Adam and J. Sánchez-Guillén. The symmetries of the Dirac– Pauli equation in two and three dimensions. Journal of Mathematical Physics, 46(5):052304, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p052304_s1.

Chen:2005:KPA

[173] Heng-Yu Chen and N. S. Manton. The Kähler potential of Abelian Higgs vortices. Journal of Mathematical Physics, 46(5):052305, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i5/p052305_s1.

Arai:2005:SPD

[174] Asao Arai, Kunimitsu Hayashi, and Itaru Sasaki. Spectral properties of a Dirac operator in the chiral quark soliton model. Journal of Mathematical Physics, 46(5):052306, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p052306_s1.

Sorensen:2005:LBP

[175] Thomas Østergaard Sørensen. The large-Z behavior of pseudorelativistic atoms. Journal of Mathematical Physics, 46(5):052307, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/
p052307_s1.

Sharples:2005:LEQ

[176] Jason Sharples. Local existence of quasispherical space–time initial data. Journal of Mathematical Physics, 46(5):052501, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p052501_s1.

Wu:2005:TNL

[177] Yabo Wu, Peng Dong, Xuemei Deng, and Guoming Zhao. The two NUTlike solutions of Ernst equation. Journal of Mathematical Physics, 46(5): 052502, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p052502_s1.

Becerril:2005:CDK

[178] Ricardo Becerril and Alfredo Herrera-Aguilar. Charging a double Kerr solution in five-dimensional Einstein–Maxwell–Kalb–Ramond theory. Journal of Mathematical Physics, 46(5):052503, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p052503_s1.

Hamanaka:2005:CFC

[179] Masashi Hamanaka. Commuting flows and conservation laws for noncommutative Lax hierarchies. Journal of Mathematical Physics, 46(5): 052701, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p052701_s1.

Carinena:2005:CPS

[180] José F. Cariñena, Manuel F. Rañada, and Mariano Santander. Central potentials on spaces of constant curvature: The Kepler problem on the two-dimensional sphere S^{2} and the hyperbolic plane H^{2}. Journal of Mathematical Physics, 46(5):052702, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i5/p052702_s1. See comment

[440] and response [441].

Rey:2005:GFS

[181] Angel M. Rey, Narciso Román-Roy, and Modesto Salgado. Günther’s formalism (k-symplectic formalism) in classical field theory: Skinner– Rusk approach and the evolution operator. Journal of Mathematical Physics, 46(5):052901, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p052901_s1.

Pucacco:2005:CIH

[182] Giuseppe Pucacco and Kjell Rosquist. Configurational invariants of Hamiltonian systems. Journal of Mathematical Physics, 46(5):052902, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i5/p052902_s1.

Brajercik:2005:VPL

[183] J. Brajercík and D. Krupka. Variational principles for locally variational forms. Journal of Mathematical Physics, 46(5):052903, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/
p052903_s1.

Lahno:2005:GCN

[184] V. Lahno and R. Zhdanov. Group classification of nonlinear wave equations. Journal of Mathematical Physics, 46(5):053301, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/
p053301_s1.

Thebaldi:2005:CED

[185] Rafael S. Thebaldi, Emmanuel Pereira, and Aldo Procacci. A cluster expansion for the decay of correlations of light-mass quantum crystals and some stochastic models under intense noise. Journal of Mathematical Physics, 46(5):053302, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p053302_s1.

Asch:2005:PWA

[186] J. Asch, I. Hradecký, and P. Stovícek. Propagators weakly associated to a family of Hamiltonians and the adiabatic theorem for the Landau Hamiltonian with a time-dependent Aharonov–Bohm flux. Journal of Mathematical Physics, 46(5):053303, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i5/p053303_s1.

Meson:2005:APP

[187] Alejandro Mesón and Fernando Vericat. An approach to the problem of phase transitions in the continuum. Journal of Mathematical Physics, 46 (5):053304, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p053304_s1.

Cassandro:2005:GCP

[188] Marzio Cassandro, Pablo Augusto Ferrari, Immacolata Merola, and Errico Presutti. Geometry of contours and Peierls estimates in d = 1 Ising models with long range interactions. Journal of Mathematical Physics, 46 (5):053305, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p053305_s1.

Alimohammadi:2005:MES

[189] M. Alimohammadi and Y. Naimi. Multispecies extension of the solvable partially asymmetric reaction-diffusion processes. Journal of Mathematical Physics, 46(5):053306, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i5/p053306_s1.

Duan:2005:GEB

[190] Renjun Duan, Tong Yang, and Changjiang Zhu. Global existence to Boltzmann equation with external force in infinite vacuum. Journal of Mathematical Physics, 46(5):053307, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i5/p053307_s1.

Frappat:2005:HKM

[191] L. Frappat and A. Sciarrino. Hyperbolic Kac–Moody superalgebras. Journal of Mathematical Physics, 46(5):053501, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p053501_s1.

Kubieniec:2005:UIR

[192] J. Kubieniec. On unitary, irreducible representations of the proper, orthochronous Lorentz group. Journal of Mathematical Physics, 46(5):

053502, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i5/p053502_s1.

Cingolani:2005:SSN

[193] Silvia Cingolani and Simone Secchi. Semiclassical states for NLS equations with magnetic potentials having polynomial growths. Journal of Mathematical Physics, 46(5):053503, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i5/p053503_s1.

Doikou:2005:RAT

[194] Anastasia Doikou. On reflection algebras and twisted Yangians. Journal of Mathematical Physics, 46(5):053504, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i5/p053504_s1.

Linnaeus:2005:SCS

[195] Staffan Linnæus. Stokes constants for a singular wave equation. Journal of Mathematical Physics, 46(5):053505, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i5/p053505_s1.

Bagarello:2005:RBM

[196] F. Bagarello. Relations between multiresolution analysis and quantum mechanics. Journal of Mathematical Physics, 46(5):053506, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/
p053506_s1.

Cordero:2005:ASG

[197] Patricio Cordero and Jamil Daboul. Analysis of the spectrum generating algebra method for obtaining energy spectra. Journal of Mathematical Physics, 46(5):053507, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p053507_s1.

Boulanger:2005:WCT

[198] Nicolas Boulanger. A Weyl-covariant tensor calculus. Journal of Mathematical Physics, 46(5):053508, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i5/p053508_s1.

Kalnins:2005:SOSa

[199] E. G. Kalnins, J. M. Kress, and W. Miller. Second-order superintegrable systems in conformally flat spaces. I. Two-dimensional classical structure theory. Journal of Mathematical Physics, 46(5):053509, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i5/p053509_s1.

Kalnins:2005:SOSb

[200] E. G. Kalnins, J. M. Kress, and W. Miller. Second order superintegrable systems in conformally flat spaces. II. The classical two-dimensional Stäckel transform. Journal of Mathematical Physics, 46(5):053510, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i5/p053510_s1.

Becker:2005:ESH

[201] Peter A. Becker. Exact solution for the hypergeometric Green’s function describing spectral formation in X-ray pulsars. Journal of Mathematical Physics, 46(5):053511, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p053511_s1.

Biswas:2005:HTG

[202] Indranil Biswas and A. K. Raina. Hecke transformation and the generalized theta function. Journal of Mathematical Physics, 46(5):053512, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i5/p053512_s1.

DelCastillo:2005:PFS

[203] Pierre Del Castillo. The Parr formula for the superheating field in a semi-infinite film. Journal of Mathematical Physics, 46(5):053513, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v46/i5/p053513_s1. See publisher’s note [302].

Patera:2005:DCCa

[204] J. Patera and A. Zaratsyan. Discrete and continuous cosine transform generalized to Lie groups SU(2) × SU(2) and O(5). Journal of Mathematical Physics, 46(5):053514, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i5/p053514_s1.

Khan:2005:CSR

[205] Abu M. Khan and Pierre Ramond. Continuous spin representations from group contraction. Journal of Mathematical Physics, 46(5):053515, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v46/i5/p053515_s1. See publisher’s note [301].

Duval:2005:QIQ

[206] C. Duval and G. Valent. Quantum integrability of quadratic Killing tensors. Journal of Mathematical Physics, 46(5):053516, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p053516_s1.

Bashkirov:2005:NST

[207] D. Bashkirov, G. Giachetta, L. Mangiarotti, and G. Sardanashvily. Noether’s second theorem for BRST symmetries. Journal of Mathematical Physics, 46(5):053517, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i5/p053517_s1.

Ali:2005:SPA

[208] S. Twareque Ali and F. Bagarello. Some physical appearances of vector coherent states and coherent states related to degenerate Hamiltonians. Journal of Mathematical Physics, 46(5):053518, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p053518_s1.

Blohmann:2005:RUD

[209] Christian Blohmann. Reconstruction of universal Drinfeld twists from representations. Journal of Mathematical Physics, 46(5):053519, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i5/p053519_s1.

Cuccagna:2005:BEE

[210] Scipio Cuccagna and Dmitry Pelinovsky. Bifurcations from the endpoints of the essential spectrum in the linearized nonlinear Schrödinger problem. Journal of Mathematical Physics, 46(5):053520, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i5/p053520_s1.

Hall:2005:ECS

[211] Brian C. Hall and Jeffrey J. Mitchell. Erratum: “Coherent states on spheres”, [J. Math. Phys. 43, 1211 (2002)]. Journal of Mathematical Physics, 46(5):059901, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i5/p059901_s1. See [1, 2208].

Giovannetti:2005:PNC

[212] Vittorio Giovannetti, Seth Lloyd, and Mary Beth Ruskai. Publisher’s note: “Conditions for multiplicativity of maximal l_{p}-norms of channels for fixed integer p”, [J. Math. Phys. 46, 042105 (2005)]. Journal of Mathematical Physics, 46(5):059902, May 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i5/p059902_s1. See [109].

Meolans:2005:SKP

[213] J. Gilbert Méolans and S. Kokou Dadzie. Scattering kernel for polyatomic molecules. Journal of Mathematical Physics, 46(6):062101, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p062101_s1.

Gheondea:2005:IQE

[214] Aurelian Gheondea, Stanley Gudder, and Peter Jonas. On the infimum of quantum effects. Journal of Mathematical Physics, 46(6):062102, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p062102_s1.

Nathanson:2005:DBO

[215] Michael Nathanson. Distinguishing bipartitite orthogonal states using LOCC: Best and worst cases. Journal of Mathematical Physics, 46(6): 062103, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p062103_s1.

Bullock:2005:TRQ

[216] Stephen S. Bullock, Gavin K. Brennen, and Dianne P. O’Leary. Time reversal and n-qubit canonical decompositions. Journal of Mathematical Physics, 46(6):062104, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i6/p062104_s1.

Exner:2005:IPL

[217] Pavel Exner. An isoperimetric problem for leaky loops and related meanchord inequalities. Journal of Mathematical Physics, 46(6):062105, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p062105_s1.

Belavkin:2005:ODF

[218] Viacheslav P. Belavkin, Giacomo Mauro D’Ariano, and Maxim Raginsky. Operational distance and fidelity for quantum channels. Journal of Mathematical Physics, 46(6):062106, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i6/p062106_s1.

Ayers:2005:GDF

[219] Paul W. Ayers. Generalized density functional theories using the kelectron densities: Development of kinetic energy functionals. Journal of Mathematical Physics, 46(6):062107, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i6/p062107_s1.

Plante:2005:ASS

[220] Guillaume Plante and Adel F. Antippa. Analytic solution of the Schrödinger equation for the Coulomb-plus-linear potential. I. The wave functions. Journal of Mathematical Physics, 46(6):062108, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i6/
p062108_s1.

Znojil:2005:SSM

[221] Miloslav Znojil. Solvable P T -symmetric model with a tunable interspersion of nonmerging levels. Journal of Mathematical Physics, 46(6): 062109, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p062109_s1.

Agop:2005:SMS

[222] M. Agop, P. D. Ioannou, and P. Nica. Superconductivity by means of the subquantum medium coherence. Journal of Mathematical Physics, 46(6): 062110, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p062110_s1.

Khrennikov:2005:RCS

[223] Andrei Khrennikov. Representation of the contextual statistical model by hyperbolic amplitudes. Journal of Mathematical Physics, 46(6):062111, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p062111_s1.

Pironio:2005:LBI

[224] Stefano Pironio. Lifting Bell inequalities. Journal of Mathematical Physics, 46(6):062112, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i6/p062112_s1.

Ghoshal:2005:NSF

[225] Debashis Ghoshal, Dileep P. Jatkar, and Maximilian Kreuzer. Neveu– Schwarz fivebrane and tachyon condensation. Journal of Mathematical Physics, 46(6):062301, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i6/p062301_s1.

Braga:2005:NOB

[226] Nelson R. F. Braga, Hector L. Carrion, and Cresus F. L. Godinho. Normal ordering and boundary conditions in open bosonic strings. Journal of Mathematical Physics, 46(6):062302, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i6/p062302_s1.

Moretti:2005:BEC

[227] Valter Moretti and Nicola Pinamonti. Bose–Einstein condensate and spontaneous breaking of conformal symmetry on Killing horizons. Journal of Mathematical Physics, 46(6):062303, June 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i6/p062303_s1.

Lange:2005:EGA

[228] Alexander Lange. The Epstein–Glaser approach to perturbative quantum field theory: graphs and Hopf algebras. Journal of Mathematical Physics, 46(6):062304, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i6/p062304_s1.

Brenier:2005:DPS

[229] Yann Brenier and Wen-An Yong. Derivation of particle, string, and membrane motions from the Born–Infeld electromagnetism. Journal of Mathematical Physics, 46(6):062305, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i6/p062305_s1.

G:2005:ARF

[230] J. Fernando Barbero G., Guillermo A. Mena Marugán, and Eduardo J. S. Villaseñor. Asymptotics of regulated field commutators for Einstein– Rosen waves. Journal of Mathematical Physics, 46(6):062306, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p062306_s1.

Shklyarov:2005:CDO

[231] Dmitry Shklyarov and Genkai Zhang. Covariant q-differential operators and unitary highest weight representations for U_{q}∫⊓_{(\,\)}. Journal of Mathematical Physics, 46(6):062307, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i6/p062307_s1.

Krisch:2005:ETV

[232] J. P. Krisch and E. N. Glass. Energy transport in the Vaidya system. Journal of Mathematical Physics, 46(6):062501, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i6/p062501_s1.

Ash:2005:RPO

[233] Avner Ash and Patrick McDonald. Random partial orders, posts, and the causal set approach to discrete quantum gravity. Journal of Mathematical Physics, 46(6):062502, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i6/p062502_s1.

Oliveira-Neto:2005:WSS

[234] G. Oliveira-Neto and F. I. Takakura. Wyman’s solution, self-similarity, and critical behavior. Journal of Mathematical Physics, 46(6):062503, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p062503_s1.

Liko:2005:ESF

[235] Tomás Liko and Paul S. Wesson. An exact solution of the fivedimensional Einstein equations with four-dimensional de Sitter-like expansion. Journal of Mathematical Physics, 46(6):062504, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i6/
p062504_s1.

Alba:2005:MEC

[236] David Alba, Luca Lusanna, and Massimo Pauri. Multipolar expansions for closed and open systems of relativistic particles. Journal of Mathematical Physics, 46(6):062505, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i6/p062505_s1.

Deryabin:2005:AHI

[237] Mikhail V. Deryabin. On asymptotic Hopf invariant for Hamiltonian systems. Journal of Mathematical Physics, 46(6):062701, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i6/
p062701_s1.

Benatti:2005:CLD

[238] Fabio Benatti and Valerio Cappellini. Continuous limit of discrete sawtooth maps and its algebraic framework. Journal of Mathematical Physics, 46(6):062702, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i6/p062702_s1.

Carinena:2005:LFN

[239] José F. Cariñena, Manuel F. Rañada, and Mariano Santander. Lagrangian formalism for nonlinear second-order Riccati systems: Onedimensional integrability and two-dimensional superintegrability. Journal of Mathematical Physics, 46(6):062703, June 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i6/p062703_s1.

Gentile:2005:SDS

[240] Guido Gentile, Michele V. Bartuccelli, and Jonathan H. B. Deane. Summation of divergent series and Borel summability for strongly dissipative differential equations with periodic or quasiperiodic forcing terms. Journal of Mathematical Physics, 46(6):062704, June 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i6/p062704_s1.

Ujino:2005:SID

[241] Hideaki Ujino, Luc Vinet, and Haruo Yoshida. A super-integrable discretization of the Calogero model. Journal of Mathematical Physics, 46 (6):062705, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p062705_s1.

Dai:2005:PHS

[242] Bo Dai and Chuu-Lian Terng. Periodic and homoclinic solutions of the modified 2 + 1 chiral model. Journal of Mathematical Physics, 46(6): 062706, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p062706_s1.

Maciejewski:2005:DPI

[243] Andrzej J. Maciejewski and Maria Przybylska. Darboux points and integrability of Hamiltonian systems with homogeneous polynomial potential. Journal of Mathematical Physics, 46(6):062901, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i6/p062901_s1.

Guo:2005:NVV

[244] Yong-Xin Guo, Yong Wang, G. Y. Chee, and Feng-Xiang Mei. Nonholonomic versus vakonomic dynamics on a Riemann–Cartan manifold. Journal of Mathematical Physics, 46(6):062902, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i6/p062902_s1.

Costeniuc:2005:CAP

[245] Marius Costeniuc, Richard S. Ellis, and Hugo Touchette. Complete analysis of phase transitions and ensemble equivalence for the Curie– Weiss–Potts model. Journal of Mathematical Physics, 46(6):063301, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p063301_s1.

Kozyrev:2005:NUD

[246] S. V. Kozyrev, V. Al. Osipov, and V. A. Avetisov. Nondegenerate ultrametric diffusion. Journal of Mathematical Physics, 46(6):063302, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p063302_s1.

Cornille:2005:PMD

[247] Henri Cornille. Planar (x, z) mixtures discrete Boltzmann equations with species either in odd or even |x| + |z| values. Journal of Mathematical Physics, 46(6):063303, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i6/p063303_s1.

Zhang:2005:ECC

[248] Fuxi Zhang. Exponential convergence of coupled diffusion processes. Journal of Mathematical Physics, 46(6):063304, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i6/p063304_s1.

Field:2005:OSC

[249] Timothy R. Field. Observability of the scattering cross-section through phase decoherence. Journal of Mathematical Physics, 46(6):063305, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p063305_s1.

Pelavas:2005:VST

[250] N. Pelavas, A. Coley, R. Milson, V. Pravda, and A. Pravdová. VSI_{i} space–times and the ϵ-property. Journal of Mathematical Physics, 46(6): 063501, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p063501_s1.

Gao:2005:MCS

[251] Jining Gao. The Maurer–Cartan structure of BRST differentials. Journal of Mathematical Physics, 46(6):063502, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i6/p063502_s1.

Pecina-Cruz:2005:CGA

[252] J. N. Pecina-Cruz. On the Casimir of the group ISL(n, R) and its algebraic decomposition. Journal of Mathematical Physics, 46(6):063503, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p063503_s1.

Gunther:2005:MDS

[253] Uwe Günther, Frank Stefani, and Miloslav Znojil. MHD α^{2}-dynamo, Squire equation and P T -symmetric interpolation between square well and harmonic oscillator. Journal of Mathematical Physics, 46(6):063504, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p063504_s1.

deFaria:2005:FIN

[254] Margarida de Faria, Maria João Oliveira, and Ludwig Streit. Feynman integrals for nonsmooth and rapidly growing potentials. Journal of Mathematical Physics, 46(6):063505, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i6/p063505_s1.

Loginov:2005:MIS

[255] E. K. Loginov. Multi-instantons in seven dimensions. Journal of Mathematical Physics, 46(6):063506, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i6/p063506_s1.

Ochiai:2005:IMP

[256] Tomoshiro Ochiai and Jose C. Nacher. Inversible Max-Plus algebras and integrable systems. Journal of Mathematical Physics, 46(6):063507, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p063507_s1.

Chakrabarti:2005:NSP

[257] A. Chakrabarti. A nested sequence of projectors and corresponding braid matrices R̂(θ). 1. Odd dimensions. Journal of Mathematical Physics, 46 (6):063508, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p063508_s1.

Chakrabarti:2005:ANC

[258] A. Chakrabarti. Aspects of a new class of braid matrices: Roots of unity and hyperelliptic q for triangularity, L-algebra, link-invariants, noncommutative spaces. Journal of Mathematical Physics, 46(6):063509, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i6/p063509_s1.

Transtrum:2005:CRF

[259] Mark K. Transtrum and Jean-François S. Van Huele. Commutation relations for functions of operators. Journal of Mathematical Physics, 46(6): 063510, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p063510_s1.

Derezinski:2005:SVH

[260] Jan Dereziński and Rafal Früboes. Stationary van Hove limit. Journal of Mathematical Physics, 46(6):063511, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i6/p063511_s1.

Keraani:2005:WMD

[261] Sahbi Keraani. Wigner measures dynamics in a Coulomb potential. Journal of Mathematical Physics, 46(6):063512, June 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i6/p063512_s1.

Odake:2005:EPS

[262] Satoru Odake and Ryu Sasaki. Equilibrium positions, shape invariance and Askey–Wilson polynomials. Journal of Mathematical Physics, 46(6): 063513, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p063513_s1.

Saenz:2005:EQM

[263] A. W. Sáenz. Erratum: “Quantum-mechanical scattering in exterior domains with impenetrable boundaries and short-range potentials”, [J. Math. Phys. 45, 1426 (2004)]. Journal of Mathematical Physics, 46(6): 069901, June 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i6/p069901_s1. See [5].

Billionnet:2005:HST

[264] Claude Billionnet. Hybrid states of two and three level atoms. Journal of Mathematical Physics, 46(7):072101, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i7/p072101_s1.

Viennot:2005:PBS

[265] David Viennot. Principal bundle structure of quantum adiabatic dynamics with a Berry phase which does not commute with the dynamical phase. Journal of Mathematical Physics, 46(7):072102, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/
p072102_s1.

Jurciukonis:2005:CQS

[266] D. Jurciukonis, E. Norvaisas, and D. O. Riska. Canonical quantization of SU(3) Skyrme model in a general representation. Journal of Mathematical Physics, 46(7):072103, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i7/p072103_s1.

March:2005:CNP

[267] N. H. March, I. A. Howard, I. Nagy, and P. M. Echenique. Changes in nonlinear potential scattering theory in electron gases brought about by reducing dimensionality. Journal of Mathematical Physics, 46(7): 072104, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i7/p072104_s1.

deOliveira:2005:SLP

[268] César R. de Oliveira and Roberto A. Prado. Spectral and localization properties for the one-dimensional Bernoulli discrete Dirac operator. Journal of Mathematical Physics, 46(7):072105, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/p072105_s1.

Liu:2005:CDC

[269] X.-M. Liu, C. Quesne, and F. Song. k-component q-deformed charge coherent states and their nonclassical properties. Journal of Mathematical Physics, 46(7):072106, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i7/p072106_s1.

Dias:2005:SGP

[270] Nuno Costa Dias and João Nuno Prata. Stargenfunctions, generally parametrized systems and a causal formulation of phase space quantum mechanics. Journal of Mathematical Physics, 46(7):072107, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i7/p072107_s1.

Pogosyan:2005:QPS

[271] George S. Pogosyan, Luis Edgar Vicent, and Kurt Bernardo Wolf. Quantum phase space for the one-dimensional hydrogen atom on the hyperbola. Journal of Mathematical Physics, 46(7):072108, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/p072108_s1.

Batalin:2005:NAC

[272] I. Batalin, M. Grigoriev, and S. Lyakhovich. Non-Abelian conversion and quantization of nonscalar second-class constraints. Journal of Mathematical Physics, 46(7):072301, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i7/p072301_s1.

Gitman:2005:LSL

[273] D. M. Gitman, P. Yu. Moshin, and A. A. Reshetnyak. Local superfield Lagrangian BRST quantization. Journal of Mathematical Physics, 46(7): 072302, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i7/p072302_s1.

Jureit:2005:CIAb

[274] Jan-Hendrik Jureit, Chtistoph Stephan, and Thomas Schücker. On a classification of irreducible almost commutative geometries. III. Journal of Mathematical Physics, 46(7):072303, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i7/p072303_s1.

Aleynikov:2005:SWE

[275] Dmitriy V. Aleynikov. Seiberg–Witten equations in R^{4}: Lie symmetries, particular solutions, integrability. Journal of Mathematical Physics, 46 (7):072304, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i7/p072304_s1.

Byun:2005:GCK

[276] Jungyoon Byun. A generalization of Connes–Kreimer Hopf algebra. Journal of Mathematical Physics, 46(7):072305, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/p072305_s1.

Wei:2005:BCF

[277] Yi Wei and Tilo Wettig. Bosonic color-flavor transformation for the special unitary group. Journal of Mathematical Physics, 46(7):072306, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i7/p072306_s1.

Andrianopoli:2005:IMS

[278] L. Andrianopoli, S. Ferrara, M. A. Lledó, and O. Maciá. Integration of massive states as contractions of nonlinear σ models. Journal of Mathematical Physics, 46(7):072307, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i7/p072307_s1.

Mendels:2005:NTE

[279] E. Mendels. Near threshold expansion of Feynman diagrams. Journal of Mathematical Physics, 46(7):072308, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i7/p072308_s1.

Bruna:2005:LHS

[280] Lluís Bruna and Joan Girbau. Laplacian in the hyperbolic space H_{n} and linearization stability of the Einstein equation for Robertson–Walker models. Journal of Mathematical Physics, 46(7):072501, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/
p072501_s1.

Zegers:2005:BTL

[281] Robin Zegers. Birkhoff’s theorem in Lovelock gravity. Journal of Mathematical Physics, 46(7):072502, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i7/p072502_s1.

Diacu:2005:KPA

[282] Florin Diacu, Ernesto Pérez-Chavela, and Manuele Santoprete. The Kepler problem with anisotropic perturbations. Journal of Mathematical Physics, 46(7):072701, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i7/p072701_s1.

Dobrovolschi:2005:BIA

[283] Dan Dobrovolschi and Jaume Llibre. Boundedness, invariant algebraic surfaces and global dynamics for a spectral model of large-scale atmospheric circulation. Journal of Mathematical Physics, 46(7):072702, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i7/p072702_s1.

Llibre:2005:IBI

[284] Jaume Llibre and Clàudia Valls. Integrability of the Bianchi IX system. Journal of Mathematical Physics, 46(7):072901, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/p072901_s1.

Du:2005:MSH

[285] Y. Du and X. B. Pan. Multiple states and hysteresis for type I superconductors. Journal of Mathematical Physics, 46(7):073301, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i7/p073301_s1.

Malyshev:2005:DCI

[286] V. A. Malyshev. Dynamical clusters of infinite particle dynamics. Journal of Mathematical Physics, 46(7):073302, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i7/p073302_s1.

Melnichuk:2005:CDW

[287] O. P. Melnichuk and I. Yu. Popov. Coupled dielectric waveguides: variational estimations. Journal of Mathematical Physics, 46(7):073501, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i7/p073501_s1.

Yang:2005:WHA

[288] Shilin Yang. Weak Hopf algebras corresponding to Cartan matrices. Journal of Mathematical Physics, 46(7):073502, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/p073502_s1.

Cronstrom:2005:USG

[289] Christofer Cronström. On the uniqueness of solutions to gauge covariant Poisson equations with compact Lie algebras. Journal of Mathematical Physics, 46(7):073503, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i7/p073503_s1.

Karabulut:2005:WFM

[290] Hasan Karabulut. A Wannier function made from distributed Gaussians. Journal of Mathematical Physics, 46(7):073504, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/p073504_s1.

Garcia-Rio:2005:FDI

[291] Eduardo García-Río, Zoran Rakić, and M. E. Vázquez-Abal. Fourdimensional indefinite Kähler Osserman manifolds. Journal of Mathematical Physics, 46(7):073505, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i7/p073505_s1.

Popov:2005:EPM

[292] Alexander D. Popov. On explicit point multimonopoles in SU(2) gauge theory. Journal of Mathematical Physics, 46(7):073506, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/
p073506_s1.

Reyes:2005:NSK

[293] Enrique G. Reyes. Nonlocal symmetries and the Kaup–Kupershmidt equation. Journal of Mathematical Physics, 46(7):073507, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/
p073507_s1.

Nunez:2005:SRE

[294] J. Fernández Núñez, W. García Fuertes, and A. M. Perelomov. Some results on the eigenfunctions of the quantum trigonometric Calogero– Sutherland model related to the Lie algebra E_{6}. Journal of Mathematical Physics, 46(7):073508, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i7/p073508_s1.

Kawamura:2005:ERC

[295] Katsunori Kawamura. Extensions of representations of the CAR algebra to the Cuntz algebra O_{2} — the Fock and the infinite wedge. Journal of Mathematical Physics, 46(7):073509, July 2005. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i7/p073509_s1.

Guo:2005:BPQ

[296] Boling Guo, Jianqing Chen, and Fengqiu Su. The “Blow up” problem for a quasilinear Schrödinger equation. Journal of Mathematical Physics, 46 (7):073510, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i7/p073510_s1.

Yue:2005:LCR

[297] Jin Yue. The 1856 lemma of Cayley revisited. I. Infinitesimal generators. Journal of Mathematical Physics, 46(7):073511, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/p073511_s1.

vonKeviczky:2005:GFS

[298] Attila B. von Keviczky, Nasser Saad, and Richard L. Hall. Green’s function for a Schrödinger operator and some related summation formulas. Journal of Mathematical Physics, 46(7):073512, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/p073512_s1.

Bluman:2005:CSC

[299] G. W. Bluman and Temuerchaolu. Comparing symmetries and conservation laws of nonlinear telegraph equations. Journal of Mathematical Physics, 46(7):073513, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i7/p073513_s1.

Guo:2005:ESW

[300] Boling Guo and Daiwen Huang. Existence of solitary waves for a simplified two fluid system of equations in plasma. Journal of Mathematical Physics, 46(7):073514, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i7/p073514_s1.

Khan:2005:PNC

[301] Abu M. Khan and Pierre Ramond. Publisher’s note: “Continuous spin representations from group contraction”, [J. Math. Phys. 46, 053515 (2005)]. Journal of Mathematical Physics, 46(7):079901, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/
p079901_s1. See [205].

DelCastillo:2005:PNP

[302] Pierre Del Castillo. Publisher’s note: “The Parr formula for the superheating field in a semi-infinite film”, [J. Math. Phys. 46, 053513 (2005)]. Journal of Mathematical Physics, 46(7):079902, July 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i7/
p079902_s1. See [203].

Brody:2005:FTS

[303] Dorje C. Brody and Lane P. Hughston. Finite-time stochastic reduction models. Journal of Mathematical Physics, 46(8):082101, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i8/p082101_s1.

Li:2005:NCN

[304] Dafa Li, Xiangrong Li, Hongtao Huang, and Xinxin Li. No-cloning of nonorthogonal states does not require inner product preserving. Journal of Mathematical Physics, 46(8):082102, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p082102_s1.

Ritter:2005:QCR

[305] William Gordon Ritter. Quantum channels and representation theory. Journal of Mathematical Physics, 46(8):082103, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/p082103_s1.

Klappenecker:2005:ASI

[306] Andreas Klappenecker, Martin Rötteler, Igor E. Shparlinski, and Arne Winterhof. On approximately symmetric informationally complete positive operator-valued measures and related systems of quantum states. Journal of Mathematical Physics, 46(8):082104, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/p082104_s1.

Vollinger:2005:GSE

[307] Stephan Völlinger. Geometry of the Schrödinger equation and stochastic mass transportation. Journal of Mathematical Physics, 46(8):082105, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i8/p082105_s1.

Khare:2005:PIP

[308] Avinash Khare and Uday Sukhatme. PT-invariant periodic potentials with a finite number of band gaps. Journal of Mathematical Physics, 46(8):082106, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i8/p082106_s1.

Synak-Radtke:2005:BLI

[309] Barbara Synak-Radtke, Karol Horodecki, and Michal Horodecki. Bounds on localizable information via semidefinite programming. Journal of Mathematical Physics, 46(8):082107, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p082107_s1.

Earp:2005:CAC

[310] Henrique N. Sá Earp and Jiannis K. Pachos. A constructive algorithm for the Cartan decomposition of SU(2^{N}). Journal of Mathematical Physics, 46(8):082108, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i8/p082108_s1.

Buscemi:2005:CPO

[311] Francesco Buscemi, Michael Keyl, Giacomo Mauro D’Ariano, Paolo Perinotti, and Reinhard F. Werner. Clean positive operator valued measures. Journal of Mathematical Physics, 46(8):082109, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/
p082109_s1.

Shin:2005:PGR

[312] Kwang C. Shin. The potential (iz)^{m} generates real eigenvalues only, under symmetric rapid decay boundary conditions. Journal of Mathematical Physics, 46(8):082110, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i8/p082110_s1.

Teh:2005:MAV

[313] Rosy Teh and Khai-Ming Wong. Monopole-antimonopole and vortex rings. Journal of Mathematical Physics, 46(8):082301, August 2005.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/
p082301_s1.

Forbes:2005:EPF

[314] Brian Forbes and Masao Jinzenji. Extending the Picard–Fuchs system of local mirror symmetry. Journal of Mathematical Physics, 46(8): 082302, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i8/p082302_s1.

Baumgaertel:2005:SPC

[315] Hellmut Baumgaertel and Hendrik Grundling. Superselection in the presence of constraints. Journal of Mathematical Physics, 46(8):082303, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i8/p082303_s1.

Bostelmann:2005:OPE

[316] Henning Bostelmann. Operator product expansions as a consequence of phase space properties. Journal of Mathematical Physics, 46(8): 082304, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i8/p082304_s1.

Walcher:2005:SLG

[317] Johannes Walcher. Stability of Landau–Ginzburg branes. Journal of Mathematical Physics, 46(8):082305, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p082305_s1.

Costa:2005:HCQ

[318] João Costa and José Natário. Homogeneous cosmologies from the quasi-Maxwell formalism. Journal of Mathematical Physics, 46(8):082501, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i8/p082501_s1.

Miritzis:2005:RPC

[319] John Miritzis. The recollapse problem of closed Friedmann–Robertson– Walker models in higher-order gravity theories. Journal of Mathematical Physics, 46(8):082502, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i8/p082502_s1.

Laine-Pearson:2005:IMW

[320] F. E. Laine-Pearson. Instabilities of multiphase wave trains in coupled nonlinear Schrödinger equations: A bisymplectic framework. Journal of Mathematical Physics, 46(8):082701, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p082701_s1.

Das:2005:SNG

[321] Ashok Das and Z. Popowicz. Supersymmetric nonlocal gas equation. Journal of Mathematical Physics, 46(8):082702, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/p082702_s1.

Nunez:2005:ETT

[322] Manuel Núñez. Existence theorems for two-fluid magnetohydrodynamics. Journal of Mathematical Physics, 46(8):083101, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/p083101_s1.

Nishiyama:2005:RMC

[323] Takahiro Nishiyama. A relaxation method for constructing a Beltrami flow in a bounded domain. Journal of Mathematical Physics, 46(8): 083102, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i8/p083102_s1.

Pule:2005:NBC

[324] Joseph V. Pulé, André F. Verbeure, and Valentin A. Zagrebnov. On nonhomogeneous Bose condensation. Journal of Mathematical Physics, 46(8):083301, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i8/p083301_s1.

Bellettini:2005:ELN

[325] G. Bellettini, A. De Masi, and E. Presutti. Energy levels of a nonlocal functional. Journal of Mathematical Physics, 46(8):083302, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i8/p083302_s1.

Chenaghlou:2005:DSG

[326] A. Chenaghlou and H. Fakhri. Dynamical symmetry group and quantum splittings for a free particle on the group manifold AdS 3. Journal of Mathematical Physics, 46(8):083501, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p083501_s1.

Baggett:2005:CPW

[327] L. W. Baggett, P. E. T. Jorgensen, K. D. Merrill, and J. A. Packer. Construction of Parseval wavelets from redundant filter systems. Journal of Mathematical Physics, 46(8):083502, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p083502_s1.

Franco:2005:NDC

[328] Daniel H. T. Franco and Caio M. M. Polito. A new derivation of the CPT and spin-statistics theorems in noncommutative field theories. Journal of Mathematical Physics, 46(8):083503, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p083503_s1.

Bu:2005:NSE

[329] Charles Bu, Kimitoshi Tsutaya, and Chenying Zhang. Nonlinear Schrödinger Equation with Inhomogeneous Dirichlet Boundary Data. Journal of Mathematical Physics, 46(8):083504, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/p083504_s1.

Kamvissis:2005:ERE

[330] Spyridon Kamvissis and Evguenii A. Rakhmanov. Existence and regularity for an energy maximization problem in two dimensions. Journal of Mathematical Physics, 46(8):083505, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p083505_s1. See comment

[2095].

Lenzi:2005:NFD

[331] E. K. Lenzi, R. S. Mendes, Kwok Sau Fa, L. S. Moraes, L. R. da Silva, and L. S. Lucena. Nonlinear fractional diffusion equation: Exact results. Journal of Mathematical Physics, 46(8):083506, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/p083506_s1.

Melgaard:2005:TPM

[332] Michael Melgaard. Threshold properties of matrix-valued Schrödinger operators. Journal of Mathematical Physics, 46(8):083507, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/
p083507_s1.

Grundland:2005:DSA

[333] A. M. Grundland and L. Snobl. Description of surfaces associated with Grassmannian sigma models on Minkowski space. Journal of Mathematical Physics, 46(8):083508, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i8/p083508_s1.

McPhedran:2005:NSL

[334] R. C. McPhedran, N. A. Nicorovici, and L. C. Botten. Neumann series and lattice sums. Journal of Mathematical Physics, 46(8):083509, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i8/p083509_s1.

Maj:2005:RBW

[335] Omar Maj. The relationship between the Wigner–Weyl kinetic formalism and the complex geometrical optics method. Journal of Mathematical Physics, 46(8):083510, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i8/p083510_s1. See publisher’s note [347].

Mendez:2005:CAG

[336] M. A. Méndez, P. Blasiak, and K. A. Penson. Combinatorial approach to generalized Bell and Stirling numbers and boson normal ordering problem. Journal of Mathematical Physics, 46(8):083511, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/
p083511_s1.

Cacciatori:2005:EAG

[337] Sergio L. Cacciatori, Bianca L. Cerchiai, Alberto Della Vedova, Giovanni Ortenzi, and Antonio Scotti. Euler angles for G_{2}. Journal of Mathematical Physics, 46(8):083512, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i8/p083512_s1.

Yan:2005:ETI

[338] Litan Yan, Ligang Lu, and Zhiqiang Xu. L^{p} estimates on a timeinhomogeneous diffusion process. Journal of Mathematical Physics, 46(8):083513, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i8/p083513_s1.

Kawamura:2005:PFO

[339] Katsunori Kawamura. The Perron–Frobenius operators, invariant measures and representations of the Cuntz–Krieger algebras. Journal of Mathematical Physics, 46(8):083514, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p083514_s1.

Humi:2005:DTS

[340] Mayer Humi. Darboux transformations for Schrödinger equations in two variables. Journal of Mathematical Physics, 46(8):083515, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i8/p083515_s1.

Nagy:2005:CQT

[341] Zoltán Nagy, Jean Avan, Anastasia Doikou, and Geneviève Rollet. Commuting quantum traces for quadratic algebras. Journal of Mathematical Physics, 46(8):083516, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i8/p083516_s1.

Gao:2005:GSS

[342] Hongjun Gao and Anyou Zhu. The global strong solutions of Hasegawa– Mima–Charney–Obukhov equation. Journal of Mathematical Physics, 46(8):083517, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i8/p083517_s1.

Blaauboer:2005:AST

[343] M. Blaauboer. Analytical solution of the time evolution of an entangled electron spin pair in a double quantum dot nanostructure. Journal of Mathematical Physics, 46(8):083518, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i8/p083518_s1.

Li:2005:SGW

[344] Fang Li and Haijun Cao. Semilattice graded weak Hopf algebra and its related quantum G-double. Journal of Mathematical Physics, 46(8): 083519, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i8/p083519_s1.

Cacciatori:2005:SPG

[345] Sergio L. Cacciatori. A simple parametrization for G_{2}. Journal of Mathematical Physics, 46(8):083520, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i8/p083520_s1.

Benito:2005:DVM

[346] Roberto Benito and David Martín de Diego. Discrete vakonomic mechanics. Journal of Mathematical Physics, 46(8):083521, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i8/
p083521_s1.

Maj:2005:PNR

[347] Omar Maj. Publisher’s note: “The relationship between the Wigner– Weyl kinetic formalism and the complex geometrical optics method”, [J. Math. Phys. 46, 083510 (2005)]. Journal of Mathematical Physics, 46(8): 089901, August 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i8/p089901_s1. See [335].

Rubinstein:2005:ISG

[348] Jacob Rubinstein and Peter Sternberg. Introduction: Superconductivity and the Ginzburg–Landau model. Journal of Mathematical Physics, 46(9):095101, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i9/p095101_s1.

Alama:2005:PET

[349] Stan Alama and Lia Bronsard. Pinning effects and their breakdown for a Ginzburg–Landau model with normal inclusions. Journal of Mathematical Physics, 46(9):095102, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i9/p095102_s1.

Almog:2005:ETL

[350] Y. Almog. Emergence of the triangular lattice near a flat wall. Journal of Mathematical Physics, 46(9):095103, September 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i9/p095103_s1.

Bauman:2005:CST

[351] Patricia Bauman, Hala Jadallah, and Daniel Phillips. Classical solutions to the time-dependent Ginzburg–Landau equations for a bounded superconducting body in a vacuum. Journal of Mathematical Physics, 46(9):095104, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i9/p095104_s1.

Berdiyorov:2005:SBF

[352] G. R. Berdiyorov, L. R. E. Cabral, and F. M. Peeters. Surface barrier for flux entry and exit in mesoscopic superconducting systems. Journal of Mathematical Physics, 46(9):095105, September 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i9/p095105_s1.

Berger:2005:TDG

[353] Jorge Berger. Time-dependent Ginzburg–Landau equations with charged boundaries. Journal of Mathematical Physics, 46(9):095106, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i9/p095106_s1.

Berlyand:2005:GLM

[354] L. Berlyand and E. Khruslov. Ginzburg–Landau model of a liquid crystal with random inclusions. Journal of Mathematical Physics, 46(9):095107, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i9/p095107_s1.

Chibotaru:2005:GLD

[355] L. F. Chibotaru, A. Ceulemans, M. Morelle, G. Teniers, C. Carballeira, and V. V. Moshchalkov. Ginzburg–Landau description of confinement and quantization effects in mesoscopic superconductors. Journal of Mathematical Physics, 46(9):095108, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i9/p095108_s1.

Du:2005:NAG

[356] Qiang Du. Numerical approximations of the Ginzburg–Landau models for superconductivity. Journal of Mathematical Physics, 46(9):095109, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i9/p095109_s1.

Kim:2005:TDL

[357] Yun hui Kim, Jacob Rubinstein, and Peter Sternberg. Topologically driven local minimizers of the Oseen–Frank energy. Journal of Mathematical Physics, 46(9):095110, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i9/p095110_s1.

Kosugi:2005:GBS

[358] Satoshi Kosugi, Yoshihisa Morita, and Shoji Yotsutani. Global bifurcation structure of a one-dimensional Ginzburg–Landau model. Journal of Mathematical Physics, 46(9):095111, September 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i9/p095111_s1.

Ma:2005:BSS

[359] Tian Ma and Shouhong Wang. Bifurcation and stability of superconductivity. Journal of Mathematical Physics, 46(9):095112, September 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i9/p095112_s1.

Nagata:2005:IET

[360] Koji Nagata. Inequalities for experimental tests of the Kochen–Specker theorem. Journal of Mathematical Physics, 46(10):102101, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p102101_s1.

Novaes:2005:RTS

[361] Marcel Novaes. Real trajectories in the semiclassical coherent state propagator. Journal of Mathematical Physics, 46(10):102102, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p102102_s1.

Fichtner:2005:TSI

[362] Karl-Heinz Fichtner, Wolfgang Freudenberg, and Masanori Ohya. Teleportation schemes in infinite dimensional Hilbert spaces. Journal of Mathematical Physics, 46(10):102103, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i10/p102103_s1.

Audenaert:2005:CBQ

[363] Koenraad M. R. Audenaert and Jens Eisert. Continuity bounds on the quantum relative entropy. Journal of Mathematical Physics, 46(10): 102104, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p102104_s1.

Lang:2005:FSC

[364] Andrew S. I. D. Lang. The frequency spectrum of the Casimir effect. Journal of Mathematical Physics, 46(10):102105, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p102105_s1.

Lyons:2005:MOD

[365] David W. Lyons and Scott N. Walck. Minimum orbit dimension for local unitary action on n-qubit pure states. Journal of Mathematical Physics, 46(10):102106, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i10/p102106_s1.

Sasaki:2005:GSM

[366] Itaru Sasaki. Ground state of the massless Nelson model in a non-Fock representation. Journal of Mathematical Physics, 46(10):102107, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p102107_s1.

Mostafazadeh:2005:APH

[367] Ali Mostafazadeh. Application of pseudo-Hermitian quantum mechanics to a P T -symmetric Hamiltonian with a continuum of scattering states. Journal of Mathematical Physics, 46(10):102108, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p102108_s1.

Strauss:2005:SDT

[368] Y. Strauss. On the semigroup decomposition of the time evolution of quantum mechanical resonances. Journal of Mathematical Physics, 46(10):102109, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i10/p102109_s1.

Martinez:2005:GFC

[369] Jorge Martínez, Claudio Meneses, and José A. Zapata. Geometry of C-flat connections, coarse graining and the continuum limit. Journal of Mathematical Physics, 46(10):102301, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i10/p102301_s1.

Salmela:2005:FGA

[370] Antti Salmela. Function group approach to unconstrained Hamiltonian Yang–Mills theory. Journal of Mathematical Physics, 46(10):102302, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p102302_s1.

Fuster:2005:NBQ

[371] A. Fuster and J. W. van Holten. A note on BRST quantization of SU(2) Yang–Mills mechanics. Journal of Mathematical Physics, 46(10): 102303, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p102303_s1.

Clark:2005:XXX

[372] T. E. Clark, S. T. Love, Muneto Nitta, and T. ter Veldhuis. AdS_{d+1} → AdS_{d}. Journal of Mathematical Physics, 46(10):102304, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p102304_s1.

Bertini:2005:RDR

[373] Massimo Bertini, Diego Noja, and Andrea Posilicano. Rigorous dynamics and radiation theory for a Pauli–Fierz model in the ultraviolet limit. Journal of Mathematical Physics, 46(10):102305, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p102305_s1.

Wospakrik:2005:SMY

[374] H. J. Wospakrik. Solutions of the massive SU(N) Yang–Mills equations by harmonic maps. Journal of Mathematical Physics, 46(10):102306, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p102306_s1.

Sasaki:2005:GSE

[375] Itaru Sasaki. Ground state energy of the polaron in the relativistic quantum electrodynamics. Journal of Mathematical Physics, 46(10): 102307, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p102307_s1.

Ringermacher:2005:IMC

[376] Harry I. Ringermacher and Lawrence R. Mead. Induced matter: Curved N-manifolds encapsulated in Riemann-flat N + 1 dimensional space. Journal of Mathematical Physics, 46(10):102501, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p102501_s1.

Garat:2005:TG

[377] Alcides Garat. Tetrads in geometrodynamics. Journal of Mathematical Physics, 46(10):102502, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i10/p102502_s1. See [2212].

Lau:2005:ASR

[378] Stephen R. Lau. Analytic structure of radiation boundary kernels for blackhole perturbations. Journal of Mathematical Physics, 46(10): 102503, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p102503_s1.

Bicak:2005:FAS

[379] Jirí Bicák and Pavel Krtous. Fields of accelerated sources: Born in de Sitter. Journal of Mathematical Physics, 46(10):102504, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p102504_s1.

Antonio:2005:GMJ

[380] N. Cirilo António and N. Manojlović. ∫↕_{∈} Gaudin model with jordanian twist. Journal of Mathematical Physics, 46(10):102701, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p102701_s1.

Gentile:2005:BLE

[381] Guido Gentile and Titus S. van Erp. Breakdown of Lindstedt expansion for chaotic maps. Journal of Mathematical Physics, 46(10):102702, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p102702_s1.

Haas:2005:JSM

[382] F. Haas. Jacobi structures in R^{3}. Journal of Mathematical Physics, 46(10):102703, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i10/p102703_s1.

Santos:2005:WMD

[383] J. P. Santos and L. O. Silva. Wigner–Moyal description of free variable mass Klein–Gordon fields. Journal of Mathematical Physics, 46(10): 102901, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p102901_s1.

Bhand:2005:RBM

[384] Ajit Bhand and Andrew D. Lewis. Rigid body mechanics in Galilean spacetimes. Journal of Mathematical Physics, 46(10):102902, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p102902_s1.

Garoni:2005:ACE

[385] T. M. Garoni, P. J. Forrester, and N. E. Frankel. Asymptotic corrections to the eigenvalue density of the GUE and LUE. Journal of Mathematical Physics, 46(10):103301, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i10/p103301_s1.

Ali:2005:FTD

[386] Abdulmuhsen H. Ali. A field theoretic derivation of the Max Born hypothesis. Journal of Mathematical Physics, 46(10):103302, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p103302_s1.

Rigotti:2005:TGR

[387] M. Rigotti and F. Debbasch. An H-theorem for the general relativistic Ornstein–Uhlenbeck process. Journal of Mathematical Physics, 46(10): 103303, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p103303_s1.

Scoufis:2005:AIS

[388] George Scoufis and Christopher M. Cosgrove. An application of the inverse scattering transform to the modified intermediate long wave equation. Journal of Mathematical Physics, 46(10):103501, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p103501_s1.

Borowiec:2005:BTQ

[389] Andrzej Borowiec, Jerzy Lukierski, Vladimir Lyakhovsky, Marek Mozrzymas, and Valeriy N. Tolstoy. Basic twist quantization of the exceptional Lie algebra g_{2}. Journal of Mathematical Physics, 46(10): 103502, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p103502_s1.

McCaw:2005:CSC

[390] James McCaw and B. H. J. McKellar. On the continuous spectral component of the Floquet operator for a periodically kicked quantum system. Journal of Mathematical Physics, 46(10):103503, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p103503_s1.

Poletaeva:2005:ESA

[391] Elena Poletaeva. Exceptional N = 6 superconformal algebra. Journal of Mathematical Physics, 46(10):103504, October 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i10/p103504_s1. See

publisher’s note [501].

Nunez:2005:QTC

[392] J. Fernández Núñez, W. García Fuertes, and A. M. Perelomov. Quantum trigonometric Calogero–Sutherland model, irreducible characters and Clebsch–Gordan series for the exceptional algebra E_{7}. Journal of Mathematical Physics, 46(10):103505, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i10/p103505_s1.

Messirdi:2005:RPM

[393] B. Messirdi, A. Senoussaoui, and G. Djellouli. Resonances of polyatomic molecules in the Born–Oppenheimer approximation. Journal of Mathematical Physics, 46(10):103506, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i10/p103506_s1.

Kalnins:2005:SOSc

[394] E. G. Kalnins, J. M. Kress, and W. Miller. Second order superintegrable systems in conformally flat spaces. III. Three-dimensional classical structure theory. Journal of Mathematical Physics, 46(10):103507, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p103507_s1.

Jarlskog:2005:RPU

[395] C. Jarlskog. A recursive parametrization of unitary matrices. Journal of Mathematical Physics, 46(10):103508, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i10/p103508_s1.

Gordoa:2005:NEV

[396] P. R. Gordoa, A. Pickering, and Z. N. Zhu. A nonisospectral extension of the Volterra hierarchy to 2 + 1 dimensions. Journal of Mathematical Physics, 46(10):103509, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i10/p103509_s1.

Aizawa:2005:QS

[397] N. Aizawa and R. Chakrabarti. Quantum spheres for OSp_{q}(1/2). Journal of Mathematical Physics, 46(10):103510, October 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i10/p103510_s1.

Kwas:2005:OMC

[398] Marek Kwas. An optimal Monte Carlo algorithm for multivariate Feynman–Kac path integrals. Journal of Mathematical Physics, 46(10): 103511, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i10/p103511_s1.

Trzetrzelewski:2005:RLT

[399] Maciej Trzetrzelewski. Reduction of su(N) loop tensors to trees. Journal of Mathematical Physics, 46(10):103512, October 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i10/p103512_s1.

Bashkirov:2005:AKT

[400] D. Bashkirov, G. Giachetta, L. Mangiarotti, and G. Sardanashvily. The antifield Koszul–Tate complex of reducible Noether identities. Journal of Mathematical Physics, 46(10):103513, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i10/p103513_s1.

Li:2005:RHO

[401] Zhi-Feng Li, Jin-Jin Liu, Wolfgang Lucha, Wen-Gan Ma, and Franz F. Schöberl. Relativistic harmonic oscillator. Journal of Mathematical Physics, 46(10):103514, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i10/p103514_s1.

Calogero:2005:NSE

[402] F. Calogero and S. Iona. Novel solvable extensions of the goldfish manybody model. Journal of Mathematical Physics, 46(10):103515, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p103515_s1.

Lizzi:2005:ACS

[403] Fedele Lizzi, Allen Stern, and Patrizia Vitale. On the absence of continuous symmetries for noncommutative 3-spheres. Journal of Mathematical Physics, 46(10):103516, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i10/p103516_s1.

Andrea:2005:GCN

[404] S. Andrea, A. Restuccia, and A. Sotomayor. The Gardner category and nonlocal conservation laws for N = 1 Super KdV. Journal of Mathematical Physics, 46(10):103517, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i10/p103517_s1.

Kubieniec:2005:SSS

[405] J. Kubieniec. On the structure of the supplementary series of unitary, irreducible representations of the proper, orthochronous Lorentz group. Journal of Mathematical Physics, 46(10):103518, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i10/
p103518_s1.

Majid:2005:GTN

[406] S. Majid. Gauge theory on nonassociative spaces. Journal of Mathematical Physics, 46(10):103519, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i10/p103519_s1.

Majid:2005:NMS

[407] S. Majid. Noncommutative model with spontaneous time generation and Planckian bound. Journal of Mathematical Physics, 46(10):103520, October 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i10/p103520_s1.

Lei:2005:ACR

[408] C. Lei, A. Vourdas, and A. Wünsche. Analytic and contour representations in the unit disk based on SU(1, 1) coherent states. Journal of Mathematical Physics, 46(11):112101, November 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i11/p112101_s1.

Lassen:2005:GTD

[409] B. Lassen, M. Willatzen, R. Melnik, and L. C. Lew Yan Voon. A general treatment of deformation effects in Hamiltonians for inhomogeneous crystalline materials. Journal of Mathematical Physics, 46(11):112102, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p112102_s1.

Chae:2005:RRW

[410] Myeongju Chae, Seung-Yeal Ha, and Hyungjin Huh. Remarks on the repulsive Wigner–Poisson system. Journal of Mathematical Physics, 46(11):112103, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p112103_s1.

Tumulka:2005:SWF

[411] Roderich Tumulka and Nino Zanghì. Smoothness of wave functions in thermal equilibrium. Journal of Mathematical Physics, 46(11):112104, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p112104_s1.

Brandt:2005:LEV

[412] Sebastian F. Brandt and Axel Pelster. Large-D expansion from variational perturbation theory. Journal of Mathematical Physics, 46 (11):112105, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p112105_s1.

Wang:2005:ETC

[413] Xue Ping Wang and Yuefei Wang. Existence of two-cluster threshold resonances and the N-body Efimov effect. Journal of Mathematical Physics, 46(11):112106, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p112106_s1.

Kan:2005:DQE

[414] Nahomi Kan and Kiyoshi Shiraishi. Divergences in quantum electrodynamics on a graph. Journal of Mathematical Physics, 46(11):112301, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p112301_s1.

Nita:2005:PGR

[415] Bogdan G. Nita, Paul MacAlevey, and Phillip T. Downes. Pure gravitational radiation with twisting rays in the linear approximation. Journal of Mathematical Physics, 46(11):112501, November 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i11/p112501_s1.

Mosna:2005:VET

[416] Ricardo A. Mosna and Alberto Saa. Volume elements and torsion. Journal of Mathematical Physics, 46(11):112502, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i11/
p112502_s1.

Armenta:2005:SBB

[417] J. Armenta and J. A. Nieto. Soldered bundle background for the de Sitter top. Journal of Mathematical Physics, 46(11):112503, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i11/p112503_s1.

Harnad:2005:CRT

[418] J. Harnad, I. Loutsenko, and O. Yermolayeva. Constrained reductions of two-dimensional dispersionless Toda hierarchy, Hamiltonian structure, and interface dynamics. Journal of Mathematical Physics, 46(11):112701, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p112701_s1.

Cimmelli:2005:ENE

[419] V. A. Cimmelli and P. Ván. The effects of nonlocality on the evolution of higher order fluxes in nonequilibrium thermodynamics. Journal of Mathematical Physics, 46(11):112901, November 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i11/p112901_s1.

Muzsnay:2005:IVP

[420] Zoltán Muzsnay. An invariant variational principle for canonical flows on Lie groups. Journal of Mathematical Physics, 46(11):112902, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i11/p112902_s1.

Forger:2005:HMF

[421] Michael Forger, Cornelius Paufler, and Hartmann Römer. Hamiltonian multivector fields and Poisson forms in multisymplectic field theory. Journal of Mathematical Physics, 46(11):112903, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i11/
p112903_s1.

Pasquero:2005:ICU

[422] Stefano Pasquero. Ideality criterion for unilateral constraints in timedependent impulsive mechanics. Journal of Mathematical Physics, 46 (11):112904, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p112904_s1.

Barros:2005:GLH

[423] Manuel Barros, Alfonso Romero, José L. Cabrerizo, and Manuel Fernández. The Gauss–Landau–Hall problem on Riemannian surfaces. Journal of Mathematical Physics, 46(11):112905, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i11/
p112905_s1.

Bisi:2005:FDE

[424] M. Bisi and G. Spiga. Fluid-dynamic equations for granular particles in a host medium. Journal of Mathematical Physics, 46(11):113301, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p113301_s1. See erratum [2206].

DeSanctis:2005:RMO

[425] Luca De Sanctis. Random multi-overlap structures for optimization problems. Journal of Mathematical Physics, 46(11):113302, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i11/
p113302_s1.

Otway:2005:HFE

[426] Thomas H. Otway. Harmonic fields on the extended projective disk and a problem in optics. Journal of Mathematical Physics, 46(11):113501, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p113501_s1. See erratum [1117].

Kodama:2005:IQD

[427] Y. Kodama, B. Konopelchenko, L. Martínez Alonso, and E. Medina. Integrable quasiclassical deformations of cubic curves. Journal of Mathematical Physics, 46(11):113502, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i11/p113502_s1.

Figueiredo:2005:ILC

[428] B. D. Bonorino Figueiredo. Ince’s limits for confluent and doubleconfluent Heun equations. Journal of Mathematical Physics, 46 (11):113503, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p113503_s1.

Stoilova:2005:CGQb

[429] N. I. Stoilova and J. Van der Jeugt. A classification of generalized quantum statistics associated with basic classical Lie superalgebras. Journal of Mathematical Physics, 46(11):113504, November 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i11/p113504_s1.

Abe:2005:NRC

[430] Yuuki Abe and Toshiki Nakashima. Nilpotent representations of classical quantum groups at roots of unity. Journal of Mathematical Physics, 46(11):113505, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p113505_s1.

Patera:2005:DCCb

[431] J. Patera and A. Zaratsyan. Discrete and continuous cosine transform generalized to Lie groups SU(3) and G(2). Journal of Mathematical Physics, 46(11):113506, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i11/p113506_s1.

Park:2005:EPM

[432] Yong Moon Park. Ergodic property of Markovian semigroups on standard forms of von Neumann algebras. Journal of Mathematical Physics, 46(11):113507, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p113507_s1.

Bruning:2005:DSG

[433] Jochen Brüning, Vladimir Geyler, and Konstantin Pankrashkin. Ondiagonal singularities of the Green functions for Schrödinger operators. Journal of Mathematical Physics, 46(11):113508, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i11/
p113508_s1.

Musso:2005:CPI

[434] E. Musso and L. Nicolodi. On the Cauchy problem for the integrable system of Lie minimal surfaces. Journal of Mathematical Physics, 46 (11):113509, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p113509_s1.

Gurses:2005:IET

[435] Metin Gürses, Gusein Sh. Guseinov, and Burcu Silindir. Integrable equations on time scales. Journal of Mathematical Physics, 46(11):113510, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p113510_s1.

Morse:2005:ECC

[436] Edward C. Morse. Eigenfunctions of the curl in cylindrical geometry. Journal of Mathematical Physics, 46(11):113511, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i11/
p113511_s1.

Pierantozzi:2005:IBW

[437] T. Pierantozzi and L. Vázquez. An interpolation between the wave and diffusion equations through the fractional evolution equations Dirac like. Journal of Mathematical Physics, 46(11):113512, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i11/
p113512_s1.

Neidhardt:2005:UDS

[438] H. Neidhardt and J. Rehberg. Uniqueness for Dissipative Schrödinger– Poisson Systems. Journal of Mathematical Physics, 46(11):113513, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p113513_s1.

Hoshino:2005:PRC

[439] Ayumu Hoshino. Polyhedral realizations of crystal bases for quantum algebras of finite types. Journal of Mathematical Physics, 46(11):113514, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i11/p113514_s1.

Shchepetilov:2005:CCP

[440] A. V. Shchepetilov. Comment on “Central potentials on spaces of constant curvature: The Kepler problem on the two-dimensional sphere S^{2} and the hyperbolic plane H^{2}”, [J. Math. Phys. 46, 052702 (2005)]. Journal of Mathematical Physics, 46(11):114101, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i11/p114101_s1. See [180, 441].

Carinena:2005:RCC

[441] J. F. Cariñena, M. F. Ranada, and M. Santander. Response to “Comment on ‘Central potentials on spaces of constant curvature: The Kepler problem on the two-dimensional sphere S^{2} and the hyperbolic plane H^{2}’ [J. Math. Phys. 46, 052702 (2005)]”. Journal of Mathematical Physics, 46(11):114102, November 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i11/p114102_s1. See [440, 180].

Chen:2005:CMB

[442] Hao Chen. Constraints on the mixing of bipartite states. Journal of Mathematical Physics, 46(12):122101, December 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i12/p122101_s1.

Hussin:2005:LOC

[443] V. Hussin and L. M. Nieto. Ladder operators and coherent states for the Jaynes–Cummings model in the rotating-wave approximation. Journal of Mathematical Physics, 46(12):122102, December 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i12/p122102_s1.

Barbaroux:2005:QEE

[444] Jean-Marie Barbaroux, Helmut Linde, and Semjon Vugalter. Quantitative estimates on the enhanced binding for the Pauli–Fierz operator. Journal of Mathematical Physics, 46(12):122103, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i12/
p122103_s1.

Garcia:2005:ACM

[445] Jaime Zaratiegui García. Analytic Coulomb matrix elements in a threedimensional geometry. Journal of Mathematical Physics, 46(12):122104, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i12/p122104_s1.

Singh:2005:CAS

[446] Sudhir Kumar Singh, Sudebkumar Prasant Pal, Somesh Kumar, and R. Srikanth. A combinatorial approach for studying local operations and classical communication transformations of multipartite states. Journal of Mathematical Physics, 46(12):122105, December 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i12/p122105_s1.

Teismann:2005:GCS

[447] Holger Teismann. Generalized coherent states and the control of quantum systems. Journal of Mathematical Physics, 46(12):122106, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i12/p122106_s1.

Janzing:2005:DTC

[448] Dominik Janzing. Decomposition of time-covariant operations on quantum systems with continuous and/or discrete energy spectrum. Journal of Mathematical Physics, 46(12):122107, December 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i12/p122107_s1.

Konig:2005:FRF

[449] Robert König and Renato Renner. A de Finetti representation for finite symmetric quantum states. Journal of Mathematical Physics, 46 (12):122108, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i12/p122108_s1.

Si:2005:NAC

[450] Tieyan Si. Non-Abelian Chern–Simons action is topological invariant on 3 simple knot. Journal of Mathematical Physics, 46(12):122301, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i12/p122301_s1.

Naot:2005:CST

[451] Gad Naot. On Chern–Simons theory with an inhomogeneous gauge group and BF theory knot invariants. Journal of Mathematical Physics, 46(12):122302, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i12/p122302_s1.

Heller:2005:NUG

[452] Michael Heller, Leszek Pysiak, and Wieslaw Sasin. Noncommutative unification of general relativity and quantum mechanics. Journal of Mathematical Physics, 46(12):122501, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i12/p122501_s1.

Christensen:2005:CSQ

[453] J. Daniel Christensen and Louis Crane. Causal sites as quantum geometry. Journal of Mathematical Physics, 46(12):122502, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i12/
p122502_s1.

Droste:2005:UHC

[454] Manfred Droste. Universal homogeneous causal sets. Journal of Mathematical Physics, 46(12):122503, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i12/p122503_s1.

Munoz-Lecanda:2005:NCC

[455] Miguel-C. Muñoz-Lecanda, Modesto Salgado, and Silvia Vilariño. Nonstandard connections in k-cosympletic field theory. Journal of Mathematical Physics, 46(12):122901, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v46/i12/p122901_s1.

Farkas:2005:SZE

[456] S. Farkas and Z. Zimborás. On the sharpness of the zero-entropy-density conjecture. Journal of Mathematical Physics, 46(12):123301, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i12/p123301_s1.

vanEerten:2005:EGS

[457] Hendrik van Eerten. Extensive ground state entropy in supersymmetric lattice models. Journal of Mathematical Physics, 46(12):123302, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i12/p123302_s1.

Assis:2005:NDE

[458] P. C. Assis, L. R. da Silva, E. K. Lenzi, L. C. Malacarne, and R. S. Mendes. Nonlinear diffusion equation, Tsallis formalism and exact solutions. Journal of Mathematical Physics, 46(12):123303, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v46/i12/
p123303_s1.

Dancer:2005:ECI

[459] K. A. Dancer, M. D. Gould, and J. Links. Eigenvalues of Casimir invariants for U_{q}[osp(m|n)]. Journal of Mathematical Physics, 46(12):123501, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i12/p123501_s1.

Chen:2005:MNB

[460] Jianqing Chen and Boling Guo. Multiple nodal bound states for a quasilinear Schrödinger equation. Journal of Mathematical Physics, 46(12):123502, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i12/p123502_s1.

Masson:2005:INC

[461] Thierry Masson and Emmanuel Serié. Invariant noncommutative connections. Journal of Mathematical Physics, 46(12):123503, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i12/p123503_s1.

Yomba:2005:CNS

[462] Emmanuel Yomba. Construction of new solutions to the fully nonlinear generalized Camassa–Holm equations by an indirect F function method. Journal of Mathematical Physics, 46(12):123504, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v46/i12/p123504_s1.

Rybkin:2005:SCL

[463] Alexei Rybkin. On the spectral L_{2} conjecture, 3/2-Lieb–Thirring inequality and distributional potentials. Journal of Mathematical Physics, 46(12):123505, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v46/i12/p123505_s1.

Bluman:2005:FPS

[464] George Bluman and Alexei F. Cheviakov. Framework for potential systems and nonlocal symmetries: Algorithmic approach. Journal of Mathematical Physics, 46(12):123506, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i12/p123506_s1.

Brunelli:2005:SPH

[465] J. C. Brunelli. The short pulse hierarchy. Journal of Mathematical Physics, 46(12):123507, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v46/i12/p123507_s1.

Rozenblyum:2005:RDA

[466] Alexander Rozenblyum. Representations of the q-deformed algebras U_{q}(so_{3}) and U_{q}(so_{5}) and q-orthogonal polynomials. Journal of Mathematical Physics, 46(12):123508, December 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v46/i12/p123508_s1.

Khuri:2005:AUL

[467] N. N. Khuri, André Martin, Pierre C. Sabatier, and Tai Tsun Wu. Addendum: “Universality of low-energy scattering in 2+1 dimensions: The nonsymmetric case”, [J. Math. Phys. 46, 032103 (2005)]. Journal of Mathematical Physics, 46(12):129901, December 2005. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v46/i12/p129901_s1. See

[64].

Altschul:2006:NHI

[468] B. Altschul. Non-Hermitian interactions between harmonic oscillators, with applications to stable, Lorentz-violating quantum electrodynamics. Journal of Mathematical Physics, 47(1):012101, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i1/p012101_s1.

Jacobs:2006:BMI

[469] Kurt Jacobs. A bound on the mutual information, and properties of entropy reduction, for quantum channels with inefficient measurements.

Journal of Mathematical Physics, 47(1):012102, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i1/p012102_s1.

Heydari:2006:EMM

[470] Hoshang Heydari. Entanglement monotones for multi-qubit states based on geometric invariant theory. Journal of Mathematical Physics, 47(1): 012103, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i1/p012103_s1.

Yonezawa:2006:IQC

[471] Nobuhiro Yonezawa and Izumi Tsutsui. Inequivalent quantizations of the N = 3 Calogero model with scale and mirror-S_{3} symmetry. Journal of Mathematical Physics, 47(1):012104, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i1/p012104_s1.

Ananos:2006:SFT

[472] G. N. J. Añaños. Scalar field theory at finite temperature in D = 2 + 1. Journal of Mathematical Physics, 47(1):012301, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i1/p012301_s1.

deMello:2006:ESG

[473] E. R. Bezerra de Mello. The Euclidean scalar Green function in the fivedimensional Kaluza–Klein magnetic monopole space–time. Journal of Mathematical Physics, 47(1):012302, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i1/p012302_s1.

Sako:2006:PFS

[474] Akifumi Sako and Toshiya Suzuki. Partition functions of supersymmetric gauge theories in noncommutative R^{2D} and their unified perspective. Journal of Mathematical Physics, 47(1):012303, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i1/p012303_s1.

Cook:2006:GWZ

[475] James S. Cook. Gauged Wess–Zumino model in noncommutative Minkowski superspace. Journal of Mathematical Physics, 47(1):012304, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i1/p012304_s1.

Leal:2006:MML

[476] Lorenzo Leal and Alexander López. Magnetic monopole in the loop representation. Journal of Mathematical Physics, 47(1):012305, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i1/p012305_s1.

Popov:2006:QGT

[477] Alexander D. Popov and Richard J. Szabo. Quiver gauge theory of non-Abelian vortices and noncommutative instantons in higher dimensions. Journal of Mathematical Physics, 47(1):012306, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i1/p012306_s1.

Bayin:2006:FBF

[478] S. Bayin, E. N. Glass, and J. P. Krisch. Fractional boundaries for fluid spheres. Journal of Mathematical Physics, 47(1):012501, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i1/p012501_s1.

Halilsoy:2006:CWS

[479] E. Halilsoy, M. Halilsoy, and O. Unver. Colliding wave solutions from five-dimensional black holes and black p-branes. Journal of Mathematical Physics, 47(1):012502, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i1/p012502_s1.

Sasaki:2006:ESC

[480] R. Sasaki and K. Takasaki. Explicit solutions of the classical Calogero and Sutherland systems for any root system. Journal of Mathematical Physics, 47(1):012701, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i1/p012701_s1.

Gallavotti:2006:FLS

[481] G. Gallavotti, G. Gentile, and A. Giuliani. Fractional Lindstedt series. Journal of Mathematical Physics, 47(1):012702, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i1/p012702_s1.

Sundkvist:2006:CIP

[482] David Sundkvist. Covariant irreducible parametrization of electromagnetic fields in arbitrary space–time. Journal of Mathematical Physics, 47(1):012901, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i1/p012901_s1. See publisher’s note [629].

Tip:2006:SMP

[483] A. Tip. Some mathematical properties of Maxwell’s equations for macroscopic dielectrics. Journal of Mathematical Physics, 47(1):012902, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i1/p012902_s1.

Rossani:2006:AEG

[484] A. Rossani and G. Spiga. Auger effect in the generalized kinetic theory of electrons and holes. Journal of Mathematical Physics, 47(1): 013301, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i1/p013301_s1.

Zhao:2006:DRC

[485] Shao-You Zhao, Wen-Li Yang, and Yao-Zhong Zhang. Determinant representation of correlation functions for the U_{q}(gl(1|1)) free Fermion model. Journal of Mathematical Physics, 47(1):013302, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i1/
p013302_s1.

Dridi:2006:EPF

[486] Raouf Dridi and Sylvain Neut. The equivalence problem for fourth order differential equations under fiber preserving diffeomorphisms. Journal of Mathematical Physics, 47(1):013501, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i1/p013501_s1.

Gromov:2006:PQK

[487] N. A. Gromov and V. V. Kuratov. Possible quantum kinematics. Journal of Mathematical Physics, 47(1):013502, January 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i1/p013502_s1.

Kong:2006:ETL

[488] De-Xing Kong, Qing-You Sun, and Yi Zhou. The equation for time-like extremal surfaces in Minkowski space R^{2+n}. Journal of Mathematical Physics, 47(1):013503, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i1/p013503_s1.

Wu:2006:BBR

[489] Shuyin Wu and Zhaoyang Yin. Blow-up, blow-up rate and decay of the solution of the weakly dissipative Camassa–Holm equation. Journal of Mathematical Physics, 47(1):013504, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i1/p013504_s1.

Boumenir:2006:ATK

[490] Amin Boumenir. The approximation of the transmutation kernel. Journal of Mathematical Physics, 47(1):013505, January 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i1/p013505_s1.

Odesskii:2006:CLB

[491] A. V. Odesskii and V. V. Sokolov. Compatible Lie brackets related to elliptic curve. Journal of Mathematical Physics, 47(1):013506, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i1/p013506_s1.

Jarlskog:2006:RPI

[492] C. Jarlskog. Recursive parametrization and invariant phases of unitary matrices. Journal of Mathematical Physics, 47(1):013507, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i1/p013507_s1.

Karabulut:2006:DGP

[493] Hasan Karabulut. Distributed Gaussian polynomials as q-oscillator eigenfunctions. Journal of Mathematical Physics, 47(1):013508, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i1/p013508_s1.

Aragon-Gonzalez:2006:DOT

[494] G. Aragón-González, J. L. Aragón, and M. A. Rodríguez-Andrade. The decomposition of an orthogonal transformation as a product of reflections. Journal of Mathematical Physics, 47(1):013509, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i1/
p013509_s1.

Gungor:2006:GDS

[495] F. Güngör and Ö. Aykanat. The generalized Davey–Stewartson equations, its Kac–Moody–Virasoro symmetry algebra and relation to Davey– Stewartson equations. Journal of Mathematical Physics, 47(1):013510, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i1/p013510_s1.

Cornean:2006:FER

[496] Horia D. Cornean, Gheorghe Nenciu, and Thomas G. Pedersen. The Faraday effect revisited: General theory. Journal of Mathematical Physics, 47(1):013511, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i1/p013511_s1.

Pelantova:2006:FGG

[497] Edita Pelantová, Milena Svobodová, and Sébastien Tremblay. Fine grading of sl(p^{2}, C) generated by tensor product of generalized Pauli matrices and its symmetries. Journal of Mathematical Physics, 47(1):013512, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i1/p013512_s1.

Forte:2006:SNE

[498] L. A. Forte and A. Sciarrino. Standard and nonstandard extensions of Lie algebras. Journal of Mathematical Physics, 47(1):013513, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i1/p013513_s1.

Mosna:2006:SSS

[499] Ricardo A. Mosna. Singular solutions to the Seiberg–Witten and Freund equations on flat space from an iterative method. Journal of Mathematical Physics, 47(1):013514, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i1/p013514_s1.

Najar:2006:SMR

[500] Hatem Najar. The spectrum minimum for random Schrödinger operators with indefinite sign potentials. Journal of Mathematical Physics, 47(1): 013515, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i1/p013515_s1.

Poletaeva:2006:PNE

[501] Elena Poletaeva. Publisher’s note: “On the exceptional N = 6 superconformal algebra”, [J. Math. Phys. 46, 103504 (2005)]. Journal of Mathematical Physics, 47(1):019901, January 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i1/p019901_s1. See [391].

Langmann:2006:RIR

[502] Edwin Langmann. Remarkable identities related to the (quantum) elliptic Calogero–Sutherland model. Journal of Mathematical Physics, 47(2):022101, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i2/p022101_s1.

DAriano:2006:OEQ

[503] Giacomo Mauro D’Ariano, Vittorio Giovannetti, and Paolo Perinotti. Optimal estimation of quantum observables. Journal of Mathematical Physics, 47(2):022102, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i2/p022102_s1.

Kurochkin:2006:RTC

[504] Yu. Kurochkin and Dz. Shoukavy. Regge trajectories of the Coulomb potential in the space of constant negative curvature ^{1}S_{3}. Journal of Mathematical Physics, 47(2):022103, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i2/p022103_s1.

Slater:2006:QFI

[505] Paul B. Slater. Quantum and Fisher information from the Husimi and related distributions. Journal of Mathematical Physics, 47(2):022104, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i2/p022104_s1.

Bergeron:2006:RBK

[506] H. Bergeron. Rigorous bra-ket formalism and wave function operator for one particle quantum mechanics. Journal of Mathematical Physics, 47(2):022105, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i2/p022105_s1.

Zamastil:2006:LOB

[507] J. Zamastil and L. Skála. Large-order behavior of the perturbation energies for the hydrogen atom in magnetic field. Journal of Mathematical Physics, 47(2):022106, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i2/p022106_s1.

Chen:2006:MAB

[508] Goong Chen, Zhonghai Ding, Sze-Bi Hsu, Moochan Kim, and Jianxin Zhou. Mathematical analysis of a Bohr atom model. Journal of Mathematical Physics, 47(2):022107, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i2/p022107_s1.

Ruuge:2006:NFF

[509] Artur E. Ruuge and Freddy Van Oystaeyen. New families of finite coherent orthoalgebras without bivaluations. Journal of Mathematical Physics, 47(2):022108, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i2/p022108_s1.

Mouchet:2006:ULB

[510] Amaury Mouchet. Upper and lower bounds for an eigenvalue associated with a positive eigenvector. Journal of Mathematical Physics, 47(2):022109, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i2/p022109_s1.

Mondragon:2006:CCF

[511] Mauricio Mondragón and Merced Montesinos. Covariant canonical formalism for four-dimensional BF theory. Journal of Mathematical Physics, 47(2):022301, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i2/p022301_s1.

Koutvitsky:2006:ICS

[512] Vladimir A. Koutvitsky and Eugene M. Maslov. Instability of coherent states of a real scalar field. Journal of Mathematical Physics, 47(2):022302, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i2/p022302_s1.

Adam:2006:IAS

[513] C. Adam, J. Sánchez-Guillén, and A. Wereszczyński. Integrability from an Abelian subgroup of the diffeomorphisms group. Journal of Mathematical Physics, 47(2):022303, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i2/p022303_s1.

Giambo:2006:GVN

[514] Roberto Giambò. Global visibility of naked singularities. Journal of Mathematical Physics, 47(2):022501, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i2/p022501_s1.

Chrusciel:2006:KVA

[515] Piotr T. Chruściel and Daniel Maerten. Killing vectors in asymptotically flat space–times. II. Asymptotically translational Killing vectors and the rigid positive energy theorem in higher dimensions. Journal of Mathematical Physics, 47(2):022502, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i2/p022502_s1.

Maran:2006:SDR

[516] Suresh K. Maran. A systematic derivation of the Riemannian Barrett– Crane intertwiner. Journal of Mathematical Physics, 47(2):022503, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i2/p022503_s1.

Takagi:2006:APA

[517] Taichiro Takagi. Analysis of a particle antiparticle description of a soliton cellular automaton. Journal of Mathematical Physics, 47(2):022701, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i2/p022701_s1.

Bjerklov:2006:PLE

[518] Kristian Bjerklöv. Positive Lyapunov exponents for continuous quasiperiodic Schrödinger equations. Journal of Mathematical Physics, 47 (2):022702, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i2/p022702_s1. See publisher’s note [628].

Bruschi:2006:NSV

[519] M. Bruschi and F. Calogero. Novel solvable variants of the goldfish manybody model. Journal of Mathematical Physics, 47(2):022703, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i2/p022703_s1.

Llibre:2006:FAI

[520] Jaume Llibre and Clàudìa Valls. Formal and analytical integrability of the Bianchi IX system. Journal of Mathematical Physics, 47(2):022704, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i2/p022704_s1.

Benatti:2006:DDS

[521] Fabio Benatti, Alberto Verjovsky, and Federico Zertuche. Discrete dynamical systems embedded in Cantor sets. Journal of Mathematical Physics, 47(2):022705, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i2/p022705_s1.

Hernandez-Bermejo:2006:NFD

[522] Benito Hernández-Bermejo. New four-dimensional solutions of the Jacobi equations for Poisson structures. Journal of Mathematical Physics, 47(2):022901, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i2/p022901_s1.

Cendra:2006:GNM

[523] H. Cendra and S. Grillo. Generalized nonholonomic mechanics, servomechanisms and related brackets. Journal of Mathematical Physics, 47(2):022902, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i2/p022902_s1.

Benn:2006:GKT

[524] I. M. Benn. Geodesics and Killing tensors in mechanics. Journal of Mathematical Physics, 47(2):022903, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i2/p022903_s1.

Toller:2006:TPA

[525] M. Toller. Test particles with acceleration-dependent Lagrangian. Journal of Mathematical Physics, 47(2):022904, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i2/p022904_s1.

Forrester:2006:AFD

[526] P. J. Forrester, N. E. Frankel, and T. M. Garoni. Asymptotic form of the density profile for Gaussian and Laguerre random matrix ensembles with orthogonal and symplectic symmetry. Journal of Mathematical Physics, 47(2):023301, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i2/p023301_s1.

Furuichi:2006:ITP

[527] Shigeru Furuichi. Information theoretical properties of Tsallis entropies. Journal of Mathematical Physics, 47(2):023302, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i2/p023302_s1.

Lehmann:2006:CES

[528] E. Lehmann. Covariant equilibrium statistical mechanics. Journal of Mathematical Physics, 47(2):023303, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i2/p023303_s1.

Alimohammadi:2006:SRD

[529] Masoud Alimohammadi. Solvable reaction-diffusion processes without exclusion. Journal of Mathematical Physics, 47(2):023304, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i2/p023304_s1.

Bachmaier:2006:SND

[530] Fabian Bachmaier and Christian Blohmann. Separation of noncommutative differential calculus on quantum Minkowski space. Journal of Mathematical Physics, 47(2):023501, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i2/p023501_s1.

Hu:2006:GQE

[531] Xinhua Hu and Zhongxing Ye. Generalized quantum entropy. Journal of Mathematical Physics, 47(2):023502, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i2/p023502_s1.

DeMicheli:2006:HGO

[532] Enrico De Micheli, Irene Scorza, and Giovanni Alberto Viano. Hyperbolic geometrical optics: Hyperbolic glass. Journal of Mathematical Physics, 47(2):023503, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i2/p023503_s1.

Bogoyavlenskij:2006:STB

[533] Oleg I. Bogoyavlenskij. Schouten tensor and bi-Hamiltonian systems of hydrodynamic type. Journal of Mathematical Physics, 47(2):023504, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i2/p023504_s1.

Palmkvist:2006:RLA

[534] Jakob Palmkvist. A realization of the Lie algebra associated to a Kantor triple system. Journal of Mathematical Physics, 47(2):023505, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i2/p023505_s1.

Kajiura:2006:OCH

[535] Hiroshige Kajiura and Jim Stasheff. Open-closed homotopy algebra in mathematical physics. Journal of Mathematical Physics, 47(2):023506, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i2/p023506_s1.

Turner:2006:VCS

[536] P. S. Turner, D. J. Rowe, and J. Repka. Vector coherent state theory of the generic representations of ∫≀(▽) in an ∫≀(∋) basis. Journal of Mathematical Physics, 47(2):023507, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i2/p023507_s1.

Chandrasekar:2006:SUA

[537] V. K. Chandrasekar, S. N. Pandey, M. Senthilvelan, and M. Lakshmanan. A simple and unified approach to identify integrable nonlinear oscillators and systems. Journal of Mathematical Physics, 47(2):023508, February 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i2/p023508_s1.

Velasco:2006:SED

[538] Claudio de Jesús Pita Ruiz Velasco and Stephen Bruce Sontz. On Shannon entropies in μ-deformed Segal–Bargmann analysis. Journal of Mathematical Physics, 47(3):032101, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i3/p032101_s1.

Combescure:2006:QFT

[539] Monique Combescure. The quantum fidelity for the time-periodic singular harmonic oscillator. Journal of Mathematical Physics, 47(3): 032102, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p032102_s1.

Brummelhuis:2006:EHA

[540] Raymond Brummelhuis and Pierre Duclos. Effective Hamiltonians for atoms in very strong magnetic fields. Journal of Mathematical Physics, 47(3):032103, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i3/p032103_s1.

Altafini:2006:RSM

[541] Claudio Altafini and Timothy F. Havel. Reflection symmetries for multiqubit density operators. Journal of Mathematical Physics, 47(3): 032104, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p032104_s1.

Shiri-Garakani:2006:FQK

[542] Mohsen Shiri-Garakani and David Ritz Finkelstein. Finite quantum kinematics of the harmonic oscillator. Journal of Mathematical Physics, 47(3):032105, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i3/p032105_s1.

Katsura:2006:ESR

[543] Hosho Katsura and Hideo Aoki. Exact supersymmetry in the relativistic hydrogen atom in general dimensions — supercharge and the generalized Johnson–Lippmann operator. Journal of Mathematical Physics, 47(3): 032301, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p032301_s1.

Moretti:2006:BEC

[544] Valter Moretti. Bose–Einstein condensate and spontaneous breaking of conformal symmetry on Killing horizons II. Journal of Mathematical Physics, 47(3):032302, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i3/p032302_s1.

Nottale:2006:NAG

[545] Laurent Nottale, Marie-Noëlle Célérier, and Thierry Lehner. Non-Abelian gauge field theory in scale relativity. Journal of Mathematical Physics, 47(3):032303, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i3/p032303_s1.

Gegenberg:2006:GRF

[546] J. Gegenberg and G. Kunstatter. The generalized Ricci flow for threedimensional manifolds with one Killing vector. Journal of Mathematical Physics, 47(3):032304, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i3/p032304_s1.

Darabi:2006:CET

[547] F. Darabi and M. R. Setare. Casimir effect in a two-dimensional signature changing space–time. Journal of Mathematical Physics, 47(3): 032501, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p032501_s1.

Stachowiak:2006:NDP

[548] Tomasz Stachowiak, Marek Szydlowski, and Andrzej J. Maciejewski. Nonintegrability of density perturbations in the Friedmann–Robertson– Walker universe. Journal of Mathematical Physics, 47(3):032502, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i3/p032502_s1.

Karmakar:2006:AIP

[549] P. K. Karmakar and C. B. Dwivedi. Analytical integrability and physical solutions of d-KdV equation. Journal of Mathematical Physics, 47(3): 032901, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p032901_s1. See comment [899].

Conca:2006:BCP

[550] Carlos Conca, Rafael Orive, and Muthusamy Vanninathan. On Burnett coefficients in periodic media. Journal of Mathematical Physics, 47(3): 032902, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p032902_s1.

Carozzi:2006:RIS

[551] T. D. Carozzi and J. E. S. Bergman. Real irreducible sesquilinearquadratic tensor concomitants of complex bivectors. Journal of Mathematical Physics, 47(3):032903, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v47/i3/p032903_s1. See pub-

lisher’s note [630].

Blaszak:2006:SSI

[552] Maciej Blaszak and Krzysztof Marciniak. From Stäckel systems to integrable hierarchies of PDE’s: Benenti class of separation relations. Journal of Mathematical Physics, 47(3):032904, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i3/p032904_s1.

Hassaine:2006:EKG

[553] Mokhtar Hassaïne. Extended Klein–Gordon action, gravity and nonrelativistic fluid. Journal of Mathematical Physics, 47(3):033101, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i3/p033101_s1. See [553].

Asch:2006:STR

[554] J. Asch, M. A. Astaburuaga, P. Briet, V. H. Cortés, P. Duclos, and C. Fernández. Sojourn time for rank one perturbations. Journal of Mathematical Physics, 47(3):033501, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i3/p033501_s1.

Kalnins:2006:EQS

[555] E. G. Kalnins, W. Miller, and G. S. Pogosyan. Exact and quasiexact solvability of second-order superintegrable quantum systems: I. Euclidean space preliminaries. Journal of Mathematical Physics, 47(3): 033502, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p033502_s1.

Engman:2006:ISG

[556] Martin Engman and Ricardo Cordero-Soto. Intrinsic spectral geometry of the Kerr–Newman event horizon. Journal of Mathematical Physics, 47(3):033503, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i3/p033503_s1.

Misra:2006:FEU

[557] Aalok Misra. Flow equations for uplifting half-flat to Spin(7) manifolds. Journal of Mathematical Physics, 47(3):033504, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i3/p033504_s1.

Scholz:2006:NZP

[558] Daniel Scholz and Michael Weyrauch. A note on the Zassenhaus product formula. Journal of Mathematical Physics, 47(3):033505, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i3/p033505_s1.

Maceda:2006:BSP

[559] M. Maceda. On the Birman–Schwinger principle applied to \sqrt{−Δ + m^{2}}− m. Journal of Mathematical Physics, 47(3):033506, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i3/p033506_s1.

Pechen:2006:MCF

[560] A. N. Pechen. The multitime correlation functions, free white noise, and the generalized Poisson statistics in the low density limit. Journal of Mathematical Physics, 47(3):033507, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i3/p033507_s1.

Negro:2006:AGT

[561] J. Negro, M. A. del Olmo, and J. Tosiek. Anyons, group theory and planar physics. Journal of Mathematical Physics, 47(3):033508, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i3/p033508_s1.

Nardone:2006:CHS

[562] Mario Nardone. Certain hidden symmetries of crystals. Journal of Mathematical Physics, 47(3):033509, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i3/p033509_s1.

Moriya:2006:MPS

[563] Hajime Moriya. Markov property and strong additivity of von Neumann entropy for graded quantum systems. Journal of Mathematical Physics, 47(3):033510, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i3/p033510_s1.

Skrypnyk:2006:GQG

[564] T. Skrypnyk. Generalized quantum Gaudin spin chains, involutive automorphisms and “twisted” classical r-matrices. Journal of Mathematical Physics, 47(3):033511, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i3/p033511_s1.

Sudo:2006:SDH

[565] Takahiro Sudo. Symmetries on the discrete Heisenberg group C^{∗}-algebra. Journal of Mathematical Physics, 47(3):033512, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i3/p033512_s1.

Krasnov:2006:RCT

[566] Kirill Krasnov and Jorma Louko. SO_{0}(1, d + 1) Racah coefficients: Type I representations. Journal of Mathematical Physics, 47(3):033513, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i3/p033513_s1.

Biondini:2006:SSK

[567] Gino Biondini and S. Chakravarty. Soliton solutions of the Kadomtsev– Petviashvili II equation. Journal of Mathematical Physics, 47(3):033514, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p033514_s1.

Iacono:2006:CEA

[568] R. Iacono. Comment on “Exact analytic solutions of nonlinear boundary value problems in fluid mechanics (Blasius equations)”, [J. Math. Phys. 46, 033101 (2005)]. Journal of Mathematical Physics, 47(3): 034101, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i3/p034101_s1. See [83].

Bergeron:2006:WWT

[569] H. Bergeron. The Wigner–Weyl transformation and the quantum path integral. Journal of Mathematical Physics, 47(4):042101, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/
p042101_s1.

Cassanas:2006:RGF

[570] Roch Cassanas. Reduced Gutzwiller formula with symmetry: Case of a finite group. Journal of Mathematical Physics, 47(4):042102, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i4/p042102_s1.

Aguila:2006:RID

[571] Carlos Angulo Aguila and Stephen Bruce Sontz. Reverse inequalities in μ-deformed Segal–Bargmann analysis. Journal of Mathematical Physics, 47(4):042103, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p042103_s1.

Yngve:2006:PIC

[572] Staffan Yngve. Phase-integral calculation of the phase of the regular Coulomb wave function F_{L}(η, ρ). Journal of Mathematical Physics, 47(4):

042104, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i4/p042104_s1.

Draganescu:2006:MCP

[573] G. E. Draganescu. Motion of a charged particle in a periodic Aharonov– Bohm potential. Journal of Mathematical Physics, 47(4):042105, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i4/p042105_s1.

King:2006:NMR

[574] Christopher King and Nilufer Koldan. New multiplicativity results for qubit maps. Journal of Mathematical Physics, 47(4):042106, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/
p042106_s1.

Arvidsson:2006:SSI

[575] Pär Arvidsson. Superconformal symmetry in the interacting theory of (2, 0) tensor multiplets and self-dual strings. Journal of Mathematical Physics, 47(4):042301, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p042301_s1.

Lin:2006:FTT

[576] De-Hone Lin. Friedel theorem for two dimensional relativistic spin-1/2 systems. Journal of Mathematical Physics, 47(4):042302, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/
p042302_s1.

Kronthaler:2006:CPW

[577] Johann Kronthaler. The Cauchy problem for the wave equation in the Schwarzschild geometry. Journal of Mathematical Physics, 47(4):042501, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i4/p042501_s1.

Anabitarte:2006:STM

[578] Mariano Anabitarte and Mauricio Bellini. Space–time matter inflation. Journal of Mathematical Physics, 47(4):042502, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/p042502_s1.

Hasse:2006:MTA

[579] Wolfgang Hasse and Volker Perlick. A Morse-theoretical analysis of gravitational lensing by a Kerr–Newman black hole. Journal of Mathematical Physics, 47(4):042503, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p042503_s1.

Lemos:2006:GMM

[580] José P. S. Lemos and Vilson T. Zanchin. Gravitational magnetic monopoles and Majumdar–Papapetrou stars. Journal of Mathematical Physics, 47(4):042504, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p042504_s1.

Christodoulakis:2006:DGS

[581] T. Christodoulakis, Th. Grammenos, Ch. Helias, P. G. Kevrekidis, and A. Spanou. Decoupling of the general scalar field mode and the solution space for Bianchi type I and V cosmologies coupled to perfect fluid sources. Journal of Mathematical Physics, 47(4):042505, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i4/p042505_s1.

Calogero:2006:IPI

[582] F. Calogero and F. Leyvraz. Isochronous and partially isochronous Hamiltonian systems are not rare. Journal of Mathematical Physics, 47(4):042901, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p042901_s1.

Margetis:2006:LFC

[583] Dionisios Margetis and Nikos Savva. Low-frequency currents induced in adjacent spherical cells. Journal of Mathematical Physics, 47(4):042902, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i4/p042902_s1.

Yavari:2006:SMC

[584] Arash Yavari, Jerrold E. Marsden, and Michael Ortiz. On spatial and material covariant balance laws in elasticity. Journal of Mathematical Physics, 47(4):042903, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p042903_s1.

Daskaloyannis:2006:UTC

[585] C. Daskaloyannis and K. Ypsilantis. Unified treatment and classification of superintegrable systems with integrals quadratic in momenta on a two-dimensional manifold. Journal of Mathematical Physics, 47(4): 042904, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i4/p042904_s1.

Cortes:2006:HTC

[586] Jorge Cortés and Alexandre M. Vinogradov. Hamiltonian theory of constrained impulsive motion. Journal of Mathematical Physics, 47(4): 042905, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i4/p042905_s1.

Deryabin:2006:FAL

[587] Mikhail V. Deryabin and Morten Willatzen. Flow acoustics and linearized equations for ideal barotropic fluid flows. Journal of Mathematical Physics, 47(4):043101, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p043101_s1.

Freidlin:2006:MRS

[588] M. I. Freidlin and A. I. Komech. On metastable regimes in stochastic Lamb system. Journal of Mathematical Physics, 47(4):043301, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i4/p043301_s1.

Bourgade:2006:QEC

[589] J.-P. Bourgade, P. Degond, F. Méhats, and C. Ringhofer. On quantum extensions to classical spherical harmonics expansion/Fokker–Planck models. Journal of Mathematical Physics, 47(4):043302, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/
p043302_s1.

Dai:2006:NVS

[590] Chao qing Dai and Jie fang Zhang. Novel variable separation solutions and exotic localized excitations via the ETM in nonlinear soliton systems. Journal of Mathematical Physics, 47(4):043501, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/p043501_s1.

Budinich:2006:SFM

[591] Marco Budinich and Paolo Budinich. A spinorial formulation of the maximum clique problem of a graph. Journal of Mathematical Physics, 47(4): 043502, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i4/p043502_s1.

Milhorat:2006:FFE

[592] Jean-Louis Milhorat. A formula for the first eigenvalue of the Dirac operator on compact spin symmetric spaces. Journal of Mathematical Physics, 47(4):043503, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p043503_s1.

Hansen:2006:TFL

[593] Frank Hansen. Trace functions as Laplace transforms. Journal of Mathematical Physics, 47(4):043504, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i4/p043504_s1.

Degiovanni:2006:THE

[594] Luca Degiovanni. Tri-Hamiltonian extensions of separable systems in the plane. Journal of Mathematical Physics, 47(4):043505, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/
p043505_s1.

Kirsten:2006:VUP

[595] Klaus Kirsten, Paul Loya, and Jinsung Park. The very unusual properties of the resolvent, heat kernel, and zeta function for the operator −d^{2}/dr^{2}−1/(4r^{2}). Journal of Mathematical Physics, 47(4):043506, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i4/p043506_s1.

Koca:2006:QRS

[596] Mehmet Koca, Muataz Al-Barwani, and Ramazan Koç. Quaternionic root systems and subgroups of the Aut(F_{4}). Journal of Mathematical Physics, 47(4):043507, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p043507_s1.

Lyakhovich:2006:BQQ

[597] S. L. Lyakhovich and A. A. Sharapov. BRST quantization of quasisymplectic manifolds and beyond. Journal of Mathematical Physics, 47(4): 043508, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i4/p043508_s1.

Levi:2006:DRP

[598] Decio Levi and Matteo Petrera. Discrete reductive perturbation technique. Journal of Mathematical Physics, 47(4):043509, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/
p043509_s1.

Bertini:2006:EA

[599] Stefano Bertini, Sergio L. Cacciatori, and Bianca L. Cerchiai. On the Euler angles for SU(N). Journal of Mathematical Physics, 47(4):043510, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i4/p043510_s1.

Calzada:2006:SQS

[600] J. A. Calzada, J. Negro, and M. A. del Olmo. Superintegrable quantum u(3) systems and higher rank factorizations. Journal of Mathematical Physics, 47(4):043511, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p043511_s1.

Patera:2006:DCS

[601] J. Patera and A. Zaratsyan. Discrete and continuous sine transform generalized to semisimple Lie groups of rank two. Journal of Mathematical Physics, 47(4):043512, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i4/p043512_s1.

Boyallian:2006:RSF

[602] Carina Boyallian, Victor G. Kac, Jose I. Liberati, and Alexei Rudakov. Representations of simple finite Lie conformal superalgebras of type W and S. Journal of Mathematical Physics, 47(4):043513, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i4/
p043513_s1.

Kalnins:2006:SOSa

[603] E. G. Kalnins, J. M. Kress, and W. Miller. Second order superintegrable systems in conformally flat spaces. IV. The classical 3D Stäckel transform and 3D classification theory. Journal of Mathematical Physics, 47(4): 043514, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i4/p043514_s1.

Yilmaz:2006:POM

[604] Bülent Yilmaz. The performance of the OSRC method for concentric penetrable circular cylinder. Journal of Mathematical Physics, 47(4): 043515, April 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i4/p043515_s1.

Chen:2006:QEG

[605] Hao Chen. Quantum entanglement and geometry of determinantal varieties. Journal of Mathematical Physics, 47(5):052101, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i5/
p052101_s1.

Tillman:2006:OAM

[606] Philip Tillman and George Sparling. Observables of angular momentum as observables on the Fedosov quantized sphere. Journal of Mathematical Physics, 47(5):052102, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i5/p052102_s1.

Pulvirenti:2006:SEW

[607] M. Pulvirenti. Semiclassical expansion of Wigner functions. Journal of Mathematical Physics, 47(5):052103, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i5/p052103_s1.

Dehesa:2006:FID

[608] J. S. Dehesa, S. López-Rosa, B. Olmos, and R. J. Yáñez. Fisher information of D-dimensional hydrogenic systems in position and momentum spaces. Journal of Mathematical Physics, 47(5):052104, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i5/p052104_s1.

Mestre:2006:CPF

[609] Ângela Mestre and Robert Oeckl. Combinatorics of n-point functions via Hopf algebra in quantum field theory. Journal of Mathematical Physics, 47(5):052301, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i5/p052301_s1.

Adam:2006:INM

[610] C. Adam, J. Sánchez-Guillén, R. A. Vázquez, and A. Wereszczyński. Investigation of the Nicole model. Journal of Mathematical Physics, 47 (5):052302, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i5/p052302_s1.

Alcalde:2006:TLM

[611] M. Aparicio Alcalde, G. Flores Hidalgo, and N. F. Svaiter. The two-loop massless (λ/4!)φ^{4} model in nontranslational invariant domain. Journal of Mathematical Physics, 47(5):052303, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i5/p052303_s1.

Hu:2006:QMO

[612] Yapeng Hu, Weigang Qiu, and Hongbao Zhang. On the quantum mechanics for one photon. Journal of Mathematical Physics, 47(5):052304, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i5/p052304_s1.

Barrett:2006:NGS

[613] John W. Barrett and Rachel A. Dawe Martins. Noncommutative geometry and the standard model vacuum. Journal of Mathematical Physics, 47 (5):052305, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i5/p052305_s1.

Kahil:2006:MKK

[614] M. E. Kahil. Motion in Kaluza–Klein type theories. Journal of Mathematical Physics, 47(5):052501, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i5/p052501_s1.

Herrera:2006:GRV

[615] L. Herrera, N. O. Santos, and J. Carot. Gravitational radiation, vorticity and the electric and magnetic part of Weyl tensor. Journal of Mathematical Physics, 47(5):052502, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i5/p052502_s1.

Melnyk:2006:LCP

[616] Fabrice Melnyk. Local Cauchy problem for the nonlinear Dirac and Dirac–Klein–Gordon equations on Kerr space–time. Journal of Mathematical Physics, 47(5):052503, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i5/p052503_s1.

Carminati:2006:DAS

[617] J. Carminati and A. E. K. Lim. The determination of all syzygies for the dependent polynomial invariants of the Riemann tensor. II. Mixed invariants of even degree in the Ricci spinor. Journal of Mathematical Physics, 47(5):052504, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i5/p052504_s1.

Zhao:2006:UAN

[618] Caidi Zhao, Shengfan Zhou, and Xinyuan Liao. Uniform attractors for nonautonomous incompressible non-Newtonian fluid with locally uniform integrable external forces. Journal of Mathematical Physics, 47(5): 052701, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i5/p052701_s1.

Yang:2006:NEM

[619] Tong Yang and Hui-Jiang Zhao. A new energy method for the Boltzmann equation. Journal of Mathematical Physics, 47(5):053301, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i5/p053301_s1.

Ma:2006:SSL

[620] Wen-Xiu Ma, Xi-Xiang Xu, and Yufeng Zhang. Semidirect sums of Lie algebras and discrete integrable couplings. Journal of Mathematical Physics, 47(5):053501, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i5/p053501_s1.

Wang:2006:TTD

[621] Erxiao Wang. Tzitzéica transformation is a dressing action. Journal of Mathematical Physics, 47(5):053502, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i5/p053502_s1.

Agop:2006:GTF

[622] M. Agop and I. Gottlieb. Gravitation theory in a fractal space–time. Journal of Mathematical Physics, 47(5):053503, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i5/p053503_s1.

Dutkay:2006:HSB

[623] Dorin Ervin Dutkay and Palle E. T. Jorgensen. Hilbert spaces built on a similarity and on dynamical renormalization. Journal of Mathematical Physics, 47(5):053504, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i5/p053504_s1.

Bakalov:2006:JIV

[624] Bojko Bakalov and Nikolay M. Nikolov. Jacobi identity for vertex algebras in higher dimensions. Journal of Mathematical Physics, 47(5): 053505, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i5/p053505_s1.

Douglas:2006:FDR

[625] Andrew Douglas. Finite dimensional representations of the Euclidean algebra ⌉(∈) having two generators. Journal of Mathematical Physics, 47 (5):053506, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i5/p053506_s1.

Mada:2006:ECB

[626] Jun Mada, Makoto Idzumi, and Tetsuji Tokihiro. The exact correspondence between conserved quantities of a periodic box-ball system and string solutions of the Bethe ansatz equations. Journal of Mathematical Physics, 47(5):053507, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i5/p053507_s1.

Abdesselam:2006:NSP

[627] B. Abdesselam and A. Chakrabarti. Nested sequence of projectors. II. Multiparameter multistate statistical models, Hamiltonians, S-matrices. Journal of Mathematical Physics, 47(5):053508, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i5/p053508_s1.

Bjerklov:2006:PNP

[628] Kristian Bjerklöv. Publisher’s note: “Positive Lyapunov exponents for continuous quasiperiodic Schrödinger equations”, [J. Math. Phys. 47, 022702 (2006)]. Journal of Mathematical Physics, 47(5):059901, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i5/p059901_s1. See [518].

Sundkvist:2006:PNC

[629] David Sundkvist. Publisher’s note: “Covariant irreducible parametrization of electromagnetic fields in arbitrary space–time”, [J. Math. Phys. 47, 012901 (2006)]. Journal of Mathematical Physics, 47(5):059902, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i5/p059902_s1. See [482].

Carozzi:2006:PNR

[630] T. D. Carozzi and J. E. S. Bergman. Publisher’s note: “Real irreducible sesquilinear-quadratic tensor concomitants of complex bivectors”, [J. Math. Phys. 47, 032903 (2006)]. Journal of Mathematical Physics, 47(5): 059903, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i5/p059903_s1. See [551].

Boscain:2006:TMT

[631] Ugo Boscain and Paolo Mason. Time minimal trajectories for a spin 1/2 particle in a magnetic field. Journal of Mathematical Physics, 47(6): 062101, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062101_s1.

Killip:2006:ARF

[632] Rowan Killip and Robert Sims. Absence of reflection as a function of the coupling constant. Journal of Mathematical Physics, 47(6):062102, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i6/p062102_s1.

Khare:2006:CPP

[633] Avinash Khare and Uday Sukhatme. Complex periodic potentials with a finite number of band gaps. Journal of Mathematical Physics, 47(6): 062103, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062103_s1.

Beneduci:2006:GCC

[634] Roberto Beneduci. A geometrical characterization of commutative positive operator valued measures. Journal of Mathematical Physics, 47(6): 062104, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062104_s1.

DAndrea:2006:SGM

[635] Francesco D’Andrea. Spectral geometry of κ-Minkowski space. Journal of Mathematical Physics, 47(6):062105, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i6/p062105_s1.

Bravyi:2006:GEY

[636] Sergey Bravyi, David Fattal, and Daniel Gottesman. GHZ extraction yield for multipartite stabilizer states. Journal of Mathematical Physics, 47(6):062106, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i6/p062106_s1.

Albeverio:2006:PQS

[637] S. Albeverio, J. Rezende, and J.-C. Zambrini. Probability and quantum symmetries. II. The Theorem of Noether in quantum mechanics. Journal of Mathematical Physics, 47(6):062107, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i6/p062107_s1.

Zhixiang:2006:WHA

[638] Wu Zhixiang. The weak Hopf algebras related to generalized Kac– Moody algebra. Journal of Mathematical Physics, 47(6):062108, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i6/p062108_s1.

Liverts:2006:ACE

[639] E. Z. Liverts, V. B. Mandelzweig, and F. Tabakin. Analytic calculation of energies and wave functions of the quartic and pure quartic oscillators. Journal of Mathematical Physics, 47(6):062109, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i6/p062109_s1.

Villegas-Blas:2006:BTR

[640] Carlos Villegas-Blas. The Bargmann transform and regularization of the 2, 3, 5-dimensional Kepler problem. Journal of Mathematical Physics, 47 (6):062110, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062110_s1.

Boztosun:2006:AIM

[641] I. Boztosun, M. Karakoc, F. Yasuk, and A. Durmus. Asymptotic iteration method solutions to the relativistic Duffin–Kemmer–Petiau equation. Journal of Mathematical Physics, 47(6):062301, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i6/p062301_s1.

Suzuki:2006:MAF

[642] Ryo Suzuki and Yuji Tachikawa. More anomaly free models of sixdimensional gauged supergravity. Journal of Mathematical Physics, 47 (6):062302, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062302_s1.

Dolan:2006:CFP

[643] F. A. Dolan. Character formulas and partition functions in higher dimensional conformal field theory. Journal of Mathematical Physics, 47 (6):062303, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062303_s1.

Sonego:2006:OGA

[644] Sebastiano Sonego and Marek A. Abramowicz. Optical geometry analysis of the electromagnetic self-force. Journal of Mathematical Physics, 47(6): 062501, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062501_s1.

Gariel:2006:PSL

[645] J. Gariel, G. Marcilhacy, and N. O. Santos. Parametrization of solutions of the Lewis metric by a Painlevé transcendent III. Journal of Mathematical Physics, 47(6):062502, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i6/p062502_s1.

Coll:2006:GKC

[646] Bartolomé Coll, Joan Josep Ferrando, and Juan Antonio Sáez. On the geometry of Killing and conformal tensors. Journal of Mathematical Physics, 47(6):062503, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i6/p062503_s1.

Derdzinski:2006:WTC

[647] Andrzej Derdzinski and Witold Roter. Walker’s theorem without coordinates. Journal of Mathematical Physics, 47(6):062504, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i6/

p062504_s1. See publisher’s note [860].

Gailis:2006:TCC

[648] R. M. Gailis and N. E. Frankel. Two-component cosmological fluids with gravitational instabilities. Journal of Mathematical Physics, 47(6): 062505, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062505_s1.

Gailis:2006:SWA

[649] R. M. Gailis and N. E. Frankel. Short wavelength analysis of the evolution of perturbations in a two-component cosmological fluid. Journal of Mathematical Physics, 47(6):062506, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i6/p062506_s1.

Vougalter:2006:EZE

[650] Vitali Vougalter and Dimitry Pelinovsky. Eigenvalues of zero energy in the linearized NLS problem. Journal of Mathematical Physics, 47(6): 062701, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062701_s1.

Loutsenko:2006:NLG

[651] Igor Loutsenko and Oksana Yermolayeva. Non-Laplacian growth, algebraic domains, and finite reflection groups. Journal of Mathematical Physics, 47(6):062702, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i6/p062702_s1.

Diaz-Cano:2006:STD

[652] A. Díaz-Cano, F. González-Gascón, and D. Peralta-Salas. On scattering trajectories of dynamical systems. Journal of Mathematical Physics, 47 (6):062703, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062703_s1.

Cacciapuoti:2006:PIA

[653] C. Cacciapuoti, R. Figari, and A. Posilicano. Point interactions in acoustics: One-dimensional models. Journal of Mathematical Physics, 47(6): 062901, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062901_s1.

Jollivet:2006:ISM

[654] Alexandre Jollivet. On inverse scattering for the multidimensional relativistic Newton equation at high energies. Journal of Mathematical Physics, 47(6):062902, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i6/p062902_s1.

Modugno:2006:CIS

[655] Marco Modugno, Dirk Saller, and Jürgen Tolksdorf. Classification of infinitesimal symmetries in covariant classical mechanics. Journal of Mathematical Physics, 47(6):062903, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i6/p062903_s1.

Saenz:2006:PEA

[656] A. W. Sáenz. Proof of the existence of the axial-to-planar channeling transition in a simple crystal model. Journal of Mathematical Physics, 47 (6):062904, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p062904_s1. See erratum [1227].

Pavon:2006:EPC

[657] Michele Pavon and Francesco Ticozzi. On entropy production for controlled Markovian evolution. Journal of Mathematical Physics, 47(6): 063301, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p063301_s1.

Dumitriu:2006:GSF

[658] Ioana Dumitriu and Alan Edelman. Global spectrum fluctuations for the β-Hermite and β-Laguerre ensembles via matrix models. Journal of Mathematical Physics, 47(6):063302, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i6/p063302_s1.

Kresic-Juric:2006:BFP

[659] Sasa Kresić-Jurić. On the Birkhoff factorization problem for the Heisenberg magnet and nonlinear Schrödinger equations. Journal of Mathematical Physics, 47(6):063501, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i6/p063501_s1.

Bogoyavlenskij:2006:BDP

[660] Oleg I. Bogoyavlenskij. Block-diagonalizability problem for hydrodynamic type systems. Journal of Mathematical Physics, 47(6):063502, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i6/p063502_s1.

Shu:2006:NSE

[661] J. Shu and J. Zhang. Nonlinear Schrödinger equation with harmonic potential. Journal of Mathematical Physics, 47(6):063503, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i6/
p063503_s1.

Chamseddine:2006:SIS

[662] Ali H. Chamseddine and Alain Connes. Scale invariance in the spectral action. Journal of Mathematical Physics, 47(6):063504, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i6/
p063504_s1.

Lima:2006:CGC

[663] I. Sánchez Lima and P. O. Hess. Clebsch–Gordan coefficients for U(8) ⊃ O(8) ⊃ SU(3). Journal of Mathematical Physics, 47(6):063505, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i6/p063505_s1.

Szmytkowski:2006:CFG

[664] Radoslaw Szmytkowski. Closed form of the generalized Green’s function for the Helmholtz operator on the two-dimensional unit sphere. Journal of Mathematical Physics, 47(6):063506, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i6/p063506_s1.

Ruzzi:2006:JFD

[665] M. Ruzzi. Jacobi ϑ-functions and discrete Fourier transforms. Journal of Mathematical Physics, 47(6):063507, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i6/p063507_s1.

Prinari:2006:IST

[666] Barbara Prinari, Mark J. Ablowitz, and Gino Biondini. Inverse scattering transform for the vector nonlinear Schrödinger equation with nonvanishing boundary conditions. Journal of Mathematical Physics, 47(6): 063508, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p063508_s1.

Skrypnyk:2006:MNA

[667] T. Skrypnyk. Modified non-Abelian Toda field equations and twisted quasigraded Lie algebras. Journal of Mathematical Physics, 47(6): 063509, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i6/p063509_s1.

Kong:2006:EET

[668] De-Xing Kong, Qing-You Sun, and Yi Zhou. Erratum: “The equation for time-like extremal surfaces in Minkowski space R^{2+n}”, [J. Math. Phys. 47, 013503 (2006)]. Journal of Mathematical Physics, 47(6):069901, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i6/p069901_s1. See erratum [?].

Torrome:2006:FVH

[669] Ricardo Gallego Torrome. A Finslerian version of ’t Hooft deterministic quantum models. Journal of Mathematical Physics, 47(7):072101, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p072101_s1.

DellAntonio:2006:QGH

[670] Gianfausto Dell’Antonio and Lucattilio Tenuta. Quantum graphs as holonomic constraints. Journal of Mathematical Physics, 47(7):072102, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p072102_s1.

Mostafazadeh:2006:DRP

[671] Ali Mostafazadeh. Differential realization of pseudo-Hermiticity: a quantum mechanical analog of Einstein’s field equation. Journal of Mathematical Physics, 47(7):072103, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i7/p072103_s1.

Kiukas:2006:PSQ

[672] J. Kiukas, P. Lahti, and K. Ylinen. Phase space quantization and the operator moment problem. Journal of Mathematical Physics, 47(7):072104, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p072104_s1.

Aerts:2006:RSP

[673] D. Aerts and S. Pulmannová. Representation of state property systems. Journal of Mathematical Physics, 47(7):072105, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i7/p072105_s1.

Kachmar:2006:GSE

[674] Ayman Kachmar. On the ground state energy for a magnetic Schrödinger operator and the effect of the DeGennes boundary condition. Journal of Mathematical Physics, 47(7):072106, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i7/p072106_s1. See erratum

[900].

Bingol:2006:SGT

[675] Elif Bingol, Burak Tevfik Kaynak, and O. Teoman Turgut. Scalar gauge theory on a cylinder. Journal of Mathematical Physics, 47(7):072301, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p072301_s1.

Korner:2006:LSE

[676] J. G. Körner, Z. Merebashvili, and M. Rogal. Laurent series expansion of a class of massive scalar one-loop integrals up to O(ε^{2}) in terms of multiple polylogarithms. Journal of Mathematical Physics, 47(7):072302, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p072302_s1.

Poplawski:2006:MIF

[677] Nikodem J. Poplawski. A Michelson interferometer in the field of a plane gravitational wave. Journal of Mathematical Physics, 47(7):072501, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p072501_s1.

Montanus:2006:OPA

[678] J. M. C. Montanus. The orbital precession around oblate spheroids. Journal of Mathematical Physics, 47(7):072502, July 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i7/p072502_s1.

Kim:2006:PTC

[679] Do-Hyung Kim. The path topology and the causal completion. Journal of Mathematical Physics, 47(7):072503, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i7/p072503_s1.

Marques:2006:GAB

[680] Geusa de A. Marques, Sandro G. Fernandes, and V. B. Bezerra. Gravitational Aharonov–Bohm effect due to weak fields. Journal of Mathematical Physics, 47(7):072504, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i7/p072504_s1.

Bernardi:2006:PBP

[681] Olga Bernardi and Franco Cardin. Poincaré–Birkhoff periodic orbits for mechanical Hamiltonian systems on T^{∗}T^{n}. Journal of Mathematical Physics, 47(7):072701, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i7/p072701_s1.

Gentile:2006:QAB

[682] Guido Gentile, Michele V. Bartuccelli, and Jonathan H. B. Deane. Quasiperiodic attractors, Borel summability and the Bryuno condition for strongly dissipative systems. Journal of Mathematical Physics, 47(7): 072702, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i7/p072702_s1.

Dejak:2006:LTD

[683] S. I. Dejak and B. L. G. Jonsson. Long-time dynamics of variable coefficient modified Korteweg–de Vries solitary waves. Journal of Mathematical Physics, 47(7):072703, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i7/p072703_s1.

Juang:2006:EPT

[684] Jonq Juang and Chin-Lung Li. Eigenvalue problems and their application to the wavelet method of chaotic control. Journal of Mathematical Physics, 47(7):072704, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i7/p072704_s1.

Drezet:2006:DLP

[685] A. Drezet, J. C. Woehl, and S. Huant. Diffraction of light by a planar aperture in a metallic screen. Journal of Mathematical Physics, 47(7): 072901, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i7/p072901_s1.

Ferreira:2006:EAE

[686] P. Castelo Ferreira. Explicit actions for electromagnetism with two gauge fields with only one electric and one magnetic physical field. Journal of Mathematical Physics, 47(7):072902, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i7/p072902_s1.

Bezuglyy:2006:GOU

[687] V. Bezuglyy, B. Mehlig, M. Wilkinson, K. Nakamura, and E. Arvedson. Generalized Ornstein–Uhlenbeck processes. Journal of Mathematical Physics, 47(7):073301, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i7/p073301_s1.

Haba:2006:HED

[688] Z. Haba. Heat equilibrium distribution in a turbulent flow. Journal of Mathematical Physics, 47(7):073302, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i7/p073302_s1.

DeRoeck:2006:QME

[689] Wojciech De Roeck, Christian Maes, and Karel Netocný. Quantum macrostates, equivalence of ensembles, and an H-theorem. Journal of Mathematical Physics, 47(7):073303, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i7/p073303_s1.

Torre:2006:USH

[690] C. G. Torre. Uniqueness of solutions to the helically reduced wave equation with Sommerfeld boundary conditions. Journal of Mathematical Physics, 47(7):073501, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i7/p073501_s1.

Furuya:2006:FPI

[691] Kiyoko Furuya. Feynman path integral of Riemann type. Journal of Mathematical Physics, 47(7):073502, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i7/p073502_s1.

Nikolov:2006:DRD

[692] Petko A. Nikolov and Gergana R. Ruseva. Dimensional reduction of Dirac operator. Journal of Mathematical Physics, 47(7):073503, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p073503_s1.

Gordoa:2006:SFP

[693] Pilar R. Gordoa, Nalini Joshi, and Andrew Pickering. Second and fourth Painlevé hierarchies and Jimbo–Miwa linear problems. Journal of Mathematical Physics, 47(7):073504, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i7/p073504_s1.

Maltsev:2006:WSD

[694] A. Ya. Maltsev. Whitham systems and deformations. Journal of Mathematical Physics, 47(7):073505, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i7/p073505_s1.

Chanu:2006:GCK

[695] C. Chanu, L. Degiovanni, and R. G. McLenaghan. Geometrical classification of Killing tensors on bidimensional flat manifolds. Journal of Mathematical Physics, 47(7):073506, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i7/p073506_s1.

Menezes:2006:SQT

[696] G. Menezes and N. F. Svaiter. Stochastic quantization of topological field theory: Generalized Langevin equation with memory kernel. Journal of Mathematical Physics, 47(7):073507, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i7/p073507_s1.

Chang-Young:2006:SQT

[697] Ee Chang-Young and Hoil Kim. Symmetry of quantum torus with crossed product algebra. Journal of Mathematical Physics, 47(7):073508, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p073508_s1.

Hernandez:2006:GMO

[698] Antonio E. Cárcamo Hernández. The Gell-Mann–Okubo and Coleman– Glashow relations for octet and decuplet baryons in the SU_{q}(3) quantum algebra. Journal of Mathematical Physics, 47(7):073509, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i7/
p073509_s1.

Thylwe:2006:SVP

[699] Karl-Erik Thylwe, Staffan Yngve, and Per Olof Fröman. Study of the validity of the phase-integral connection formula for potential barriers of arbitrary thickness. Journal of Mathematical Physics, 47(7):073510, July 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i7/p073510_s1.

Dias:2006:CSE

[700] N. C. Dias, A. Miković, and J. N. Prata. Coherent states expectation values as semiclassical trajectories. Journal of Mathematical Physics, 47(8): 082101, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i8/p082101_s1.

Janzing:2006:MDH

[701] Dominik Janzing and Thomas Decker. Minimally disturbing Heisenberg– Weyl symmetric measurements using hard-core collisions of Schrödinger particles. Journal of Mathematical Physics, 47(8):082102, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/
p082102_s1.

Miyamoto:2006:ZER

[702] Manabu Miyamoto. Zero energy resonance and the logarithmically slow decay of unstable multilevel systems. Journal of Mathematical Physics, 47(8):082103, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i8/p082103_s1.

Guo:2006:SPA

[703] Xiaoyi Guo and Mingyu Xu. Some physical applications of fractional Schrödinger equation. Journal of Mathematical Physics, 47(8):082104, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i8/p082104_s1.

Golovnev:2006:TLQ

[704] Alexey V. Golovnev. Thin layer quantization in higher dimensions and codimensions. Journal of Mathematical Physics, 47(8):082105, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p082105_s1.

DeLeo:2006:APW

[705] Stefano De Leo, Gisele C. Ducati, and Tiago M. Madureira. Analytic plane wave solutions for the quaternionic potential step. Journal of Mathematical Physics, 47(8):082106, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i8/p082106_s1.

Damak:2006:TAS

[706] Mondher Damak, Marius Mantoiu, and Rafael Tiedra de Aldecoa. Toeplitz algebras and spectral results for the one-dimensional Heisenberg model. Journal of Mathematical Physics, 47(8):082107, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p082107_s1.

Kerimov:2006:GTA

[707] G. A. Kerimov and A. Ventura. Group-theoretical approach to reflectionless potentials. Journal of Mathematical Physics, 47(8):082108, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p082108_s1.

DAlessandro:2006:DUE

[708] Domenico D’Alessandro and Raffaele Romano. Decompositions of unitary evolutions and entanglement dynamics of bipartite quantum systems. Journal of Mathematical Physics, 47(8):082109, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/
p082109_s1.

Altunkaynak:2006:FMD

[709] Baris Altunkaynak, Fatih Erman, and O. Teoman Turgut. Finitely many Dirac-delta interactions on Riemannian manifolds. Journal of Mathematical Physics, 47(8):082110, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i8/p082110_s1.

Batra:2006:AFS

[710] Puneet Batra, Bogdan A. Dobrescu, and David Spivak. Anomaly-free sets of fermions. Journal of Mathematical Physics, 47(8):082301, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p082301_s1.

Yasuk:2006:EAS

[711] F. Yasuk, A. Durmus, and I. Boztosun. Exact analytical solution to the relativistic Klein–Gordon equation with noncentral equal scalar and vector potentials. Journal of Mathematical Physics, 47(8):082302, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p082302_s1.

Fewster:2006:QEI

[712] Christopher J. Fewster and Michael J. Pfenning. Quantum energy inequalities and local covariance. I. Globally hyperbolic spacetimes. Journal of Mathematical Physics, 47(8):082303, August 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i8/p082303_s1.

Hadasz:2006:LTU

[713] Leszek Hadasz and Zbigniew Jaskólski. Liouville theory and uniformization of four-punctured sphere. Journal of Mathematical Physics, 47(8): 082304, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i8/p082304_s1.

Tolksdorf:2006:GTD

[714] Jürgen Tolksdorf and Torsten Thumstädter. Gauge theories of Dirac type. Journal of Mathematical Physics, 47(8):082305, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/
p082305_s1.

Prikas:2006:SST

[715] Athanasios Prikas. q stars in scalar-tensor gravitational theories in extra dimensions. Journal of Mathematical Physics, 47(8):082501, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p082501_s1.

Shieh:2006:MAW

[716] Shih-Feng Shieh, Yiqian Wang, Guowei Wei, and Choy-Heng Lai. Mathematical analysis of the wavelet method of chaos control. Journal of Mathematical Physics, 47(8):082701, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i8/p082701_s1.

Calogero:2006:GGT

[717] F. Calogero and E. Langmann. Goldfishing by gauge theory. Journal of Mathematical Physics, 47(8):082702, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i8/p082702_s1.

Cuccu:2006:OPE

[718] F. Cuccu, B. Emamizadeh, and G. Porru. Optimization problems for an elastic plate. Journal of Mathematical Physics, 47(8):082901, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p082901_s1.

Draganescu:2006:AVI

[719] G. E. Draganescu. Application of a variational iteration method to linear and nonlinear viscoelastic models with fractional derivatives. Journal of Mathematical Physics, 47(8):082902, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i8/p082902_s1.

Pasquero:2006:SPU

[720] Stefano Pasquero. On the simultaneous presence of unilateral and kinetic constraints in time-dependent impulsive mechanics. Journal of Mathematical Physics, 47(8):082903, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i8/p082903_s1.

Wang:2006:NSB

[721] Weike Wang, Tong Yang, and Xiongfeng Yang. Nonlinear stability of boundary layers of the Boltzmann equation for cutoff hard potentials. Journal of Mathematical Physics, 47(8):083301, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/p083301_s1.

Ali:2006:SQP

[722] Abdulmuhsen H. Ali. Statistical quantum probabilities in coherent bosonic and fermionic states. Journal of Mathematical Physics, 47(8): 083302, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i8/p083302_s1.

Celik:2006:CCQ

[723] Salih Çelik. Cartan calculi on the quantum superplane. Journal of Mathematical Physics, 47(8):083501, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i8/p083501_s1.

Papageorgiou:2006:YBM

[724] Vassilios G. Papageorgiou, Anastasios G. Tongas, and Alexander P. Veselov. Yang–Baxter maps and symmetries of integrable equations on quad-graphs. Journal of Mathematical Physics, 47(8):083502, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p083502_s1.

Jiang:2006:NMS

[725] Bin Jiang, Yushun Wang, and Jiaxiang Cai. A new multisymplectic scheme for generalized Kadomtsev–Petviashvili equation. Journal of Mathematical Physics, 47(8):083503, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i8/p083503_s1.

Barletta:2006:YMF

[726] Elisabetta Barletta, Sorin Dragomir, and Hajime Urakawa. Yang– Mills fields on CR manifolds. Journal of Mathematical Physics, 47(8): 083504, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i8/p083504_s1.

Keller:2006:UQF

[727] Jaime Keller and Peter Weinberger. The use of quadratic forms in the calculation of ground state electronic structures. Journal of Mathematical Physics, 47(8):083505, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i8/p083505_s1.

Perez-Garcia:2006:CPT

[728] David Pérez-García, Michael M. Wolf, Denes Petz, and Mary Beth Ruskai. Contractivity of positive and trace-preserving maps under L_{p} norms. Journal of Mathematical Physics, 47(8):083506, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/
p083506_s1.

Daboul:2006:CBS

[729] Jamil Daboul. Contraction of broken symmetries via Kac–Moody formalism. Journal of Mathematical Physics, 47(8):083507, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/
p083507_s1.

Guo:2006:EWS

[730] Boling Guo and Daiwen Huang. Existence of weak solutions and trajectory attractors for the moist atmospheric equations in geophysics. Journal of Mathematical Physics, 47(8):083508, August 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i8/p083508_s1.

Strumia:2006:WPF

[731] Alberto Strumia. Waves, particles, and field dynamics. Journal of Mathematical Physics, 47(8):083509, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i8/p083509_s1.

Dita:2006:SUM

[732] Petre Dita. Separation of unistochastic matrices from the double stochastic ones: Recovery of a 3 × 3 unitary matrix from experimental data. Journal of Mathematical Physics, 47(8):083510, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/p083510_s1.

Briet:2006:DEP

[733] Philippe Briet, Horia D. Cornean, and Delphine Louis. Diamagnetic expansions for perfect quantum gases. Journal of Mathematical Physics, 47(8):083511, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i8/p083511_s1.

Alonso:2006:SEW

[734] Luis Martínez Alonso, Elena Medina, and Manuel Mañas. String equations in Whitham hierarchies: τ-functions and Virasoro constraints. Journal of Mathematical Physics, 47(8):083512, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/p083512_s1.

Wu:2006:STC

[735] Chao-Zhong Wu. On solutions of the two-component Camassa–Holm system. Journal of Mathematical Physics, 47(8):083513, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i8/
p083513_s1.

Abe:2006:ERQ

[736] Yuuki Abe and Toshiki Nakashima. Evaluation representations of quantum affine algebras at roots of unity. Journal of Mathematical Physics, 47(8):083514, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i8/p083514_s1.

Cao:2006:MTA

[737] Xifang Cao, Hongyou Wu, and Chuanyou Xu. On Miura transformations among nonlinear partial differential equations. Journal of Mathematical Physics, 47(8):083515, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i8/p083515_s1.

Cognola:2006:HKE

[738] Guido Cognola, Emilio Elizalde, and Sergio Zerbini. Heat-kernel expansion on noncompact domains and a generalized zeta-function regularization procedure. Journal of Mathematical Physics, 47(8):083516, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i8/p083516_s1.

Mostafazadeh:2006:WPH

[739] Ali Mostafazadeh. Is weak pseudo-Hermiticity weaker than pseudo-Hermiticity? Journal of Mathematical Physics, 47(9):092101, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p092101_s1.

Nagaj:2006:OQE

[740] Daniel Nagaj and Iordanis Kerenidis. On the optimality of quantum encryption schemes. Journal of Mathematical Physics, 47(9):092102, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i9/p092102_s1.

Zafiris:2006:STR

[741] Elias Zafiris. Sheaf-theoretic representation of quantum measure algebras. Journal of Mathematical Physics, 47(9):092103, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i9/
p092103_s1.

Vourdas:2006:GQS

[742] A. Vourdas. Galois quantum systems, irreducible polynomials and Riemann surfaces. Journal of Mathematical Physics, 47(9):092104, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p092104_s1.

Viennot:2006:MMQ

[743] David Viennot. Magnetic monopoles in quantum adiabatic dynamics and the immersion property of the control manifold. Journal of Mathematical Physics, 47(9):092105, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i9/p092105_s1.

Nasiri:2006:PSQ

[744] S. Nasiri, Y. Sobouti, and F. Taati. Phase space quantum mechanics — Direct. Journal of Mathematical Physics, 47(9):092106, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p092106_s1.

Chiribella:2006:ECM

[745] Giulio Chiribella and Giacomo Mauro D’Ariano. Extremal covariant measurements. Journal of Mathematical Physics, 47(9):092107, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p092107_s1.

Yamashita:2006:PAS

[746] Hideyasu Yamashita. Pointwise analysis of scalar fields: A nonstandard approach. Journal of Mathematical Physics, 47(9):092301, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v47/i9/p092301_s1. See comment [1157].

Tureanu:2006:ASA

[747] Anca Tureanu. Analyticity of the scattering amplitude, causality, and high-energy bounds in quantum field theory on noncommutative space– time. Journal of Mathematical Physics, 47(9):092302, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i9/
p092302_s1.

Duan:2006:HDK

[748] Yi shi Duan, Shao feng Wu, and Peng ming Zhang. Higher-dimensional knotlike topological defects in local non-Abelian topological tensor currents. Journal of Mathematical Physics, 47(9):092303, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i9/
p092303_s1.

Braun:2006:BCG

[749] Volker Braun and Sakura Schäfer-Nameki. D-brane charges in Gepner models. Journal of Mathematical Physics, 47(9):092304, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p092304_s1.

Unal:2006:EGI

[750] Nuri Ünal. Electromagnetic and gravitational interactions of the spinning particle. Journal of Mathematical Physics, 47(9):092501, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p092501_s1.

Papadopoulos:2006:ECR

[751] G. O. Papadopoulos. On the essential constants in Riemannian geometries. Journal of Mathematical Physics, 47(9):092502, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p092502_s1.

Gravanis:2006:IHA

[752] Elias Gravanis and Steven Willison. Intersecting hypersurfaces in antide Sitter and Lovelock gravity. Journal of Mathematical Physics, 47 (9):092503, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i9/p092503_s1.

Szablikowski:2006:MLR

[753] Blazej M. Szablikowski and Maciej Blaszak. Meromorphic Lax representations of (1 + 1)-dimensional multi-Hamiltonian dispersionless systems. Journal of Mathematical Physics, 47(9):092701, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i9/
p092701_s1.

Tarasov:2006:MDS

[754] Vasily E. Tarasov. Map of discrete system into continuous. Journal of Mathematical Physics, 47(9):092901, September 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i9/p092901_s1.

Khare:2006:DWP

[755] Avinash Khare and Avadh Saxena. Domain wall and periodic solutions of coupled ϕ^{4} models in an external field. Journal of Mathematical Physics, 47(9):092902, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i9/p092902_s1.

Humi:2006:SSS

[756] Mayer Humi. Steady states of self-gravitating incompressible fluid in two dimensions. Journal of Mathematical Physics, 47(9):093101, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p093101_s1.

Vafeas:2006:SFE

[757] Panayiotis Vafeas and George Dassios. Stokes flow in ellipsoidal geometry. Journal of Mathematical Physics, 47(9):093102, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i9/
p093102_s1.

Kalnins:2006:SOSb

[758] E. G. Kalnins, J. M. Kress, and W. Miller. Second-order superintegrable systems in conformally flat spaces. V. Two- and three-dimensional quantum systems. Journal of Mathematical Physics, 47(9):093501, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i9/p093501_s1.

Atakishiyev:2006:MQO

[759] M. N. Atakishiyev, N. M. Atakishiyev, and A. U. Klimyk. On su_{q}(1, 1)- models of quantum oscillator. Journal of Mathematical Physics, 47 (9):093502, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i9/p093502_s1.

Michev:2006:FI

[760] Iordan P. Michev. On Fay identity. Journal of Mathematical Physics, 47(9):093503, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i9/p093503_s1.

Vacaru:2006:CFA

[761] Sergiu I. Vacaru. Clifford–Finsler algebroids and nonholonomic Einstein– Dirac structures. Journal of Mathematical Physics, 47(9):093504, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i9/p093504_s1.

Smirnov:2006:CSS

[762] Roman G. Smirnov and Pavel Winternitz. A class of superintegrable systems of Calogero type. Journal of Mathematical Physics, 47(9):093505, September 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i9/p093505_s1. See erratum [1118].

Stuchi:2006:NTB

[763] T. J. Stuchi, A. López-Castillo, and M. A. Almeida. Nonintegrability of the three-body problems for the classical helium atom. Journal of Mathematical Physics, 47(9):093506, September 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i9/p093506_s1.

Gonul:2006:ESD

[764] B. Gönül and M. Koçak. Explicit solutions for N-dimensional Schrödinger equations with position-dependent mass. Journal of Mathematical Physics, 47(10):102101, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i10/p102101_s1.

Odake:2006:UTA

[765] Satoru Odake and Ryu Sasaki. Unified theory of annihilation-creation operators for solvable (“discrete”) quantum mechanics. Journal of Mathematical Physics, 47(10):102102, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i10/p102102_s1.

Yuan:2006:NIP

[766] Li Yuan and Hong-Ke Du. A note on the infimum problem of Hilbert space effects. Journal of Mathematical Physics, 47(10):102103, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p102103_s1. See comment [1236].

DeLeo:2006:QDP

[767] Stefano De Leo and Gisele C. Ducati. Quaternionic diffusion by a potential step. Journal of Mathematical Physics, 47(10):102104, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p102104_s1.

Donnelly:2006:RCF

[768] Harold Donnelly. Resonance counting function in black box scattering. Journal of Mathematical Physics, 47(10):102105, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i10/
p102105_s1.

Hikami:2006:QIM

[769] Kazuhiro Hikami. Quantum invariants, modular forms, and lattice points II. Journal of Mathematical Physics, 47(10):102301, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i10/
p102301_s1.

Bowman:2006:CQL

[770] Daniel A. Bowman and John L. Challifour. Canonical quantization of lattice Higgs–Maxwell–Chern–Simons fields: Krein self-adjointness. Journal of Mathematical Physics, 47(10):102302, October 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i10/p102302_s1.

Bandelloni:2006:LCD

[771] G. Bandelloni and S. Lazzarini. Large chiral diffeomorphisms on Riemann surfaces and W -algebras. Journal of Mathematical Physics, 47(10): 102303, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i10/p102303_s1.

Noui:2006:TDL

[772] Karim Noui. Three-dimensional loop quantum gravity: Particles and the quantum double. Journal of Mathematical Physics, 47(10):102501, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p102501_s1.

Christodoulakis:2006:ACS

[773] T. Christodoulakis and Petros A. Terzis. Automorphisms and a cartography of the solution space for vacuum Bianchi cosmologies: The Type III case. Journal of Mathematical Physics, 47(10):102502, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i10/
p102502_s1.

Winklmeier:2006:SAR

[774] Monika Winklmeier and Osanobu Yamada. Spectral analysis of radial Dirac operators in the Kerr–Newman metric and its applications to timeperiodic solutions. Journal of Mathematical Physics, 47(10):102503, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p102503_s1.

Natorf:2006:SRT

[775] W. Natorf and J. Tafel. Symmetries of the Robinson–Trautman equation. Journal of Mathematical Physics, 47(10):102504, October 2006.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i10/
p102504_s1.

Bruschi:2006:GNS

[776] M. Bruschi and F. Calogero. Goldfishing: A new solvable many-body problem. Journal of Mathematical Physics, 47(10):102701, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i10/
p102701_s1.

Chen:2006:KFD

[777] Yu-Tung Chen and Ming-Hsien Tu. On kernel formulas and dispersionless Hirota equations of the extended dispersionless BKP hierarchy. Journal of Mathematical Physics, 47(10):102702, October 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i10/p102702_s1.

Pitkonen:2006:PDD

[778] Mikko Pitkonen. Polarizability of the dielectric double-sphere. Journal of Mathematical Physics, 47(10):102901, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i10/p102901_s1.

Chen:2006:CRM

[779] Yong Chen, Min-Ping Qian, and Jian-Sheng Xie. On characterization of reversible Markov processes by monotonicity of the fluctuation spectral density. Journal of Mathematical Physics, 47(10):103301, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p103301_s1.

Assis:2006:NDE

[780] P. C. Assis, P. C. da Silva, L. R. da Silva, E. K. Lenzi, and M. K. Lenzi. Nonlinear diffusion equation and nonlinear external force: Exact solution. Journal of Mathematical Physics, 47(10):103302, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i10/
p103302_s1.

Pastur:2006:LLL

[781] L. Pastur. Limiting laws of linear eigenvalue statistics for Hermitian matrix models. Journal of Mathematical Physics, 47(10):103303, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p103303_s1.

Lueck:2006:ECR

[782] T. Lueck, H.-J. Sommers, and M. R. Zirnbauer. Energy correlations for a random matrix model of disordered bosons. Journal of Mathematical Physics, 47(10):103304, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i10/p103304_s1.

Barra:2006:OFB

[783] Adriano Barra and Luca De Sanctis. Overlap fluctuations from the Boltzmann random overlap structure. Journal of Mathematical Physics, 47(10):103305, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i10/p103305_s1.

Dey:2006:GPM

[784] Rukmini Dey. Geometric prequantization of the moduli space of the vortex equations on a Riemann surface. Journal of Mathematical Physics, 47(10):103501, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i10/p103501_s1. See erratum [2141].

Li:2006:RSS

[785] Wei Li. Removable singularities for solutions of coupled Yang–Mills– Dirac equations. Journal of Mathematical Physics, 47(10):103502, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p103502_s1. See comment [1116].

Baleanu:2006:FHA

[786] Dumitru Baleanu, Sami I. Muslih, and Kenan Tas. Fractional Hamiltonian analysis of higher order derivatives systems. Journal of Mathematical Physics, 47(10):103503, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i10/p103503_s1.

Romera:2006:URF

[787] Elvira Romera, P. Sánchez-Moreno, and J. S. Dehesa. Uncertainty relation for Fisher information of D-dimensional single-particle systems with central potentials. Journal of Mathematical Physics, 47(10):103504, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p103504_s1.

Matsuno:2006:LSS

[788] Yoshimasa Matsuno. The Lyapunov stability of the N-soliton solutions in the Lax hierarchy of the Benjamin–Ono equation. Journal of Mathematical Physics, 47(10):103505, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i10/p103505_s1.

Langer:2006:VPE

[789] Matthias Langer and Christiane Tretter. Variational principles for eigenvalues of the Klein–Gordon equation. Journal of Mathematical Physics, 47(10):103506, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i10/p103506_s1.

Ferapontov:2006:CIH

[790] E. V. Ferapontov, K. R. Khusnutdinova, D. G. Marshall, and M. V. Pavlov. Classification of integrable Hamiltonian hydrodynamic chains associated with Kupershmidt’s brackets. Journal of Mathematical Physics, 47(10):103507, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i10/p103507_s1.

Aldridge:2006:EAI

[791] J. E. Aldridge, G. E. Prince, W. Sarlet, and G. Thompson. An EDS approach to the inverse problem in the calculus of variations. Journal of Mathematical Physics, 47(10):103508, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i10/p103508_s1.

Winternitz:2006:ISS

[792] Pavel Winternitz and Ismet Yurdusen. Integrable and superintegrable systems with spin. Journal of Mathematical Physics, 47(10):103509, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i10/p103509_s1.

Kasman:2006:WNP

[793] Alex Kasman and Stéphane Lafortune. When is negativity not a problem for the ultradiscrete limit? Journal of Mathematical Physics, 47(10): 103510, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i10/p103510_s1.

Dancer:2006:RQD

[794] K. A. Dancer, P. S. Isac, and J. Links. Representations of the quantum doubles of finite group algebras and spectral parameter dependent solutions of the Yang–Baxter equation. Journal of Mathematical Physics, 47(10):103511, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i10/p103511_s1.

Park:2006:IBT

[795] Choonkil Park. Isomorphisms between C^{∗}-ternary algebras. Journal of Mathematical Physics, 47(10):103512, October 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i10/p103512_s1.

Brunner:2006:BTI

[796] Nicolas Brunner, Valerio Scarani, and Nicolas Gisin. Bell-type inequalities for nonlocal resources. Journal of Mathematical Physics, 47 (11):112101, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i11/p112101_s1.

Bracci:2006:WAS

[797] L. Bracci and L. E. Picasso. On the Weyl algebras for systems with semibounded and bounded configuration space. Journal of Mathematical Physics, 47(11):112102, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i11/p112102_s1.

Borges:2006:SBS

[798] P. F. Borges, H. Boschi-Filho, and A. N. Vaidya. Scattering and bound state Green’s functions on a plane via so(2, 1) Lie algebra. Journal of Mathematical Physics, 47(11):112103, November 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i11/p112103_s1.

Bolte:2006:STF

[799] Jens Bolte and Hans-Michael Stiepan. The Selberg trace formula for Dirac operators. Journal of Mathematical Physics, 47(11):112104, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i11/p112104_s1.

Pankrashkin:2006:LEP

[800] Konstantin Pankrashkin. Localization effects in a periodic quantum graph with magnetic field and spin-orbit interaction. Journal of Mathematical Physics, 47(11):112105, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i11/p112105_s1.

Hislop:2006:BPC

[801] Peter D. Hislop and Olivier Lenoble. Basic properties of the currentcurrent correlation measure for random Schrödinger operators. Journal of Mathematical Physics, 47(11):112106, November 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i11/p112106_s1.

Camus:2006:EEE

[802] Brice Camus. Equilibrium and eigenfunctions estimates in the semiclassical regime. Journal of Mathematical Physics, 47(11):112107, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i11/p112107_s1.

Castro:2006:CSS

[803] Carlos Castro. On Chern–Simons (super) gravity, E_{8} Yang–Mills and polyvector-valued gauge theories in Clifford spaces. Journal of Mathematical Physics, 47(11):112301, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i11/p112301_s1.

Seidewitz:2006:FSP

[804] Ed Seidewitz. Foundations of a spacetime path formalism for relativistic quantum mechanics. Journal of Mathematical Physics, 47(11):112302, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i11/p112302_s1.

Miralles:2006:OOS

[805] David Miralles and José M. Pozo. Observers, observables, spinors, and the confusion of tongues. Journal of Mathematical Physics, 47 (11):112303, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i11/p112303_s1.

Sako:2006:DRS

[806] Akifumi Sako and Toshiya Suzuki. Dimensional reduction of Seiberg– Witten monopole equations, N = 2 noncommutative supersymmetric field theories and Young diagrams. Journal of Mathematical Physics, 47(11):112304, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i11/p112304_s1.

Ferrando:2006:TVS

[807] Joan Josep Ferrando and Juan Antonio Sáez. Type I vacuum solutions with aligned Papapetrou fields: An intrinsic characterization. Journal of Mathematical Physics, 47(11):112501, November 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i11/p112501_s1.

Shiromizu:2006:EMO

[808] Tetsuya Shiromizu, Sumio Yamada, and Hirotaka Yoshino. On existence of matter outside a static black hole. Journal of Mathematical Physics, 47(11):112502, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i11/p112502_s1.

Prikas:2006:DWS

[809] Athanasios Prikas. Domain walls on the surface of q-stars. Journal of Mathematical Physics, 47(11):112503, November 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i11/p112503_s1.

Poplawski:2006:PTE

[810] Nikodem J. Poplawski. Propagating torsion in the Einstein frame. Journal of Mathematical Physics, 47(11):112504, November 2006. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i11/
p112504_s1.

Giambo:2006:VAR

[811] Roberto Giambò, Fabio Giannoni, and Giulio Magli. Variational approach to Robertson–Walker spacetimes with homogeneous scalar fields. Journal of Mathematical Physics, 47(11):112505, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i11/
p112505_s1.

Qiao:2006:NIE

[812] Zhijun Qiao. A new integrable equation with cuspons and W/M-shape-peaks solitons. Journal of Mathematical Physics, 47(11):112701, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i11/p112701_s1. See comment [766].

Rawashdeh:2006:IPS

[813] M. Rawashdeh and G. Thompson. The inverse problem for sixdimensional codimension two nilradical Lie algebras. Journal of Mathematical Physics, 47(11):112901, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i11/p112901_s1.

Gerdt:2006:HRG

[814] V. Gerdt, R. Horan, A. Khvedelidze, M. Lavelle, D. McMullan, and Yu. Palii. On the Hamiltonian reduction of geodesic motion on SU(3) to SU(3)/SU(2). Journal of Mathematical Physics, 47(11):112902, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i11/p112902_s1.

Kondratiev:2006:NGT

[815] Yuri G. Kondratiev, Oleksandr V. Kutoviy, and Elena Zhizhina. Nonequilibrium Glauber-type dynamics in continuum. Journal of Mathematical Physics, 47(11):113501, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i11/p113501_s1.

Borisov:2006:DPA

[816] D. Borisov and P. Exner. Distant perturbation asymptotics in windowcoupled waveguides. I. The nonthreshold case. Journal of Mathematical Physics, 47(11):113502, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i11/p113502_s1.

Singer:2006:CBF

[817] R. Singer and M. Fähnle. Construction of basis functions for the spincluster expansion of the magnetic energy on the atomic scale in rotationally invariant systems. Journal of Mathematical Physics, 47(11):113503, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i11/p113503_s1.

Lievens:2006:HOC

[818] S. Lievens, N. I. Stoilova, and J. Van der Jeugt. Harmonic oscillators coupled by springs: Discrete solutions as a Wigner quantum system. Journal of Mathematical Physics, 47(11):113504, November 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i11/p113504_s1.

Bluman:2006:FNR

[819] George Bluman, Alexei F. Cheviakov, and Nataliya M. Ivanova. Framework for nonlocally related partial differential equation systems and nonlocal symmetries: Extension, simplification, and examples. Journal of Mathematical Physics, 47(11):113505, November 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i11/p113505_s1.

Heil:2006:GSB

[820] Christopher Heil and Alexander M. Powell. Gabor Schauder bases and the Balian–Low theorem. Journal of Mathematical Physics, 47 (11):113506, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i11/p113506_s1.

Martins:2006:NGT

[821] R. A. Dawe Martins. Noncommutative geometry, topology, and the standard model vacuum. Journal of Mathematical Physics, 47(11):113507, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i11/p113507_s1.

Wang:2006:SDC

[822] Jing Ping Wang. On the structure of (2 + 1)-dimensional commutative and noncommutative integrable equations. Journal of Mathematical Physics, 47(11):113508, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i11/p113508_s1.

Gough:2006:QSC

[823] John Gough. Quantum Stratonovich calculus and the quantum Wong– Zakai theorem. Journal of Mathematical Physics, 47(11):113509, November 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i11/p113509_s1.

Marlow:2006:BAG

[824] Thomas Marlow. Bayesian analog of Gleason’s theorem. Journal of Mathematical Physics, 47(12):122101, December 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i12/p122101_s1.

Nouicer:2006:PHP

[825] Khireddine Nouicer. Pauli–Hamiltonian in the presence of minimal lengths. Journal of Mathematical Physics, 47(12):122102, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i12/p122102_s1.

Herbut:2006:BPS

[826] Fedor Herbut. On bipartite pure-state entanglement structure in terms of disentanglement. Journal of Mathematical Physics, 47(12):122103, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i12/p122103_s1.

Organista:2006:SIT

[827] José Orlando Organista, Marek Nowakowski, and H. C. Rosu. Shape invariance through Crum transformation. Journal of Mathematical Physics, 47(12):122104, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i12/p122104_s1.

Unal:2006:PTC

[828] Nuri Ünal. Parametric-time coherent states for the generalized MIC– Kepler system. Journal of Mathematical Physics, 47(12):122105, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i12/p122105_s1.

Anastopoulos:2006:TAP

[829] Charis Anastopoulos and Ntina Savvidou. Time-of-arrival probabilities and quantum measurements. Journal of Mathematical Physics, 47(12):122106, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p122106_s1.

Gross:2006:HTF

[830] D. Gross. Hudson’s theorem for finite-dimensional quantum systems. Journal of Mathematical Physics, 47(12):122107, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i12/
p122107_s1.

Daoud:2006:FSH

[831] M. Daoud and M. R. Kibler. Fractional supersymmetry and hierarchy of shape invariant potentials. Journal of Mathematical Physics, 47(12):122108, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p122108_s1.

Cacciari:2006:PFR

[832] Ilaria Cacciari and Paolo Moretti. Propagator for finite range potentials. Journal of Mathematical Physics, 47(12):122109, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i12/
p122109_s1.

Schwartz:2006:RQW

[833] Charles Schwartz. Relativistic quaternionic wave equation. Journal of Mathematical Physics, 47(12):122301, December 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i12/p122301_s1. See

erratum [2213].

Mestre:2006:GLG

[834] Ângela Mestre and Robert Oeckl. Generating loop graphs via Hopf algebra in quantum field theory. Journal of Mathematical Physics, 47 (12):122302, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p122302_s1.

Svegstrup:2006:EHS

[835] Rolf Dyre Svegstrup. Endomorphisms on half-sided modular inclusions. Journal of Mathematical Physics, 47(12):122303, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i12/
p122303_s1.

Parvizi:2006:MRO

[836] Shahrokh Parvizi and Alireza Tavanfar. Minimal redefinition of the OSV ensemble. Journal of Mathematical Physics, 47(12):122304, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i12/p122304_s1.

Krisch:2006:FIL

[837] J. P. Krisch. Fractional Israel layers. Journal of Mathematical Physics, 47(12):122501, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p122501_s1.

Corbera:2006:IMP

[838] Montserrat Corbera and Jaume Llibre. Infinitely many periodic orbits for the rhomboidal five-body problem. Journal of Mathematical Physics, 47(12):122701, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p122701_s1.

Juang:2006:PBC

[839] Jonq Juang, Chin-Lung Li, and Jing-Wei Chang. Perturbed block circulant matrices and their application to the wavelet method of chaotic control. Journal of Mathematical Physics, 47(12):122702, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i12/
p122702_s1.

Berti:2006:EUR

[840] Alessia Berti. Existence and uniqueness for the reflection and transmission problem in stratified electromagnetic media. Journal of Mathematical Physics, 47(12):122901, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i12/p122901_s1.

Aharonovich:2006:GFW

[841] I. Aharonovich and L. P. Horwitz. Green functions for wave propagation on a five-dimensional manifold and the associated gauge fields generated by a uniformly moving point source. Journal of Mathematical Physics, 47(12):122902, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p122902_s1.

Abdesselam:2006:NCS

[842] B. Abdesselam and A. Chakrabarti. New class ô_{N} of statistical models: Transfer matrix eigenstates, chain Hamiltonians, factorizable S matrix. Journal of Mathematical Physics, 47(12):123301, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i12/
p123301_s1.

Bach:2006:LLT

[843] Volker Bach and Jacob Schach Møller. Localization at low temperature and infrared bounds. Journal of Mathematical Physics, 47(12):123302, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i12/p123302_s1.

Ueltschi:2006:FCB

[844] Daniel Ueltschi. Feynman cycles in the Bose gas. Journal of Mathematical Physics, 47(12):123303, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v47/i12/p123303_s1.

Hayashi:2006:FSP

[845] Nakao Hayashi and Pavel I. Naumkin. Final state problem for Korteweg– de Vries type equations. Journal of Mathematical Physics, 47(12):123501, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i12/p123501_s1.

Sbihi:2006:STN

[846] Mohammed Sbihi. Spectral theory of neutron transport semigroups with partly elastic collision operators. Journal of Mathematical Physics, 47(12):123502, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p123502_s1.

Holschneider:2006:ECO

[847] Matthias Holschneider and Gerd Teschke. Existence and computation of optimally localized coherent states. Journal of Mathematical Physics, 47(12):123503, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p123503_s1.

Botman:2006:GEC

[848] David M. Botman and William P. Joyce. Geometric equivalence of Clifford algebras. Journal of Mathematical Physics, 47(12):123504, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i12/p123504_s1.

Strauss:2006:ARS

[849] Y. Strauss, L. P. Horwitz, and A. Volovick. Approximate resonance states in the semigroup decomposition of resonance evolution. Journal of Mathematical Physics, 47(12):123505, December 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i12/p123505_s1.

deGosson:2006:ECZ

[850] Maurice A. de Gosson and Serge M. de Gosson. Extension of the Conley– Zehnder index, a product formula, and an application to the Weyl representation of metaplectic operators. Journal of Mathematical Physics, 47(12):123506, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p123506_s1.

Czaja:2006:UPG

[851] Wojciech Czaja and Jacek Zienkiewicz. Uncertainty principle for Gabor systems and the Zak transform. Journal of Mathematical Physics, 47(12):123507, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p123507_s1.

Cai:2006:NMS

[852] Jiaxiang Cai, Yushun Wang, Bin Wang, and Bin Jiang. New multisymplectic self-adjoint scheme and its composition scheme for the time-domain Maxwell’s equations. Journal of Mathematical Physics, 47(12):123508, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p123508_s1.

Wehefritz-Kaufmann:2006:CRI

[853] Birgit Wehefritz-Kaufmann. Clifford representations in integrable systems. Journal of Mathematical Physics, 47(12):123509, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i12/
p123509_s1.

Boiti:2006:STN

[854] M. Boiti, F. Pempinelli, and A. K. Pogrebkov. Scattering transform for nonstationary Schrödinger equation with bidimensionally perturbed N-soliton potential. Journal of Mathematical Physics, 47(12):123510, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i12/p123510_s1.

Aizawa:2006:UMR

[855] N. Aizawa, R. Chakrabarti, S. S. Naina Mohammed, and J. Segar. Universal T -matrix, representations of OSp_{q}(1/2) and little Q-Jacobi polynomials. Journal of Mathematical Physics, 47(12):123511, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v47/i12/p123511_s1.

Izaurieta:2006:ELS

[856] Fernando Izaurieta, Eduardo Rodríguez, and Patricio Salgado. Expanding Lie (super)algebras through Abelian semigroups. Journal of Mathematical Physics, 47(12):123512, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v47/i12/p123512_s1.

Bering:2006:NSA

[857] K. Bering. A note on semidensities in antisymplectic geometry. Journal of Mathematical Physics, 47(12):123513, December 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i12/p123513_s1.

Pavlov:2006:HCC

[858] Maxim V. Pavlov. Hydrodynamic chains and the classification of their Poisson brackets. Journal of Mathematical Physics, 47(12):123514, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v47/i12/p123514_s1.

Nesterenko:2006:CLD

[859] Maryna Nesterenko and Roman Popovych. Contractions of lowdimensional Lie algebras. Journal of Mathematical Physics, 47(12): 123515, December 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i12/p123515_s1.

Derdzinski:2006:PNW

[860] Andrzej Derdzinski and Witold Roter. Publisher’s note: “Walker’s theorem without coordinates”, [J. Math. Phys. 47, 062504 (2006)]. Journal of Mathematical Physics, 47(12):129901, December 2006. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v47/i12/p129901_s1. See

[647].

Carley:2007:SCL

[861] Holly Carley. Strong-coupling limit for the ground state of a particle harmonic oscillator interaction. Journal of Mathematical Physics, 48(1): 012101, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i1/p012101_s1.

Akhtarshenas:2007:CCP

[862] S. J. Akhtarshenas. Canonical coset parametrization and the Bures metric of the three-level quantum systems. Journal of Mathematical Physics, 48(1):012102, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p012102_s1.

Derezinski:2007:EWC

[863] Jan Dereziński and Wojciech De Roeck. Extended weak coupling limit for Friedrichs Hamiltonians. Journal of Mathematical Physics, 48(1): 012103, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i1/p012103_s1.

Amosov:2007:WCB

[864] Grigori G. Amosov. On Weyl channels being covariant with respect to the maximum commutative group of unitaries. Journal of Mathematical Physics, 48(1):012104, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p012104_s1.

Charest:2007:QVS

[865] F. Charest, C. Hudon, and P. Winternitz. Quasiseparation of variables in the Schrödinger equation with a magnetic field. Journal of Mathematical Physics, 48(1):012105, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p012105_s1.

Blanchard:2007:QDS

[866] Ph. Blanchard, M. Hellmich, P. Lugiewicz, and R. Olkiewicz. Quantum dynamical semigroups for finite and infinite Bose systems. Journal of Mathematical Physics, 48(1):012106, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i1/p012106_s1.

Petz:2007:CRT

[867] Dénes Petz and Jonas Kahn. Complementary reductions for two qubits. Journal of Mathematical Physics, 48(1):012107, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i1/p012107_s1.

Gour:2007:DMI

[868] Gilad Gour, Somshubhro Bandyopadhyay, and Barry C. Sanders. Dual monogamy inequality for entanglement. Journal of Mathematical Physics, 48(1):012108, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p012108_s1.

Dias:2007:FMT

[869] Nuno Costa Dias and João Nuno Prata. Features of Moyal trajectories. Journal of Mathematical Physics, 48(1):012109, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i1/p012109_s1.

Sbarba:2007:DDE

[870] L. Dello Sbarba and V. Hussin. Degenerate discrete energy spectra and associated coherent states. Journal of Mathematical Physics, 48(1): 012110, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i1/p012110_s1.

Zoli:2007:FSE

[871] Marco Zoli. Finite size effects in bistable ϕ^{4} models. Journal of Mathematical Physics, 48(1):012111, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i1/p012111_s1.

Almeida:2007:HGC

[872] José B. Almeida. Hidden geometric character of relativistic quantum mechanics. Journal of Mathematical Physics, 48(1):012301, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i1/p012301_s1.

Trzetrzelewski:2007:LBT

[873] Maciej Trzetrzelewski. Large N behavior of two dimensional supersymmetric Yang–Mills quantum mechanics. Journal of Mathematical Physics, 48(1):012302, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p012302_s1.

Barrett:2007:LVN

[874] John W. Barrett. Lorentzian version of the noncommutative geometry of the standard model of particle physics. Journal of Mathematical Physics, 48(1):012303, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p012303_s1.

Giribet:2007:NSK

[875] Gaston E. Giribet. Note on Z_{2} symmetries of the Knizhnik– Zamolodchikov equation. Journal of Mathematical Physics, 48(1): 012304, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i1/p012304_s1.

Harte:2007:MSC

[876] Abraham I. Harte. Multipole structure of current vectors in curved space–time. Journal of Mathematical Physics, 48(1):012501, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i1/p012501_s1.

Antippa:2007:IDS

[877] Adel F. Antippa and Daniel M. Dubois. Incursive discretization, system bifurcation, and energy conservation. Journal of Mathematical Physics, 48(1):012701, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p012701_s1.

DeLuca:2007:GIE

[878] Jayme De Luca. Geometric integration of the electromagnetic two-body problem. Journal of Mathematical Physics, 48(1):012702, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i1/
p012702_s1.

Zhang:2007:DBT

[879] Ruifeng Zhang and Boling Guo. Dynamical behavior for the threedimensional generalized Hasegawa–Mima equations. Journal of Mathematical Physics, 48(1):012703, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i1/p012703_s1.

Cabrerizo:2007:MPT

[880] Jose L. Cabrerizo, Manuel Fernandez, and Miguel Ortega. Massless particles in three-dimensional Lorentzian warped products. Journal of Mathematical Physics, 48(1):012901, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i1/p012901_s1.

Marquette:2007:PPA

[881] Ian Marquette and Pavel Winternitz. Polynomial Poisson algebras for classical superintegrable systems with a third-order integral of motion. Journal of Mathematical Physics, 48(1):012902, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i1/

p012902_s1. See erratum [1329].

Mukhamedov:2007:CAT

[882] Farrukh Mukhamedov, Utkir Rozikov, and José F. F. Mendes. On contour arguments for the three state Potts model with competing interactions on a semi-infinite Cayley tree. Journal of Mathematical Physics, 48(1):013301, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p013301_s1.

Tautz:2007:NKI

[883] R. C. Tautz, R. Schlickeiser, and I. Lerche. Nonresonant kinetic instabilities of a relativistic plasma in a uniform magnetic field: Longitudinal and transverse mode coupling effects. Journal of Mathematical Physics, 48(1): 013302, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i1/p013302_s1.

Wang:2007:DBR

[884] Hongyun Wang and Hong Qian. On detailed balance and reversibility of semi-Markov processes and single-molecule enzyme kinetics. Journal of Mathematical Physics, 48(1):013303, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i1/p013303_s1.

DeRoeck:2007:TCS

[885] W. De Roeck and E. L. Lakshtanov. Total cross section exceeds transport cross section for quantum scattering from hard bodies at low and high wave numbers. Journal of Mathematical Physics, 48(1):013501, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i1/p013501_s1.

Fokas:2007:GDN

[886] A. S. Fokas and B. Pelloni. Generalized Dirichlet to Neumann map for moving initial-boundary value problems. Journal of Mathematical Physics, 48(1):013502, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p013502_s1.

Figueiredo:2007:GSW

[887] B. D. Bonorino Figueiredo. Generalized spheroidal wave equation and limiting cases. Journal of Mathematical Physics, 48(1):013503, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i1/p013503_s1. See erratum [987].

Chen:2007:DEE

[888] Huaitang Chen and Huicheng Yin. Double elliptic equation method and new exact solutions of the (n + 1)-dimensional sinh-Gordon equation. Journal of Mathematical Physics, 48(1):013504, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i1/p013504_s1.

Tek:2007:MKV

[889] Süleyman Tek. Modified Korteweg–de Vries surfaces. Journal of Mathematical Physics, 48(1):013505, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i1/p013505_s1. See erratum [961].

Quevedo:2007:G

[890] Hernando Quevedo. Geometrothermodynamics. Journal of Mathematical Physics, 48(1):013506, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p013506_s1.

Ieda:2007:ISM

[891] Jun’ichi Ieda, Masaru Uchiyama, and Miki Wadati. Inverse scattering method for square matrix nonlinear Schrödinger equation under nonvanishing boundary conditions. Journal of Mathematical Physics, 48(1): 013507, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i1/p013507_s1.

Kengne:2007:SSE

[892] E. Kengne, Boris A. Malomed, S. T. Chui, and W. M. Liu. Solitary signals in electrical nonlinear transmission line. Journal of Mathematical Physics, 48(1):013508, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p013508_s1.

Chen:2007:CMT

[893] Bang-Yen Chen and Franki Dillen. Classification of marginally trapped Lagrangian surfaces in Lorentzian complex space forms. Journal of Mathematical Physics, 48(1):013509, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i1/p013509_s1. See erratum [1499].

Zhou:2007:NSP

[894] Ruguang Zhou. Nonlinearizations of spectral problems of the nonlinear Schrödinger equation and the real-valued modified Korteweg–de Vries equation. Journal of Mathematical Physics, 48(1):013510, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i1/
p013510_s1.

Bagarello:2007:BVB

[895] F. Bagarello. Bounded version of bosonic creation and annihilation operators and their related quasicoherent states. Journal of Mathematical Physics, 48(1):013511, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p013511_s1.

Khrennikov:2007:QAT

[896] Andrei Khrennikov. To quantum averages through asymptotic expansion of classical averages on infinite-dimensional space. Journal of Mathematical Physics, 48(1):013512, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i1/p013512_s1.

Doliwa:2007:ILTa

[897] A. Doliwa, P. Grinevich, M. Nieszporski, and P. M. Santini. Integrable lattices and their sublattices: From the discrete Moutard (discrete Cauchy–Riemann) 4-point equation to the self-adjoint 5-point scheme. Journal of Mathematical Physics, 48(1):013513, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i1/p013513_s1.

Nucci:2007:FSP

[898] M. C. Nucci and P. G. L. Leach. Fuchs’ solution of Painlevé VI equation by means of Jacobi’s last multiplier. Journal of Mathematical Physics, 48(1):013514, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i1/p013514_s1.

Yan:2007:CAI

[899] Zhenya Yan. Comment on “Analytical integrability and physical solutions of d-KdV equation”, [J. Math. Phys. 47, 032901 (2006)]. Journal of Mathematical Physics, 48(1):014101, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i1/p014101_s1. See [549].

Kachmar:2007:EGS

[900] Ayman Kachmar. Erratum: “On the ground state energy for a magnetic Schrödinger operator and the effect of the DeGennes boundary condition”, [J. Math. Phys. 47, 072106 (2006)]. Journal of Mathematical Physics, 48(1):019901, January 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i1/p019901_s1. See [674].

Bruneau:2007:BHO

[901] Laurent Bruneau and Jan Dereziński. Bogoliubov Hamiltonians and oneparameter groups of Bogoliubov transformations. Journal of Mathematical Physics, 48(2):022101, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i2/p022101_s1.

Beneduci:2007:NOS

[902] Roberto Beneduci. Neumark operators and sharp reconstructions: The finite dimensional case. Journal of Mathematical Physics, 48(2):022102, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i2/p022102_s1.

Gerstenhaber:2007:LUP

[903] Murray Gerstenhaber. Least uncertainty principle in deformation quantization. Journal of Mathematical Physics, 48(2):022103, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i2/
p022103_s1.

Munger:2007:IBS

[904] Charles Thomas Munger. Ideal basis sets for the Dirac Coulomb problem: Eigenvalue bounds and convergence proofs. Journal of Mathematical Physics, 48(2):022301, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i2/p022301_s1.

Mastropietro:2007:NAB

[905] Vieri Mastropietro. Nonperturbative Adler–Bardeen theorem. Journal of Mathematical Physics, 48(2):022302, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i2/p022302_s1.

Redington:2007:SFD

[906] N. Redington and M. A. K. Lodhi. Simple five-dimensional wave equation for a Dirac particle. Journal of Mathematical Physics, 48(2):022303, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i2/p022303_s1.

Sen:2007:LRH

[907] Dipak K. Sen. Left- and right-handed neutrinos and Baryon–Lepton masses. Journal of Mathematical Physics, 48(2):022304, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i2/
p022304_s1.

Adam:2007:SSS

[908] C. Adam, J. Sánchez-Guillén, and A. Wereszczyński. Soliton stability in some knot soliton models. Journal of Mathematical Physics, 48(2):022305, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i2/p022305_s1.

Kaura:2007:SPF

[909] Payal Kaura, Aalok Misra, and Pramod Shukla. Super Picard–Fuchs equation and monodromies for supermanifolds. Journal of Mathematical Physics, 48(2):022306, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i2/p022306_s1.

Massa:2007:VTG

[910] Enrico Massa and Stefano Vignolo. Variational techniques in general relativity: A metric-affine approach to Kaluza’s theory. Journal of Mathematical Physics, 48(2):022501, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i2/p022501_s1.

Batic:2007:SMD

[911] D. Batic. Scattering for massive Dirac fields on the Kerr metric. Journal of Mathematical Physics, 48(2):022502, February 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i2/p022502_s1.

Hardin:2007:SLE

[912] D. P. Hardin, E. B. Saff, and H. Stahl. Support of the logarithmic equilibrium measure on sets of revolution in R^{3}. Journal of Mathematical Physics, 48(2):022901, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i2/p022901_s1.

Buitrago:2007:SEL

[913] J. Buitrago and S. Hajjawi. Spinor extended Lorentz-force-like equations as consequence of a spinorial structure of space–time. Journal of Mathematical Physics, 48(2):022902, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i2/p022902_s1.

Hernandez-Bermejo:2007:NSF

[914] Benito Hernández-Bermejo. New solution family of the Jacobi equations: Characterization, invariants, and global Darboux analysis. Journal of Mathematical Physics, 48(2):022903, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i2/p022903_s1.

Tamura:2007:RPF

[915] Hiroshi Tamura and Keiichi R. Ito. Random point fields for paraparticles of any order. Journal of Mathematical Physics, 48(2):023301, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i2/p023301_s1.

Giuliani:2007:GSE

[916] Alessandro Giuliani. Ground state energy of the low density Hubbard model: An upper bound. Journal of Mathematical Physics, 48(2):023302, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i2/p023302_s1.

Koma:2007:SGD

[917] Tohru Koma. Spectral gap and decay of correlations in U(1)-symmetric lattice systems in dimensions D < 2. Journal of Mathematical Physics, 48(2):023303, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i2/p023303_s1.

Chevalier:2007:FDT

[918] C. Chevalier and F. Debbasch. Fluctuation-dissipation theorems in an expanding universe. Journal of Mathematical Physics, 48(2):023304, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i2/p023304_s1.

Romano:2007:ASQ

[919] V. Romano, M. Torrisi, and R. Tracinà. Approximate solutions to the quantum drift-diffusion model of semiconductors. Journal of Mathematical Physics, 48(2):023501, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i2/p023501_s1.

Ye:2007:WQB

[920] Lixia Ye and Fang Li. Weak quantum Borcherds superalgebras and their representations. Journal of Mathematical Physics, 48(2):023502, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i2/p023502_s1.

Kalnins:2007:EQS

[921] E. G. Kalnins, W. Miller, and G. S. Pogosyan. Exact and quasiexact solvability of second order superintegrable quantum systems. II. Relation to separation of variables. Journal of Mathematical Physics, 48(2):023503, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i2/p023503_s1.

Zhao:2007:SVM

[922] Shao-You Zhao and Yao-Zhong Zhang. Supersymmetric vertex models with domain wall boundary conditions. Journal of Mathematical Physics, 48(2):023504, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i2/p023504_s1.

Nakamula:2007:VAC

[923] Atsushi Nakamula. Variant of ADHMN construction associated with q-analysis. Journal of Mathematical Physics, 48(2):023505, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i2/p023505_s1.

Skrypnyk:2007:QIS

[924] T. Skrypnyk. Quantum integrable systems, non-skew-symmetric rmatrices and algebraic Bethe ansatz. Journal of Mathematical Physics, 48(2):023506, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i2/p023506_s1.

Dolivet:2007:CSM

[925] Yacine Dolivet and Miguel Tierz. Chern–Simons matrix models and Stieltjes–Wigert polynomials. Journal of Mathematical Physics, 48 (2):023507, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i2/p023507_s1.

Golinski:2007:SSM

[926] Tomasz Goliński, Maciej Horowski, Anatol Odzijewicz, and Aneta Slizewska. sl(2, R) symmetry and solvable multiboson systems. Journal of Mathematical Physics, 48(2):023508, February 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i2/p023508_s1.

Dito:2007:SMP

[927] Giuseppe Dito and Rémi Léandre. Stochastic Moyal product on the Wiener space. Journal of Mathematical Physics, 48(2):023509, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i2/p023509_s1.

Fakhri:2007:NLO

[928] H. Fakhri, A. Dehghani, and A. Jafari. New ladder operators for the monopole harmonics. Journal of Mathematical Physics, 48(2):023510, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i2/p023510_s1.

Crampe:2007:QTY

[929] Nicolas Crampé and Anastasia Doikou. (quantum) twisted Yangians: Symmetry, Baxterisation, and centralizers. Journal of Mathematical Physics, 48(2):023511, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i2/p023511_s1.

Ramm:2007:MBWa

[930] A. G. Ramm. Many-body wave scattering by small bodies. Journal of Mathematical Physics, 48(2):023512, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i2/p023512_s1.

Stephan:2007:MNA

[931] Christoph A. Stephan. Massive neutrinos in almost-commutative geometry. Journal of Mathematical Physics, 48(2):023513, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i2/
p023513_s1.

Hestenes:2007:CSG

[932] David Hestenes and Jeremy W. Holt. Crystallographic space groups in geometric algebra. Journal of Mathematical Physics, 48(2):023514, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i2/p023514_s1.

Gordoa:2007:NDN

[933] P. R. Gordoa, A. Pickering, and Z. N. Zhu. New 2 + 1 dimensional nonisospectral Toda lattice hierarchy. Journal of Mathematical Physics, 48(2):023515, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i2/p023515_s1.

Torre:2007:EHR

[934] C. G. Torre. Erratum: “The helically-reduced wave equation as a symmetric-positive system”, [J. Math. Phys. 44, 6223 (2003)]. Journal of Mathematical Physics, 48(2):029901, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i2/p029901_s1. See [4].

Aschbacher:2007:TPQ

[935] W. Aschbacher, V. Jaksić, Y. Pautrat, and C.-A. Pillet. Transport properties of quasi-free fermions. Journal of Mathematical Physics, 48(3): 032101, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p032101_s1.

Benamira:2007:EPI

[936] F. Benamira, L. Guechi, S. Mameri, and M. A. Sadoun. Exact path integral treatment of a diatomic molecule potential. Journal of Mathematical Physics, 48(3):032102, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i3/p032102_s1.

Albeverio:2007:CSL

[937] Sergio Albeverio, Claudio Cacciapuoti, and Domenico Finco. Coupling in the singular limit of thin quantum waveguides. Journal of Mathematical Physics, 48(3):032103, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i3/p032103_s1.

Braun:2007:SCSa

[938] Carol Braun and Anupam Garg. Semiclassical coherent-state propagator for many particles. Journal of Mathematical Physics, 48(3): 032104, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p032104_s1.

Meng:2007:MKP

[939] Guowu Meng. MICZ-kepler problems in all dimensions. Journal of Mathematical Physics, 48(3):032105, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i3/p032105_s1.

Anastopoulos:2007:QPT

[940] Charis Anastopoulos and Ntina Savvidou. Quantum probabilities for time-extended alternatives. Journal of Mathematical Physics, 48(3): 032106, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p032106_s1.

Geloun:2007:DVC

[941] Joseph Ben Geloun, Jan Govaerts, and M. Norbert Hounkonnou. (p, q) deformations and (p, q)-vector coherent states of the Jaynes–Cummings model in the rotating wave approximation. Journal of Mathematical Physics, 48(3):032107, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i3/p032107_s1.

Cirilo-Lombardo:2007:MSH

[942] Diego Julio Cirilo-Lombardo. On the mathematical structure and hidden symmetries of the Born–Infeld field equations. Journal of Mathematical Physics, 48(3):032301, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i3/p032301_s1.

Adam:2007:CLS

[943] C. Adam, J. Sánchez-Guillén, and A. Wereszczyński. Conservation laws in Skyrme-type models. Journal of Mathematical Physics, 48(3): 032302, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p032302_s1.

Feigin:2007:KLD

[944] B. L. Feigin, A. M. Gainutdinov, A. M. Semikhatov, and I. Yu. Tipunin. Kazhdan–Lusztig-dual quantum group for logarithmic extensions of Virasoro minimal models. Journal of Mathematical Physics, 48(3):032303, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p032303_s1.

Major:2007:RCT

[945] Seth Major, David Rideout, and Sumati Surya. On recovering continuum topology from a causal set. Journal of Mathematical Physics, 48(3): 032501, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p032501_s1.

Chandrasekar:2007:LHD

[946] V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan. On the Lagrangian and Hamiltonian description of the damped linear harmonic oscillator. Journal of Mathematical Physics, 48(3):032701, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i3/
p032701_s1.

Zhao:2007:CUA

[947] Caidi Zhao and Shengfan Zhou. L^{2}-compact uniform attractors for a nonautonomous incompressible non-Newtonian fluid with locally uniformly integrable external forces in distribution space. Journal of Mathematical Physics, 48(3):032702, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i3/p032702_s1.

Fiorani:2007:GAA

[948] E. Fiorani and G. Sardanashvily. Global action-angle coordinates for completely integrable systems with noncompact invariant submanifolds. Journal of Mathematical Physics, 48(3):032901, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i3/p032901_s1.

Vankerschaver:2007:EPR

[949] Joris Vankerschaver. Euler–Poincaré reduction for discrete field theories. Journal of Mathematical Physics, 48(3):032902, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i3/p032902_s1.

Bibbona:2007:CVV

[950] Enrico Bibbona, Lorenzo Fatibene, and Mauro Francaviglia. Chetaev versus vakonomic prescriptions in constrained field theories with parametrized variational calculus. Journal of Mathematical Physics, 48(3):032903, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i3/p032903_s1.

Giardina:2007:DEC

[951] Cristian Giardinà, Jorge Kurchan, and Frank Redig. Duality and exact correlations for a model of heat conduction. Journal of Mathematical Physics, 48(3):033301, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i3/p033301_s1.

Nenciu:2007:IEM

[952] Gheorghe Nenciu. Independent electron model for open quantum systems: Landauer–Büttiker formula and strict positivity of the entropy production. Journal of Mathematical Physics, 48(3):033302, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i3/
p033302_s1.

Refregier:2007:SCT

[953] Philippe Réfrégier. Symmetries in coherence theory of partially polarized light. Journal of Mathematical Physics, 48(3):033303, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i3/p033303_s1.

McPhedran:2007:SIT

[954] R. C. McPhedran, L. C. Botten, N. A. Nicorovici, and I. J. Zucker. Systematic investigation of two-dimensional static array sums. Journal of Mathematical Physics, 48(3):033501, March 2007. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i3/p033501_s1.

Andruskiewitsch:2007:PHA

[955] Nicolás Andruskiewitsch and Fernando Fantino. On pointed Hopf algebras associated with unmixed conjugacy classes in s_{m}. Journal of Mathematical Physics, 48(3):033502, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i3/p033502_s1.

Guil:2007:ESL

[956] Francisco Guil. Electronic states on a lattice in periodic magnetic fields. Journal of Mathematical Physics, 48(3):033503, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i3/p033503_s1.

Cresson:2007:FED

[957] Jacky Cresson. Fractional embedding of differential operators and Lagrangian systems. Journal of Mathematical Physics, 48(3):033504, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i3/p033504_s1.

Guerra:2007:MMP

[958] Ediel Guerra, Jacqueline Rojas, and Ramón Mendoza. A mathematical model for the partition function. Journal of Mathematical Physics, 48(3): 033505, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p033505_s1.

Bolorizadeh:2007:PFT

[959] M. A. Bolorizadeh, M. J. Brunger, T. Maddern, and E. Ghanbari Adivi. Positronium formation as a three-body reaction. I. The secondorder positron-electron interaction amplitude. Journal of Mathematical Physics, 48(3):033506, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i3/p033506_s1.

Llibre:2007:GAF

[960] Jaume Llibre and Clàudia Valls. Global analytic first integrals for the real planar Lotka–Volterra system. Journal of Mathematical Physics, 48(3):

033507, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p033507_s1.

Tek:2007:EMK

[961] Süleyman Tek. Erratum: “Modified Korteweg–de Vries surfaces”, [J. Math. Phys. 48, 013505 (2007)]. Journal of Mathematical Physics, 48(3): 039901, March 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i3/p039901_s1. See [889].

Adachi:2007:MIS

[962] Tadayoshi Adachi and Katsuhiro Maehara. On multidimensional inverse scattering for Stark Hamiltonians. Journal of Mathematical Physics, 48(4):042101, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i4/p042101_s1.

Nakano:2007:FVA

[963] Fumihiko Nakano. Finite volume approximation of the Anderson model. Journal of Mathematical Physics, 48(4):042102, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i4/p042102_s1.

Cacciari:2007:PFR

[964] Ilaria Cacciari and Paolo Moretti. Propagator for finite range potentials: The case of reflection. Journal of Mathematical Physics, 48(4): 042103, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i4/p042103_s1.

Correggi:2007:RRBa

[965] M. Correggi, T. Rindler-Daller, and J. Yngvason. Rapidly rotating Bose– Einstein condensates in strongly anharmonic traps. Journal of Mathematical Physics, 48(4):042104, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i4/p042104_s1.

Xia:2007:GWR

[966] Qiaoling Xia. Generalized Weierstrass representations of surfaces with the constant Gauss curvature in pseudo-Riemannian three-dimensional space forms. Journal of Mathematical Physics, 48(4):042301, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i4/
p042301_s1.

Komathiraj:2007:TSS

[967] K. Komathiraj and S. D. Maharaj. Tikekar superdense stars in electric fields. Journal of Mathematical Physics, 48(4):042501, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i4/
p042501_s1.

Batic:2007:HET

[968] D. Batic and H. Schmid. Heun equation, Teukolsky equation, and type-D metrics. Journal of Mathematical Physics, 48(4):042502, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i4/
p042502_s1.

Khasin:2007:RRR

[969] Michael Khasin and Lazar Friedland. Removal of resonances by rotation in linearly degenerate two-dimensional oscillator systems. Journal of Mathematical Physics, 48(4):042701, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i4/p042701_s1.

Sergyeyev:2007:SSC

[970] Artur Sergyeyev and Dmitry Demskoi. Sasa–Satsuma (complex modified Korteweg–de Vries II) and the complex sine-Gordon II equation revisited: Recursion operators, nonlocal symmetries, and more. Journal of Mathematical Physics, 48(4):042702, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i4/p042702_s1.

Bender:2007:SBC

[971] Carl M. Bender and Daniel W. Darg. Spontaneous breaking of classical PT symmetry. Journal of Mathematical Physics, 48(4):042703, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i4/p042703_s1.

Birtea:2007:EEM

[972] P. Birtea and M. Puta. Equivalence of energy methods in stability theory. Journal of Mathematical Physics, 48(4):042704, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i4/p042704_s1.

Hamza:2007:LEU

[973] Eman Hamza and Günter Stolz. Lyapunov exponents for unitary Anderson models. Journal of Mathematical Physics, 48(4):043301, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i4/p043301_s1.

Khare:2007:DWP

[974] Avinash Khare and Avadh Saxena. Domain wall and periodic solutions of coupled asymmetric double well models. Journal of Mathematical Physics, 48(4):043302, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i4/p043302_s1.

Gottschalk:2007:HDL

[975] Hanno Gottschalk and Boubaker Smii. How to determine the law of the solution to a stochastic partial differential equation driven by a Lévy space–time noise? Journal of Mathematical Physics, 48(4):043303, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i4/p043303_s1.

Albeverio:2007:RRD

[976] S. Albeverio, R. Hryniv, and Ya. Mykytyuk. Reconstruction of radial Dirac operators. Journal of Mathematical Physics, 48(4):043501, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i4/p043501_s1.

Wang:2007:GFS

[977] Shaowei Wang and Mingyu Xu. Generalized fractional Schrödinger equation with space–time fractional derivatives. Journal of Mathematical Physics, 48(4):043502, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i4/p043502_s1.

Dehesa:2007:ITM

[978] J. S. Dehesa, S. López-Rosa, and R. J. Yáñez. Information-theoretic measures of hyperspherical harmonics. Journal of Mathematical Physics, 48(4):043503, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i4/p043503_s1.

Stoilova:2007:CGQ

[979] N. I. Stoilova and J. Van der Jeugt. Classification of generalized quantum statistics associated with the exceptional Lie (super)algebras. Journal of Mathematical Physics, 48(4):043504, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i4/p043504_s1.

Bagarello:2007:IAO

[980] F. Bagarello and S. Triolo. Invariant analytic orthonormalization procedure with an application to coherent states. Journal of Mathematical Physics, 48(4):043505, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i4/p043505_s1.

Gomez:2007:NST

[981] Fernando Gómez. Nonunitary similarity transformation of conservative to dissipative evolutions: Intertwining without time operator. Journal of Mathematical Physics, 48(4):043506, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i4/p043506_s1.

Cvijovic:2007:CFS

[982] Djurdje Cvijović and H. M. Srivastava. Closed-form summation of the Dowker and related sums. Journal of Mathematical Physics, 48(4): 043507, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i4/p043507_s1.

Fu:2007:CMI

[983] Jiayuan Fu, Qifen Jiang, and Yucai Su. Classification of modules of the intermediate series over Ramond N = 2 superconformal algebras. Journal of Mathematical Physics, 48(4):043508, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i4/p043508_s1.

Yamada:2007:SRS

[984] Daisuke Yamada. Scattering rule in soliton cellular automaton associated with crystal base of U_{q}(D_{4}^{(3)}). Journal of Mathematical Physics, 48(4):

043509, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i4/p043509_s1.

Bershtein:2007:ASR

[985] Olga Bershtein. On a q-analog of a Sahi result. Journal of Mathematical Physics, 48(4):043510, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i4/p043510_s1.

Coquereaux:2007:ODM

[986] R. Coquereaux and G. Schieber. Orders and dimensions for sl(2) or sl(3) module categories and boundary conformal field theories on a torus. Journal of Mathematical Physics, 48(4):043511, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i4/p043511_s1.

Figueiredo:2007:EGS

[987] B. D. Bonorino Figueiredo. Erratum: “Generalized spheroidal wave equation and limiting cases”, [J. Math. Phys. 48, 013503 (2007)]. Journal of Mathematical Physics, 48(4):049901, April 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i4/p049901_s1. See [887].

Jurciukonis:2007:QSM

[988] D. Jurciukonis and E. Norvaisas. Quantum SU(3) Skyrme model with noncanonical embedded SO(3) soliton. Journal of Mathematical Physics, 48(5):052101, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p052101_s1.

Viennot:2007:GQA

[989] David Viennot. Geometry of quantum active subspaces and of effective Hamiltonians. Journal of Mathematical Physics, 48(5):052102, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p052102_s1.

Benatti:2007:ERC

[990] Fabio Benatti, Roberto Floreanini, and Alexandra M. Liguori. Extended reduction criterion and lattice states. Journal of Mathematical Physics, 48(5):052103, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p052103_s1.

Gross:2007:EDU

[991] D. Gross, K. Audenaert, and J. Eisert. Evenly distributed unitaries: On the structure of unitary designs. Journal of Mathematical Physics, 48(5): 052104, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052104_s1.

Bombin:2007:HEC

[992] H. Bombin and M. A. Martin-Delgado. Homological error correction: Classical and quantum codes. Journal of Mathematical Physics, 48(5): 052105, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052105_s1.

Bengtsson:2007:MUB

[993] Ingemar Bengtsson, Wojciech Bruzda, Åsa Ericsson, Jan-Åke Larsson, Wojciech Tadej, and Karol Zyczkowski. Mutually unbiased bases and Hadamard matrices of order six. Journal of Mathematical Physics, 48 (5):052106, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052106_s1.

Krivoruchenko:2007:WSH

[994] M. I. Krivoruchenko and Amand Faessler. Weyl’s symbols of Heisenberg operators of canonical coordinates and momenta as quantum characteristics. Journal of Mathematical Physics, 48(5):052107, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p052107_s1.

Slawny:2007:PLO

[995] J. Slawny. Position and length operators in a theory with minimal length. Journal of Mathematical Physics, 48(5):052108, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p052108_s1.

Pavlyukh:2007:ASH

[996] Y. Pavlyukh and W. Hübner. Analytic solution of Hedin’s equations in zero dimensions. Journal of Mathematical Physics, 48(5):052109, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p052109_s1.

Cappellini:2007:SST

[997] Valerio Cappellini, Hans-Jürgen Sommers, and Karol Zyczkowski. Subnormalized states and trace-nonincreasing maps. Journal of Mathematical Physics, 48(5):052110, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i5/p052110_s1.

DeLeo:2007:QWP

[998] Stefano De Leo and Gisele C. Ducati. Quaternionic wave packets. Journal of Mathematical Physics, 48(5):052111, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i5/p052111_s1.

Lisovyy:2007:ABE

[999] Oleg Lisovyy. Aharonov–Bohm effect on the Poincaré disk. Journal of Mathematical Physics, 48(5):052112, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i5/p052112_s1.

Molinari:2007:APG

[1000] Luca Guido Molinari. Another proof of Gell-Mann and Low’s theorem. Journal of Mathematical Physics, 48(5):052113, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p052113_s1.

Sergyeyev:2007:ESS

[1001] Artur Sergyeyev. Exact solvability of superintegrable Benenti systems. Journal of Mathematical Physics, 48(5):052114, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p052114_s1.

Reyes:2007:GPD

[1002] Marat C. Reyes. Generalized path dependent representations for gauge theories. Journal of Mathematical Physics, 48(5):052301, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/
p052301_s1.

Gangopadhyay:2007:LAN

[1003] Sunandan Gangopadhyay. Lie algebraic noncommuting structures from reparametrization symmetry. Journal of Mathematical Physics, 48(5): 052302, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052302_s1.

Schwartz:2007:RQW

[1004] Charles Schwartz. Relativistic quaternionic wave equation. II. Journal of Mathematical Physics, 48(5):052303, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i5/p052303_s1.

Mack:2007:SST

[1005] Gerhard Mack and Mathias de Riese. Simple space–time symmetries: Generalizing conformal field theory. Journal of Mathematical Physics, 48 (5):052304, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052304_s1.

Agostini:2007:MRH

[1006] Alessandra Agostini. κ-Minkowski representations on Hilbert spaces. Journal of Mathematical Physics, 48(5):052305, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p052305_s1.

Tsamis:2007:MSV

[1007] N. C. Tsamis and R. P. Woodard. Maximally symmetric vector propagator. Journal of Mathematical Physics, 48(5):052306, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p052306_s1.

Moses:2007:MFC

[1008] Harry E. Moses. Magnetic forces on current circuits when there are no photons in the electromagnetic field. Journal of Mathematical Physics, 48 (5):052307, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052307_s1.

Dimock:2007:TAS

[1009] J. Dimock. Transition amplitudes and sewing properties for bosons on the Riemann sphere. Journal of Mathematical Physics, 48(5):052308, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p052308_s1.

Werner:2007:LFP

[1010] Marcus C. Werner. A Lefschetz fixed point theorem in gravitational lensing. Journal of Mathematical Physics, 48(5):052501, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p052501_s1.

Anco:2007:MCF

[1011] Stephen C. Anco. Mean-curvature flow and quasilocal mass for 2-surfaces in Hamiltonian general relativity. Journal of Mathematical Physics, 48 (5):052502, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052502_s1.

Sucu:2007:ESD

[1012] Yusuf Sucu and Nuri Ünal. Exact solution of Dirac equation in 2 + 1 dimensional gravity. Journal of Mathematical Physics, 48(5):052503, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p052503_s1.

Vrbik:2007:KPT

[1013] Jan Vrbik. Kepler problem with time-dependent and resonant perturbations. Journal of Mathematical Physics, 48(5):052701, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/
p052701_s1.

Dooling:2007:LSS

[1014] David C. Dooling and James E. Hammerberg. Lindstedt series solutions of the Fermi–Pasta–Ulam lattice. Journal of Mathematical Physics, 48 (5):052702, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052702_s1.

Asch:2007:DCH

[1015] J. Asch and P. Stovícek. Dynamics of a classical Hall system driven by a time-dependent Aharonov–Bohm flux. Journal of Mathematical Physics, 48(5):052901, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p052901_s1.

Schoberl:2007:CFG

[1016] Markus Schöberl and Kurt Schlacher. Covariant formulation of the governing equations of continuum mechanics in an Eulerian description. Journal of Mathematical Physics, 48(5):052902, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p052902_s1.

Bejancu:2007:GNM

[1017] Aurel Bejancu. On the geometry of nonholonomic mechanical systems with vertical distribution. Journal of Mathematical Physics, 48(5): 052903, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p052903_s1.

Cendra:2007:LSH

[1018] H. Cendra and S. D. Grillo. Lagrangian systems with higher order constraints. Journal of Mathematical Physics, 48(5):052904, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/
p052904_s1.

Koralov:2007:IPG

[1019] Leonid Koralov. An inverse problem for Gibbs fields with hard core potential. Journal of Mathematical Physics, 48(5):053301, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/
p053301_s1.

Ge:2007:GJE

[1020] Hao Ge and Min Qian. Generalized Jarzynski’s equality in inhomogeneous Markov chains. Journal of Mathematical Physics, 48(5):053302, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p053302_s1.

Rilke:2007:TOH

[1021] Florian Rilke. Transfer operators for hard-rod-type models. Journal of Mathematical Physics, 48(5):053303, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i5/p053303_s1.

Kota:2007:WRA

[1022] V. K. B. Kota. U(2Ω) ⊃ U(Ω) ⊃ ⊗SU(2) Wigner–Racah algebra for embedded Gaussian unitary ensemble of random matrices with spin. Journal of Mathematical Physics, 48(5):053304, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i5/p053304_s1.

Sukiasian:2007:RBC

[1023] H. S. Sukiasian and Wilfried Gille. Relation between the chord length distribution of an infinitely long cylinder and that of its base. Journal of Mathematical Physics, 48(5):053305, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i5/p053305_s1.

Maes:2007:MEP

[1024] Christian Maes and Karel Netocný. Minimum entropy production principle from a dynamical fluctuation law. Journal of Mathematical Physics, 48(5):053306, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p053306_s1.

Dong:2007:TTM

[1025] Huanhe Dong and Wei Yuan. Two types of multicomponent matrix loop algebras and its application. Journal of Mathematical Physics, 48(5): 053501, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p053501_s1.

Nishiyama:2007:SCF

[1026] Seiya Nishiyama, João da Providência, and Takao Komatsu. Selfconsistent-field method and τ-functional method on group manifold in soliton theory. II. Laurent coefficients of soliton solutions for ŝl_{n} and for ŝu_{n}. Journal of Mathematical Physics, 48(5):053502, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p053502_s1.

Mihai:2007:NDC

[1027] Dana Mihai and George A. J. Sparling. Null decomposition of conformal algebras. Journal of Mathematical Physics, 48(5):053503, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p053503_s1.

Ali:2007:MVB

[1028] S. Twareque Ali and M. Englis. Matrix-valued Berezin–Toeplitz quantization. Journal of Mathematical Physics, 48(5):053504, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/
p053504_s1.

Gunara:2007:BDW

[1029] Bobby E. Gunara, Freddy P. Zen, and Arianto. BPS domain walls and vacuum structure of N = 1 supergravity coupled to a chiral multiplet. Journal of Mathematical Physics, 48(5):053505, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p053505_s1.

Gomez:2007:ESC

[1030] A. Gómez, M. Torres, and J. L. Aragón. Eutactic star closest to a given star. Journal of Mathematical Physics, 48(5):053506, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i5/p053506_s1.

Matolcsi:2007:ATD

[1031] T. Matolcsi and P. Ván. Absolute time derivatives. Journal of Mathematical Physics, 48(5):053507, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i5/p053507_s1.

Abdesselam:2007:HDU

[1032] B. Abdesselam, A. Chakrabarti, V. K. Dobrev, and S. G. Mihov. Higher dimensional unitary braid matrices: Construction, associated structures, and entanglements. Journal of Mathematical Physics, 48(5):053508, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p053508_s1.

Zhou:2007:PET

[1033] Gang Zhou. Perturbation expansion and N th order Fermi golden rule of the nonlinear Schrödinger equations. Journal of Mathematical Physics, 48(5):053509, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p053509_s1.

Watanabe:2007:ISP

[1034] Michiyuki Watanabe. Inverse scattering problem for time dependent Hartree–Fock equations in the three-body case. Journal of Mathematical Physics, 48(5):053510, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p053510_s1.

Kleppe:2007:SSG

[1035] A. F. Kleppe and Chris Wainwright. Supercoset space geometry. Journal of Mathematical Physics, 48(5):053511, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i5/p053511_s1.

Henneaux:2007:SCR

[1036] Marc Henneaux, Mauricio Leston, Daniel Persson, and Philippe Spindel. A special class of rank 10 and 11 Coxeter groups. Journal of Mathematical Physics, 48(5):053512, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p053512_s1. See erratum [1608].

Wu:2007:GPC

[1037] Dongsheng Wu. Generalized Pickands constants. Journal of Mathematical Physics, 48(5):053513, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i5/p053513_s1.

Yang:2007:FFR

[1038] Wen-Li Yang, Yao-Zhong Zhang, and Xin Liu. Free field realization of current superalgebra gl(m|n)_{k}. Journal of Mathematical Physics, 48(5): 053514, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p053514_s1.

Slawny:2007:BSH

[1039] J. Slawny. Bound states of hydrogen atom in a theory with minimal length uncertainty relations. Journal of Mathematical Physics, 48(5): 053515, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p053515_s1.

Yurova:2007:ADM

[1040] Maria Yurova. Application of dressing method for long wave-short wave resonance interaction equation. Journal of Mathematical Physics, 48(5):

053516, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i5/p053516_s1.

Kajiura:2007:CHL

[1041] Hiroshige Kajiura. Categories of holomorphic line bundles on higher dimensional noncommutative complex tori. Journal of Mathematical Physics, 48(5):053517, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p053517_s1.

Lin:2007:HSS

[1042] Tai-Chia Lin and Juncheng Wei. Half-Skyrmions and spike-vortex solutions of two-component nonlinear Schrödinger systems. Journal of Mathematical Physics, 48(5):053518, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i5/p053518_s1.

Ma:2007:BSN

[1043] Li Ma and Lin Zhao. Blow up of solutions to the nonlinear Schrödinger equations on manifolds. Journal of Mathematical Physics, 48(5):053519, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p053519_s1.

Gal:2007:DFV

[1044] Tamás Gál. Differentiation of functionals with variables coupled by conservation constraints: Analysis through a fluid-dynamical model. Journal of Mathematical Physics, 48(5):053520, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i5/p053520_s1.

Pan:2007:NSS

[1045] Xing-Bin Pan. Nodal sets of solutions of equations involving magnetic Schrödinger operator in three dimensions. Journal of Mathematical Physics, 48(5):053521, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p053521_s1.

Lin:2007:QLS

[1046] Christopher Lin and Zhiqin Lu. Quantum layers over surfaces ruled outside a compact set. Journal of Mathematical Physics, 48(5):053522, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i5/p053522_s1.

Ibragimov:2007:CLM

[1047] Nail H. Ibragimov, Raisa Khamitova, and Bo Thidé. Conservation laws for the Maxwell–Dirac equations with dual Ohm’s law. Journal of Mathematical Physics, 48(5):053523, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i5/p053523_s1.

Doering:2007:ISI

[1048] Charles R. Doering and Paul K. Newton. Introduction to special issue: Mathematical Fluid Dynamics. Journal of Mathematical Physics, 48(6): 065101, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i6/p065101_s1.

Dascaliuc:2007:UBA

[1049] R. Dascaliuc, C. Foias, and M. S. Jolly. Universal bounds on the attractor of the Navier–Stokes equation in the energy, enstrophy plane. Journal of Mathematical Physics, 48(6):065201, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i6/p065201_s1.

Gibbon:2007:ETD

[1050] J. D. Gibbon and G. A. Pavliotis. Estimates for the two-dimensional Navier–Stokes equations in terms of the Reynolds number. Journal of Mathematical Physics, 48(6):065202, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i6/p065202_s1.

Kukavica:2007:NSE

[1051] Igor Kukavica and Mohammed Ziane. Navier–Stokes equations with regularity in one direction. Journal of Mathematical Physics, 48(6):065203, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i6/p065203_s1.

Chernyshenko:2007:PRT

[1052] Sergei I. Chernyshenko, Peter Constantin, James C. Robinson, and Edriss S. Titi. A posteriori regularity of the three-dimensional Navier– Stokes equations from numerical computations. Journal of Mathematical Physics, 48(6):065204, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i6/p065204_s1.

Ohkitani:2007:BPC

[1053] K. Ohkitani. A blow-up problem of a class of axisymmetric Navier–Stokes equations with infinite energy. Journal of Mathematical Physics, 48(6): 065205, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i6/p065205_s1.

Cordoba:2007:ABT

[1054] Diego Córdoba, Francisco Gancedo, and Rafael Orive. Analytical behavior of two-dimensional incompressible flow in porous media. Journal of Mathematical Physics, 48(6):065206, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i6/p065206_s1.

Galdi:2007:FPS

[1055] Giovanni P. Galdi. Further properties of steady-state solutions to the Navier–Stokes problem past a three-dimensional obstacle. Journal of Mathematical Physics, 48(6):065207, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i6/p065207_s1.

Cheskidov:2007:EDF

[1056] Alexey Cheskidov, Charles R. Doering, and Nikola P. Petrov. Energy dissipation in fractal-forced flow. Journal of Mathematical Physics, 48 (6):065208, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i6/p065208_s1.

Birnir:2007:DVM

[1057] Björn Birnir, Songming Hou, and Niklas Wellander. Derivation of the viscous Moore–Greitzer equation for aeroengine flow. Journal of Mathematical Physics, 48(6):065209, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i6/p065209_s1.

Jung:2007:AAS

[1058] Chang-Yeol Jung and Roger Temam. Asymptotic analysis for singularly perturbed convection-diffusion equations with a turning point. Journal of Mathematical Physics, 48(6):065301, June 2007. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i6/p065301_s1.

Marchioro:2007:VVL

[1059] Carlo Marchioro. Vanishing viscosity limit for an incompressible fluid with concentrated vorticity. Journal of Mathematical Physics, 48(6): 065302, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i6/p065302_s1.

Cheng:2007:DKT

[1060] Wenfang (Wendy) Cheng and Xiaoming Wang. Discrete Kato-type theorem on inviscid limit of Navier–Stokes flows. Journal of Mathematical Physics, 48(6):065303, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i6/p065303_s1.

Aref:2007:PVD

[1061] Hassan Aref. Point vortex dynamics: A classical mathematics playground. Journal of Mathematical Physics, 48(6):065401, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i6/
p065401_s1.

Blackmore:2007:SPT

[1062] Denis Blackmore, Lu Ting, and Omar Knio. Studies of perturbed three vortex dynamics. Journal of Mathematical Physics, 48(6):065402, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i6/p065402_s1.

Borisov:2007:DIP

[1063] Alexey V. Borisov, Ivan S. Mamaev, and Sergey M. Ramodanov. Dynamic interaction of point vortices and a two-dimensional cylinder. Journal of Mathematical Physics, 48(6):065403, June 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i6/p065403_s1.

Lekien:2007:LCS

[1064] François Lekien, Shawn C. Shadden, and Jerrold E. Marsden. Lagrangian coherent structures in n-dimensional systems. Journal of Mathematical Physics, 48(6):065404, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i6/p065404_s1.

Julien:2007:RMF

[1065] Keith Julien and Edgar Knobloch. Reduced models for fluid flows with strong constraints. Journal of Mathematical Physics, 48(6):065405, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i6/p065405_s1.

Majda:2007:NEL

[1066] Andrew Majda and Xiaoming Wang. Note on the emergence of large scale coherent structure under small scale random bombardments: The discrete case. Journal of Mathematical Physics, 48(6):065501, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i6/
p065501_s1.

Oliver:2007:TBC

[1067] Marcel Oliver and Oliver Bühler. Transparent boundary conditions as dissipative subgrid closures for the spectral representation of scalar advection by shear flows. Journal of Mathematical Physics, 48(6):065502, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i6/p065502_s1.

Cheskidov:2007:IDM

[1068] Alexey Cheskidov, Susan Friedlander, and Natasa Pavlović. Inviscid dyadic model of turbulence: The fixed point and Onsager’s conjecture. Journal of Mathematical Physics, 48(6):065503, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i6/p065503_s1.

Linshiz:2007:ASC

[1069] Jasmine S. Linshiz and Edriss S. Titi. Analytical study of certain magnetohydrodynamic-α models. Journal of Mathematical Physics, 48 (6):065504, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i6/p065504_s1.

Jamaloodeen:2007:TLQ

[1070] Mohamed I. Jamaloodeen and Paul K. Newton. Two-layer quasigeostrophic potential vorticity model. Journal of Mathematical Physics, 48(6):065601, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i6/p065601_s1.

Hsia:2007:SRB

[1071] Chun-Hsiung Hsia, Tian Ma, and Shouhong Wang. Stratified rotating Boussinesq equations in geophysical fluid dynamics: Dynamic bifurcation and periodic solutions. Journal of Mathematical Physics, 48(6): 065602, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i6/p065602_s1.

Lim:2007:EEN

[1072] Chjan C. Lim. Energy extremals and nonlinear stability in a variational theory of a coupled barotropic flow — Rotating solid sphere system. Journal of Mathematical Physics, 48(6):065603, June 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i6/p065603_s1.

Fukuda:2007:SAP

[1073] Motohisa Fukuda and Michael M. Wolf. Simplifying additivity problems using direct sum constructions. Journal of Mathematical Physics, 48(7): 072101, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i7/p072101_s1.

Ogura:2007:ASF

[1074] Akihiro Ogura and Motoo Sekiguchi. Algebraic structure of the Feynman propagator and a new correspondence for canonical transformations. Journal of Mathematical Physics, 48(7):072102, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p072102_s1.

Selesnick:2007:TSC

[1075] S. A. Selesnick. Type structure and chiral breaking in the standard model. Journal of Mathematical Physics, 48(7):072103, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/
p072103_s1.

Nagaj:2007:NCQ

[1076] Daniel Nagaj and Shay Mozes. New construction for a QMA complete three-local Hamiltonian. Journal of Mathematical Physics, 48(7):072104, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i7/p072104_s1.

Dong:2007:SSS

[1077] Jianping Dong and Mingyu Xu. Some solutions to the space fractional Schrödinger equation using momentum representation method. Journal of Mathematical Physics, 48(7):072105, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i7/p072105_s1.

Burch:2007:PSR

[1078] Aidan Burch. Phase space representations and perturbation theory for continuous-time histories. Journal of Mathematical Physics, 48(7): 072106, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i7/p072106_s1.

Khrennikov:2007:QRE

[1079] Andrei Khrennikov. Quantumlike representation of extensive form games: Probabilistic aspects. Journal of Mathematical Physics, 48(7): 072107, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i7/p072107_s1.

Daskaloyannis:2007:QSS

[1080] C. Daskaloyannis and Y. Tanoudis. Quantum superintegrable systems with quadratic integrals on a two dimensional manifold. Journal of Mathematical Physics, 48(7):072108, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i7/p072108_s1.

Gibilisco:2007:UPQ

[1081] Paolo Gibilisco, Daniele Imparato, and Tommaso Isola. Uncertainty principle and quantum Fisher information. II. Journal of Mathematical Physics, 48(7):072109, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p072109_s1.

Roy:2007:WCP

[1082] Aidan Roy and A. J. Scott. Weighted complex projective 2-designs from bases: Optimal state determination by orthogonal measurements. Journal of Mathematical Physics, 48(7):072110, July 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i7/p072110_s1.

Urias:2007:SDB

[1083] Jesús Urías. Spectral decomposition of Bell operators for multiqubit systems. Journal of Mathematical Physics, 48(7):072111, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/
p072111_s1.

Hu:2007:MFK

[1084] Po Hu and Igor Kriz. A mathematical formalism for the Kondo effect in Wess–Zumino–Witten branes. Journal of Mathematical Physics, 48(7): 072301, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i7/p072301_s1.

Bengtsson:2007:SHS

[1085] Anders K. H. Bengtsson. Structure of higher spin gauge interactions. Journal of Mathematical Physics, 48(7):072302, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p072302_s1.

Dahia:2007:WPS

[1086] F. Dahia, Lúcio F. P. da Silva, Carlos Romero, and Reza Tavakol. Warped product spaces and geodesic motion in the neighborhood of branes. Journal of Mathematical Physics, 48(7):072501, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/
p072501_s1.

Amaricci:2007:ASR

[1087] A. Amaricci, F. Bonetto, and P. Falco. Analyticity of the Sinai–Ruelle– Bowen measure for a class of simple Anosov flows. Journal of Mathematical Physics, 48(7):072701, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i7/p072701_s1.

Chen:2007:PGD

[1088] Yu-Tung Chen, Niann-Chern Lee, and Ming-Hsien Tu. Polytropic gas dynamics revisited. Journal of Mathematical Physics, 48(7):072702, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i7/p072702_s1.

Cresson:2007:SED

[1089] Jacky Cresson and Sébastien Darses. Stochastic embedding of dynamical systems. Journal of Mathematical Physics, 48(7):072703, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/
p072703_s1.

Chen:2007:IIO

[1090] Yong Chen, Ming-Ping Qian, and Jian-Sheng Xie. Irreversibility implies the occurrence of nonmonotonic power spectra. Journal of Mathematical Physics, 48(7):073301, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p073301_s1.

Ho:2007:DMF

[1091] Choon-Lin Ho and Ryu Sasaki. Deformed multivariable Fokker–Planck equations. Journal of Mathematical Physics, 48(7):073302, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/
p073302_s1.

Ha:2007:RSF

[1092] Seung-Yeal Ha and Se Eun Noh. Remarks on the stability of the frictionless Vlasov–Poisson–Fokker–Planck system. Journal of Mathematical Physics, 48(7):073303, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p073303_s1.

Wang:2007:EBL

[1093] Weike Wang, Tong Yang, and Xiongfeng Yang. Existence of boundary layers to the Boltzmann equation with cutoff soft potentials. Journal of Mathematical Physics, 48(7):073304, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i7/p073304_s1.

Luhn:2007:FG

[1094] Christoph Luhn, Salah Nasri, and Pierre Ramond. Flavor group Δ(3n^{2}). Journal of Mathematical Physics, 48(7):073501, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p073501_s1.

Ren:2007:ITS

[1095] Ji rong Ren, Ran Li, and Yi shi Duan. Inner topological structure of Hopf invariant. Journal of Mathematical Physics, 48(7):073502, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i7/p073502_s1.

Adamovic:2007:LIO

[1096] Drazen Adamović and Antun Milas. Logarithmic intertwining operators and W (2, 2p − 1) algebras. Journal of Mathematical Physics, 48(7): 073503, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i7/p073503_s1.

Maillard:2007:ESP

[1097] Jean-Marie Maillard. Explicit star products on orbits of nilpotent Lie groups with square integrable representations. Journal of Mathematical Physics, 48(7):073504, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p073504_s1.

Cederwall:2007:OI

[1098] Martin Cederwall and Jakob Palmkvist. The octic E_{8} invariant. Journal of Mathematical Physics, 48(7):073505, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i7/p073505_s1.

Kisil:2007:TDC

[1099] Vladimir V. Kisil. Two-dimensional conformal models of space–time and their compactification. Journal of Mathematical Physics, 48(7):073506, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i7/p073506_s1.

Huang:2007:GAE

[1100] Ding jiang Huang and Nataliya M. Ivanova. Group analysis and exact solutions of a class of variable coefficient nonlinear telegraph equations. Journal of Mathematical Physics, 48(7):073507, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p073507_s1.

Yang:2007:RPA

[1101] Zai-Chun Yang and Bo Tian. Remarks on the Painlevé analysis for a (2+ 1)-dimensional generalization of nonlinear-Schrödinger-type equations. Journal of Mathematical Physics, 48(7):073508, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p073508_s1.

Chen:2007:SLS

[1102] Bang-Yen Chen and Joeri Van der Veken. Spatial and Lorentzian surfaces in Robertson–Walker space times. Journal of Mathematical Physics, 48 (7):073509, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i7/p073509_s1.

Schief:2007:LES

[1103] W. K. Schief, A. Szereszewski, and C. Rogers. The Lamé equation in shell membrane theory. Journal of Mathematical Physics, 48(7):073510, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i7/p073510_s1.

Wu:2007:DRM

[1104] Dongsheng Wu. Dimension results of multifractional Brownian sheets. Journal of Mathematical Physics, 48(7):073511, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p073511_s1.

Huhtanen:2007:ECU

[1105] Marko Huhtanen. Energy conservation and unitary approximation numbers. Journal of Mathematical Physics, 48(7):073512, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p073512_s1.

Chen:2007:ECG

[1106] Guanggan Chen, Jian Zhang, and Yunyun Wei. Energy criterion of global existence for supercritical nonlinear Schrödinger equation with harmonic potential. Journal of Mathematical Physics, 48(7):073513, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i7/p073513_s1.

Sinelshchikov:2007:DCS

[1107] S. Sinel’shchikov, A. Stolin, and L. Vaksman. Differential calculi on some quantum prehomogeneous vector spaces. Journal of Mathematical Physics, 48(7):073514, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p073514_s1.

Merad:2007:WFD

[1108] M. Merad and S. Bensaid. Wave functions for a Duffin–Kemmer– Petiau particle in a time-dependent potential. Journal of Mathematical Physics, 48(7):073515, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p073515_s1. See comment [2196].

Albeverio:2007:MWA

[1109] S. Albeverio, P. E. T. Jorgensen, and A. M. Paolucci. Multiresolution wavelet analysis of integer scale Bessel functions. Journal of Mathematical Physics, 48(7):073516, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p073516_s1.

Castro:2007:NNG

[1110] Carlos Castro. On the noncommutative and nonassociative geometry of octonionic space time, modified dispersion relations and grand unification. Journal of Mathematical Physics, 48(7):073517, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p073517_s1.

Bender:2007:OSB

[1111] Carl M. Bender and Michael A. Bender. Optimal shape of a blob. Journal of Mathematical Physics, 48(7):073518, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i7/p073518_s1.

Degiovanni:2007:CVS

[1112] Luca Degiovanni and Giovanni Rastelli. Complex variables for separation of the Hamilton–Jacobi equation on real pseudo-Riemannian manifolds. Journal of Mathematical Physics, 48(7):073519, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/p073519_s1.

Kengne:2007:ESG

[1113] E. Kengne and R. Vaillancourt. Exact solutions of the Gross–Pitaevskii equation in periodic potential in the presence of external source. Journal of Mathematical Physics, 48(7):073520, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i7/p073520_s1.

Wang:2007:CPG

[1114] Shubin Wang, Guixiang Xu, and Guowang Chen. Cauchy problem for the generalized Benney–Luke equation. Journal of Mathematical Physics, 48 (7):073521, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i7/p073521_s1.

Loginov:2007:CGT

[1115] E. K. Loginov. On a class of gauge theories. Journal of Mathematical Physics, 48(7):073522, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i7/p073522_s1.

Otway:2007:CRS

[1116] Thomas H. Otway. Comment on “Removable singularities for solutions of coupled Yang–Mills–Dirac equations”, [J. Math. Phys. 47, 103502 (2006)]. Journal of Mathematical Physics, 48(7):074101, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/
p074101_s1. See [785].

Otway:2007:EHF

[1117] Thomas T. Otway. Erratum: “Harmonic fields on the extended projective disc and a problem in optics”, [J. Math. Phys. 46, 113501 (2005)]. Journal of Mathematical Physics, 48(7):079901, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/
p079901_s1. See [426].

Smirnov:2007:ECS

[1118] Roman G. Smirnov and Pavel Winternitz. Erratum: “A class of superintegrable systems of Calogero type”, [J. Math. Phys. 47, 093505 (2006)]. Journal of Mathematical Physics, 48(7):079902, July 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i7/
p079902_s1. See [762].

Zoli:2007:DRF

[1119] Marco Zoli. On the decay rate of the false vacuum. Journal of Mathematical Physics, 48(8):082101, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i8/p082101_s1.

Strauch:2007:RER

[1120] Frederick W. Strauch. Relativistic effects and rigorous limits for discreteand continuous-time quantum walks. Journal of Mathematical Physics, 48(8):082102, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p082102_s1.

Busch:2007:UJM

[1121] P. Busch and D. B. Pearson. Universal joint-measurement uncertainty relation for error bars. Journal of Mathematical Physics, 48(8):082103, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i8/p082103_s1.

Mangiarotti:2007:QMR

[1122] L. Mangiarotti and G. Sardanashvily. Quantum mechanics with respect to different reference frames. Journal of Mathematical Physics, 48(8): 082104, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i8/p082104_s1.

Ulfbeck:2007:SDV

[1123] Ole Ulfbeck. Simple derivation of the vector addition coefficients. Journal of Mathematical Physics, 48(8):082105, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i8/p082105_s1.

Odake:2007:EHO

[1124] Satoru Odake and Ryu Sasaki. Exact Heisenberg operator solutions for multiparticle quantum mechanics. Journal of Mathematical Physics, 48(8):082106, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p082106_s1.

Dupuis:2007:NNB

[1125] Frédéric Dupuis, Nicolas Gisin, Avinatan Hasidim, André Allan Méthot, and Haran Pilpel. No nonlocal box is universal. Journal of Mathematical Physics, 48(8):082107, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p082107_s1.

Blanchard:2007:PIH

[1126] Ph. Blanchard, R. Figari, and A. Mantile. Point interaction Hamiltonians in bounded domains. Journal of Mathematical Physics, 48(8): 082108, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i8/p082108_s1.

Eab:2007:FTC

[1127] C. H. Eab, S. C. Lim, and L. P. Teo. Finite temperature Casimir effect for a massless fractional Klein–Gordon field with fractional Neumann conditions. Journal of Mathematical Physics, 48(8):082301, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i8/p082301_s1.

Soylu:2007:ASD

[1128] A. Soylu, O. Bayrak, and I. Boztosun. An approximate solution of Dirac– Hulthén problem with pseudospin and spin symmetry for any κ state. Journal of Mathematical Physics, 48(8):082302, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i8/p082302_s1.

Shirzad:2007:FPG

[1129] A. Shirzad. Full and partial gauge fixing. Journal of Mathematical Physics, 48(8):082303, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p082303_s1.

Aspinwall:2007:LGC

[1130] Paul S. Aspinwall. Landau–Ginzburg to Calabi–Yau dictionary for Dbranes. Journal of Mathematical Physics, 48(8):082304, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i8/
p082304_s1.

Bagrov:2007:CPC

[1131] V. G. Bagrov, M. C. Baldiotti, and D. M. Gitman. Charged particles in crossed and longitudinal electromagnetic fields and beam guides. Journal of Mathematical Physics, 48(8):082305, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i8/p082305_s1.

Buchbinder:2007:CBR

[1132] I. L. Buchbinder and P. M. Lavrov. Classical Becchi–Rouet–Stora–Tyutin charge for nonlinear algebras. Journal of Mathematical Physics, 48(8): 082306, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i8/p082306_s1.

Clausen:2007:AFA

[1133] Adam Clausen and James Isenberg. Areal foliation and asymptotically velocity-term dominated behavior in T^{2} symmetric space–times with positive cosmological constant. Journal of Mathematical Physics, 48(8): 082501, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i8/p082501_s1.

Bessa:2007:SSM

[1134] Carlos H. G. Béssa and V. B. Bezerra. On some solutions of Maxwell’s equations in a Schwarzschild–de Sitter black hole. Journal of Mathematical Physics, 48(8):082502, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i8/p082502_s1.

Qiao:2007:NIH

[1135] Zhijun Qiao. New integrable hierarchy, its parametric solutions, cuspons, one-peak solitons, and M/W -shape peak solitons. Journal of Mathematical Physics, 48(8):082701, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i8/p082701_s1.

Cao:2007:DLS

[1136] Jinlong Cao and Xiang Zhang. Dynamics of the Lorenz system having an invariant algebraic surface. Journal of Mathematical Physics, 48(8): 082702, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i8/p082702_s1.

Birtea:2007:ASC

[1137] Petre Birtea, Mihai Boleantu, Mircea Puta, and Razvan Micu Tudoran. Asymptotic stability for a class of metriplectic systems. Journal of Mathematical Physics, 48(8):082703, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i8/p082703_s1.

Guo:2007:INC

[1138] Yong-Xin Guo, Shi-Xing Liu, Chang Liu, Shao-Kai Luo, and Yong Wang. Influence of nonholonomic constraints on variations, symplectic structure, and dynamics of mechanical systems. Journal of Mathematical Physics, 48(8):082901, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p082901_s1.

Yehia:2007:ATD

[1139] Hamad M. Yehia. Atlas of two-dimensional irreversible conservative Lagrangian mechanical systems with a second quadratic integral. Journal of Mathematical Physics, 48(8):082902, August 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i8/p082902_s1.

Vaisman:2007:WHD

[1140] Izu Vaisman. Weak-Hamiltonian dynamical systems. Journal of Mathematical Physics, 48(8):082903, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i8/p082903_s1.

Barros:2007:MVF

[1141] Manuel Barros, José L. Cabrerizo, Manuel Fernández, and Alfonso Romero. Magnetic vortex filament flows. Journal of Mathematical Physics, 48(8):082904, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p082904_s1.

Harland:2007:LSL

[1142] Derek Harland. Large scale and large period limits of symmetric calorons. Journal of Mathematical Physics, 48(8):082905, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i8/p082905_s1.

Angst:2007:CLT

[1143] Jürgen Angst and Jacques Franchi. Central limit theorem for a class of relativistic diffusions. Journal of Mathematical Physics, 48(8):083101, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i8/p083101_s1.

Zenchuk:2007:NMP

[1144] A. I. Zenchuk. New multidimensional partially integrable generalization of S-integrable N-wave equation. Journal of Mathematical Physics, 48(8):083501, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p083501_s1.

Petalidou:2007:GQT

[1145] Fani Petalidou. On the geometric quantization of twisted Poisson manifolds. Journal of Mathematical Physics, 48(8):083502, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i8/

p083502_s1. See addendum [1414].

Lim:2007:DAS

[1146] A. E. K. Lim and J. Carminati. Determination of all syzygies for the dependent polynomial invariants of the Riemann tensor. III. Mixed invariants of arbitrary degree in the Ricci spinor. Journal of Mathematical Physics, 48(8):083503, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p083503_s1.

Morse:2007:ECA

[1147] Edward C. Morse. Eigenfunctions of the curl in annular cylindrical and rectangular geometry. Journal of Mathematical Physics, 48(8): 083504, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i8/p083504_s1.

Vogeler:2007:PGS

[1148] Kirsten Vogeler and Michael Flohr. Pure gauge SU(2) Seiberg–Witten theory and modular forms. Journal of Mathematical Physics, 48(8): 083505, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i8/p083505_s1.

Romero:2007:NDB

[1149] David Romero and Federico Zertuche. Number of different binary functions generated by NK-Kauffman networks and the emergence of genetic robustness. Journal of Mathematical Physics, 48(8):083506, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i8/p083506_s1.

Audenaert:2007:SE

[1150] Koenraad M. R. Audenaert. Subadditivity of q-entropies for q > 1. Journal of Mathematical Physics, 48(8):083507, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i8/p083507_s1.

Trzetrzelewski:2007:SLG

[1151] Maciej Trzetrzelewski. Supersymmetry and Lie groups. Journal of Mathematical Physics, 48(8):083508, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i8/p083508_s1.

Martins:2007:FTC

[1152] R. A. Martins. Finite temperature corrections and embedded strings in noncommutative geometry and the standard model with neutrino mixing. Journal of Mathematical Physics, 48(8):083509, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v48/i8/p083509_s1.

Huang:2007:EMR

[1153] Shou-Jun Huang and De-Xing Kong. Equations for the motion of relativistic torus in the Minkowski space R^{1+n}. Journal of Mathematical Physics, 48(8):083510, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i8/p083510_s1.

Jorgensen:2007:HAI

[1154] Palle E. T. Jorgensen, Keri Kornelson, and Karen Shuman. Harmonic analysis of iterated function systems with overlap. Journal of Mathematical Physics, 48(8):083511, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i8/p083511_s1.

Serov:2007:PRP

[1155] Valery Serov and Markus Harju. Partial recovery of potentials in generalized nonlinear Schrödinger equations on the line. Journal of Mathematical Physics, 48(8):083512, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i8/p083512_s1.

Iglesias:2007:GFN

[1156] D. Iglesias, Juan C. Marrero, D. Martín de Diego, and Diana Sosa. General framework for nonholonomic mechanics: Nonholonomic systems on Lie affgebroids. Journal of Mathematical Physics, 48(8):083513, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i8/p083513_s1.

Yamashita:2007:CPA

[1157] Hideyasu Yamashita. Comment on “Pointwise analysis of scalar fields: A nonstandard approach”, [J. Math. Phys. 47, 092301 (2006)]. Journal of Mathematical Physics, 48(8):084101, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i8/p084101_s1. See [746].

Granstrom:2007:IKS

[1158] Helena Granström. Incomplete Kochen–Specker coloring. Journal of Mathematical Physics, 48(9):092101, September 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i9/p092101_s1.

Csiszar:2007:LRQ

[1159] Imre Csiszár, Fumio Hiai, and Dénes Petz. Limit relation for quantum entropy and channel capacity per unit cost. Journal of Mathematical Physics, 48(9):092102, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i9/p092102_s1.

Spohn:2007:NCB

[1160] Herbert Spohn. Notes on coherent backscattering from a random potential. Journal of Mathematical Physics, 48(9):092103, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i9/
p092103_s1.

Exner:2007:CRT

[1161] Pavel Exner and Olaf Post. Convergence of resonances on thin branched quantum waveguides. Journal of Mathematical Physics, 48(9):092104, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i9/p092104_s1.

Lim:2007:NGQ

[1162] Adrian P. C. Lim. A nonstandard geometric quantization of the harmonic oscillator. Journal of Mathematical Physics, 48(9):092105, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i9/p092105_s1.

Curtright:2007:BQS

[1163] Thomas Curtright and Luca Mezincescu. Biorthogonal quantum systems. Journal of Mathematical Physics, 48(9):092106, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i9/
p092106_s1.

Helm:2007:LWE

[1164] Mario Helm and Ivan Veselić. Linear Wegner estimate for alloy-type Schrödinger operators on metric graphs. Journal of Mathematical Physics, 48(9):092107, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i9/p092107_s1.

Curtright:2007:SBQ

[1165] Thomas Curtright, Luca Mezincescu, and David Schuster. Supersymmetric biorthogonal quantum systems. Journal of Mathematical Physics, 48(9):092108, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i9/p092108_s1.

Birkandan:2007:DEB

[1166] T. Birkandan and M. Hortaçsu. Dirac equation in the background of the Nutku helicoid metric. Journal of Mathematical Physics, 48(9):092301, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i9/p092301_s1. See comment [1498].

Kharlanov:2007:CLV

[1167] O. G. Kharlanov and V. Ch. Zhukovsky. CPT and Lorentz violation effects in hydrogenlike atoms. Journal of Mathematical Physics, 48(9):092302, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i9/p092302_s1.

Pitelli:2007:QSS

[1168] Paulo M. Pitelli and Patricio S. Letelier. Quantum singularities in space– times with spherical and cylindrical topological defects. Journal of Mathematical Physics, 48(9):092501, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i9/p092501_s1.

Ori:2007:SMM

[1169] Amos Ori and Dan Gorbonos. Simplified mathematical model for the formation of null singularities inside black holes. I. Basic formulation and a conjecture. Journal of Mathematical Physics, 48(9):092502, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i9/p092502_s1.

Gorbonos:2007:SMM

[1170] Dan Gorbonos and Gershon Wolansky. Simplified mathematical model for the formation of null singularities inside black holes. II. Proof for a particular case. Journal of Mathematical Physics, 48(9):092503, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i9/p092503_s1.

Heller:2007:AMS

[1171] Michael Heller, Zdzislaw Odrzygóźdź, Leszek Pysiak, and Wieslaw Sasin. Anatomy of malicious singularities. Journal of Mathematical Physics, 48(9):092504, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i9/p092504_s1.

Yaremko:2007:RRE

[1172] Yurij Yaremko. Radiation reaction in 2 + 1 electrodynamics. Journal of Mathematical Physics, 48(9):092901, September 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i9/p092901_s1.

Fasso:2007:IAS

[1173] Francesco Fassò and Nicola Sansonetto. Integrable almost-symplectic Hamiltonian systems. Journal of Mathematical Physics, 48(9):092902, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i9/p092902_s1.

Calogero:2007:IEH

[1174] F. A. Calogero and F. Leyvraz. Isochronous extension of the Hamiltonian describing free motion in the Poincaré half-plane: Classical and quantum treatments. Journal of Mathematical Physics, 48(9):092903, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i9/p092903_s1.

Poghosyan:2007:AEL

[1175] Suren Poghosyan and Hans Zessin. Asymptotic expansion of the logpartition function for a gas of interacting Brownian loops. Journal of Mathematical Physics, 48(9):093301, September 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i9/p093301_s1.

Mudrov:2007:RET

[1176] A. I. Mudrov. Reflection equation and twisted Yangians. Journal of Mathematical Physics, 48(9):093501, September 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i9/p093501_s1.

Dunajski:2007:MIS

[1177] Maciej Dunajski, James D. E. Grant, and Ian A. B. Strachan. Multidimensional integrable systems and deformations of Lie algebra homomorphisms. Journal of Mathematical Physics, 48(9):093502, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i9/p093502_s1.

Patera:2007:FGG

[1178] Jirí Patera, Edita Pelantová, and Milena Svobodová. Fine group gradings of the real forms of sl(4, C), sp(4, C), and o(4, C). Journal of Mathematical Physics, 48(9):093503, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i9/p093503_s1.

Klimyk:2007:ASM

[1179] A. Klimyk and J. Patera. (anti)symmetric multivariate trigonometric functions and corresponding Fourier transforms. Journal of Mathematical Physics, 48(9):093504, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i9/p093504_s1.

Geloun:2007:CNV

[1180] Joseph Ben Geloun and M. Norbert Hounkonnou. Canonical and nonlinear vector coherent states of generalized models with spin-orbit interaction. Journal of Mathematical Physics, 48(9):093505, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i9/p093505_s1.

Eveson:2007:MDQ

[1181] Simon P. Eveson and Christopher J. Fewster. Mass dependence of quantum energy inequality bounds. Journal of Mathematical Physics, 48(9):093506, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i9/p093506_s1.

Goze:2007:PAO

[1182] M. Goze, M. Rausch de Traubenberg, and A. Tanasa. Poincaré and sl(2) algebras of order 3. Journal of Mathematical Physics, 48(9):093507, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i9/p093507_s1.

Barrett:2007:OTV

[1183] John W. Barrett, João Faria Martins, and J. Manuel García-Islas. Observables in the Turaev–Viro and Crane–Yetter models. Journal of Mathematical Physics, 48(9):093508, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i9/p093508_s1.

DeLeo:2007:WPS

[1184] M. De Leo and D. Rial. Well posedness and smoothing effect of Schrödinger–Poisson equation. Journal of Mathematical Physics, 48 (9):093509, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i9/p093509_s1.

Kawamura:2007:RBS

[1185] Katsunori Kawamura. Recursive boson system in the Cuntz algebra O_{∞}. Journal of Mathematical Physics, 48(9):093510, September 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i9/
p093510_s1.

Li:2007:GIS

[1186] Yuan Li, Xiu-Hong Sun, and Zheng-Li Chen. Generalized infimum and sequential product of quantum effects. Journal of Mathematical Physics, 48(10):102101, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i10/p102101_s1.

Michelangeli:2007:RSL

[1187] Alessandro Michelangeli. Role of scaling limits in the rigorous analysis of Bose–Einstein condensation. Journal of Mathematical Physics, 48(10): 102102, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i10/p102102_s1.

Correggi:2007:RRBb

[1188] M. Correggi, T. Rindler-Daller, and J. Yngvason. Rapidly rotating Bose– Einstein condensates in homogeneous traps. Journal of Mathematical Physics, 48(10):102103, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i10/p102103_s1.

Braun:2007:SCSb

[1189] Carol Braun and Anupam Garg. Semiclassical coherent-state propagator for many spins. Journal of Mathematical Physics, 48(10):102104, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p102104_s1.

Hao:2007:GWP

[1190] Chengchun Hao, Ling Hsiao, and Hai-Liang Li. Global well posedness for the Gross–Pitaevskii equation with an angular momentum rotational term in three dimensions. Journal of Mathematical Physics, 48(10): 102105, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i10/p102105_s1.

Carinena:2007:QHO

[1191] José F. Cariñena, Manuel F. Rañada, and Mariano Santander. The quantum harmonic oscillator on the sphere and the hyperbolic plane: κ-dependent formalism, polar coordinates, and hypergeometric functions. Journal of Mathematical Physics, 48(10):102106, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i10/
p102106_s1.

Avron:2007:VTQ

[1192] J. E. Avron, G. Bisker, and O. Kenneth. Visualizing two qubits. Journal of Mathematical Physics, 48(10):102107, October 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i10/p102107_s1.

Hildebrand:2007:CR

[1193] Roland Hildebrand. Concurrence revisited. Journal of Mathematical Physics, 48(10):102108, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i10/p102108_s1.

Albeverio:2007:RAF

[1194] Sergio Albeverio, Laura Cattaneo, Sonia Mazzucchi, and Luca Di Persio. A rigorous approach to the Feynman–Vernon influence functional and its applications. I. Journal of Mathematical Physics, 48(10):102109, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p102109_s1.

Farkas:2007:NEA

[1195] S. Farkas and Z. Zimborás. The von Neumann entropy asymptotics in multidimensional fermionic systems. Journal of Mathematical Physics, 48(10):102110, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i10/p102110_s1.

Jansen:2007:BAA

[1196] Sabine Jansen, Mary-Beth Ruskai, and Ruedi Seiler. Bounds for the adiabatic approximation with applications to quantum computation. Journal of Mathematical Physics, 48(10):102111, October 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i10/p102111_s1.

Curtright:2007:QHQ

[1197] Thomas Curtright and Andrzej Veitia. Quasi-Hermitian quantum mechanics in phase space. Journal of Mathematical Physics, 48(10):102112, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i10/p102112_s1.

Fiore:2007:QDI

[1198] Gaetano Fiore. q-quaternions and q-deformed su(2) instantons. Journal of Mathematical Physics, 48(10):102301, October 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i10/p102301_s1.

Kawakami:2007:CEP

[1199] N. A. Kawakami, M. C. Nemes, and Walter F. Wreszinski. The Casimir effect for parallel plates revisited. Journal of Mathematical Physics, 48(10):102302, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i10/p102302_s1.

Mastropietro:2007:RGW

[1200] Vieri Mastropietro. Renormalization group and Ward identities for infrared QED4. Journal of Mathematical Physics, 48(10):102303, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p102303_s1.

Haesen:2007:PC

[1201] Stefan Haesen and Leopold Verstraelen. Pseudosymmetry collineations. Journal of Mathematical Physics, 48(10):102501, October 2007. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i10/
p102501_s1.

Couch:2007:LPB

[1202] W. E. Couch and C. L. Holder. Liouvillian perturbations of black holes. Journal of Mathematical Physics, 48(10):102502, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i10/
p102502_s1.

Bilge:2007:SFC

[1203] A. H. Bilge and D. Daghan. Scalar fields coupled to gravity in (3 + 1) dimensions: Static spherically symmetric stiff matter solutions. Journal of Mathematical Physics, 48(10):102503, October 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i10/p102503_s1.

Ferrando:2007:ISM

[1204] Joan Josep Ferrando and Juan Antonio Sáez. On the invariant symmetries of the D-metrics. Journal of Mathematical Physics, 48(10): 102504, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i10/p102504_s1.

Wang:2007:DAS

[1205] Wei Wang and Jinqiao Duan. A dynamical approximation for stochastic partial differential equations. Journal of Mathematical Physics, 48(10): 102701, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i10/p102701_s1.

Gurses:2007:IBV

[1206] Metin Gürses, Ismagil Habibullin, and Kostyantyn Zheltukhin. Integrable boundary value problems for elliptic type Toda lattice in a disk. Journal of Mathematical Physics, 48(10):102702, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i10/
p102702_s1.

Low:2007:RGN

[1207] Stephen G. Low. Relativity group for noninertial frames in Hamilton’s mechanics. Journal of Mathematical Physics, 48(10):102901, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p102901_s1.

Horwood:2007:HOF

[1208] Joshua T. Horwood. Higher order first integrals in classical mechanics. Journal of Mathematical Physics, 48(10):102902, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i10/
p102902_s1.

Ghanam:2007:BIN

[1209] R. Ghanam, F. Hindeleh, and G. Thompson. Bi-invariant and noninvariant metrics on Lie groups. Journal of Mathematical Physics, 48(10): 102903, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i10/p102903_s1.

Wang:2007:NDR

[1210] Hongyun Wang. A new derivation of the randomness parameter. Journal of Mathematical Physics, 48(10):103301, October 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i10/p103301_s1.

Perez-Madrid:2007:SMT

[1211] A. Pérez-Madrid. Statistical mechanical theory of an oscillating isolated system: The relaxation to equilibrium. Journal of Mathematical Physics, 48(10):103302, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i10/p103302_s1.

Konno:2007:MLE

[1212] Hidetoshi Konno and Fumitoshi Watanabe. Maximum likelihood estimators for generalized Cauchy processes. Journal of Mathematical Physics, 48(10):103303, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i10/p103303_s1.

Adams:2007:AAB

[1213] Stefan Adams and Tony Dorlas. C^{∗}-algebraic approach to the Bose– Hubbard model. Journal of Mathematical Physics, 48(10):103304, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p103304_s1.

Choe:2007:ABC

[1214] Kwangseok Choe. Asymptotic behavior of condensate solutions in the Chern–Simons–Higgs theory. Journal of Mathematical Physics, 48(10): 103501, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i10/p103501_s1.

Heo:2007:SSP

[1215] Jaeseong Heo. Stationary stochastic processes in a group system. Journal of Mathematical Physics, 48(10):103502, October 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i10/p103502_s1.

Jorgensen:2007:EEH

[1216] Palle E. T. Jorgensen and Myung-Sin Song. Entropy encoding, Hilbert space, and Karhunen–Loève transforms. Journal of Mathematical Physics, 48(10):103503, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i10/p103503_s1.

Poletaeva:2007:ELS

[1217] Elena Poletaeva. Embedding of the Lie superalgebra D(2, 1; α) into the Lie superalgebra of pseudodifferential symbols on S^{1|2}. Journal of Mathematical Physics, 48(10):103504, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i10/p103504_s1.

Abdesselam:2007:HDM

[1218] B. Abdesselam, A. Chakrabarti, V. K. Dobrev, and S. G. Mihov. Higher dimensional multiparameter unitary and nonunitary braid matrices: Even dimensions. Journal of Mathematical Physics, 48(10):103505, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p103505_s1.

Jorgensen:2007:MM

[1219] Palle E. T. Jorgensen. The measure of a measurement. Journal of Mathematical Physics, 48(10):103506, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i10/p103506_s1.

MacKay:2007:RMC

[1220] N. J. MacKay and A. Taylor. Rational R-matrices, centralizer algebras and tensor identities for e_{6} and e_{7} exceptional families of Lie algebras. Journal of Mathematical Physics, 48(10):103507, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i10/
p103507_s1.

Kong:2007:WCM

[1221] De-Xing Kong and Kefeng Liu. Wave character of metrics and hyperbolic geometric flow. Journal of Mathematical Physics, 48(10):103508, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p103508_s1.

Ji:2007:CLB

[1222] Lina Ji and Changzheng Qu. Conditional Lie Bäcklund symmetries and solutions to (n+1)-dimensional nonlinear diffusion equations. Journal of Mathematical Physics, 48(10):103509, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i10/p103509_s1.

Zhou:2007:IRD

[1223] Ruguang Zhou. Integrable Rosochatius deformations of the restricted soliton flows. Journal of Mathematical Physics, 48(10):103510, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p103510_s1.

Ramm:2007:MBWb

[1224] A. G. Ramm. Many-body wave scattering by small bodies and applications. Journal of Mathematical Physics, 48(10):103511, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i10/
p103511_s1.

Joshi:2007:LPI

[1225] N. Joshi, A. V. Kitaev, and P. A. Treharne. On the linearization of the Painlevé III–VI equations and reductions of the three-wave resonant system. Journal of Mathematical Physics, 48(10):103512, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i10/
p103512_s1.

Abdollahpour:2007:IAO

[1226] M. R. Abdollahpour, F. Bagarello, and S. Triolo. An invariant analytic orthonormalization procedure with applications. Journal of Mathematical Physics, 48(10):103513, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i10/p103513_s1.

Saenz:2007:EPE

[1227] A. W. Sáenz. Erratum: “Proof of the existence of the axial-to-planar channeling transition in a simple crystal model”, [J. Math. Phys. 47, 062904 (2006)]. Journal of Mathematical Physics, 48(10):109901, October 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i10/p109901_s1. See [656].

Kopf:2007:GCQ

[1228] Tomás Kopf and Mario Paschke. Generally covariant quantum mechanics on noncommutative configuration spaces. Journal of Mathematical Physics, 48(11):112101, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i11/p112101_s1.

Huang:2007:GEP

[1229] Siendong Huang. Generalized Einstein–Podolsky–Rosen states. Journal of Mathematical Physics, 48(11):112102, November 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i11/p112102_s1.

Ribeiro:2007:TFS

[1230] A. D. Ribeiro, M. A. M. de Aguiar, and A. F. R. de Toledo Piza. Trace formula for systems with spin from the coherent state propagator. Journal of Mathematical Physics, 48(11):112103, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i11/
p112103_s1.

Nouicer:2007:CPO

[1231] Khireddine Nouicer. Coulomb potential in one dimension with minimal length: A path integral approach. Journal of Mathematical Physics, 48(11):112104, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p112104_s1.

Galiautdinov:2007:SSC

[1232] Andrei Galiautdinov. Single-step controlled-NOT logic from any exchange interaction. Journal of Mathematical Physics, 48(11):112105, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p112105_s1.

Chenaghlou:2007:BGK

[1233] A. Chenaghlou and O. Faizy. Barut–Girardello and Klauder–Perelomov coherent states for the Kravchuk functions. Journal of Mathematical Physics, 48(11):112106, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i11/p112106_s1. See comment [1415].

Esteban:2007:SAD

[1234] Maria J. Esteban and M. Loss. Self-adjointness for Dirac operators via Hardy–Dirac inequalities. Journal of Mathematical Physics, 48 (11):112107, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p112107_s1.

Bezak:2007:TTI

[1235] Viktor Bezák. Tunneling through inhomogeneous delta barriers as diffraction and scattering. Journal of Mathematical Physics, 48 (11):112108, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p112108_s1.

Gheondea:2007:CNI

[1236] Aurelian Gheondea. Comment on “A note on the infimum problem of Hilbert space effects”, [J. Math. Phys. 47, 102103 (2006)]. Journal of Mathematical Physics, 48(11):112109, November 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i11/p112109_s1. See

[766].

Pitrik:2007:MQS

[1237] József Pitrik. Markovian quasifree states on canonical anticommutation relation algebras. Journal of Mathematical Physics, 48(11):112110, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p112110_s1.

Loran:2007:KQM

[1238] Farhang Loran. Kähler quantization of H^{∗}(T^{2}, R) and modular forms. Journal of Mathematical Physics, 48(11):112301, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i11/
p112301_s1.

deAMarques:2007:SEQ

[1239] Geusa de A.Marques, J. G. de Assis, and V. B. Bezerra. Some effects on quantum systems due to the gravitational field of a cosmic string. Journal of Mathematical Physics, 48(11):112501, November 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i11/p112501_s1.

Echeverria-Enriquez:2007:EHS

[1240] Arturo Echeverría-Enríquez, Manuel de León, Miguel C. Muñoz-Lecanda, and Narciso Román-Roy. Extended Hamiltonian systems in multisymplectic field theories. Journal of Mathematical Physics, 48 (11):112901, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p112901_s1.

Bruno:2007:CCS

[1241] Danilo Bruno. Constructing a class of solutions for the Hamilton–Jacobi equation in field theory. Journal of Mathematical Physics, 48(11):112902, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p112902_s1.

Pucacco:2007:DSV

[1242] Giuseppe Pucacco and Kjell Rosquist. (1 + 1)-dimensional separation of variables. Journal of Mathematical Physics, 48(11):112903, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i11/p112903_s1.

Schwammle:2007:TPG

[1243] Veit Schwämmle and Constantino Tsallis. Two-parameter generalization of the logarithm and exponential functions and Boltzmann–Gibbs– Shannon entropy. Journal of Mathematical Physics, 48(11):113301, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p113301_s1.

Aschbacher:2007:ESD

[1244] Walter H. Aschbacher and Jean-Marie Barbaroux. Exponential spatial decay of spin-spin correlations in translation invariant quasifree states. Journal of Mathematical Physics, 48(11):113302, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i11/
p113302_s1.

Iguchi:2007:GCD

[1245] Kazumoto Iguchi and Hiroaki S. Yamada. General connectivity distribution functions for growing networks with preferential attachment of fractional power. Journal of Mathematical Physics, 48(11):113303, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p113303_s1.

Zhang:2007:SCP

[1246] Xianwen Zhang and Shigeng Hu. L^{p} solutions to the Cauchy problem of the BGK equation. Journal of Mathematical Physics, 48(11):113304, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p113304_s1.

Kaynak:2007:SCZ

[1247] Burak Tevfik Kaynak and O. Teoman Turgut. Symbol calculus and zetafunction regularized determinants. Journal of Mathematical Physics, 48(11):113501, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p113501_s1.

Gao:2007:EDC

[1248] Ya-Jun Gao. Extended double-complex linear systems and new multiple infinite-dimensional hidden symmetries for the general symplectic gravity models. Journal of Mathematical Physics, 48(11):113502, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i11/p113502_s1.

Yahiaoui:2007:EPR

[1249] S.-A. Yahiaoui, O. Cherroud, and M. Bentaiba. Effective potential and resummation procedure to multidimensional complex cubic potentials for weak and strong coupling. Journal of Mathematical Physics, 48(11):113503, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p113503_s1.

Ko:2007:RDD

[1250] Chul Ki Ko. Remarks on the decomposition of Dirichlet forms on standard forms of von Neumann algebras. Journal of Mathematical Physics, 48(11):113504, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p113504_s1.

Fioresi:2007:MCS

[1251] R. Fioresi, M. A. Lledó, and V. S. Varadarajan. The Minkowski and conformal superspaces. Journal of Mathematical Physics, 48(11):113505, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p113505_s1. See erratum [2215].

Doliwa:2007:ILTb

[1252] A. Doliwa, M. Nieszporski, and P. M. Santini. Integrable lattices and their sublattices. II. From the B-quadrilateral lattice to the self-adjoint schemes on the triangular and the honeycomb lattices. Journal of Mathematical Physics, 48(11):113506, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i11/p113506_s1.

Dossena:2007:SRZ

[1253] Giacomo Dossena. Some results on the Zeeman topology. Journal of Mathematical Physics, 48(11):113507, November 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i11/p113507_s1.

Sadeghi:2007:DOM

[1254] J. Sadeghi. Dirac oscillator with minimal lengths and free particle on AdS_{2} and S^{2}. Journal of Mathematical Physics, 48(11):113508, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i11/p113508_s1.

Kaikina:2007:PON

[1255] Elena Kaikina. Pseudodifferential operator with a nonanalytic symbol on a half-line. Journal of Mathematical Physics, 48(11):113509, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i11/p113509_s1.

Cimmelli:2007:ELP

[1256] Vito Antonio Cimmelli. An extension of Liu procedure in weakly nonlocal thermodynamics. Journal of Mathematical Physics, 48(11):113510, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p113510_s1.

Klinker:2007:DSB

[1257] Frank Klinker. The decomposition of the spinor bundle of Grassmann manifolds. Journal of Mathematical Physics, 48(11):113511, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i11/p113511_s1.

Freidel:2007:SDR

[1258] L. Freidel, K. Noui, and Ph. Roche. 6j symbols duality relations. Journal of Mathematical Physics, 48(11):113512, November 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i11/p113512_s1.

Tian:2007:OCV

[1259] Lixin Tian and Chunyu Shen. Optimal control of the viscous Degasperis– Procesi equation. Journal of Mathematical Physics, 48(11):113513, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i11/p113513_s1.

Koca:2007:POC

[1260] Mehmet Koca, Mudhahir Al-Ajmi, and Ramazan Koç. Polyhedra obtained from Coxeter groups and quaternions. Journal of Mathematical Physics, 48(11):113514, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i11/p113514_s1.

Chilla:2007:SIR

[1261] Vincenzo Chilla. Selection and identity rules for subductions of type A quantum Iwahori–Hecke algebras. Journal of Mathematical Physics, 48(11):113515, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p113515_s1.

Kanakoglou:2007:BLG

[1262] K. Kanakoglou and C. Daskaloyannis. A braided look at Green ansatz for parabosons. Journal of Mathematical Physics, 48(11):113516, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i11/p113516_s1.

Chen:2007:LTD

[1263] Q. Chen, J. Jost, and G. Wang. Liouville theorems for Dirac-harmonic maps. Journal of Mathematical Physics, 48(11):113517, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i11/
p113517_s1.

Kalnins:2007:NTD

[1264] E. G. Kalnins, J. M. Kress, and W. Miller. Nondegenerate threedimensional complex Euclidean superintegrable systems and algebraic varieties. Journal of Mathematical Physics, 48(11):113518, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i11/p113518_s1.

He:2007:GTC

[1265] Jingsong He, Zhiwei Wu, and Yi Cheng. Gauge transformations for the constrained CKP and BKP hierarchies. Journal of Mathematical Physics, 48(11):113519, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i11/p113519_s1.

Zakrzewski:2007:SBH

[1266] W. J. Zakrzewski. Surfaces in R^{N^{2}−1} based on harmonic maps S^{2} → CP^{N−1}. Journal of Mathematical Physics, 48(11):113520, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i11/p113520_s1.

Skrypnyk:2007:GGS

[1267] T. Skrypnyk. Generalized Gaudin spin chains, nonskew symmetric rmatrices, and reflection equation algebras. Journal of Mathematical Physics, 48(11):113521, November 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i11/p113521_s1.

Gerard:2007:GDT

[1268] Christian Gérard and Rafael Tiedra de Aldecoa. Generalized definition of time delay in scattering theory. Journal of Mathematical Physics, 48(12):122101, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p122101_s1.

Michalakis:2007:MMO

[1269] S. Michalakis. Multiplicativity of the maximal output 2-norm for depolarized Werner–Holevo channels. Journal of Mathematical Physics, 48(12):122102, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p122102_s1.

Erman:2007:NLM

[1270] Fatih Erman and O. Teoman Turgut. Nonrelativistic Lee model in three dimensional Riemannian manifolds. Journal of Mathematical Physics, 48(12):122103, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p122103_s1.

Sasaki:2007:QSD

[1271] Ryu Sasaki. Quasiexactly solvable difference equations. Journal of Mathematical Physics, 48(12):122104, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i12/p122104_s1.

Odake:2007:MQS

[1272] Satoru Odake and Ryu Sasaki. Multiparticle quasiexactly solvable difference equations. Journal of Mathematical Physics, 48(12):122105, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i12/p122105_s1.

Boos:2007:SRF

[1273] William Boos. Stochastic representations of Feynman integration. Journal of Mathematical Physics, 48(12):122106, December 2007. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i12/
p122106_s1.

Unal:2007:PTC

[1274] Nuri Ünal. Parametric-time coherent states for Smorodinsky–Winternitz potentials. Journal of Mathematical Physics, 48(12):122107, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i12/p122107_s1.

Halpern:2007:OPT

[1275] M. B. Halpern. The orbifolds of permutation type as physical string systems at multiples of c = 26. V. Cyclic permutation orbifolds. Journal of Mathematical Physics, 48(12):122301, December 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i12/p122301_s1.

Morgan:2007:LFR

[1276] Peter Morgan. Lie fields revisited. Journal of Mathematical Physics, 48(12):122302, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p122302_s1.

Ibison:2007:CFR

[1277] M. Ibison. On the conformal forms of the Robertson–Walker metric. Journal of Mathematical Physics, 48(12):122501, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i12/
p122501_s1.

Schuch:2007:CBQ

[1278] Dieter Schuch. Connection between quantum mechanical and classical time evolution of certain dissipative systems via a dynamical invariant. Journal of Mathematical Physics, 48(12):122701, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i12/
p122701_s1.

Brauchart:2007:SLD

[1279] J. S. Brauchart, D. P. Hardin, and E. B. Saff. The support of the limit distribution of optimal Riesz energy points on sets of revolution in R^{3}. Journal of Mathematical Physics, 48(12):122901, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i12/
p122901_s1.

Freitas:2007:CSM

[1280] P. Freitas, R. S. Laugesen, and G. F. Liddell. On convex surfaces with minimal moment of inertia. Journal of Mathematical Physics, 48(12):122902, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p122902_s1.

Carrisi:2007:EFM

[1281] M. C. Carrisi and S. Pennisi. An exact fluid model for relativistic electron beams: The many moment case. Journal of Mathematical Physics, 48(12):123101, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p123101_s1.

Li:2007:SRO

[1282] Dening Li. Shock reflection and oblique shock waves. Journal of Mathematical Physics, 48(12):123102, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i12/p123102_s1.

Hiai:2007:LDC

[1283] Fumio Hiai, Milán Mosonyi, and Tomohiro Ogawa. Large deviations and Chernoff bound for certain correlated states on a spin chain. Journal of Mathematical Physics, 48(12):123301, December 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i12/p123301_s1.

Dai:2007:IQG

[1284] Wu-Sheng Dai and Mi Xie. Interacting quantum gases in confined space: Two- and three-dimensional equations of state. Journal of Mathematical Physics, 48(12):123302, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i12/p123302_s1.

Khelifi:2007:DSA

[1285] Abdessatar Khelifi. Determination of small amplitude perturbations for the electric permittivity from partial dynamic boundary measurements. Journal of Mathematical Physics, 48(12):123501, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i12/
p123501_s1.

Klaus:2007:CCB

[1286] Martin Klaus and Boris Mityagin. Coupling constant behavior of eigenvalues of Zakharov–Shabat systems. Journal of Mathematical Physics, 48(12):123502, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p123502_s1.

Sodin:2007:RMN

[1287] Sasha Sodin. Random matrices, nonbacktracking walks, and orthogonal polynomials. Journal of Mathematical Physics, 48(12):123503, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v48/i12/p123503_s1.

Romano:2007:QCS

[1288] V. Romano. Quantum corrections to the semiclassical hydrodynamical model of semiconductors based on the maximum entropy principle. Journal of Mathematical Physics, 48(12):123504, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i12/
p123504_s1.

Pickl:2007:GED

[1289] Peter Pickl. Generalized eigenfunctions for Dirac operators near criticality. Journal of Mathematical Physics, 48(12):123505, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v48/i12/
p123505_s1.

Lo:2007:EDP

[1290] Assane Lo. On the exponential decay of the n-point correlation functions and the analyticity of the pressure. Journal of Mathematical Physics, 48(12):123506, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p123506_s1.

Schmidt:2007:PCS

[1291] Matthias Schmidt and Mike R. Jeffrey. Peel or coat spheres by convolution, repeatedly. Journal of Mathematical Physics, 48(12):123507, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i12/p123507_s1.

Ha:2007:GWP

[1292] Seung-Yeal Ha and Ho Lee. Global well posedness of the relativistic Vlasov–Yukawa system with small data. Journal of Mathematical Physics, 48(12):123508, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i12/p123508_s1.

Vacaru:2007:DQA

[1293] Sergiu I. Vacaru. Deformation quantization of almost Kähler models and Lagrange–Finsler spaces. Journal of Mathematical Physics, 48(12):123509, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p123509_s1.

Nucci:2007:LG

[1294] M. C. Nucci and P. G. L. Leach. Lagrangians galore. Journal of Mathematical Physics, 48(12):123510, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i12/p123510_s1.

Starkl:2007:NES

[1295] Reinhard Starkl. New explicit solutions of Einstein equations in the framework of GAP Theory. Journal of Mathematical Physics, 48 (12):123511, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p123511_s1.

Shpot:2007:MFI

[1296] M. A. Shpot. A massive Feynman integral and some reduction relations for Appell functions. Journal of Mathematical Physics, 48(12):123512, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i12/p123512_s1.

Freyre:2007:NA

[1297] Sebastián Freyre, Matías Graña, and Leandro Vendramin. On Nichols algebras over SL(2, F_{q}) and GL(2, F_{q}). Journal of Mathematical Physics, 48(12):123513, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p123513_s1.

Odzijewicz:2007:CSM

[1298] Anatol Odzijewicz, Tomasz Goliński, and Agnieszka Tereszkiewicz. Coherent state maps related to the bounded positive operators. Journal of Mathematical Physics, 48(12):123514, December 2007. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v48/i12/p123514_s1.

Manojlovic:2007:ABA

[1299] Nenad Manojlović and Zoltán Nagy. Algebraic Bethe ansatz for the elliptic quantum group E_{τ,η}(A_{2}^{(2)}). Journal of Mathematical Physics, 48(12):123515, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p123515_s1.

Nikolov:2007:DAC

[1300] Petko Nikolov and Tihomir Valchev. Description of all conformally invariant differential operators, acting on scalar functions. Journal of Mathematical Physics, 48(12):123516, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v48/i12/p123516_s1.

daRocha:2007:DSF

[1301] R. da Rocha and J. M. Hoff da Silva. From Dirac spinor fields to eigenspinoren des ladungskonjugationsoperators. Journal of Mathematical Physics, 48(12):123517, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v48/i12/p123517_s1.

Calvaruso:2007:TDH

[1302] Giovanni Calvaruso. Three-dimensional homogeneous Lorentzian metrics with prescribed Ricci tensor. Journal of Mathematical Physics, 48(12):123518, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i12/p123518_s1.

Luhn:2007:SFN

[1303] Christoph Luhn, Salah Nasri, and Pierre Ramond. Simple finite non-Abelian flavor groups. Journal of Mathematical Physics, 48(12):123519, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v48/i12/p123519_s1.

Miki:2007:AA

[1304] Kei Miki. A (q, γ) analog of the W_{1+∞} algebra. Journal of Mathematical Physics, 48(12):123520, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v48/i12/p123520_s1.

Nachtergaele:2008:ASIa

[1305] Bruno L. Z. Nachtergaele. Announcement: Special issue on “Statistical Mechanics on Random Structures”. Journal of Mathematical Physics, 49(1):010201, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p010201_s1.

Pitowsky:2008:NBI

[1306] Itamar Pitowsky. New Bell inequalities for the singlet state: Going beyond the Grothendieck bound. Journal of Mathematical Physics, 49(1): 012101, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p012101_s1.

Appleby:2008:SPP

[1307] D. M. Appleby, Ingemar Bengtsson, and S. Chaturvedi. Spectra of phase point operators in odd prime dimensions and the extended Clifford group. Journal of Mathematical Physics, 49(1):012102, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i1/p012102_s1.

Chen:2008:EUE

[1308] Jianxin Chen, Runyao Duan, Zhengfeng Ji, Mingsheng Ying, and Jun Yu. Existence of universal entangler. Journal of Mathematical Physics, 49(1): 012103, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p012103_s1.

Suslov:2008:HPN

[1309] Sergei K. Suslov and Benjamin Trey. The Hahn polynomials in the nonrelativistic and relativistic Coulomb problems. Journal of Mathematical Physics, 49(1):012104, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p012104_s1.

Hassan:2008:DCS

[1310] Ali Saif M. Hassan and Pramod S. Joag. On the degree conjecture for separability of multipartite quantum states. Journal of Mathematical Physics, 49(1):012105, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p012105_s1.

Andai:2008:UPQ

[1311] Attila Andai. Uncertainty principle with quantum Fisher information. Journal of Mathematical Physics, 49(1):012106, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i1/p012106_s1.

Bernardoni:2008:MGG

[1312] Fabio Bernardoni, Sergio L. Cacciatori, Bianca L. Cerchiai, and Antonio Scotti. Mapping the geometry of the E_{6} group. Journal of Mathematical Physics, 49(1):012107, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p012107_s1.

Yilmaz:2008:DPS

[1313] Nejat T. Yilmaz. Dualization of the principal sigma model. Journal of Mathematical Physics, 49(1):012301, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i1/p012301_s1.

Benguria:2008:SAM

[1314] Rafael D. Benguria, Michael Loss, and Heinz Siedentop. Stability of atoms and molecules in an ultrarelativistic Thomas–Fermi–Weizsäcker model. Journal of Mathematical Physics, 49(1):012302, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i1/
p012302_s1.

Rajah:2008:RER

[1315] S. S. Rajah and S. D. Maharaj. A Riccati equation in radiative stellar collapse. Journal of Mathematical Physics, 49(1):012501, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i1/p012501_s1.

Liu:2008:HCM

[1316] Dong Liu and Cuipo Jiang. Harish-Chandra modules over the twisted Heisenberg–Virasoro algebra. Journal of Mathematical Physics, 49(1): 012901, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p012901_s1.

Bolivar:2008:GCK

[1317] A. O. Bolivar. Generalization of the classical Kramers rate for non-Markovian open systems out of equilibrium. Journal of Mathematical Physics, 49(1):013301, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p013301_s1.

Xu:2008:TTG

[1318] Tao Xu, Hai-Qiang Zhang, Ya-Xing Zhang, Juan Li, Qian Feng, and Bo Tian. Two types of generalized integrable decompositions and new solitary-wave solutions for the modified Kadomtsev–Petviashvili equation with symbolic computation. Journal of Mathematical Physics, 49(1): 013501, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p013501_s1.

Lipatov:2008:TCA

[1319] L. N. Lipatov, M. Rausch de Traubenberg, and G. G. Volkov. On the ternary complex analysis and its applications. Journal of Mathematical Physics, 49(1):013502, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p013502_s1.

Gorin:2008:MIC

[1320] T. Gorin and G. V. López. Monomial integrals on the classical groups. Journal of Mathematical Physics, 49(1):013503, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i1/p013503_s1.

Albini:2008:RBC

[1321] Paolo Albini, Alessandro Toigo, and Veronica Umanità. Relations between convergence rates in Schatten p-norms. Journal of Mathematical Physics, 49(1):013504, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p013504_s1.

Nesterov:2008:IDR

[1322] Alexander I. Nesterov and Fermín Aceves de la Cruz. Infinite-dimensional representations of the rotation group and Dirac monopole problem. Journal of Mathematical Physics, 49(1):013505, January 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i1/p013505_s1.

Hu:2008:SGD

[1323] Juan Hu, Hong-Yan Wang, and Hon-Wah Tam. Source generation of the Davey–Stewartson equation. Journal of Mathematical Physics, 49(1): 013506, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p013506_s1.

Du:2008:AGL

[1324] Hong-Ke Du, Yue-Qing Wang, and Jun-Lian Xu. Applications of the generalized Lüders theorem. Journal of Mathematical Physics, 49(1): 013507, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p013507_s1.

Arroyo:2008:USS

[1325] J. Arroyo, O. J. Garay, and J. J. Mencía. Unit speed stationary points of the acceleration. Journal of Mathematical Physics, 49(1):013508, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i1/p013508_s1.

Lim:2008:SPP

[1326] S. C. Lim and L. P. Teo. Sample path properties of fractional Riesz– Bessel field of variable order. Journal of Mathematical Physics, 49(1): 013509, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p013509_s1.

Athanasiadis:2008:RRE

[1327] Christodoulos Athanasiadis and Nikolaos L. Tsitsas. Radiation relations for electromagnetic excitation of a layered chiral medium by an interior dipole. Journal of Mathematical Physics, 49(1):013510, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i1/p013510_s1.

Chanachowicz:2008:ICR

[1328] Mark Chanachowicz, Claudia M. Chanu, and Raymond G. McLenaghan. Invariant classification of the rotationally symmetric R-separable webs for the Laplace equation in Euclidean space. Journal of Mathematical Physics, 49(1):013511, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p013511_s1.

Marquette:2008:EPP

[1329] I. Marquette and P. Winternitz. Erratum: “Polynomial Poisson Algebras for Classical Superintegrable Systems with a Third Order Integral of Motion”, [J. Math. Phys. 48, 012902 (2007)]. Journal of Mathematical Physics, 49(1):019901, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i1/p019901_s1. See [881].

Caves:2008:EUQ

[1330] Carlton M. Caves, Christopher A. Fuchs, and Rüdiger Schack. Erratum: “Unknown quantum states: The quantum de Finetti representation”, [J. Math. Phys. 44, 4537 (2002)]. Journal of Mathematical Physics, 49(1): 019902, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i1/p019902_s1. See [3].

Anastopoulos:2008:TAPa

[1331] Charis Anastopoulos and Ntina Savvidou. Time-of-arrival probabilities and quantum measurements. II. Application to tunneling times. Journal of Mathematical Physics, 49(2):022101, February 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i2/p022101_s1.

Masanes:2008:UEC

[1332] Lluís Masanes. Useful entanglement can be extracted from all nonseparable states. Journal of Mathematical Physics, 49(2):022102, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i2/p022102_s1.

Anastopoulos:2008:TAPb

[1333] Charis Anastopoulos. Time-of-arrival probabilities and quantum measurements. III. Decay of unstable states. Journal of Mathematical Physics, 49(2):022103, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i2/p022103_s1.

Chenaghlou:2008:GKC

[1334] A. Chenaghlou and O. Faizy. Gazeau–Klauder coherent states for trigonometric Rosen–Morse potential. Journal of Mathematical Physics, 49(2):022104, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i2/p022104_s1. See comments [1456].

Eskin:2008:IPS

[1335] G. Eskin. Inverse problems for the Schrödinger equations with timedependent electromagnetic potentials and the Aharonov–Bohm effect. Journal of Mathematical Physics, 49(2):022105, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i2/p022105_s1.

Quesne:2008:OMC

[1336] C. Quesne. Oscillator–Morse–Coulomb mappings and algebras for constant or position-dependent mass. Journal of Mathematical Physics, 49(2):022106, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i2/p022106_s1.

Cubitt:2008:DSB

[1337] Toby Cubitt, Ashley Montanaro, and Andreas Winter. On the dimension of subspaces with bounded Schmidt rank. Journal of Mathematical Physics, 49(2):022107, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i2/p022107_s1.

Wu:2008:CLO

[1338] Rebing Wu, Alexander Pechen, Herschel Rabitz, Michael Hsieh, and Benjamin Tsou. Control landscapes for observable preparation with open quantum systems. Journal of Mathematical Physics, 49(2):022108, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i2/p022108_s1.

Michel:2008:DDC

[1339] N. Michel. Direct demonstration of the completeness of the eigenstates of the Schrödinger equation with local and nonlocal potentials bearing a Coulomb tail. Journal of Mathematical Physics, 49(2):022109, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i2/p022109_s1.

Nagy:2008:SLO

[1340] Gabriel Nagy. On spectra of Lüders operations. Journal of Mathematical Physics, 49(2):022110, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i2/p022110_s1.

Strich:2008:PSE

[1341] Robert Strich. Passive states for essential observers. Journal of Mathematical Physics, 49(2):022301, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i2/p022301_s1.

Boumali:2008:EOD

[1342] A. Boumali. On the eigensolutions of the one-dimensional Duffin– Kemmer–Petiau oscillator. Journal of Mathematical Physics, 49 (2):022302, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i2/p022302_s1.

Gariel:2008:SAS

[1343] J. Gariel, G. Marcilhacy, and N. O. Santos. Stationary axisymmetric solutions involving a third order equation irreducible to Painlevé transcendents. Journal of Mathematical Physics, 49(2):022501, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i2/p022501_s1.

Hall:2008:PEC

[1344] G. S. Hall and D. P. Lonie. The principle of equivalence and cosmological metrics. Journal of Mathematical Physics, 49(2):022502, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i2/p022502_s1.

Marvan:2008:LEP

[1345] M. Marvan and O. Stolín. On local equivalence problem of space–times with two orthogonally transitive commuting Killing fields. Journal of Mathematical Physics, 49(2):022503, February 2008. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i2/p022503_s1.

Ding:2008:PMT

[1346] Lu Ding. Positive mass theorems for higher dimensional Lorentzian manifolds. Journal of Mathematical Physics, 49(2):022504, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i2/p022504_s1.

Przybylska:2008:DGO

[1347] Maria Przybylska. Differential Galois obstructions for integrability of homogeneous Newton equations. Journal of Mathematical Physics, 49(2):022701, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i2/p022701_s1.

Maccari:2008:CFC

[1348] Attilio Maccari. Chaotic, fractal, and coherent solutions for a new integrable system of equations in 2 + 1 dimensions. Journal of Mathematical Physics, 49(2):022702, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i2/p022702_s1.

Castilho:2008:VPS

[1349] César Castilho and Hélio Machado. The N-vortex problem on a symmetric ellipsoid: A perturbation approach. Journal of Mathematical Physics, 49(2):022703, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i2/p022703_s1.

Yavari:2008:GDE

[1350] Arash Yavari. On geometric discretization of elasticity. Journal of Mathematical Physics, 49(2):022901, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i2/p022901_s1.

Verrier:2008:NSH

[1351] P. E. Verrier and N. W. Evans. A new superintegrable Hamiltonian. Journal of Mathematical Physics, 49(2):022902, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i2/p022902_s1.

Acatrinei:2008:DRG

[1352] Ciprian Sorin Acatrinei. Dimensional reduction for generalized Poisson brackets. Journal of Mathematical Physics, 49(2):022903, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i2/p022903_s1.

Qin:2008:GES

[1353] Yuming Qin and Shenglan Wen. Global existence of spherically symmetric solutions for nonlinear compressible Navier–Stokes equations. Journal of Mathematical Physics, 49(2):023101, February 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i2/p023101_s1.

Horwood:2008:OSV

[1354] Joshua T. Horwood and Raymond G. McLenaghan. Orthogonal separation of variables for the Hamilton–Jacobi and wave equations in three-dimensional Minkowski space. Journal of Mathematical Physics, 49(2):023501, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i2/p023501_s1.

Ruijsenaars:2008:EVP

[1355] S. Ruijsenaars and L. Stodolsky. On the Euclidean version of the photon number integral. Journal of Mathematical Physics, 49(2):023502, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i2/p023502_s1.

Kengne:2008:ESE

[1356] E. Kengne, X. X. Liu, B. A. Malomed, S. T. Chui, and W. M. Liu. Explicit solutions to an effective Gross–Pitaevskii equation: Onedimensional Bose–Einstein condensate in specific traps. Journal of Mathematical Physics, 49(2):023503, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i2/p023503_s1.

DeTurck:2008:EGL

[1357] Dennis DeTurck and Herman Gluck. Electrodynamics and the Gauss linking integral on the 3-sphere and in hyperbolic 3-space. Journal of Mathematical Physics, 49(2):023504, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i2/p023504_s1.

Boyallian:2008:QRL

[1358] Carina Boyallian and Vanesa Meinardi. Quasifinite representations of the Lie superalgebra of quantum pseudodifferential operators. Journal of Mathematical Physics, 49(2):023505, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i2/p023505_s1.

Helein:2008:RFM

[1359] Frédéric Hélein. A representation formula for maps on supermanifolds. Journal of Mathematical Physics, 49(2):023506, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i2/p023506_s1.

delRio:2008:SAT

[1360] Rafael del Rio, Carmen Martinez, and Hermann Schulz-Baldes. Spectral averaging techniques for Jacobi matrices. Journal of Mathematical Physics, 49(2):023507, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i2/p023507_s1.

Daoud:2008:NSF

[1361] M. Daoud and A. Hamama. Noncommutative scalar fields from symplectic deformation. Journal of Mathematical Physics, 49(2):023508, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i2/p023508_s1.

Geloun:2008:BQE

[1362] Joseph Ben Geloun and M. Norbert Hounkonnou. Bosonic quasideterminants and eigenvalue problems of generalized spin-orbit operators. Journal of Mathematical Physics, 49(2):023509, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i2/p023509_s1.

Kulish:2008:QSA

[1363] P. P. Kulish, N. Manojlovic, and Z. Nagy. Quantum symmetry algebras of spin systems related to Temperley–Lieb R-matrices. Journal of Mathematical Physics, 49(2):023510, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i2/p023510_s1.

Celik:2008:HAS

[1364] Salih Celik and Ergün Yasar. The Hopf algebra structure of the Z^{3}-graded quantum supergroup GL_{q,j}(1|1). Journal of Mathematical Physics, 49(2):023511, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i2/p023511_s1.

Gal:2008:EDF

[1365] Tamás Gál. Erratum: “Differentiation of functionals with variables coupled by conservation constraints: Analysis through a fluid-dynamical model”, [J. Math. Phys. 48, 053520 (2007)]. Journal of Mathematical Physics, 49(2):029901, February 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i2/p029901_s1. See [?].

Salem:2008:SWD

[1366] Walid K. Abou Salem. Solitary wave dynamics in time-dependent potentials. Journal of Mathematical Physics, 49(3):032101, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i3/
p032101_s1.

Durran:2008:DCB

[1367] Richard Durran, Andrew Neate, and Aubrey Truman. The divine clockwork: Bohr’s correspondence principle and Nelson’s stochastic mechanics for the atomic elliptic state. Journal of Mathematical Physics, 49(3): 032102, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p032102_s1.

Rubensson:2008:ROI

[1368] Emanuel H. Rubensson, Elias Rudberg, and Pawel Salek. Rotations of occupied invariant subspaces in self-consistent field calculations. Journal of Mathematical Physics, 49(3):032103, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p032103_s1.

Stulpe:2008:SCE

[1369] Werner Stulpe and Paul Busch. The structure of classical extensions of quantum probability theory. Journal of Mathematical Physics, 49(3): 032104, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p032104_s1.

Boileau:2008:QRF

[1370] J.-C. Boileau, L. Sheridan, M. Laforest, and S. D. Bartlett. Quantum reference frames and the classification of rotationally invariant maps. Journal of Mathematical Physics, 49(3):032105, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i3/p032105_s1.

Carbone:2008:SPT

[1371] R. Carbone, F. Fagnola, J. C. García, and R. Quezada. Spectral properties of the two-photon absorption and emission process. Journal of Mathematical Physics, 49(3):032106, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p032106_s1.

Beylkin:2008:AWU

[1372] Gregory Beylkin, Martin J. Mohlenkamp, and Fernando Pérez. Approximating a wavefunction as an unconstrained sum of Slater determinants. Journal of Mathematical Physics, 49(3):032107, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i3/p032107_s1.

Lenzi:2008:SSE

[1373] E. K. Lenzi, B. F. de Oliveira, L. R. da Silva, and L. R. Evangelista. Solutions for a Schrödinger equation with a nonlocal term. Journal of Mathematical Physics, 49(3):032108, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p032108_s1.

Dierckx:2008:FQS

[1374] B. Dierckx, M. Fannes, and M. Pogorzelska. Fermionic quasifree states and maps in information theory. Journal of Mathematical Physics, 49(3): 032109, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p032109_s1.

Ali:2008:SAV

[1375] S. Twareque Ali and F. Bagarello. Supersymmetric associated vector coherent states and generalized Landau levels arising from two-dimensional supersymmetry. Journal of Mathematical Physics, 49(3):032110, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i3/p032110_s1.

Exner:2008:HPS

[1376] P. Exner and S. Kondej. Hiatus perturbation for a singular Schrödinger operator with an interaction supported by a curve in R^{3}. Journal of Mathematical Physics, 49(3):032111, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p032111_s1.

Hiai:2008:EEH

[1377] Fumio Hiai, Milán Mosonyi, and Tomohiro Ogawa. Error exponents in hypothesis testing for correlated states on a spin chain. Journal of Mathematical Physics, 49(3):032112, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p032112_s1.

Szarek:2008:GSQ

[1378] Stanislaw J. Szarek, Elisabeth Werner, and Karol Zyczkowski. Geometry of sets of quantum maps: A generic positive map acting on a highdimensional system is not completely positive. Journal of Mathematical Physics, 49(3):032113, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p032113_s1.

Molnar:2008:MSP

[1379] Lajos Molnár. Maps on states preserving the relative entropy. Journal of Mathematical Physics, 49(3):032114, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p032114_s1.

Albertsson:2008:PLP

[1380] Cecilia Albertsson, Ladislav Hlavatý, and Libor Snobl. On the Poisson– Lie T -plurality of boundary conditions. Journal of Mathematical Physics, 49(3):032301, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p032301_s1.

Douglas:2008:NCY

[1381] Michael R. Douglas, Robert L. Karp, Sergio Lukic, and René Reinbacher. Numerical Calabi–Yau metrics. Journal of Mathematical Physics, 49(3): 032302, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p032302_s1.

Tiwari:2008:TP

[1382] S. C. Tiwari. Topological photon. Journal of Mathematical Physics, 49(3):032303, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p032303_s1.

Finster:2008:RFP

[1383] Felix Finster. On the regularized fermionic projector of the vacuum. Journal of Mathematical Physics, 49(3):032304, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i3/p032304_s1.

Jakubassa-Amundsen:2008:HKE

[1384] D. H. Jakubassa-Amundsen. Heat kernel estimates and spectral properties of a pseudorelativistic operator with magnetic field. Journal of Mathematical Physics, 49(3):032305, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p032305_s1.

Garidi:2008:MVF

[1385] T. Garidi, J.-P. Gazeau, S. Rouhani, and M. V. Takook. “massless” vector field in de Sitter universe. Journal of Mathematical Physics, 49(3): 032501, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p032501_s1.

Bali:2008:BTC

[1386] Raj Bali and Naresh Kumar Chandnani. Bianchi type-I cosmological model for perfect fluid distribution in Lyra geometry. Journal of Mathematical Physics, 49(3):032502, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i3/p032502_s1.

Girelli:2008:THG

[1387] F. Girelli, H. Pfeiffer, and E. M. Popescu. Topological higher gauge theory: From BF to BF CG theory. Journal of Mathematical Physics, 49(3): 032503, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p032503_s1.

Antippa:2008:HDH

[1388] Adel F. Antippa and Daniel M. Dubois. Hyperincursive discrete harmonic oscillator. Journal of Mathematical Physics, 49(3):032701, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i3/p032701_s1.

Llibre:2008:DIA

[1389] Jaume Llibre and Clàudia Valls. Darboux integrability and algebraic invariant surfaces for the Rikitake system. Journal of Mathematical Physics, 49(3):032702, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p032702_s1.

Carinena:2008:HOR

[1390] José F. Cariñena, Manuel F. Rañada, and Mariano Santander. The harmonic oscillator on Riemannian and Lorentzian configuration spaces of constant curvature. Journal of Mathematical Physics, 49(3):032703, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p032703_s1.

Anzaldo-Meneses:2008:DSN

[1391] A. Anzaldo-Meneses and F. Monroy-Pérez. Dynamical systems and nilpotent sub-Riemannian geometry. Journal of Mathematical Physics, 49(3):032704, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p032704_s1.

Crampin:2008:RPN

[1392] M. Crampin and T. Mestdag. Routh’s procedure for non-Abelian symmetry groups. Journal of Mathematical Physics, 49(3):032901, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i3/p032901_s1.

Adams:2008:PCR

[1393] A. J. Adams. Postbuckling of circular rings: An analytical solution. Journal of Mathematical Physics, 49(3):032902, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i3/p032902_s1.

Ochiai:2008:NDD

[1394] T. Ochiai, U. Leonhardt, and J. C. Nacher. A novel design of dielectric perfect invisibility devices. Journal of Mathematical Physics, 49(3):

032903, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p032903_s1.

Shcherbina:2008:DSL

[1395] Mariya Shcherbina. Double scaling limit for matrix models with nonanalytic potentials. Journal of Mathematical Physics, 49(3):033501, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i3/p033501_s1.

Jureit:2008:CIA

[1396] Jan-Hendrik Jureit and Christoph A. Stephan. On a classification of irreducible almost-commutative geometries IV. Journal of Mathematical Physics, 49(3):033502, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p033502_s1.

Yong:2008:ICP

[1397] Wen-An Yong. An interesting class of partial differential equations. Journal of Mathematical Physics, 49(3):033503, March 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i3/p033503_s1.

Ferreira:2008:RID

[1398] Erasmo Ferreira and Javier Sesma. Representation of integral dispersion relations by local forms. Journal of Mathematical Physics, 49(3): 033504, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p033504_s1.

Camargo:2008:DFO

[1399] R. Figueiredo Camargo, Ary O. Chiacchio, and E. Capelas de Oliveira. Differentiation to fractional orders and the fractional telegraph equation. Journal of Mathematical Physics, 49(3):033505, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i3/p033505_s1.

Becker:2008:SFX

[1400] Peter A. Becker. Spectral formation in X-ray pulsars and associated identities involving the Laguerre polynomials. Journal of Mathematical Physics, 49(3):033506, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p033506_s1.

Weimar-Woods:2008:CIL

[1401] Evelyn Weimar-Woods. Contractions of invariants of Lie algebras with applications to classical inhomogeneous Lie algebras. Journal of Mathematical Physics, 49(3):033507, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i3/p033507_s1.

Choup:2008:EEL

[1402] Leonard N. Choup. Edgeworth expansion of the largest eigenvalue distribution function of Gaussian unitary ensemble revisited. Journal of Mathematical Physics, 49(3):033508, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p033508_s1.

Sitarz:2008:TDO

[1403] Andrzej Sitarz. Twisted Dirac operators over quantum spheres. Journal of Mathematical Physics, 49(3):033509, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p033509_s1.

Iguri:2008:SRG

[1404] S. M. Iguri and M. A. Castagnino. Some remarks on the Gelfand– Naimark–Segal representations of topological ^{∗}-algebras. Journal of Mathematical Physics, 49(3):033510, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p033510_s1.

Ma:2008:SIA

[1405] Wen-Xiu Ma, Jing-Song He, and Zhen-Yun Qin. A supertrace identity and its applications to superintegrable systems. Journal of Mathematical Physics, 49(3):033511, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i3/p033511_s1.

Akant:2008:ESG

[1406] Levent Akant. Equivariant symplectic geometry of gauge fixing in Yang– Mills theory. Journal of Mathematical Physics, 49(3):033512, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i3/p033512_s1.

Wang:2008:GNQ

[1407] Hai-Jun Wang. General nonlocality in quantum fields. Journal of Mathematical Physics, 49(3):033513, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v49/i3/p033513_s1. See comment

[2202] and erratum [2214].

Minami:2008:FES

[1408] Kazuhiko Minami. The free energies of six-vertex models and the nequivalence relation. Journal of Mathematical Physics, 49(3):033514, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i3/p033514_s1.

Batalin:2008:OSC

[1409] Igor A. Batalin and Klaus Bering. Odd scalar curvature in field-antifield formalism. Journal of Mathematical Physics, 49(3):033515, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i3/
p033515_s1.

Guo:2008:BGE

[1410] Zhengguang Guo. Blow up, global existence, and infinite propagation speed for the weakly dissipative Camassa–Holm equation. Journal of Mathematical Physics, 49(3):033516, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p033516_s1.

Sahadevan:2008:BTF

[1411] R. Sahadevan and S. Rajakumar. Bilinear, trilinear forms, and exact solution of certain fourth order integrable difference equations. Journal of Mathematical Physics, 49(3):033517, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p033517_s1.

Kocabova:2008:GBA

[1412] P. Kocábová and P. Stovícek. Generalized Bloch analysis and propagators on Riemannian manifolds with a discrete symmetry. Journal of Mathematical Physics, 49(3):033518, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i3/p033518_s1.

Chmielinski:2008:PWE

[1413] Jacek Chmieliński, Dijana Ilisević, Mohammad Sal Moslehian, and Ghadir Sadeghi. Perturbation of the Wigner equation in inner product C^{∗}-modules. Journal of Mathematical Physics, 49(3):033519, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i3/p033519_s1.

Petalidou:2008:AGQ

[1414] Fani Petalidou. Addendum: “On the geometric quantization of twisted Poisson manifolds”, [J. Math. Phys. 48, 083502 (2007)]. Journal of Mathematical Physics, 49(3):033520, March 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i3/p033520_s1. See [1145].

Fakhri:2008:CBG

[1415] H. Fakhri and A. Dehghani. Comment on “Barut–Girardello and Klauder–Perelomov coherent states for the Kravchuk functions”, [J. Math. Phys. 48, 112106 (2007)]. Journal of Mathematical Physics, 49(4): 042101, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i4/p042101_s1. See [1233].

Brandao:2008:CED

[1416] F. G. S. L. Brandão and J. Eisert. Correlated entanglement distillation and the structure of the set of undistillable states. Journal of Mathematical Physics, 49(4):042102, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i4/p042102_s1.

Benatti:2008:EIB

[1417] Fabio Benatti, Alexandra M. Liguori, and Adam Nagy. Environment induced bipartite entanglement. Journal of Mathematical Physics, 49(4): 042103, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i4/p042103_s1.

Robert:2008:SVG

[1418] D. Robert and A. V. Smilga. Supersymmetry versus ghosts. Journal of Mathematical Physics, 49(4):042104, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i4/p042104_s1.

DellAntonio:2008:JEP

[1419] Gianfausto Dell’Antonio, Rodolfo Figari, and Alessandro Teta. Joint excitation probability for two harmonic oscillators in one dimension and the Mott problem. Journal of Mathematical Physics, 49(4):042105, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i4/p042105_s1.

Rotteler:2008:DDS

[1420] Martin Rötteler. Dynamical decoupling schemes derived from Hamilton cycles. Journal of Mathematical Physics, 49(4):042106, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i4/
p042106_s1.

Bouguerra:2008:TDP

[1421] Y. Bouguerra, A. Bounames, M. Maamache, and Y. Saadi. Timedependent Pauli equation in the presence of the Aharonov–Bohm effect. Journal of Mathematical Physics, 49(4):042107, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i4/p042107_s1.

Dzhunushaliev:2008:NSH

[1422] Vladimir Dzhunushaliev. Nonassociativity, supersymmetry, and hidden variables. Journal of Mathematical Physics, 49(4):042108, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i4/
p042108_s1.

Suzuki:2008:PSM

[1423] Akito Suzuki. Physical subspace in a model of the quantized electromagnetic field coupled to an external field with an indefinite metric. Journal of Mathematical Physics, 49(4):042301, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i4/p042301_s1.

Chaichian:2008:NFA

[1424] M. Chaichian, P. P. Kulish, A. Tureanu, R. B. Zhang, and Xiao Zhang. Noncommutative fields and actions of twisted Poincaré algebra. Journal of Mathematical Physics, 49(4):042302, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i4/p042302_s1.

daVeiga:2008:EWD

[1425] Paulo A. Faria da Veiga and Michael O’Carroll. Eightfold way from dynamical first principles in strongly coupled lattice quantum chromodynamics. Journal of Mathematical Physics, 49(4):042303, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i4/
p042303_s1.

Castro:2008:EGA

[1426] Carlos Castro. The Euclidean gravitational action as black hole entropy, singularities, and spacetime voids. Journal of Mathematical Physics, 49(4):042501, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i4/p042501_s1. See comments [1834].

Garraffo:2008:GSV

[1427] C. Garraffo, G. Giribet, E. Gravanis, and S. Willison. Gravitational solitons and C^{0} vacuum metrics in five-dimensional Lovelock gravity. Journal of Mathematical Physics, 49(4):042502, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i4/p042502_s1.

Clement:2008:SFD

[1428] Gérard Clément. The symmetries of five-dimensional minimal supergravity reduced to three dimensions. Journal of Mathematical Physics, 49(4): 042503, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i4/p042503_s1. See erratum [1566].

Giambo:2008:GBH

[1429] Roberto Giambò, Fabio Giannoni, and Giulio Magli. Genericity of black hole formation in the gravitational collapse of homogeneous selfinteracting scalar fields. Journal of Mathematical Physics, 49(4):042504, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i4/p042504_s1.

Sugny:2008:FHM

[1430] D. Sugny, P. Mardesić, M. Pelletier, A. Jebrane, and H. R. Jauslin. Fractional Hamiltonian monodromy from a Gauss–Manin monodromy. Journal of Mathematical Physics, 49(4):042701, April 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i4/p042701_s1.

Zak:2008:QIM

[1431] Michail Zak. Quantum-inspired maximizer. Journal of Mathematical Physics, 49(4):042702, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i4/p042702_s1.

Barnich:2008:SCA

[1432] Glenn Barnich and Geoffrey Compère. Surface charge algebra in gauge theories and thermodynamic integrability. Journal of Mathematical Physics, 49(4):042901, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i4/p042901_s1.

Pasquero:2008:NUC

[1433] Stefano Pasquero. Nonideal unilateral constraints in impulsive mechanics: A geometric approach. Journal of Mathematical Physics, 49(4): 042902, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i4/p042902_s1.

Santoprete:2008:GHO

[1434] Manuele Santoprete. Gravitational and harmonic oscillator potentials on surfaces of revolution. Journal of Mathematical Physics, 49(4):042903, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i4/p042903_s1.

Heibig:2008:RSS

[1435] Arnaud Heibig and Liviu Iulian Palade. On the rest state stability of an objective fractional derivative viscoelastic fluid model. Journal of Mathematical Physics, 49(4):043101, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i4/p043101_s1.

Scardicchio:2008:EOD

[1436] A. Scardicchio, F. H. Stillinger, and S. Torquato. Estimates of the optimal density of sphere packings in high dimensions. Journal of Mathematical Physics, 49(4):043301, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i4/p043301_s1.

Pule:2008:SBM

[1437] Joe V. Pulé, André F. Verbeure, and Valentin A. Zagrebnov. On solvable boson models. Journal of Mathematical Physics, 49(4):043302, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i4/p043302_s1.

Chevalier:2008:RDU

[1438] C. Chevalier and F. Debbasch. Relativistic diffusions: A unifying approach. Journal of Mathematical Physics, 49(4):043303, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i4/
p043303_s1.

Abe:2008:BLE

[1439] Mitsuo Abe and Katsunori Kawamura. Branching laws for endomorphisms of fermions and the Cuntz algebra O_{2}. Journal of Mathematical Physics, 49(4):043501, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i4/p043501_s1.

Grundland:2008:SVH

[1440] A. M. Grundland and A. J. Hariton. Supersymmetric version of a hydrodynamic system in Riemann invariants and its solutions. Journal of Mathematical Physics, 49(4):043502, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i4/p043502_s1.

Rumanov:2008:CBT

[1441] Igor Rumanov. The correspondence between Tracy–Widom and Adler– Shiota–van Moerbeke approaches in random matrix theory: The Gaussian case. Journal of Mathematical Physics, 49(4):043503, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i4/
p043503_s1.

Vacaru:2008:NRF

[1442] Sergiu I. Vacaru. Nonholonomic Ricci flows. II. Evolution equations and dynamics. Journal of Mathematical Physics, 49(4):043504, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i4/p043504_s1.

Iwai:2008:RVE

[1443] Toshihiro Iwai and Hidetaka Yamaoka. Rotational-vibrational energy spectra of triatomic molecules near relative equilibria. Journal of Mathematical Physics, 49(4):043505, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i4/p043505_s1.

Donnelly:2008:MNE

[1444] Harold Donnelly. Maxima of Neumann eigenfunctions. Journal of Mathematical Physics, 49(4):043506, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i4/p043506_s1.

Liu:2008:IBV

[1445] Jianli Liu and Yi Zhou. Initial-boundary value problem for the equation of timelike extremal surfaces in Minkowski space. Journal of Mathematical Physics, 49(4):043507, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i4/p043507_s1.

Zhang:2008:FOF

[1446] Zhiyue Zhang and Dingwen Deng. A fourth-order finite difference solver for nerve conduction equation in rectangular domains. Journal of Mathematical Physics, 49(4):043508, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i4/p043508_s1.

Marucho:2008:NBS

[1447] M. Marucho. Non-Borel summable Φ^{4} theory in zero dimension: A toy model for testing numerical and analytical methods. Journal of Mathematical Physics, 49(4):043509, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i4/p043509_s1.

Coffey:2008:TDS

[1448] Mark W. Coffey. On a three-dimensional symmetric Ising tetrahedron and contributions to the theory of the dilogarithm and Clausen functions. Journal of Mathematical Physics, 49(4):043510, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i4/p043510_s1.

Kim:2008:SCC

[1449] Ki hong Kim and Jack J. Yoh. Shock compression of condensed matter using multimaterial reactive ghost fluid method. Journal of Mathematical Physics, 49(4):043511, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i4/p043511_s1.

Xu:2008:RNS

[1450] Runzhang Xu and Yacheng Liu. Remarks on nonlinear Schrödinger equation with harmonic potential. Journal of Mathematical Physics, 49(4): 043512, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i4/p043512_s1.

Chang:2008:VOE

[1451] Wen-Jing Chang and Xiang-Mao Ding. On the vertex operators of the elliptic quantum algebra U_{q,p}( ˆ∫↕_{2})_{k}. Journal of Mathematical Physics, 49(4):043513, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i4/p043513_s1.

Balinsky:2008:DSI

[1452] A. Balinsky, W. D. Evans, and Y. Saito. Dirac–Sobolev inequalities and estimates for the zero modes of massless Dirac operators. Journal of Mathematical Physics, 49(4):043514, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i4/p043514_s1.

Donninger:2008:MSS

[1453] Roland Donninger and Peter C. Aichelburg. On the mode stability of a self-similar wave map. Journal of Mathematical Physics, 49(4):043515, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i4/p043515_s1.

Bering:2008:SSC

[1454] K. Bering. Semidensities, second-class constraints, and conversion in anti-Poisson geometry. Journal of Mathematical Physics, 49(4):043516, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i4/p043516_s1.

Lee:2008:ICL

[1455] Chiun-Chang Lee and Tai-Chia Lin. Incompressible and compressible limits of two-component Gross–Pitaevskii equations with rotating fields and trap potentials. Journal of Mathematical Physics, 49(4):043517, April 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i4/p043517_s1.

Fakhri:2008:CGK

[1456] H. Fakhri and A. Dehghani. Comments on “Gazeau–Klauder coherent states for trigonometric Rosen–Morse potential”, [J. Math. Phys. 49, 022104 (2008)]. Journal of Mathematical Physics, 49(5):052101, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p052101_s1. See [1334].

Khrennikov:2008:RNL

[1457] Andrei Khrennikov. The role of von Neumann and Lüders postulates in the Einstein, Podolsky, and Rosen considerations: Comparing measurements with degenerate and nondegenerate spectra. Journal of Mathematical Physics, 49(5):052102, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i5/p052102_s1.

Clark:2008:DRG

[1458] Jeremy Clark. Decoherence rates for Galilean covariant dynamics. Journal of Mathematical Physics, 49(5):052103, May 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i5/p052103_s1.

Shikano:2008:OCM

[1459] Yutaka Shikano and Akio Hosoya. Optimal covariant measurement of momentum on a half line in quantum mechanics. Journal of Mathematical Physics, 49(5):052104, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p052104_s1.

Dong:2008:ACD

[1460] Jianping Dong and Mingyu Xu. Applications of continuity and discontinuity of a fractional derivative of the wave functions to fractional quantum mechanics. Journal of Mathematical Physics, 49(5):052105, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p052105_s1.

Gudder:2008:CSP

[1461] Stan Gudder and Frédéric Latrémolière. Characterization of the sequential product on quantum effects. Journal of Mathematical Physics, 49(5): 052106, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i5/p052106_s1.

Jaffe:2008:RPM

[1462] Arthur Jaffe and Gordon Ritter. Reflection positivity and monotonicity. Journal of Mathematical Physics, 49(5):052301, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/p052301_s1.

Kelnhofer:2008:AGT

[1463] Gerald Kelnhofer. Abelian gauge theories on compact manifolds and the Gribov ambiguity. Journal of Mathematical Physics, 49(5):052302, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p052302_s1.

Brouder:2008:RSS

[1464] Christian Brouder and Michael Dütsch. Relating on-shell and off-shell formalisms in perturbative quantum field theory. Journal of Mathematical Physics, 49(5):052303, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i5/p052303_s1.

Bruning:2008:SLM

[1465] E. Brüning and S. Nagamachi. Solution of a linearized model of Heisenberg’s fundamental equation II. Journal of Mathematical Physics, 49(5): 052304, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i5/p052304_s1.

Al-Badawi:2008:SDE

[1466] A. Al-Badawi and I. Sakalli. Solution of the Dirac equation in the rotating Bertotti–Robinson spacetime. Journal of Mathematical Physics, 49(5):

052501, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i5/p052501_s1.

Ulanovskii:2008:GPC

[1467] M. Ulanovskii. On global properties of certain homogeneous exact solutions of Einstein field equations. Journal of Mathematical Physics, 49(5): 052502, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i5/p052502_s1.

Calogero:2008:NCS

[1468] Francesco Calogero. A new class of solvable dynamical systems. Journal of Mathematical Physics, 49(5):052701, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i5/p052701_s1.

Jones:2008:CCQ

[1469] Robert T. Jones. Central configurations with a quasihomogeneous potential function. Journal of Mathematical Physics, 49(5):052901, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p052901_s1.

Abdesselam:2008:NEV

[1470] B. Abdesselam and A. Chakrabarti. A new eight vertex model and higher dimensional, multiparameter generalizations. Journal of Mathematical Physics, 49(5):053301, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p053301_s1.

Imamura:2008:CFS

[1471] T. Imamura and T. Sasamoto. Correlation function of the Schur process with a fixed final partition. Journal of Mathematical Physics, 49(5): 053302, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i5/p053302_s1.

Ruiz:2008:KLS

[1472] Jean Ruiz and Marc Wouts. On the Kertész line: Some rigorous bounds. Journal of Mathematical Physics, 49(5):053303, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/p053303_s1.

Bogoyavlenskij:2008:LVS

[1473] Oleg Bogoyavlenskij, Yoshiaki Itoh, and Tetsuyuki Yukawa. Lotka– Volterra systems integrable in quadratures. Journal of Mathematical Physics, 49(5):053501, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p053501_s1.

Arnlind:2008:RTA

[1474] Joakim Arnlind. Representation theory of C-algebras for a higher-order class of spheres and tori. Journal of Mathematical Physics, 49(5):053502, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p053502_s1.

Odake:2008:OPH

[1475] Satoru Odake and Ryu Sasaki. Orthogonal polynomials from Hermitian matrices. Journal of Mathematical Physics, 49(5):053503, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p053503_s1.

Sahai:2008:MCA

[1476] Vivek Sahai and Sarasvati Yadav. On models of certain p, q-algebra representations: The p, q-oscillator algebra. Journal of Mathematical Physics, 49(5):053504, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p053504_s1.

Tkadlec:2008:AOE

[1477] Josef Tkadlec. Atomistic and orthoatomistic effect algebras. Journal of Mathematical Physics, 49(5):053505, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i5/p053505_s1.

Ma:2008:DPT

[1478] Tian Ma and Shouhong Wang. Dynamic phase transitions for ferromagnetic systems. Journal of Mathematical Physics, 49(5):053506, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p053506_s1.

Fyodorov:2008:HHS

[1479] Y. V. Fyodorov, Y. Wei, and M. R. Zirnbauer. Hyperbolic Hubbard– Stratonovich transformation made rigorous. Journal of Mathematical Physics, 49(5):053507, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p053507_s1.

Wang:2008:LEC

[1480] Hai-Jun Wang. The Lorentz extension as consequence of the family symmetry. Journal of Mathematical Physics, 49(5):053508, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/
p053508_s1.

Dattoli:2008:HAU

[1481] G. Dattoli, D. Levi, and P. Winternitz. Heisenberg algebra, umbral calculus and orthogonal polynomials. Journal of Mathematical Physics, 49 (5):053509, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i5/p053509_s1.

Wu:2008:CDV

[1482] Xian Wu. On a class of damped vibration problems with obstacles and without direct variational structures. Journal of Mathematical Physics, 49(5):053510, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p053510_s1.

Park:2008:EAS

[1483] Jong Yeoul Park and Tae Gab Ha. Existence and asymptotic stability for the semilinear wave equation with boundary damping and source term. Journal of Mathematical Physics, 49(5):053511, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/

p053511_s1. See erratum [1940].

Li:2008:RSV

[1484] Junbo Li and Yucai Su. Representations of the Schrödinger–Virasoro algebras. Journal of Mathematical Physics, 49(5):053512, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/
p053512_s1.

Hu:2008:LHS

[1485] Liangzhong Hu. The Laplacian on homogeneous spaces. Journal of Mathematical Physics, 49(5):053513, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i5/p053513_s1.

Conca:2008:BCP

[1486] C. Conca, J. San Martín, L. Smaranda, and M. Vanninathan. On Burnett coefficients in periodic media in low contrast regime. Journal of Mathematical Physics, 49(5):053514, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i5/p053514_s1.

Doring:2008:TFTa

[1487] A. Döring and C. J. Isham. A topos foundation for theories of physics: I. Formal languages for physics. Journal of Mathematical Physics, 49(5): 053515, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i5/p053515_s1.

Doring:2008:TFTb

[1488] A. Döring and C. J. Isham. A topos foundation for theories of physics: II. Daseinisation and the liberation of quantum theory. Journal of Mathematical Physics, 49(5):053516, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i5/p053516_s1.

Doring:2008:TFTc

[1489] A. Döring and C. J. Isham. A topos foundation for theories of physics: III. The representation of physical quantities with arrows δ̆^{o}(A): \underline{Σ} → \underline{R}^{⪰}. Journal of Mathematical Physics, 49(5):053517, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/p053517_s1.

Doring:2008:TFTd

[1490] A. Döring and C. J. Isham. A topos foundation for theories of physics: IV. Categories of systems. Journal of Mathematical Physics, 49(5):053518, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p053518_s1.

Gao:2008:NID

[1491] Ya-Jun Gao. New infinite-dimensional quadruple symmetry groups for the general symplectic gravity model. Journal of Mathematical Physics, 49(5):053519, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p053519_s1.

Bogoyavlenskij:2008:FIP

[1492] O. Bogoyavlenskij and P. Reynolds. Form-invariant Poisson brackets of hydrodynamic type with several spatial variables. Journal of Mathematical Physics, 49(5):053520, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i5/p053520_s1.

El-Nabulsi:2008:FAV

[1493] Rami Ahmad El-Nabulsi and Delfim F. M. Torres. Fractional actionlike variational problems. Journal of Mathematical Physics, 49(5):053521, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i5/p053521_s1.

Bagarello:2008:AQD

[1494] F. Bagarello. O^{⋆}-algebras and quantum dynamics: Some existence results. Journal of Mathematical Physics, 49(5):053522, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/
p053522_s1.

Skorupski:2008:PIA

[1495] Andrzej A. Skorupski. Phase integral approximation for coupled ordinary differential equations of the Schrödinger type. Journal of Mathematical Physics, 49(5):053523, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p053523_s1.

Dolya:2008:QES

[1496] S. N. Dolya. Quasi-exactly solvable models based on special functions. Journal of Mathematical Physics, 49(5):053524, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/p053524_s1.

Luhn:2008:QFS

[1497] Christoph Luhn and Pierre Ramond. Quintics with finite simple symmetries. Journal of Mathematical Physics, 49(5):053525, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/p053525_s1.

Birkandan:2008:CDE

[1498] T. Birkandan and M. Hortaçsu. Comment on “Dirac equation in the background of the Nutku helicoid metric”, [J. Math. Phys. 48, 092301 (2007)]. Journal of Mathematical Physics, 49(5):054101, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i5/

p054101_s1. See [1166].

Chen:2008:ECM

[1499] Bang-Yen Chen and Franki Dillen. Erratum: “Classification of marginally trapped Lagrangian surfaces in Lorentzian complex space form”, [J. Math. Phys. 48, 013509 (2007)]. Journal of Mathematical Physics, 49(5):059901, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i5/p059901_s1. See [893].

Lundholm:2008:GSQ

[1500] D. Lundholm. On the geometry of supersymmetric quantum mechanical systems. Journal of Mathematical Physics, 49(6):062101, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i6/
p062101_s1.

Chen:2008:VDS

[1501] Goong Chen, Zhonghai Ding, Alain Perronnet, and Zhigang Zhang. Visualization and dimensional scaling for some three-body problems in atomic and molecular quantum mechanics. Journal of Mathematical Physics, 49 (6):062102, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i6/p062102_s1.

Ma:2008:STB

[1502] Li Ma and Lin Zhao. Sharp thresholds of blow-up and global existence for the coupled nonlinear Schrödinger system. Journal of Mathematical Physics, 49(6):062103, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i6/p062103_s1.

Aleixo:2008:PSS

[1503] A. N. F. Aleixo and A. B. Balantekin. Purely squeezed states for quantum deformed systems. Journal of Mathematical Physics, 49(6):062104, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v49/i6/p062104_s1. See publisher’s note [1588].

Wehner:2008:HEU

[1504] Stephanie Wehner and Andreas Winter. Higher entropic uncertainty relations for anti-commuting observables. Journal of Mathematical Physics, 49(6):062105, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i6/p062105_s1.

Tho:2008:IIC

[1505] Nguyen Vien Tho. Interaction of imaginary-charge-carrying dyon with particles. Journal of Mathematical Physics, 49(6):062301, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i6/
p062301_s1.

Dimock:2008:MTA

[1506] J. Dimock. More transition amplitudes on the Riemann sphere. Journal of Mathematical Physics, 49(6):062302, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i6/p062302_s1.

Krishnaswami:2008:SDO

[1507] Govind S. Krishnaswami. Schwinger–Dyson operators as invariant vector fields on a matrix model analog of the group of loops. Journal of Mathematical Physics, 49(6):062303, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i6/p062303_s1.

Dolan:2008:RLE

[1508] P. Dolan and A. Gerber. The Riemann–Lanczos equations in general relativity and their integrability. Journal of Mathematical Physics, 49(6): 062501, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i6/p062501_s1.

Bini:2008:MFP

[1509] D. Bini, C. Cherubini, and A. Geralico. Massless field perturbations of the spinning C metric. Journal of Mathematical Physics, 49(6):062502, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i6/p062502_s1.

Minguzzi:2008:NIC

[1510] E. Minguzzi. Non-imprisonment conditions on spacetime. Journal of Mathematical Physics, 49(6):062503, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i6/p062503_s1.

Claussen:2008:TER

[1511] Jens Christian Claussen. Time evolution of the rule 150 cellular automaton activity from a Fibonacci iteration. Journal of Mathematical Physics, 49(6):062701, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i6/p062701_s1.

Diaz:2008:ICF

[1512] Rafael Díaz and Lorenzo Leal. Invariants from classical field theory. Journal of Mathematical Physics, 49(6):062901, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i6/p062901_s1.

Barbero-Linan:2008:UFN

[1513] María Barbero-Liñán, Arturo Echeverría-Enríquez, David Martín de Diego, Miguel C. Muñoz-Lecanda, and Narciso Román-Roy. Unified formalism for nonautonomous mechanical systems. Journal of Mathematical Physics, 49(6):062902, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i6/p062902_s1.

Harnad:2008:MHS

[1514] John Harnad and Jacques Hurtubise. Multi-Hamiltonian structures for r-matrix systems. Journal of Mathematical Physics, 49(6):062903, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i6/p062903_s1.

Gryc:2008:HCC

[1515] William E. Gryc. On the holonomy of the Coulomb connection over manifolds with boundary. Journal of Mathematical Physics, 49(6):062904, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i6/p062904_s1.

Baty:2008:NAJ

[1516] Roy S. Baty, F. Farassat, and Don H. Tucker. Nonstandard analysis and jump conditions for converging shock waves. Journal of Mathematical Physics, 49(6):063101, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i6/p063101_s1.

Khare:2008:DWP

[1517] Avinash Khare and Avadh Saxena. Domain wall and periodic solutions of a coupled ϕ^{6} model. Journal of Mathematical Physics, 49(6): 063301, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i6/p063301_s1.

Bhattacharya:2008:TDC

[1518] Satyabrata Bhattacharya, Suman Kumar Banik, Sudip Chattopadhyay, and Jyotipratim Ray Chaudhuri. Time dependent current in a nonstationary environment: A microscopic approach. Journal of Mathematical Physics, 49(6):063302, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i6/p063302_s1.

Droz-Vincent:2008:SPE

[1519] Philippe Droz-Vincent. Scalar products of elementary distributions. Journal of Mathematical Physics, 49(6):063501, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i6/p063501_s1.

Zenchuk:2008:LDR

[1520] A. I. Zenchuk. Lower-dimensional reductions of GL(M, C) self-dual Yang Mills equation: Solutions with break of wave profiles. Journal of Mathematical Physics, 49(6):063502, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i6/p063502_s1.

Kwon:2008:SSA

[1521] Namhee Kwon and Daniel S. Sage. Subrepresentation semirings and an analog of 6j-symbols. Journal of Mathematical Physics, 49(6):063503, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i6/p063503_s1.

Carmeli:2008:ECP

[1522] Claudio Carmeli, Teiko Heinosaari, Juha-Pekka Pellonpää, and Alessandro Toigo. Extremal covariant positive operator valued measures: The case of a compact symmetry group. Journal of Mathematical Physics, 49 (6):063504, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i6/p063504_s1.

Zhenyun:2008:GAL

[1523] Qin Zhenyun. A generalized Ablowitz–Ladik hierarchy, multi-Hamiltonian structure and Darboux transformation. Journal of Mathematical Physics, 49(6):063505, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i6/p063505_s1.

Campoamor-Stursberg:2008:KSL

[1524] R. Campoamor-Stursberg and M. Rausch de Traubenberg. Kinematical superalgebras and Lie algebras of order 3. Journal of Mathematical Physics, 49(6):063506, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i6/p063506_s1.

Rassias:2008:AHD

[1525] John Michael Rassias and Hark-Mahn Kim. Approximate homomorphisms and derivations between C^{∗}-ternary algebras. Journal of Mathematical Physics, 49(6):063507, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i6/p063507_s1.

Ancarani:2008:DOC

[1526] L. U. Ancarani and G. Gasaneo. Derivatives of any order of the confluent hypergeometric function _{1}F_{1}(a, b, z) with respect to the parameter a or b. Journal of Mathematical Physics, 49(6):063508, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i6/p063508_s1.

Daboul:2008:GCA

[1527] Claudia Daboul, Jamil Daboul, and Marc de Montigny. Gradings and contractions of affine Kac–Moody algebras. Journal of Mathematical Physics, 49(6):063509, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i6/p063509_s1.

Lehner:2008:CEM

[1528] C. Lehner, T. Wettig, T. Guhr, and Y. Wei. Character expansion method for supergroups and extended superversions of the Leutwyler–Smilga and Berezin–Karpelevich integrals. Journal of Mathematical Physics, 49(6): 063510, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i6/p063510_s1.

Miao:2008:DSP

[1529] Dong Miao. The discrete spectrum of the periodic Maxwell operator perturbed by a decreasing potential. Journal of Mathematical Physics, 49 (6):063511, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i6/p063511_s1.

Bastos:2008:WWF

[1530] Catarina Bastos, Orfeu Bertolami, Nuno Costa Dias, and João Nuno Prata. Weyl–Wigner formulation of noncommutative quantum mechanics. Journal of Mathematical Physics, 49(7):072101, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i7/p072101_s1.

Meiler:2008:SCP

[1531] Maria Meiler, Ricardo Cordero-Soto, and Sergei K. Suslov. Solution of the Cauchy problem for a time-dependent Schrödinger equation. Journal of Mathematical Physics, 49(7):072102, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i7/p072102_s1.

Konstein:2008:DCE

[1532] S. E. Konstein and I. V. Tyutin. Deformations and central extensions of the antibracket superalgebra. Journal of Mathematical Physics, 49(7): 072103, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i7/p072103_s1.

Mosonyi:2008:ADM

[1533] Milán Mosonyi, Fumio Hiai, Tomohiro Ogawa, and Mark Fannes. Asymptotic distinguishability measures for shift-invariant quasifree states of fermionic lattice systems. Journal of Mathematical Physics, 49(7): 072104, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i7/p072104_s1.

Gudder:2008:QMC

[1534] Stanley Gudder. Quantum Markov chains. Journal of Mathematical Physics, 49(7):072105, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i7/p072105_s1.

Neto:2008:MEW

[1535] Antônio Francisco Neto, Michael O’Carroll, and Paulo A. Faria da Veiga. Mesonic eightfold way from dynamics and confinement in strongly coupled lattice quantum chromodynamics. Journal of Mathematical Physics, 49(7):072301, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i7/p072301_s1.

Rybkin:2008:ERC

[1536] Alexei Rybkin. On the evolution of a reflection coefficient under the Korteweg–de Vries flow. Journal of Mathematical Physics, 49(7):072701, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p072701_s1.

Blaszak:2008:IDS

[1537] Maciej Blaszak, Metin Gürses, Burcu Silindir, and Blazej M. Szablikowski. Integrable discrete systems on R and related dispersionless systems. Journal of Mathematical Physics, 49(7):072702, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i7/
p072702_s1.

Szydlowski:2008:PFR

[1538] Marek Szydlowski and Orest Hrycyna. Paths of Friedmann–Robertson– Walker brane models. Journal of Mathematical Physics, 49(7):072703, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p072703_s1.

Jia:2008:ICH

[1539] M. Jia, P. Liu, and S. Y. Lou. An integrable coupled hybrid lattice and its exact solutions. Journal of Mathematical Physics, 49(7):072901, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p072901_s1.

Niedermaier:2008:LEH

[1540] M. Niedermaier and E. Seiler. The large N expansion in hyperbolic sigma models. Journal of Mathematical Physics, 49(7):073301, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p073301_s1.

Lv:2008:SMF

[1541] Long-Jin Lv, Jian-Bin Xiao, Fu-Yao Ren, and Lei Gao. Solutions for multidimensional fractional anomalous diffusion equations. Journal of Mathematical Physics, 49(7):073302, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i7/p073302_s1.

Jafarizadeh:2008:CER

[1542] M. A. Jafarizadeh, R. Sufiani, and S. Jafarizadeh. Calculating effective resistances on underlying networks of association schemes. Journal of Mathematical Physics, 49(7):073303, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i7/p073303_s1.

Ma:2008:DMP

[1543] Tian Ma and Shouhong Wang. Dynamic model and phase transitions for liquid helium. Journal of Mathematical Physics, 49(7):073304, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p073304_s1.

Kang:2008:NAE

[1544] Nam Lyong Kang and Sang Don Choi. A new approach to electronspin resonance absorption line shape function. Journal of Mathematical Physics, 49(7):073501, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i7/p073501_s1.

Lievens:2008:HOC

[1545] S. Lievens, N. I. Stoilova, and J. Van der Jeugt. Harmonic oscillator chains as Wigner quantum systems: Periodic and fixed wall boundary conditions in }↕(∞|\) solutions. Journal of Mathematical Physics, 49(7): 073502, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i7/p073502_s1.

Li:2008:LTS

[1546] Hai-Liang Li, Guo-Jing Zhang, Min Zhang, and Chengchun Hao. Longtime self-similar asymptotic of the macroscopic quantum models. Journal of Mathematical Physics, 49(7):073503, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i7/p073503_s1.

Xin:2008:GSD

[1547] Jie Xin, Boling Guo, Yongqian Han, and Daiwen Huang. The global solution of the (2+1)-dimensional long wave-short wave resonance interaction equation. Journal of Mathematical Physics, 49(7):073504, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p073504_s1.

Garfinkle:2008:MDN

[1548] David Garfinkle and James Isenberg. The modeling of degenerate neck pinch singularities in Ricci flow by Bryant solitons. Journal of Mathematical Physics, 49(7):073505, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i7/p073505_s1.

Wesslen:2008:GDT

[1549] Maria S. M. Wesslén. A geometric description of tensor product decompositions in ∫⊓(∋). Journal of Mathematical Physics, 49(7):073506, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p073506_s1.

Fialowski:2008:MSL

[1550] Alice Fialowski and Michael Penkava. Moduli spaces of low dimensional real Lie algebras. Journal of Mathematical Physics, 49(7):073507, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p073507_s1.

Matsuno:2008:PSS

[1551] Yoshimasa Matsuno. Periodic solutions of the short pulse model equation. Journal of Mathematical Physics, 49(7):073508, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i7/p073508_s1.

Atakishiyev:2008:ECD

[1552] N. M. Atakishiyev and A. U. Klimyk. On eigenfunctions of continuous and discrete (anti)symmetric multivariate Fourier transforms. Journal of Mathematical Physics, 49(7):073509, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i7/p073509_s1.

Jorgensen:2008:ESA

[1553] Palle E. T. Jorgensen. Essential self-adjointness of the graph-Laplacian. Journal of Mathematical Physics, 49(7):073510, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i7/p073510_s1.

Chaichian:2008:RGN

[1554] M. Chaichian, A. Tureanu, R. B. Zhang, and Xiao Zhang. Riemannian geometry of noncommutative surfaces. Journal of Mathematical Physics, 49(7):073511, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i7/p073511_s1.

Isaev:2008:BRS

[1555] A. P. Isaev, S. O. Krivonos, and O. V. Ogievetsky. Becchi–Rouet–Stora– Tyutin operators for W algebras. Journal of Mathematical Physics, 49 (7):073512, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i7/p073512_s1.

Lim:2008:MCE

[1556] S. C. Lim and L. P. Teo. On the minima and convexity of Epstein zeta function. Journal of Mathematical Physics, 49(7):073513, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p073513_s1.

Lavaei-Yanesi:2008:PSQ

[1557] L. Lavaei-Yanesi and M. Khorrami. Phase structure of the quartic-cubic generalized two dimensional Yang–Mills U(N) on the sphere. Journal of Mathematical Physics, 49(7):073514, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i7/p073514_s1.

Ju:2008:CNS

[1558] Qiangchang Ju, Fucai Li, and Shu Wang. Convergence of the Navier– Stokes–Poisson system to the incompressible Navier–Stokes equations. Journal of Mathematical Physics, 49(7):073515, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i7/p073515_s1.

Karasu-Kalkanli:2008:NIG

[1559] Ayse Karasu-Kalkanli, Atalay Karasu, Anton Sakovich, Sergei Sakovich, and Refik Turhan. A new integrable generalization of the Korteweg–de Vries equation. Journal of Mathematical Physics, 49(7):073516, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p073516_s1.

Nucci:2008:JLM

[1560] M. C. Nucci and P. G. L. Leach. Jacobi’s last multiplier and Lagrangians for multidimensional systems. Journal of Mathematical Physics, 49(7): 073517, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i7/p073517_s1.

Kirkinis:2008:RGI

[1561] Eleftherios Kirkinis. The renormalization group and the implicit function theorem for amplitude equations. Journal of Mathematical Physics, 49 (7):073518, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i7/p073518_s1.

Ghaderipoor:2008:GSI

[1562] Alireza Ghaderipoor and Chintha Tellambura. Generalization of some integrals over unitary matrices by character expansion of groups. Journal of Mathematical Physics, 49(7):073519, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i7/p073519_s1.

Dikande:2008:BMS

[1563] Alain Moïse Dikandé. Bright-matter solitons in a uniformly feeded Bose– Einstein condensate with expulsive harmonic potential. Journal of Mathematical Physics, 49(7):073520, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i7/p073520_s1.

Weihua:2008:RTI

[1564] Liu Weihua and Wu Junde. A representation theorem of infimum of bounded quantum observables. Journal of Mathematical Physics, 49(7): 073521, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i7/p073521_s1.

Kowalski:2008:HSD

[1565] Kenneth L. Kowalski. Horn–Silver distributions in D-dimensional Euclidean spaces. Journal of Mathematical Physics, 49(7):074101, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p074101_s1.

Clement:2008:ESF

[1566] Gérard Clément. Erratum: “The symmetries of five-dimensional minimal supergravity reduced to three dimensions”, [J. Math. Phys. 49, 042503 (2008)]. Journal of Mathematical Physics, 49(7):079901, July 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i7/p079901_s1. See [1428].

Spector:2008:MLU

[1567] Donald Spector. Minimal length uncertainty relations and new shape invariant models. Journal of Mathematical Physics, 49(8):082101, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i8/p082101_s1.

Alhaidari:2008:ECS

[1568] A. D. Alhaidari and H. Bahlouli. Extending the class of solvable potentials. I. The infinite potential well with a sinusoidal bottom. Journal of Mathematical Physics, 49(8):082102, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i8/p082102_s1.

Skotiniotis:2008:EVQ

[1569] Michael Skotiniotis, Aidan Roy, and Barry C. Sanders. On the epistemic view of quantum states. Journal of Mathematical Physics, 49(8):

082103, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i8/p082103_s1.

Yanez:2008:FIS

[1570] R. J. Yáñez, P. Sánchez-Moreno, A. Zarzo, and J. S. Dehesa. Fisher information of special functions and second-order differential equations. Journal of Mathematical Physics, 49(8):082104, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i8/p082104_s1.

Mehri-Dehnavi:2008:GPN

[1571] Hossein Mehri-Dehnavi and Ali Mostafazadeh. Geometric phase for non-Hermitian Hamiltonians and its holonomy interpretation. Journal of Mathematical Physics, 49(8):082105, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i8/p082105_s1.

Castro:2008:WOT

[1572] P. G. Castro, B. Chakraborty, and F. Toppan. Wigner oscillators, twisted Hopf algebras, and second quantization. Journal of Mathematical Physics, 49(8):082106, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i8/p082106_s1.

Lucietti:2008:ACB

[1573] James Lucietti and Mukund Rangamani. Asymptotic counting of BPS operators in superconformal field theories. Journal of Mathematical Physics, 49(8):082301, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i8/p082301_s1.

Hendi:2008:RBB

[1574] S. H. Hendi. Rotating black branes in Brans–Dicke–Born–Infeld theory. Journal of Mathematical Physics, 49(8):082501, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i8/p082501_s1.

Szablikowski:2008:DAW

[1575] Blazej M. Szablikowski and Maciej Blaszak. Dispersionful analog of the Whitham hierarchy. Journal of Mathematical Physics, 49(8):082701, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i8/p082701_s1.

Inizan:2008:HFE

[1576] Pierre Inizan. Homogeneous fractional embeddings. Journal of Mathematical Physics, 49(8):082901, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i8/p082901_s1.

Hirota:2008:EHM

[1577] M. Hirota and Y. Fukumoto. Energy of hydrodynamic and magnetohydrodynamic waves with point and continuous spectra. Journal of Mathematical Physics, 49(8):083101, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i8/p083101_s1.

Hunsicker:2008:APS

[1578] Eugenie Hunsicker, Victor Nistor, and Jorge O. Sofo. Analysis of periodic Schrödinger operators: Regularity and approximation of eigenfunctions. Journal of Mathematical Physics, 49(8):083501, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i8/p083501_s1.

Cheviakov:2008:EPF

[1579] Alexei F. Cheviakov. An extended procedure for finding exact solutions of partial differential equations arising from potential symmetries. Applications to gas dynamics. Journal of Mathematical Physics, 49(8): 083502, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i8/p083502_s1.

Ghanmi:2008:LAF

[1580] A. Ghanmi and A. Intissar. Landau automorphic functions on C^{n} of magnitude ν. Journal of Mathematical Physics, 49(8):083503, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i8/p083503_s1.

Sanchez:2008:ATL

[1581] Richard Sanchez, Jean Ragusa, and Emiliano Masiello. Asymptotic theory of the linear transport equation in anisotropic media. Journal of Mathematical Physics, 49(8):083504, August 2008. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i8/p083504_s1.

Katore:2008:PGW

[1582] S. D. Katore and R. S. Rane. Plane gravitational waves of gaugeinvariant generalized field equations with asymmetric fundamental tensor in plane symmetry. Journal of Mathematical Physics, 49(8):083505, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i8/p083505_s1.

Gao:2008:ISM

[1583] Ya-Jun Gao. Inverse scattering method and soliton double solution family for the general symplectic gravity model. Journal of Mathematical Physics, 49(8):083506, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i8/p083506_s1.

Abdeljawad:2008:EUT

[1584] Thabet Maraaba (Abdeljawad), Dumitru Baleanu, and Fahd Jarad. Existence and uniqueness theorem for a class of delay differential equations with left and right Caputo fractional derivatives. Journal of Mathematical Physics, 49(8):083507, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i8/p083507_s1.

El-Jaick:2008:SCD

[1585] Léa Jaccoud El-Jaick and Bartolomeu D. B. Figueiredo. Solutions for confluent and double-confluent Heun equations. Journal of Mathematical Physics, 49(8):083508, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i8/p083508_s1.

Aydin:2008:SML

[1586] A. Aydin and B. Karasözen. Symplectic and multisymplectic Lobatto methods for the “good” Boussinesq equation. Journal of Mathematical Physics, 49(8):083509, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i8/p083509_s1.

Meusburger:2008:QPL

[1587] C. Meusburger and B. J. Schroers. Quaternionic and Poisson–Lie structures in three-dimensional gravity: The cosmological constant as deformation parameter. Journal of Mathematical Physics, 49(8):083510, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i8/p083510_s1.

Aleixo:2008:PNP

[1588] A. N. F. Aleixo and A. B. Balantekin. Publisher’s note: “Purely squeezed states for quantum deformed systems”, [J. Math. Phys. 49, 062104 (2008)]. Journal of Mathematical Physics, 49(8):089901, August 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i8/

p089901_s1. See [1503].

Dorlas:2008:DFI

[1589] T. C. Dorlas and E. G. F. Thomas. The discrete Feynman integral. Journal of Mathematical Physics, 49(9):092101, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v49/i9/p092101_s1.

Mendas:2008:CTE

[1590] Istok P. Mendas. Classification and time evolution of density matrices for a N-state system. Journal of Mathematical Physics, 49(9):092102, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i9/p092102_s1.

Ferrario:2008:PSM

[1591] Andrea Ferrario. Poisson sigma model with branes and hyperelliptic Riemann surfaces. Journal of Mathematical Physics, 49(9):092301, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i9/p092301_s1.

Minguzzi:2008:LCT

[1592] E. Minguzzi. Limit curve theorems in Lorentzian geometry. Journal of Mathematical Physics, 49(9):092501, September 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i9/p092501_s1.

Fontanelli:2008:BHA

[1593] Laura Fontanelli, Paolo Lorenzoni, Marco Pedroni, and Jorge P. Zubelli. Bi-Hamiltonian aspects of a matrix Harry Dym hierarchy. Journal of Mathematical Physics, 49(9):092901, September 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i9/p092901_s1.

Evans:2008:SCA

[1594] N. W. Evans and P. E. Verrier. Superintegrability of the caged anisotropic oscillator. Journal of Mathematical Physics, 49(9):092902, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i9/p092902_s1.

Baxter:2008:FES

[1595] Nathan P. Baxter and Sergei V. Shabanov. The flip-over effect in self-similar laser-induced plasma expansion. Journal of Mathematical Physics, 49(9):093101, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i9/p093101_s1.

Ding:2008:TPD

[1596] Qing Ding and Wei Lin. The transmission property of the discrete Heisenberg ferromagnetic spin chain. Journal of Mathematical Physics, 49(9):093501, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i9/p093501_s1.

Mazzucchi:2008:FPI

[1597] S. Mazzucchi. Feynman path integrals for the inverse quartic oscillator. Journal of Mathematical Physics, 49(9):093502, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i9/
p093502_s1.

Albeverio:2008:JPL

[1598] Sergio Albeverio and Witold Karwowski. Jump processes on leaves of multibranching trees. Journal of Mathematical Physics, 49(9):093503, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i9/p093503_s1.

Bruneau:2008:LTE

[1599] Vincent Bruneau and El Maati Ouhabaz. Lieb–Thirring estimates for non-self-adjoint Schrödinger operators. Journal of Mathematical Physics, 49(9):093504, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i9/p093504_s1.

Jiao:2008:ASR

[1600] Xiaoyu Jiao, Ruoxia Yao, and S. Y. Lou. Approximate similarity reduction for singularly perturbed Boussinesq equation via symmetry perturbation and direct method. Journal of Mathematical Physics, 49 (9):093505, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i9/p093505_s1.

Liu:2008:ETD

[1601] Xiaojun Liu, Yunbo Zeng, and Runliang Lin. An extended twodimensional Toda lattice hierarchy and two-dimensional Toda lattice with self-consistent sources. Journal of Mathematical Physics, 49 (9):093506, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i9/p093506_s1.

Lisovyy:2008:PVT

[1602] O. Lisovyy. On Painlevé VI transcendents related to the Dirac operator on the hyperbolic disk. Journal of Mathematical Physics, 49(9):093507, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i9/p093507_s1.

Coffey:2008:EIA

[1603] Mark W. Coffey. Evaluation of a ln tan integral arising in quantum field theory. Journal of Mathematical Physics, 49(9):093508, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i9/p093508_s1.

Cardoso:2008:NAS

[1604] Pedro G. S. Cardoso, Ernesto P. Borges, Thierry C. P. Lobão, and Suani T. R. Pinho. Nondistributive algebraic structures derived from nonextensive statistical mechanics. Journal of Mathematical Physics, 49(9):093509, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i9/p093509_s1.

Wu:2008:NMC

[1605] Hongxia Wu, Yunbo Zeng, and Tianyou Fan. A new multicomponent CKP hierarchy and solutions. Journal of Mathematical Physics, 49(9):093510, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i9/p093510_s1.

Fialowski:2008:LAD

[1606] Alice Fialowski and Ashis Mandal. Leibniz algebra deformations of a Lie algebra. Journal of Mathematical Physics, 49(9):093511, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i9/p093511_s1.

Foldes:2008:LGF

[1607] Stephan Foldes. The Lorentz group and its finite field analogs: Local isomorphism and approximation. Journal of Mathematical Physics, 49(9):093512, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i9/p093512_s1.

Henneaux:2008:ESC

[1608] Marc Henneaux, Mauricio Leston, Daniel Persson, and Philippe Spindel. Erratum: “A Special Class of Rank 10 and 11 Coxeter Groups”, [J. Math. Phys. 48, 053512 (2007)]. Journal of Mathematical Physics, 49(9):099901, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i9/p099901_s1. See [1036].

Galapon:2008:EST

[1609] Eric. A. Galapon. Erratum: “Shouldn’t there be an antithesis to quantization?”, [J. Math. Phys. 45, 3180 (2004)]. Journal of Mathematical Physics, 49(9):099902, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i9/p099902_s1. See [7].

Nachtergaele:2008:ASIb

[1610] Bruno L. Z. Nachtergaele. Announcement: Special issue on “Integrable Quantum Systems and Solvable Statistical Mechanics Models”. Journal of Mathematical Physics, 49(10):100201, October 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i10/p100201_s1.

Hayashi:2008:APO

[1611] Masahito Hayashi and Keiji Matsumoto. Asymptotic performance of optimal state estimation in qubit system. Journal of Mathematical Physics, 49(10):102101, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i10/p102101_s1.

Aguilar:2008:ECN

[1612] Jean-Philippe Aguilar and Nils Berglund. The effect of classical noise on a quantum two-level system. Journal of Mathematical Physics, 49(10): 102102, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i10/p102102_s1.

Durran:2008:DCS

[1613] Richard Durran, Andrew Neate, Aubrey Truman, and Feng-Yu Wang. On the divine clockwork: The spectral gap for the correspondence limit of the Nelson diffusion generator for the atomic elliptic state. Journal of Mathematical Physics, 49(10):102103, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i10/p102103_s1.

Cubitt:2008:SDQ

[1614] Toby S. Cubitt, Mary Beth Ruskai, and Graeme Smith. The structure of degradable quantum channels. Journal of Mathematical Physics, 49(10): 102104, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i10/p102104_s1.

Angelescu:2008:LSB

[1615] Nicolae Angelescu, Robert A. Minlos, Jean Ruiz, and Valentin A. Zagrebnov. Lower spectral branches of a spin-boson model. Journal of Mathematical Physics, 49(10):102105, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v49/i10/p102105_s1. See erratum

[2195].

Cornean:2008:ASE

[1616] H. D. Cornean, P. Duclos, G. Nenciu, and R. Purice. Adiabatically switched-on electrical bias and the Landauer–Büttiker formula. Journal of Mathematical Physics, 49(10):102106, October 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i10/p102106_s1.

Rosgen:2008:ADR

[1617] Bill Rosgen. Additivity and distinguishability of random unitary channels. Journal of Mathematical Physics, 49(10):102107, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i10/
p102107_s1.

Kryukov:2008:NTU

[1618] A. Kryukov. Nine theorems on the unification of quantum mechanics and relativity. Journal of Mathematical Physics, 49(10):102108, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i10/p102108_s1.

Bouten:2008:DAQ

[1619] Luc Bouten and Ramon Van Handel. Discrete approximation of quantum stochastic models. Journal of Mathematical Physics, 49(10):102109, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i10/p102109_s1.

Aki:2008:CLS

[1620] Gonca L. Aki, Peter A. Markowich, and Christof Sparber. Classical limit for semirelativistic Hartree systems. Journal of Mathematical Physics, 49(10):102110, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i10/p102110_s1.

Meng:2008:KP

[1621] Guowu Meng. The O(1)-Kepler problems. Journal of Mathematical Physics, 49(10):102111, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i10/p102111_s1.

Tarasov:2008:WQF

[1622] Vasily E. Tarasov. Weyl quantization of fractional derivatives. Journal of Mathematical Physics, 49(10):102112, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i10/p102112_s1.

Menezes:2008:SQC

[1623] G. Menezes and N. F. Svaiter. Stochastic quantization for complex actions. Journal of Mathematical Physics, 49(10):102301, October 2008.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i10/
p102301_s1.

Sakaguchi:2008:SSA

[1624] Makoto Sakaguchi and Kentaroh Yoshida. Super Schrödinger algebra in AdS/CFT. Journal of Mathematical Physics, 49(10):102302, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i10/p102302_s1.

Dahia:2008:EST

[1625] F. Dahia, G. A. T. Gomez, and C. Romero. On the embedding of space–time in five-dimensional Weyl spaces. Journal of Mathematical Physics, 49(10):102501, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i10/p102501_s1.

Llibre:2008:AIC

[1626] Jaume Llibre and Clàudia Valls. Analytic integrability of a Chua system. Journal of Mathematical Physics, 49(10):102701, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i10/
p102701_s1.

Habibullin:2008:CDI

[1627] Ismagil Habibullin, Natalya Zheltukhina, and Asli Pekcan. On the classification of Darboux integrable chains. Journal of Mathematical Physics, 49(10):102702, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i10/p102702_s1.

Chiba:2008:ACM

[1628] Hayato Chiba. Approximation of center manifolds on the renormalization group method. Journal of Mathematical Physics, 49(10):102703, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i10/p102703_s1.

Gurses:2008:HTI

[1629] Metin Gürses, Ismagil Habibullin, and Kostyantyn Zheltukhin. Hydrodynamic type integrable equations on a segment and a half-line. Journal of Mathematical Physics, 49(10):102704, October 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i10/p102704_s1.

Yan:2008:UAN

[1630] Xingjie Yan and Chengkui Zhong. L^{p}-uniform attractor for nonautonomous reaction-diffusion equations in unbounded domains. Journal of Mathematical Physics, 49(10):102705, October 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i10/p102705_s1.

Gsponer:2008:CPE

[1631] Andre Gsponer. The classical point electron in Colombeau’s theory of nonlinear generalized functions. Journal of Mathematical Physics, 49(10):102901, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i10/p102901_s1.

Fabrizio:2008:IWL

[1632] Mauro Fabrizio. Ice-water and liquid-vapor phase transitions by a Ginzburg–Landau model. Journal of Mathematical Physics, 49(10): 102902, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i10/p102902_s1.

Ao:2008:CBB

[1633] Ji jun Ao and Achilaltu Wang. The calculation of the bilinear Boltzmann operator for gas mixtures. Journal of Mathematical Physics, 49(10): 103101, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i10/p103101_s1.

Bernardin:2008:SAE

[1634] Cédric Bernardin. Superdiffusivity of asymmetric energy model in dimensions 1 and 2. Journal of Mathematical Physics, 49(10):103301, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i10/p103301_s1.

Hayashi:2008:NSS

[1635] Nakao Hayashi, Pavel I. Naumkin, and Ratno Bagus Edy Wibowo. Nonlinear scattering for a system of nonlinear Klein–Gordon equations. Journal of Mathematical Physics, 49(10):103501, October 2008.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i10/
p103501_s1.

Kawamura:2008:UIF

[1636] Katsunori Kawamura. Unitary isomorphism of Fock spaces of bosons and fermions arising from a representation of the Cuntz algebra O_{2}. Journal of Mathematical Physics, 49(10):103502, October 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i10/p103502_s1.

Li:2008:SFS

[1637] Biao Li, Wang-Chuan Ye, and Yong Chen. Symmetry, full symmetry groups, and some exact solutions to a generalized Davey–Stewartson system. Journal of Mathematical Physics, 49(10):103503, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i10/
p103503_s1.

Curro:2008:RNQ

[1638] Carmela Currò and Francesco Oliveri. Reduction of nonhomogeneous quasilinear 2×2 systems to homogeneous and autonomous form. Journal of Mathematical Physics, 49(10):103504, October 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i10/p103504_s1.

Giannoulis:2008:LHT

[1639] Johannes Giannoulis, Michael Herrmann, and Alexander Mielke. Lagrangian and Hamiltonian two-scale reduction. Journal of Mathematical Physics, 49(10):103505, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i10/p103505_s1.

Kunzinger:2008:PCL

[1640] Michael Kunzinger and Roman O. Popovych. Potential conservation laws. Journal of Mathematical Physics, 49(10):103506, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i10/
p103506_s1.

Heo:2008:RKH

[1641] Jaeseong Heo. Reproducing kernel Hilbert C^{∗}-modules and kernels associated with cocycles. Journal of Mathematical Physics, 49(10):103507, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i10/p103507_s1.

Velazquez-Arcos:2008:CTK

[1642] J. M. Velázquez-Arcos, C. A. Vargas, J. L. Fernández-Chapou, and A. L. Salas-Brito. On computing the trace of the kernel of the homogeneous Fredholm’s equation. Journal of Mathematical Physics, 49(10): 103508, October 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i10/p103508_s1.

Nachtergaele:2008:ASIc

[1643] Bruno L. Z. Nachtergaele. Announcement: Special issue on “Integrable Quantum Systems and Solvable Statistical Mechanics Models”. Journal of Mathematical Physics, 49(11):110201, November 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i11/p110201_s1.

Huang:2008:SPC

[1644] Siendong Huang. On states of perfect correlation. Journal of Mathematical Physics, 49(11):112101, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i11/p112101_s1.

Chi:2008:MEQ

[1645] Dong Pyo Chi, Jeong Woon Choi, Kabgyun Jeong, Jeong San Kim, Taewan Kim, and Soojoon Lee. Monogamy equality in 2 ⊃ ⊗2 ⊃ ⊗d quantum systems. Journal of Mathematical Physics, 49(11):112102, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i11/p112102_s1.

Kiukas:2008:SMC

[1646] Jukka Kiukas, Pekka Lahti, and Kari Ylinen. Semispectral measures as convolutions and their moment operators. Journal of Mathematical Physics, 49(11):112103, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i11/p112103_s1.

Schlingemann:2008:SCQ

[1647] Dirk-M. Schlingemann, Holger Vogts, and Reinhard F. Werner. On the structure of Clifford quantum cellular automata. Journal of Mathematical Physics, 49(11):112104, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i11/p112104_s1.

Dahia:2008:EFE

[1648] F. Dahia, C. Romero, and M. A. S. Cruz. Effective field equations of brane-induced electromagnetism. Journal of Mathematical Physics, 49(11):112501, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i11/p112501_s1.

Saha:2008:ACM

[1649] Bijan Saha and Victor Rikhvitsky. Anisotropic cosmological models with spinor and scalar fields and viscous fluid in presence of a Λ term: Qualitative solutions. Journal of Mathematical Physics, 49(11):112502, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i11/p112502_s1.

Corsi:2008:MTA

[1650] Livia Corsi and Guido Gentile. Melnikov theory to all orders and Puiseux series for subharmonic solutions. Journal of Mathematical Physics, 49 (11):112701, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i11/p112701_s1.

Chanu:2008:STB

[1651] Claudia Chanu, Luca Degiovanni, and Giovanni Rastelli. Superintegrable three-body systems on the line. Journal of Mathematical Physics, 49(11):112901, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i11/p112901_s1.

Galaktionov:2008:BSJ

[1652] V. A. Galaktionov. On blow-up space jets for the Navier–Stokes equations in R^{3} with convergence to Euler equations. Journal of Mathematical Physics, 49(11):113101, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i11/p113101_s1.

Manwai:2008:ASN

[1653] Yuen Manwai. Analytical solutions to the Navier–Stokes equations. Journal of Mathematical Physics, 49(11):113102, November 2008. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i11/

p113102_s1. See erratum [1758].

Tay:2008:BEM

[1654] B. A. Tay and T. Petrosky. Biorthonormal eigenbasis of a Markovian master equation for the quantum Brownian motion. Journal of Mathematical Physics, 49(11):113301, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i11/p113301_s1.

Faris:2008:CGI

[1655] William G. Faris. A connected graph identity and convergence of cluster expansions. Journal of Mathematical Physics, 49(11):113302, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i11/p113302_s1.

Chaudhuri:2008:TBK

[1656] Jyotipratim Ray Chaudhuri, Suman Kumar Banik, Sudip Chattopadhyay, and Pinaki Chaudhury. Transport and bistable kinetics of a Brownian particle in a nonequilibrium environment. Journal of Mathematical Physics, 49(11):113303, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i11/p113303_s1.

Pelinovsky:2008:STD

[1657] D. E. Pelinovsky and A. Stefanov. On the spectral theory and dispersive estimates for a discrete Schrödinger equation in one dimension. Journal of Mathematical Physics, 49(11):113501, November 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i11/p113501_s1.

Gagliardo:2008:HFH

[1658] Michael Gagliardo. The higher flows of harmonic maps and an application to the Virasoro action. Journal of Mathematical Physics, 49 (11):113502, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i11/p113502_s1.

Liu:2008:CIW

[1659] Dong Liu, Shoulan Gao, and Linsheng Zhu. Classification of irreducible weight modules over W -algebra W (2, 2). Journal of Mathematical Physics, 49(11):113503, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i11/p113503_s1.

Zhang:2008:BPP

[1660] Shuanghu Zhang and Zhaoyang Yin. On the blow-up phenomena of the periodic Dullin–Gottwald–Holm equation. Journal of Mathematical Physics, 49(11):113504, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i11/p113504_s1.

Bohacik:2008:FIU

[1661] J. Bohácik and P. Presnajder. The functional integral with unconditional Wiener measure for anharmonic oscillator. Journal of Mathematical Physics, 49(11):113505, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i11/p113505_s1.

Kupin:2008:ACS

[1662] S. Kupin. Absolutely continuous spectrum of a Schrödinger operator on a tree. Journal of Mathematical Physics, 49(11):113506, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i11/p113506_s1.

Wu:2008:CPW

[1663] Derchyi Wu. The Cauchy problem of the Ward equation with mixed scattering data. Journal of Mathematical Physics, 49(11):113507, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i11/p113507_s1.

Miao:2008:EBP

[1664] Dong Miao. Eigenvalue branches of the perturbed Maxwell operator M + λD in a gap of σ(M). Journal of Mathematical Physics, 49(11):113508, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i11/p113508_s1.

Figueroa-OFarrill:2008:LLA

[1665] José Figueroa-O’Farrill. Lorentzian Lie n-algebras. Journal of Mathematical Physics, 49(11):113509, November 2008. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i11/p113509_s1.

Sahadevan:2008:SRG

[1666] R. Sahadevan and S. Balakrishnan. Similarity reduction, generalized symmetries, recursion operator, and integrability of coupled Volterra system. Journal of Mathematical Physics, 49(11):113510, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i11/
p113510_s1.

Fakhri:2008:ASH

[1667] H. Fakhri and T. Ghaneh. Approach of spherical harmonics to the representation of the deformed su(1, 1) algebra. Journal of Mathematical Physics, 49(11):113511, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i11/p113511_s1.

Cruz:2008:STK

[1668] Maximino Cruz and Juan H. Arredondo R. Scattering theory for the Klein–Gordon equation with nondecreasing potentials. Journal of Mathematical Physics, 49(11):113512, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v49/i11/p113512_s1.

Dowker:2008:EAS

[1669] J. S. Dowker and Klaus Kirsten. Elliptic aspects of statistical mechanics on spheres. Journal of Mathematical Physics, 49(11):113513, November 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i11/p113513_s1.

Contucci:2008:ISI

[1670] Pierluigi Contucci and Cristian Giardinà. Introduction to special issue: Statistical Mechanics on Random Structures. Journal of Mathematical Physics, 49(12):125101, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i12/p125101_s1.

Crawford:2008:IBM

[1671] Nicholas Crawford. The interaction between multioverlaps in the high temperature phase of the Sherrington–Kirkpatrick spin glass. Journal of Mathematical Physics, 49(12):125201, December 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i12/p125201_s1.

Bovier:2008:FPF

[1672] Anton Bovier and Anton Klimovsky. Fluctuations of the partition function in the generalized random energy model with external field. Journal of Mathematical Physics, 49(12):125202, December 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i12/p125202_s1.

Morters:2008:MSS

[1673] Peter Mörters and Marcel Ortgiese. Minimal supporting subtrees for the free energy of polymers on disordered trees. Journal of Mathematical Physics, 49(12):125203, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v49/i12/p125203_s1.

Arguin:2008:RIV

[1674] Louis-Pierre Arguin. A remark on the infinite-volume Gibbs measures of spin glasses. Journal of Mathematical Physics, 49(12):125204, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i12/p125204_s1.

Bolthausen:2008:USS

[1675] Erwin Bolthausen and Nicola Kistler. Universal structures in some mean field spin glasses and an application. Journal of Mathematical Physics, 49(12):125205, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i12/p125205_s1.

Bayati:2008:RAC

[1676] Mohsen Bayati, Alfredo Braunstein, and Riccardo Zecchina. A rigorous analysis of the cavity equations for the minimum spanning tree. Journal of Mathematical Physics, 49(12):125206, December 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i12/p125206_s1.

Janson:2008:SSR

[1677] Svante Janson and Malwina J. Luczak. Susceptibility in subcritical random graphs. Journal of Mathematical Physics, 49(12):125207, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i12/p125207_s1.

vanEnter:2008:LRG

[1678] A. C. D. van Enter and W. M. Ruszel. Loss and recovery of Gibbsianness for XY models in external fields. Journal of Mathematical Physics, 49(12):125208, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i12/p125208_s1.

vanderHofstad:2008:UDP

[1679] Remco van der Hofstad and Gerard Hooghiemstra. Universality for distances in power-law random graphs. Journal of Mathematical Physics, 49(12):125209, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i12/p125209_s1.

Morita:2008:MFQ

[1680] Satoshi Morita and Hidetoshi Nishimori. Mathematical foundation of quantum annealing. Journal of Mathematical Physics, 49(12):125210, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v49/i12/p125210_s1.

Basile:2008:EMA

[1681] Chiara Basile, Dario Benedetto, Emanuele Caglioti, and Mirko Degli Esposti. An example of mathematical authorship attribution. Journal of Mathematical Physics, 49(12):125211, December 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i12/p125211_s1.

Talagrand:2008:MFM

[1682] Michel Talagrand. A mean field model for spin glasses based on a 2-level perceptron. Journal of Mathematical Physics, 49(12):125212, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v49/i12/p125212_s1.

Lenci:2008:CLT

[1683] Marco Lenci. Central limit theorem and recurrence for random walks in bistochastic random environments. Journal of Mathematical Physics, 49(12):125213, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i12/p125213_s1.

Chazottes:2008:CCI

[1684] Jean-René Chazottes, Pierre Collet, and Frank Redig. Coupling, concentration inequalities, and stochastic dynamics. Journal of Mathematical Physics, 49(12):125214, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v49/i12/p125214_s1.

Kulske:2008:CSM

[1685] Christof Külske and Alex A. Opoku. Continuous spin mean-field models: Limiting kernels and Gibbs properties of local transforms. Journal of Mathematical Physics, 49(12):125215, December 2008. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v49/i12/p125215_s1.

Parisi:2008:SPE

[1686] Giorgio Parisi. On the survey-propagation equations in random constraint satisfiability problems. Journal of Mathematical Physics, 49 (12):125216, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i12/p125216_s1.

Barra:2008:AER

[1687] Adriano Barra and Francesco Guerra. About the ergodic regime in the analogical Hopfield neural networks: Moments of the partition function. Journal of Mathematical Physics, 49(12):125217, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i12/
p125217_s1.

Bhamidi:2008:FPP

[1688] Shankar Bhamidi. First passage percolation on locally treelike networks. I. Dense random graphs. Journal of Mathematical Physics, 49 (12):125218, December 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i12/p125218_s1.

Percus:2008:PPS

[1689] Allon G. Percus, Gabriel Istrate, Bruno Gonçalves, Robert Z. Sumi, and Stefan Boettcher. The peculiar phase structure of random graph bisection. Journal of Mathematical Physics, 49(12):125219, December 2008.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i12/
p125219_s1.

Nachtergaele:2009:AII

[1690] Bruno L. Z. Nachtergaele. Announcement: ICMP09 — 16th International Congress on Mathematical Physics (3–8 August 2009, Prague, Czech Republic). Journal of Mathematical Physics, 50(1):010201, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p010201_s1.

Cassar:2009:ACJ

[1691] Mark M. Cassar. Announcement: Chinese, Japanese, and Korean characters available for author names. Journal of Mathematical Physics, 50(1): 010202, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p010202_s1.

Cassar:2009:ARF

[1692] Mark M. Cassar. Announcement: Reduced fees for color in print. Journal of Mathematical Physics, 50(1):010203, January 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i1/p010203_s1.

Marquette:2009:STOa

[1693] Ian Marquette. Superintegrability with third order integrals of motion, cubic algebras, and supersymmetric quantum mechanics. I. Rational function potentials. Journal of Mathematical Physics, 50(1):012101, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p012101_s1.

Konig:2009:ECV

[1694] Robert König and Michael M. Wolf. On exchangeable continuous variable systems. Journal of Mathematical Physics, 50(1):012102, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p012102_s1.

Chen:2009:ISS

[1695] Thomas Chen, Jürg Fröhlich, and Alessandro Pizzo. Infraparticle scattering states in nonrelativistic quantum electrodynamics. II. Mass shell properties. Journal of Mathematical Physics, 50(1):012103, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p012103_s1.

Li:2009:EMQ

[1696] Dafa Li, Xiangrong Li, Hongtao Huang, and Xinxin Li. An entanglement measure for n qubits. Journal of Mathematical Physics, 50(1): 012104, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p012104_s1.

Konig:2009:MCF

[1697] Robert König and Graeme Mitchison. A most compendious and facile quantum de Finetti theorem. Journal of Mathematical Physics, 50(1): 012105, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p012105_s1.

Salcedo:2009:PSL

[1698] L. L. Salcedo. Phase space localization of antisymmetric functions. Journal of Mathematical Physics, 50(1):012106, January 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i1/p012106_s1.

Skinner:2009:UBH

[1699] A. J. Skinner, V. A. Newell, and R. Sanchez. Unbiased bases (Hadamards) for six-level systems: Four ways from Fourier. Journal of Mathematical Physics, 50(1):012107, January 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i1/p012107_s1.

Kihara:2009:GIC

[1700] Hironobu Kihara and Muneto Nitta. Generalized instantons on complex projective spaces. Journal of Mathematical Physics, 50(1):012301, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p012301_s1.

Mironov:2009:OSO

[1701] Victor L. Mironov and Sergey V. Mironov. Octonic second-order equations of relativistic quantum mechanics. Journal of Mathematical Physics, 50(1):012302, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i1/p012302_s1.

Giambo:2009:GCH

[1702] Roberto Giambò. Gravitational collapse of homogeneous perfect fluids in higher order gravity theories. Journal of Mathematical Physics, 50(1): 012501, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p012501_s1.

Beyer:2009:NKG

[1703] Horst R. Beyer. A note on the Klein–Gordon equation in the background of a rotating black hole. Journal of Mathematical Physics, 50(1): 012502, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p012502_s1.

Dittrich:2009:SGO

[1704] Bianca Dittrich and Thomas Thiemann. Are the spectra of geometrical operators in loop quantum gravity really discrete? Journal of Mathematical Physics, 50(1):012503, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i1/p012503_s1.

Mahan:2009:CWN

[1705] G. D. Mahan. Classical waves on nonlinear lattices. II. Q^{n}. Journal of Mathematical Physics, 50(1):012701, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p012701_s1.

Ban:2009:DKS

[1706] Ailing Ban, Shengfan Zhou, and Xiaoying Han. Dimension of kernel sections for damped nonautonomous wave equations with critical exponent. Journal of Mathematical Physics, 50(1):012702, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/p012702_s1.

Tudoran:2009:LCT

[1707] Razvan M. Tudoran and Ramona A. Tudoran. On a large class of threedimensional Hamiltonian systems. Journal of Mathematical Physics, 50(1):012703, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i1/p012703_s1.

Lenells:2009:SCH

[1708] J. Lenells and O. Lechtenfeld. On the N = 2 supersymmetric Camassa– Holm and Hunter–Saxton equations. Journal of Mathematical Physics, 50(1):012704, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i1/p012704_s1.

Conti:2009:CJI

[1709] Roberto Conti and Wojciech Szymański. Computing the Jones index of quadratic permutation endomorphisms of O_{2}. Journal of Mathematical Physics, 50(1):012705, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i1/p012705_s1.

Mironov:2009:ORE

[1710] Victor L. Mironov and Sergey V. Mironov. Octonic representation of electromagnetic field equations. Journal of Mathematical Physics, 50(1): 012901, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p012901_s1.

Ozeren:2009:ITE

[1711] Serhat F. Özeren. Investigation of the time evolution of Lane–Emdentype Kanai–Caldirola oscillator. Journal of Mathematical Physics, 50(1): 012902, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p012902_s1.

Giachetta:2009:NGS

[1712] G. Giachetta, L. Mangiarotti, and G. Sardanashvily. On the notion of gauge symmetries of generic Lagrangian field theory. Journal of Mathematical Physics, 50(1):012903, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i1/p012903_s1.

Chin:2009:RMC

[1713] Siu A. Chin. Relativistic motion in a constant electromagnetic field. Journal of Mathematical Physics, 50(1):012904, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/p012904_s1.

Wei:2009:SIB

[1714] Juncheng Wei and Matthias Winter. Strongly interacting bumps for the Schrödinger–Newton equations. Journal of Mathematical Physics, 50(1): 012905, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p012905_s1.

Fu:2009:WPB

[1715] Ying Fu and Changzheng Qu. Well posedness and blow-up solution for a new coupled Camassa–Holm equations with peakons. Journal of Mathematical Physics, 50(1):012906, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p012906_s1.

Deriglazov:2009:IEH

[1716] A. A. Deriglazov. Improved extended Hamiltonian and search for local symmetries. Journal of Mathematical Physics, 50(1):012907, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p012907_s1.

Lind-Johansen:2009:PLF

[1717] Øyvind Lind-Johansen, Kristian Seip, and Johannes Skaar. The perfect lens on a finite bandwidth. Journal of Mathematical Physics, 50(1): 012908, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p012908_s1.

Gustafson:2009:SSS

[1718] Stephen Gustafson, Hideo Takaoka, and Tai-Peng Tsai. Stability in H^{1/2} of the sum of K solitons for the Benjamin–Ono equation. Journal of Mathematical Physics, 50(1):013101, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p013101_s1.

Garcia-Pelayo:2009:DMF

[1719] Ricardo García-Pelayo. Distribution of magnetization in the finite Ising chain. Journal of Mathematical Physics, 50(1):013301, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/
p013301_s1.

Preissmann:2009:BMM

[1720] E. Preissmann, J.-Ch. Anglès d’Auriac, and J.-M. Maillard. Birational mappings and matrix subalgebra from the chiral Potts model. Journal of Mathematical Physics, 50(1):013302, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p013302_s1.

Furuichi:2009:MEP

[1721] Shigeru Furuichi. On the maximum entropy principle and the minimization of the Fisher information in Tsallis statistics. Journal of Mathematical Physics, 50(1):013303, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i1/p013303_s1.

Darvish:2009:KSR

[1722] Esmat Darvish and Amir Ali Masoudi. Kinetic surface roughening for the Mullins–Herring equation. Journal of Mathematical Physics, 50(1): 013304, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p013304_s1.

Shojaei:2009:PFT

[1723] F. Shojaei, E. Ghanbari-Adivi, M. J. Brunger, and M. A. Bolorizadeh. Positronium formation as a three-body reaction. II. The second-order nuclear amplitudes. Journal of Mathematical Physics, 50(1):013501, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p013501_s1.

Zayed:2009:EMF

[1724] E. M. E. Zayed and Khaled A. Gepreel. The (G′/G)-expansion method for finding traveling wave solutions of nonlinear partial differential equations in mathematical physics. Journal of Mathematical Physics, 50(1): 013502, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p013502_s1.

Caroca:2009:EHO

[1725] R. Caroca, N. Merino, and P. Salgado. S expansion of higher-order Lie algebras. Journal of Mathematical Physics, 50(1):013503, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p013503_s1.

Brunelli:2009:INE

[1726] J. C. Brunelli. Integrability of a nonlinear evolution equation derived from isoperimetric plane curve motion. Journal of Mathematical Physics, 50(1):013504, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i1/p013504_s1.

Zhang:2009:BSD

[1727] Guoping Zhang. Breather solutions of the discrete nonlinear Schrödinger equations with unbounded potentials. Journal of Mathematical Physics, 50(1):013505, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i1/p013505_s1.

Park:2009:WPU

[1728] Jong Yeoul Park and Tae Gab Ha. Well-posedness and uniform decay rates for the Klein–Gordon equation with damping term and acoustic boundary conditions. Journal of Mathematical Physics, 50(1):013506, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p013506_s1.

Pickering:2009:BTD

[1729] A. Pickering. Bäcklund transformations for a discrete second Painlevé hierarchy. Journal of Mathematical Physics, 50(1):013507, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/
p013507_s1.

Dassios:2009:OET

[1730] George Dassios, Demetra Hadjiloizi, and Fotini Kariotou. The octapolic ellipsoidal term in magnetoencephalography. Journal of Mathematical Physics, 50(1):013508, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i1/p013508_s1.

Drissi:2009:RST

[1731] L. B. Drissi, H. Jehjouh, and E. H. Saidi. Refining the shifted topological vertex. Journal of Mathematical Physics, 50(1):013509, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p013509_s1.

Anastasiei:2009:FQL

[1732] Mihai Anastasiei and Sergiu I. Vacaru. Fedosov quantization of Lagrange–Finsler and Hamilton–Cartan spaces and Einstein gravity lifts on (co) tangent bundles. Journal of Mathematical Physics, 50(1):013510, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p013510_s1.

Alcalde:2009:VPS

[1733] M. Aparicio Alcalde, R. Kullock, and N. F. Svaiter. Virtual processes and super-radiance in spin-boson models. Journal of Mathematical Physics, 50(1):013511, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i1/p013511_s1.

Choup:2009:EEL

[1734] Leonard N. Choup. Edgeworth expansion of the largest eigenvalue distribution function of Gaussian orthogonal ensemble. Journal of Mathematical Physics, 50(1):013512, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i1/p013512_s1.

Gibbons:2009:RMF

[1735] J. Gibbons and A. P. Veselov. On the rational monodromy-free potentials with sextic growth. Journal of Mathematical Physics, 50(1): 013513, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p013513_s1.

Duarte:2009:FEF

[1736] L. G. S. Duarte and L. A. C. P. da Mota. Finding elementary first integrals for rational second order ordinary differential equations. Journal of Mathematical Physics, 50(1):013514, January 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i1/p013514_s1.

Gilkey:2009:GRG

[1737] P. Gilkey, S. Nikcević, and D. Westerman. Geometric realizations of generalized algebraic curvature operators. Journal of Mathematical Physics, 50(1):013515, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i1/p013515_s1.

Dinu:2009:NDL

[1738] Victor Dinu, Arne Jensen, and Gheorghe Nenciu. Nonexponential decay laws in perturbation theory of near threshold eigenvalues. Journal of Mathematical Physics, 50(1):013516, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p013516_s1.

Coons:2009:GMT

[1739] Michael Coons and Klaus Kirsten. General moment theorems for nondistinct unrestricted partitions. Journal of Mathematical Physics, 50(1): 013517, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p013517_s1.

Bai:2009:PRY

[1740] Cheng-Ming Bai, Mo-Lin Ge, and Naihuan Jing. Principal realization of the Yangian Y (}↕(\)). Journal of Mathematical Physics, 50(1): 013518, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i1/p013518_s1.

Ganji:2009:ETS

[1741] D. D. Ganji and M. Abdollahzadeh. Exact traveling solutions of some nonlinear evolution equation by (G′/G)-expansion method. Journal of Mathematical Physics, 50(1):013519, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p013519_s1.

Grabowski:2009:NCN

[1742] J. Grabowski, M. de León, J. C. Marrero, and D. Martín de Diego. Nonholonomic constraints: A new viewpoint. Journal of Mathematical Physics, 50(1):013520, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i1/p013520_s1.

Wigman:2009:DNS

[1743] Igor Wigman. On the distribution of the nodal sets of random spherical harmonics. Journal of Mathematical Physics, 50(1):013521, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p013521_s1.

Tetradis:2009:EFT

[1744] Nikolaos Tetradis. Effective field theory with a variable ultraviolet cutoff. Journal of Mathematical Physics, 50(1):013522, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/p013522_s1.

Schwartz:2009:CQV

[1745] Charles Schwartz. Calculus with a quaternionic variable. Journal of Mathematical Physics, 50(1):013523, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p013523_s1.

Escobar:2009:FG

[1746] J. A. Escobar and Christoph Luhn. The flavor group Δ(6n^{2}). Journal of Mathematical Physics, 50(1):013524, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p013524_s1.

Fan:2009:PNC

[1747] Engui Fan. The positive and negative Camassa–Holm-γ hierarchies, zero curvature representations, bi-Hamiltonian structures, and algebrogeometric solutions. Journal of Mathematical Physics, 50(1):013525, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p013525_s1.

Meremianin:2009:HHA

[1748] A. V. Meremianin. Hyperspherical harmonics with arbitrary arguments. Journal of Mathematical Physics, 50(1):013526, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/p013526_s1.

Gershgorin:2009:CHW

[1749] Boris Gershgorin, Yuri V. Lvov, and Sergey Nazarenko. Canonical Hamiltonians for waves in inhomogeneous media. Journal of Mathematical Physics, 50(1):013527, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i1/p013527_s1.

Kieburg:2009:IGV

[1750] Mario Kieburg, Heiner Kohler, and Thomas Guhr. Integration of Grassmann variables over invariant functions on flat superspaces. Journal of Mathematical Physics, 50(1):013528, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i1/p013528_s1.

Bertola:2009:PFT

[1751] M. Bertola and O. Marchal. The partition function of the two-matrix model as an isomonodromic τ function. Journal of Mathematical Physics, 50(1):013529, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i1/p013529_s1.

Asorey:2009:FRS

[1752] M. Asorey and P. M. Lavrov. Fedosov and Riemannian supermanifolds. Journal of Mathematical Physics, 50(1):013530, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/p013530_s1.

Kaneko:2009:LEP

[1753] Kazuo Kaneko. Local expansion of Painlevé VI transcendents around a fixed singularity. Journal of Mathematical Physics, 50(1):013531, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p013531_s1.

Palmkvist:2009:GCR

[1754] Jakob Palmkvist. Generalized conformal realizations of Kac–Moody algebras. Journal of Mathematical Physics, 50(1):013532, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/
p013532_s1.

Ju:2009:RCN

[1755] Qiangchang Ju, Yong Li, and Shu Wang. Rate of convergence from the Navier–Stokes–Poisson system to the incompressible Euler equations. Journal of Mathematical Physics, 50(1):013533, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i1/p013533_s1.

Lian:2009:DGF

[1756] Qiao-Fang Lian and Yun-Zhang Li. The duals of Gabor frames on discrete periodic sets. Journal of Mathematical Physics, 50(1):013534, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p013534_s1.

Cai:2009:NEM

[1757] Jiaxiang Cai. A new explicit multisymplectic scheme for the regularized long-wave equation. Journal of Mathematical Physics, 50(1):013535, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i1/p013535_s1.

Manwai:2009:EAS

[1758] Yuen Manwai. Erratum: “Analytical solutions to the Navier–Stokes equations”, [J. Math. Phys. 49, 113102 (2008)]. Journal of Mathematical Physics, 50(1):019901, January 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i1/p019901_s1. See [1653].

Cassar:2009:AMC

[1759] Mark M. Cassar. Announcement: Multimedia capability added to Journal of Mathematical Physics. Journal of Mathematical Physics, 50(2):020201, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i2/p020201_s1.

Belgiorno:2009:CQS

[1760] Francesco Belgiorno and Franco Gallone. Confined quantum systems and their limits. Journal of Mathematical Physics, 50(2):022101, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i2/p022101_s1.

Gheondea:2009:ACO

[1761] Aurelian Gheondea and Ali Samil Kavruk. Absolute continuity for operator valued completely positive maps on C^{∗}-algebras. Journal of Mathematical Physics, 50(2):022102, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i2/p022102_s1.

Adam:2009:PVF

[1762] C. Adam, P. Klimas, J. Sánchez-Guillén, and A. Wereszczyński. Pullback of the volume form, integrable models in higher dimensions and exotic textures. Journal of Mathematical Physics, 50(2):022301, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i2/p022301_s1.

Shapiro:2009:FML

[1763] I. Shapiro. Frobenius map on local Calabi–Yau manifolds. Journal of Mathematical Physics, 50(2):022302, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i2/p022302_s1.

Amorim:2009:FNT

[1764] Ricardo Amorim. Fermions and noncommutative theories. Journal of Mathematical Physics, 50(2):022303, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i2/p022303_s1.

Karp:2009:FB

[1765] Robert L. Karp. On the C^{n}/Z_{m} fractional branes. Journal of Mathematical Physics, 50(2):022304, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i2/p022304_s1.

Pradhan:2009:CSV

[1766] Anirudh Pradhan. Cylindrically symmetric viscous fluid universe in Lyra geometry. Journal of Mathematical Physics, 50(2):022501, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i2/p022501_s1.

Thirukkanesh:2009:RRM

[1767] S. Thirukkanesh and S. D. Maharaj. Radiating relativistic matter in geodesic motion. Journal of Mathematical Physics, 50(2):022502, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i2/p022502_s1.

Couch:2009:LQM

[1768] W. E. Couch and C. L. Holder. Liouvillian quasinormal modes of Reissner–Nordstrom black holes. Journal of Mathematical Physics, 50(2):022503, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i2/p022503_s1.

Shieh:2009:HCD

[1769] Shih-Feng Shieh. Horseshoes for coupled discrete nonlinear Schrödinger equations. Journal of Mathematical Physics, 50(2):022701, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i2/p022701_s1.

DEliseo:2009:GE

[1770] Maurizio M. D’Eliseo. The gravitational ellipse. Journal of Mathematical Physics, 50(2):022901, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i2/p022901_s1.

Ding:2009:GSO

[1771] Shijin Ding, Huanyao Wen, Lei Yao, and Changjiang Zhu. Global solutions to one-dimensional compressible Navier–Stokes–Poisson equations with density-dependent viscosity. Journal of Mathematical Physics, 50(2):023101, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i2/p023101_s1.

Tamura:2009:MFI

[1772] Hiroshi Tamura and Valentin A. Zagrebnov. Mean-field interacting boson random point fields in weak harmonic traps. Journal of Mathematical Physics, 50(2):023301, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i2/p023301_s1.

Jafarizadeh:2009:RCE

[1773] M. A. Jafarizadeh, R. Sufiani, and S. Jafarizadeh. Recursive calculation of effective resistances in distance-regular networks based on Bose– Mesner algebra and Christoffel–Darboux identity. Journal of Mathematical Physics, 50(2):023302, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i2/p023302_s1.

VanValkenburgh:2009:MSP

[1774] Michael VanValkenburgh. Manipulation of semiclassical photon states. Journal of Mathematical Physics, 50(2):023501, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i2/p023501_s1.

Qian:2009:OMM

[1775] Zhi Qian. An optimal modified method for a two-dimensional inverse heat conduction problem. Journal of Mathematical Physics, 50(2):023502, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i2/p023502_s1.

Huang:2009:PGC

[1776] Qing Huang, V. Lahno, C. Z. Qu, and R. Zhdanov. Preliminary group classification of a class of fourth-order evolution equations. Journal of Mathematical Physics, 50(2):023503, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i2/p023503_s1.

Yang:2009:SES

[1777] Jianke Yang and D. J. Kaup. Squared eigenfunctions for the Sasa– Satsuma equation. Journal of Mathematical Physics, 50(2):023504, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i2/p023504_s1.

Chen:2009:BPS

[1778] Jianqing Chen and Boling Guo. Blow-up profile to the solutions of two-coupled Schrödinger equations. Journal of Mathematical Physics, 50(2):023505, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i2/p023505_s1.

Wang:2009:LST

[1779] Jing Ping Wang. Lenard scheme for two-dimensional periodic Volterra chain. Journal of Mathematical Physics, 50(2):023506, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i2/
p023506_s1.

Du:2009:RBC

[1780] Yi Du, Yan Liu, and Zhengan Yao. Remarks on the blow-up criteria for three-dimensional ideal magnetohydrodynamics equations. Journal of Mathematical Physics, 50(2):023507, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i2/p023507_s1.

Falek:2009:BOP

[1781] M. Falek and M. Merad. Bosonic oscillator in the presence of minimal length. Journal of Mathematical Physics, 50(2):023508, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i2/p023508_s1.

Filipuk:2009:MAS

[1782] G. Filipuk and R. G. Halburd. Movable algebraic singularities of secondorder ordinary differential equations. Journal of Mathematical Physics, 50(2):023509, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i2/p023509_s1.

Yildirim:2009:HPM

[1783] Ahmet Yildirim. Homotopy perturbation method to obtain exact special solutions with solitary patterns for Boussinesq-like B(m, n) equations with fully nonlinear dispersion. Journal of Mathematical Physics, 50(2):023510, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i2/p023510_s1.

Baumgartel:2009:SPR

[1784] Hellmut Baumgärtel, Hani Kaldass, and Soliman Komy. On spectral properties of the resonances for selected potential scattering systems. Journal of Mathematical Physics, 50(2):023511, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i2/p023511_s1.

Palmer:2009:EAB

[1785] Bennett Palmer. Equilibria for anisotropic bending energies. Journal of Mathematical Physics, 50(2):023512, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i2/p023512_s1.

Albuquerque:2009:TQC

[1786] C. D. Albuquerque, R. Palazzo, Jr., and E. B. Silva. Topological quantum codes on compact surfaces with genus g ≥ 2. Journal of Mathematical Physics, 50(2):023513, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i2/p023513_s1.

Komech:2009:ACS

[1787] A. I. Komech and A. E. Merzon. On asymptotic completeness for scattering in the nonlinear lamb system. Journal of Mathematical Physics, 50(2):023514, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i2/p023514_s1.

Cvijovic:2009:DIA

[1788] Djurdje Cvijović. A dilogarithmic integral arising in quantum field theory. Journal of Mathematical Physics, 50(2):023515, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i2/
p023515_s1.

Cho:2009:SGP

[1789] Ilwoo Cho and Palle Jorgensen. C^{∗}-subalgebras generated by partial isometries. Journal of Mathematical Physics, 50(2):023516, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i2/p023516_s1.

Marinakis:2009:CNI

[1790] Vangelis Marinakis. Comment on “A new integrable equation with cuspons and W/M-shape-peaks solitons”, [J. Math. Phys. 47, 112701 (2006)]. Journal of Mathematical Physics, 50(2):024101, February 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i2/
p024101_s1. See [812].

Du:2009:SCG

[1791] Hong-Ke Du and Yan-Ni Dou. A spectral characterization for generalized quantum gates. Journal of Mathematical Physics, 50(3):032101, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i3/p032101_s1.

Buisseret:2009:SEB

[1792] Fabien Buisseret, Claude Semay, and Bernard Silvestre-Brac. Some equivalences between the auxiliary field method and envelope theory. Journal of Mathematical Physics, 50(3):032102, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i3/p032102_s1.

Lawton:2009:SRB

[1793] Wayne Lawton, Anders S. Mouritzen, Jiao Wang, and Jiangbin Gong. Spectral relationships between kicked Harper and on-resonance double kicked rotor operators. Journal of Mathematical Physics, 50(3): 032103, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i3/p032103_s1.

Combescure:2009:BCM

[1794] Monique Combescure. Block-circulant matrices with circulant blocks, Weil sums, and mutually unbiased bases. II. The prime power case. Journal of Mathematical Physics, 50(3):032104, March 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i3/p032104_s1.

Mosonyi:2009:HTG

[1795] Milán Mosonyi. Hypothesis testing for Gaussian states on bosonic lattices. Journal of Mathematical Physics, 50(3):032105, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i3/
p032105_s1.

Tyson:2009:TSE

[1796] Jon Tyson. Two-sided estimates of minimum-error distinguishability of mixed quantum states via generalized Holevo–Curlander bounds. Journal of Mathematical Physics, 50(3):032106, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i3/p032106_s1. See comment

[1907] and erratum [2097].

Suto:2009:PMC

[1797] András Sütö. A possible mechanism of concurring diagonal and offdiagonal long-range order for soft interactions. Journal of Mathematical Physics, 50(3):032107, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i3/p032107_s1.

Zhang:2009:DSK

[1798] Fu-Lin Zhang and Jing-Ling Chen. Dynamical symmetries of the Klein– Gordon equation. Journal of Mathematical Physics, 50(3):032301, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i3/p032301_s1.

Aazami:2009:UMTa

[1799] A. B. Aazami and A. O. Petters. A universal magnification theorem for higher-order caustic singularities. Journal of Mathematical Physics, 50(3):032501, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i3/p032501_s1.

Zhijian:2009:GAC

[1800] Yang Zhijian and Jin Baoxia. Global attractor for a class of Kirchhoff models. Journal of Mathematical Physics, 50(3):032701, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i3/p032701_s1.

Manton:2009:SIS

[1801] Nicholas S. Manton. Scaling identities for solitons beyond Derrick’s theorem. Journal of Mathematical Physics, 50(3):032901, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i3/
p032901_s1.

Ma:2009:BLI

[1802] Tian Ma and Shouhong Wang. Boundary-layer and interior separations in the Taylor–Couette–Poiseuille flow. Journal of Mathematical Physics, 50(3):033101, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i3/p033101_s1.

Rebenko:2009:QAS

[1803] A. L. Rebenko and M. V. Tertychnyi. Quasilattice approximation of statistical systems with strong superstable interactions: Correlation functions. Journal of Mathematical Physics, 50(3):033301, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i3/
p033301_s1.

Bassler:2009:ESS

[1804] K. E. Bassler, P. J. Forrester, and N. E. Frankel. Eigenvalue separation in some random matrix models. Journal of Mathematical Physics, 50(3): 033302, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i3/p033302_s1.

Kawamura:2009:JTE

[1805] Katsunori Kawamura, Dan Lascu, and Ion Coltescu. Jump transformations and an embedding of O_{∞} into O_{2}. Journal of Mathematical Physics, 50(3):033501, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i3/p033501_s1.

Xu:2009:EES

[1806] Xiao-Ming Xu, Hong-Ke Du, and Xiaochun Fang. An explicit expression of supremum of bounded quantum observables. Journal of Mathematical Physics, 50(3):033502, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i3/p033502_s1.

Tian:2009:GWP

[1807] Lixin Tian and Shujuan Liang. Global well posedness and limit behavior of the solutions to the viscous Degasperis–Procesi equation. Journal of Mathematical Physics, 50(3):033503, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i3/p033503_s1.

Skrypnyk:2009:NSS

[1808] T. Skrypnyk. Non-skew-symmetric classical r-matrices, algebraic Bethe ansatz, and Bardeen–Cooper–Schrieffer-type integrable systems. Journal of Mathematical Physics, 50(3):033504, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i3/p033504_s1.

Amour:2009:ISR

[1809] L. Amour, J. Faupin, and T. Raoux. Inverse spectral results for Schrödinger operators on the unit interval with partial information given on the potentials. Journal of Mathematical Physics, 50(3):033505, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i3/p033505_s1.

Tucker:2009:DFV

[1810] Robin W. Tucker. Differential form valued forms and distributional electromagnetic sources. Journal of Mathematical Physics, 50(3):033506, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i3/p033506_s1.

Fernandez:2009:BHM

[1811] M. Fernández, A. Tomassini, L. Ugarte, and R. Villacampa. Balanced Hermitian metrics from SU(2)-structures. Journal of Mathematical Physics, 50(3):033507, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i3/p033507_s1.

Pan:2009:MCM

[1812] Xing-Bin Pan. Minimizing curl in a multiconnected domain. Journal of Mathematical Physics, 50(3):033508, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i3/p033508_s1.

Dokovic:2009:PIS

[1813] D. Ž. – Doković and A. Osterloh. On polynomial invariants of several qubits. Journal of Mathematical Physics, 50(3):033509, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i3/
p033509_s1.

Cai:2009:MPS

[1814] Jiaxiang Cai, Yushun Wang, and Zhonghua Qiao. Multisymplectic Preissman scheme for the time-domain Maxwell’s equations. Journal of Mathematical Physics, 50(3):033510, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i3/p033510_s1.

Vinet:2009:AHW

[1815] Luc Vinet and Alexei Zhedanov. Automorphisms of the Heisenberg– Weyl algebra and d-orthogonal polynomials. Journal of Mathematical Physics, 50(3):033511, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i3/p033511_s1.

Dolya:2009:QLA

[1816] S. N. Dolya. Quadratic Lie algebras and quasiexact solvability of the two-photon Rabi Hamiltonian. Journal of Mathematical Physics, 50(3): 033512, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i3/p033512_s1.

Casas:2009:EAC

[1817] Fernando Casas and Ander Murua. An efficient algorithm for computing the Baker–Campbell–Hausdorff series and some of its applications. Journal of Mathematical Physics, 50(3):033513, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i3/p033513_s1.

Zhang:2009:IDI

[1818] Zhongzhi Zhang, Yichao Zhang, Shuigeng Zhou, Ming Yin, and Jihong Guan. Influences of degree inhomogeneity on average path length and random walks in disassortative scale-free networks. Journal of Mathematical Physics, 50(3):033514, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i3/p033514_s1. See publisher’s note

[1941].

Chiribella:2009:RSQ

[1819] Giulio Chiribella, Giacomo Mauro D’Ariano, and Paolo Perinotti. Realization schemes for quantum instruments in finite dimensions. Journal of Mathematical Physics, 50(4):042101, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p042101_s1.

Cintio:2009:SRD

[1820] A. Cintio and G. Morchio. Sum rules and density waves spectrum for nonrelativistic fermions. Journal of Mathematical Physics, 50(4):042102, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042102_s1.

Aaberg:2009:CMA

[1821] Johan Åberg and Graeme Mitchison. Cumulants and the moment algebra: Tools for analyzing weak measurements. Journal of Mathematical Physics, 50(4):042103, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i4/p042103_s1.

Christandl:2009:FFT

[1822] Matthias Christandl and Ben Toner. Finite de Finetti theorem for conditional probability distributions describing physical theories. Journal of Mathematical Physics, 50(4):042104, April 2009. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p042104_s1.

Ho:2009:SIP

[1823] Choon-Lin Ho. Shape invariance in prepotential approach to exactly solvable models. Journal of Mathematical Physics, 50(4):042105, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042105_s1.

Ambainis:2009:NEQ

[1824] Andris Ambainis, Jan Bouda, and Andreas Winter. Nonmalleable encryption of quantum information. Journal of Mathematical Physics, 50(4):042106, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i4/p042106_s1.

Ammari:2009:MFL

[1825] Z. Ammari and F. Nier. Mean field limit for bosons and propagation of Wigner measures. Journal of Mathematical Physics, 50(4):042107, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042107_s1.

Shirai:2009:SAP

[1826] Shin ichi Shirai. Some applications of the Pohozaev identity. Journal of Mathematical Physics, 50(4):042108, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p042108_s1.

Chuang:2009:CSQ

[1827] Isaac Chuang, Andrew Cross, Graeme Smith, John Smolin, and Bei Zeng. Codeword stabilized quantum codes: Algorithm and structure. Journal of Mathematical Physics, 50(4):042109, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p042109_s1.

Barnich:2009:DIE

[1828] Glenn Barnich and Cédric Troessaert. Duality and integrability: Electromagnetism, linearized gravity, and massless higher spin gauge fields as bi-Hamiltonian systems. Journal of Mathematical Physics, 50(4):042301, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042301_s1.

Bogner:2009:PFI

[1829] Christian Bogner and Stefan Weinzierl. Periods and Feynman integrals. Journal of Mathematical Physics, 50(4):042302, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i4/p042302_s1.

Savadkouhi:2009:ATJ

[1830] M. Bavand Savadkouhi, M. Eshaghi Gordji, J. M. Rassias, and N. Ghobadipour. Approximate ternary Jordan derivations on Banach ternary algebras. Journal of Mathematical Physics, 50(4):042303, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v50/i4/p042303_s1. See comment [2198].

Giribet:2009:CTP

[1831] Gaston Giribet and Lorena Nicolás. Comment on three-point function in AdS(3)/CFT(2). Journal of Mathematical Physics, 50(4):042304, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042304_s1.

Wickramasekara:2009:RDQa

[1832] S. Wickramasekara. Relativistic dynamics of quasistable states. I. Perturbation theory for the Poincaré group. Journal of Mathematical Physics, 50(4):042305, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i4/p042305_s1.

deBerredo-Peixoto:2009:TDG

[1833] G. de Berredo-Peixoto and M. O. Katanaev. Tube dislocations in gravity. Journal of Mathematical Physics, 50(4):042501, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i4/p042501_s1.

Mitra:2009:CEG

[1834] Abhas Mitra. Comments on “The Euclidean gravitational action as black hole entropy, singularities, and space–time voids”, [J. Math. Phys. 49, 042501 (2008)]. Journal of Mathematical Physics, 50(4):042502, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042502_s1. See [1426, 2217].

Ramm:2009:IMD

[1835] A. G. Ramm. Invariant manifolds for dissipative systems. Journal of Mathematical Physics, 50(4):042701, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p042701_s1.

Song:2009:BND

[1836] Haitao Song and Chengkui Zhong. Blow up of the nonclassical diffusion equation. Journal of Mathematical Physics, 50(4):042702, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042702_s1.

Chiba:2009:LEA

[1837] Hayato Chiba and Masatomo Iwasa. Lie equations for asymptotic solutions of perturbation problems of ordinary differential equations. Journal of Mathematical Physics, 50(4):042703, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p042703_s1.

Ivanchenko:2009:DET

[1838] E. A. Ivanchenko. Dynamic equations for three different qudits in a magnetic field. Journal of Mathematical Physics, 50(4):042704, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042704_s1.

Cronstrom:2009:NSV

[1839] Christofer Cronström and Tommi Raita. On nonholonomic systems and variational principles. Journal of Mathematical Physics, 50(4):042901, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042901_s1.

Sonego:2009:FAR

[1840] Sebastiano Sonego and Massimo Pin. Foundations of anisotropic relativistic mechanics. Journal of Mathematical Physics, 50(4):042902, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p042902_s1.

Jimenez-Lara:2009:GPP

[1841] Lidia Jiménez-Lara, Jaume Llibre, and Martín Vargas. Global phase portraits of the planar perpendicular problem of two fixed centers. Journal of Mathematical Physics, 50(4):042903, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p042903_s1.

Bauer:2009:CSL

[1842] Michel Bauer, Denis Bernard, and Tom Kennedy. Conditioning Schramm–Loewner evolutions and loop erased random walks. Journal of Mathematical Physics, 50(4):043301, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p043301_s1.

Yamano:2009:GKL

[1843] Takuya Yamano. A generalization of the Kullback–Leibler divergence and its properties. Journal of Mathematical Physics, 50(4):043302, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p043302_s1.

Tsallis:2009:EMV

[1844] C. Tsallis, A. R. Plastino, and R. F. Alvarez-Estrada. Escort mean values and the characterization of power-law-decaying probability densities. Journal of Mathematical Physics, 50(4):043303, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i4/p043303_s1.

Huang:2009:QAQ

[1845] Hua-Lin Huang. Quiver approaches to quasi-Hopf algebras. Journal of Mathematical Physics, 50(4):043501, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p043501_s1.

Sahadevan:2009:HDI

[1846] R. Sahadevan and S. Rajakumar. Higher dimensional integrable mappings derived from coupled discrete nonlinear Schrödinger equations. Journal of Mathematical Physics, 50(4):043502, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i4/p043502_s1.

Chen:2009:CSS

[1847] Bang-Yen Chen. Classification of spatial surfaces with parallel mean curvature vector in pseudo-Euclidean spaces of arbitrary dimension. Journal of Mathematical Physics, 50(4):043503, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p043503_s1.

Hlavaty:2009:PLP

[1848] Ladislav Hlavatý, Ivo Petr, and Vojtech Stepán. Poisson–Lie T -plurality with spectators. Journal of Mathematical Physics, 50(4):043504, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p043504_s1.

Geloun:2009:CSN

[1849] J. Ben Geloun and F. G. Scholtz. Coherent states in noncommutative quantum mechanics. Journal of Mathematical Physics, 50(4):043505, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p043505_s1.

Gao:2009:EUE

[1850] Zhifeng Gao and Jing Zhang. Existence, uniqueness, and equivalence theorems for magnetic monopoles in general (4 p-1)-dimensional Yang– Mills theory. Journal of Mathematical Physics, 50(4):043506, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i4/
p043506_s1.

Khochman:2009:RSS

[1851] Abdallah Khochman. Resonances and spectral shift function for a magnetic Schrödinger operator. Journal of Mathematical Physics, 50(4): 043507, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i4/p043507_s1.

Klimcik:2009:IYB

[1852] Ctirad Klimcík. On integrability of the Yang–Baxter σ-model. Journal of Mathematical Physics, 50(4):043508, April 2009. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p043508_s1.

Bagarello:2009:IOB

[1853] F. Bagarello. Intertwining operators between different Hilbert spaces: Connection with frames. Journal of Mathematical Physics, 50(4):043509, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p043509_s1.

Hoppe:2009:SAM

[1854] Jens Hoppe, Douglas Lundholm, and Maciej Trzetrzelewski. Spin(9) average of SU(N) matrix models. Journal of Mathematical Physics, 50(4): 043510, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i4/p043510_s1.

Nimmo:2009:TSV

[1855] J. J. C. Nimmo and S. N. M. Ruijsenaars. Tzitzeica solitons versus relativistic Calogero–Moser three-body clusters. Journal of Mathematical Physics, 50(4):043511, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i4/p043511_s1.

Rasmussen:2009:PFR

[1856] Jørgen Rasmussen. Polynomial fusion rings of W -extended logarithmic minimal models. Journal of Mathematical Physics, 50(4):043512, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p043512_s1.

Zertuche:2009:RKN

[1857] Federico Zertuche. On the robustness of NK-Kauffman networks against changes in their connections and Boolean functions. Journal of Mathematical Physics, 50(4):043513, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i4/p043513_s1.

Camargo:2009:SFG

[1858] R. Figueiredo Camargo, R. Charnet, and E. Capelas de Oliveira. On some fractional Green’s functions. Journal of Mathematical Physics, 50(4):

043514, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i4/p043514_s1.

Stephan:2009:KDS

[1859] Christoph A. Stephan. Krajewski diagrams and the standard model. Journal of Mathematical Physics, 50(4):043515, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i4/p043515_s1.

Alabina:2009:ASS

[1860] Y. F. Alabina, A. V. Latyshev, and A. A. Yushkanov. Analytical solution of the skin effect problem with the displacement current in Maxwell plasma by the source method. Journal of Mathematical Physics, 50(4): 043516, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i4/p043516_s1.

Ali:2009:NRC

[1861] S. Twareque Ali, Lubka Balkova, E. M. F. Curado, J. P. Gazeau, M. A. Rego-Monteiro, Ligia M. C. S. Rodrigues, and K. Sekimoto. Noncommutative reading of the complex plane through Delone sequences. Journal of Mathematical Physics, 50(4):043517, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i4/p043517_s1.

Wei:2009:MRC

[1862] Yi Wei. Moments of ratios of characteristic polynomials of a certain class of random matrices. Journal of Mathematical Physics, 50(4):043518, April 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i4/p043518_s1.

Nachtergaele:2009:EJM

[1863] Bruno L. Z. Nachtergaele. Editorial: Journal of Mathematical Physics turns 50. Journal of Mathematical Physics, 50(5):050401, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/
p050401_s1.

Viennot:2009:NFH

[1864] David Viennot. The need for a flat higher gauge structure to describe a Berry phase associated with some resonance phenomena. Journal of Mathematical Physics, 50(5):052101, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i5/p052101_s1.

Rudolph:2009:AQO

[1865] G. Rudolph and M. Schmidt. On the algebra of quantum observables for a certain gauge model. Journal of Mathematical Physics, 50(5): 052102, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p052102_s1.

Amorim:2009:TCN

[1866] Ricardo Amorim. Tensor coordinates in noncommutative mechanics. Journal of Mathematical Physics, 50(5):052103, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/p052103_s1.

Fakhri:2009:CBG

[1867] H. Fakhri and A. Dehghani. Coherency of su(1, 1)-Barut–Girardello type and entanglement for spherical harmonics. Journal of Mathematical Physics, 50(5):052104, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i5/p052104_s1.

Cannata:2009:TDS

[1868] F. Cannata, M. V. Ioffe, and D. N. Nishnianidze. Three-dimensional solutions of supersymmetrical intertwining relations and pairs of isospectral Hamiltonians. Journal of Mathematical Physics, 50(5):052105, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i5/p052105_s1.

Marino:2009:MIM

[1869] Marcos Mariño, Ricardo Schiappa, and Marlene Weiss. Multi-instantons and multicuts. Journal of Mathematical Physics, 50(5):052301, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i5/p052301_s1.

Zamani:2009:QMP

[1870] Farhad Zamani and Ali Mostafazadeh. Quantum mechanics of Proca fields. Journal of Mathematical Physics, 50(5):052302, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/
p052302_s1.

Chrusciel:2009:HDB

[1871] Piotr T. Chruściel. On higher dimensional black holes with Abelian isometry group. Journal of Mathematical Physics, 50(5):052501, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i5/p052501_s1.

Sobreira:2009:CSE

[1872] A. A. R. Sobreira, Geusa de A. Marques, J. B. Fonseca-Neto, and V. B. Bezerra. On a class of solutions of the Einstein–Maxwell field equations in scalar-tensor theories of gravity. Journal of Mathematical Physics, 50 (5):052502, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p052502_s1.

Joung:2009:TDQ

[1873] E. Joung, J. Mourad, and K. Noui. Three dimensional quantum geometry and deformed symmetry. Journal of Mathematical Physics, 50(5): 052503, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p052503_s1.

Qin:2009:DRW

[1874] Wen-Xin Qin and Pei-Lin Zhang. Discrete rotating waves in a ring of coupled mechanical oscillators with strong damping. Journal of Mathematical Physics, 50(5):052701, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i5/p052701_s1.

ElBialy:2009:CBR

[1875] Mohamed Sami ElBialy. Collective branch regularization of simultaneous binary collisions in the 3D N-body problem. Journal of Mathematical Physics, 50(5):052702, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i5/p052702_s1.

Guo:2009:DSZ

[1876] Boling Guo, Yan Lv, and Xiaoping Yang. Dynamics of stochastic Zakharov equations. Journal of Mathematical Physics, 50(5):052703, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i5/p052703_s1.

Pradeep:2009:NCH

[1877] R. Gladwin Pradeep, V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan. Nonstandard conserved Hamiltonian structures in dissipative/ damped systems: Nonlinear generalizations of damped harmonic oscillator. Journal of Mathematical Physics, 50(5):052901, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/p052901_s1.

Cimmelli:2009:NTF

[1878] V. A. Cimmelli, A. Sellitto, and V. Triani. A new thermodynamic framework for second-grade Korteweg-type viscous fluids. Journal of Mathematical Physics, 50(5):053101, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i5/p053101_s1.

Janaswamy:2009:TPF

[1879] R. Janaswamy. Transitional probabilities for the four-state random walk on a lattice in the presence of partially reflecting boundaries. Journal of Mathematical Physics, 50(5):053301, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i5/p053301_s1.

Ha:2009:GEC

[1880] Seung-Yeal Ha and Ho Lee. Global existence of classical solutions to the damped Vlasov–Klein–Gordon equations with small data. Journal of Mathematical Physics, 50(5):053302, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i5/p053302_s1.

Genovese:2009:MAM

[1881] Giuseppe Genovese and Adriano Barra. A mechanical approach to mean field spin models. Journal of Mathematical Physics, 50(5):053303, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i5/p053303_s1.

Yang:2009:SSS

[1882] Yisong Yang and Ruifeng Zhang. Steady state solutions for nonlinear Schrödinger equation arising in optics. Journal of Mathematical Physics, 50(5):053501, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i5/p053501_s1. See erratum [2200].

Luczak:2009:MPD

[1883] Andrzej Luczak. Maximizing the probability of detection for pure states. Journal of Mathematical Physics, 50(5):053502, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/p053502_s1.

Anishetty:2009:ISB

[1884] Ramesh Anishetty, Manu Mathur, and Indrakshi Raychowdhury. Irreducible SU(3) Schwinger bosons. Journal of Mathematical Physics, 50 (5):053503, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p053503_s1.

Katore:2009:GPG

[1885] S. D. Katore and R. S. Rane. Generalized plane gravitational waves of weakened field equations in general relativity. Journal of Mathematical Physics, 50(5):053504, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i5/p053504_s1.

Sahadevan:2009:SRN

[1886] R. Sahadevan and L. Nalinidevi. Similarity reduction, nonlocal and master symmetries of sixth order Korteweg–de Vries equation. Journal of Mathematical Physics, 50(5):053505, May 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i5/p053505_s1.

Liu:2009:GDA

[1887] Xiaojun Liu, Runliang Lin, Bo Jin, and Yunbo Zeng. A generalized dressing approach for solving the extended KP and the extended mKP hierarchy. Journal of Mathematical Physics, 50(5):053506, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/p053506_s1.

Horwood:2009:HJT

[1888] Joshua T. Horwood, Raymond G. McLenaghan, and Roman G. Smirnov. Hamilton–Jacobi theory in three-dimensional Minkowski space via Cartan geometry. Journal of Mathematical Physics, 50(5):053507, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i5/p053507_s1.

Yang:2009:ENV

[1889] Sze-Guang Yang. The existence of nontopological vortices in a twoparticle Chern–Simons system. Journal of Mathematical Physics, 50(5): 053508, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p053508_s1.

Poghosyan:2009:ACE

[1890] Suren Poghosyan and Daniel Ueltschi. Abstract cluster expansion with applications to statistical mechanical systems. Journal of Mathematical Physics, 50(5):053509, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i5/p053509_s1.

Hu:2009:IES

[1891] Juan Hu, Xing-Biao Hu, and Hon-Wah Tam. Ishimori-I equation with self-consistent sources. Journal of Mathematical Physics, 50(5):053510, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i5/p053510_s1.

Rossi:2009:HPT

[1892] Carlo A. Rossi. Holonomy and parallel transport in the differential geometry of loop spaces and generalized gauge transformations. Journal of Mathematical Physics, 50(5):053511, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i5/p053511_s1.

Franssens:2009:DDS

[1893] Ghislain R. Franssens. Delta distributions supported on quadratic O(p, q)-invariant surfaces. Journal of Mathematical Physics, 50(5): 053512, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p053512_s1.

Gordoa:2009:MRO

[1894] P. R. Gordoa and A. Pickering. A method of reduction of order for discrete systems. Journal of Mathematical Physics, 50(5):053513, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i5/p053513_s1.

Budinich:2009:CCC

[1895] Marco Budinich. On computational complexity of Clifford algebra. Journal of Mathematical Physics, 50(5):053514, May 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i5/p053514_s1.

Agrawal:2009:TDM

[1896] Gunjan Agrawal and Sampada Shrivastava. t-topology on the ndimensional Minkowski space. Journal of Mathematical Physics, 50(5): 053515, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p053515_s1.

Fatibene:2009:SOD

[1897] Lorenzo Fatibene, Raymond G. McLenaghan, Giovanni Rastelli, and Shane N. Smith. Symmetry operators for Dirac’s equation on twodimensional spin manifolds. Journal of Mathematical Physics, 50(5): 053516, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p053516_s1.

Bogoyavlensky:2009:MDI

[1898] Oleg Bogoyavlensky. Method of descent for integrable lattices. Journal of Mathematical Physics, 50(5):053517, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i5/p053517_s1.

Yomba:2009:MIE

[1899] Emmanuel Yomba and George R. Sell. Modulational instability and exact solutions for a three-component system of vector nonlinear Schrödinger equations. Journal of Mathematical Physics, 50(5):053518, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/
p053518_s1.

Huan-he:2009:SLA

[1900] Dong Huan-he. A subalgebra of Lie algebra A_{2} and its associated two types of loop algebras, as well as Hamiltonian structures of integrable hierarchy. Journal of Mathematical Physics, 50(5):053519, May 2009.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/
p053519_s1.

Dinlemez:2009:PFF

[1901] Ülkü Dinlemez and Gülten Torun. Properties of the far-field operator for a multilayered scatterer in acoustic scattering theory. Journal of Mathematical Physics, 50(5):053520, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i5/p053520_s1.

Kawamura:2009:UFB

[1902] Katsunori Kawamura. Universal fermionization of bosons on permutative representations of the Cuntz algebra O_{2}. Journal of Mathematical Physics, 50(5):053521, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i5/p053521_s1.

Zhdanov:2009:RBP

[1903] Renat Zhdanov. On relation between potential and contact symmetries of evolution equations. Journal of Mathematical Physics, 50(5): 053522, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p053522_s1.

Paal:2009:DDT

[1904] Eugen Paal and Jüri Virkepu. Dynamical deformations of threedimensional Lie algebras in Bianchi classification over the harmonic oscillator. Journal of Mathematical Physics, 50(5):053523, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i5/
p053523_s1.

Yang:2009:FGS

[1905] Xiao Yang and Junmin Wang. Finite genus solution of a new (2+1) Burgers equation with a discrete variable. Journal of Mathematical Physics, 50 (5):053524, May 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i5/p053524_s1.

Kotulek:2009:NSN

[1906] Jan Kotulek. Nontrivial systems and the necessity of the scalar quantum mechanics axioms. Journal of Mathematical Physics, 50(6):062101, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i6/p062101_s1.

Tyson:2009:MEQ

[1907] Jon Tyson. Minimum-error quantum distinguishability bounds from matrix monotone functions: A comment on “two-sided estimates of minimum-error distinguishability of mixed quantum states via generalized holevo–curlander bounds”, [J. math. phys. 50, 032106 (2009)]. Journal of Mathematical Physics, 50(6):062102, June 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i6/p062102_s1. See

[2097, 1796].

Cornean:2009:IEM

[1908] H. D. Cornean, J. Dereziński, and P. Ziń. On the infimum of the energymomentum spectrum of a homogeneous Bose gas. Journal of Mathematical Physics, 50(6):062103, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i6/p062103_s1.

Korbicz:2009:EQG

[1909] J. K. Korbicz, J. Wehr, and M. Lewenstein. Entanglement and quantum groups. Journal of Mathematical Physics, 50(6):062104, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/
p062104_s1.

Laing:2009:UGT

[1910] W. B. Laing, M. Dunn, and D. K. Watson. On the use of group theoretical and graphical techniques toward the solution of the general N-body problem. Journal of Mathematical Physics, 50(6):062105, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/p062105_s1.

Skowronek:2009:CPM

[1911] Lukasz Skowronek, Erling Størmer, and Karol Zyczkowski. Cones of positive maps and their duality relations. Journal of Mathematical Physics, 50(6):062106, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i6/p062106_s1.

Delbourgo:2009:IME

[1912] R. Delbourgo and D. Elliott. Inverse momentum expectation values for hydrogenic systems. Journal of Mathematical Physics, 50(6):062107, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i6/p062107_s1.

Beny:2009:QEC

[1913] Cédric Bény, Achim Kempf, and David W. Kribs. Quantum error correction on infinite-dimensional Hilbert spaces. Journal of Mathematical Physics, 50(6):062108, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i6/p062108_s1.

Amooshahi:2009:CQE

[1914] M. Amooshahi. Canonical quantization of electromagnetic field in an anisotropic polarizable and magnetizable medium. Journal of Mathematical Physics, 50(6):062301, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i6/p062301_s1.

Takaesu:2009:SAQ

[1915] Toshimitsu Takaesu. On the spectral analysis of quantum electrodynamics with spatial cutoffs. I. Journal of Mathematical Physics, 50(6): 062302, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i6/p062302_s1.

Torre:2009:SRQ

[1916] C. G. Torre. Symmetry reduction of quasi-free states. Journal of Mathematical Physics, 50(6):062303, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i6/p062303_s1.

Dappiaggi:2009:DQS

[1917] Claudio Dappiaggi, Valter Moretti, and Nicola Pinamonti. Distinguished quantum states in a class of cosmological spacetimes and their Hadamard property. Journal of Mathematical Physics, 50(6):062304, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/
p062304_s1.

Bessa:2009:BMR

[1918] Carlos H. G. Béssa, V. B. Bezerra, and L. H. Ford. Brownian motion in Robertson–Walker spacetimes from electromagnetic vacuum fluctuations. Journal of Mathematical Physics, 50(6):062501, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/
p062501_s1.

DeLuca:2009:VPW

[1919] Jayme De Luca. Variational principle for the Wheeler–Feynman electrodynamics. Journal of Mathematical Physics, 50(6):062701, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/
p062701_s1.

Macfarlane:2009:ECA

[1920] A. J. Macfarlane. On the evolution of the cellular automaton of rule 150 from some simple initial states. Journal of Mathematical Physics, 50(6): 062702, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i6/p062702_s1.

Qian:2009:IRR

[1921] Min-Ping Qian and Jian-Sheng Xie. On the imaginary-real ratio rule of power spectra. Journal of Mathematical Physics, 50(6):063301, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v50/i6/p063301_s1. See comment [2197].

Jiang:2009:ECO

[1922] Tiefeng Jiang. The entries of circular orthogonal ensembles. Journal of Mathematical Physics, 50(6):063302, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i6/p063302_s1.

Fan:2009:MES

[1923] Hong yi Fan and Xue xiang Xu. Mean energy of some interacting bosonic systems derived by virtue of the generalized Hellmann–Feynman theorem. Journal of Mathematical Physics, 50(6):063303, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/p063303_s1.

Zeitlin:2009:SFT

[1924] Anton M. Zeitlin. String field theory-inspired algebraic structures in gauge theories. Journal of Mathematical Physics, 50(6):063501, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i6/p063501_s1.

Miao:2009:CSS

[1925] Rong-Xin Miao. The covariant and on-shell statistics in κ-deformed space–time. Journal of Mathematical Physics, 50(6):063502, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/
p063502_s1.

Gade:2009:SSC

[1926] R. M. Gade. A set of sums for continuous dual q-2-Hahn polynomials. Journal of Mathematical Physics, 50(6):063503, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/p063503_s1.

Akemann:2009:GPN

[1927] G. Akemann, M. J. Phillips, and L. Shifrin. Gap probabilities in non-Hermitian random matrix theory. Journal of Mathematical Physics, 50 (6):063504, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i6/p063504_s1.

Zenchuk:2009:RBN

[1928] A. I. Zenchuk. On the relationship between nonlinear equations integrable by the method of characteristics and equations associated with commuting vector fields. Journal of Mathematical Physics, 50(6):063505, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i6/p063505_s1.

Ruggieri:2009:SRT

[1929] Marianna Ruggieri and Antonino Valenti. Symmetries and reduction techniques for dissipative models. Journal of Mathematical Physics, 50 (6):063506, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i6/p063506_s1.

Camargo:2009:SFL

[1930] R. Figueiredo Camargo, Ary O. Chiacchio, R. Charnet, and E. Capelas de Oliveira. Solution of the fractional Langevin equation and the Mittag-Leffler functions. Journal of Mathematical Physics, 50(6):063507, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i6/p063507_s1.

Goze:2009:HAT

[1931] M. Goze and M. Rausch de Traubenberg. Hopf algebras for ternary algebras. Journal of Mathematical Physics, 50(6):063508, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i6/
p063508_s1.

Calvaruso:2009:TDI

[1932] Giovanni Calvaruso. Three-dimensional Ivanov–Petrova manifolds. Journal of Mathematical Physics, 50(6):063509, June 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i6/p063509_s1.

Goto:2009:EFP

[1933] Shin itiro Goto and Robin W. Tucker. Electromagnetic fields produced by moving sources in a curved beam pipe. Journal of Mathematical Physics, 50(6):063510, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i6/p063510_s1.

Cho:2009:AAG

[1934] Ilwoo Cho and Palle E. T. Jorgensen. Applications of automata and graphs: Labeling operators in Hilbert space. II. Journal of Mathematical Physics, 50(6):063511, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i6/p063511_s1.

Hoang:2009:IPH

[1935] N. S. Hoang and A. G. Ramm. An inverse problem for a heat equation with piecewise-constant thermal conductivity. Journal of Mathematical Physics, 50(6):063512, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i6/p063512_s1.

Jardim:2009:ETD

[1936] Marcos Jardim and Rafael F. Leão. On the eigenvalues of the twisted Dirac operator. Journal of Mathematical Physics, 50(6):063513, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i6/p063513_s1.

Gunara:2009:FBP

[1937] Bobby Eka Gunara and Freddy Permana Zen. Flat Bogomolnyi–Prasad– Sommerfeld domain walls on two-dimensional Kähler–Ricci soliton. Journal of Mathematical Physics, 50(6):063514, June 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i6/p063514_s1. See

publisher’s note [2096].

Ghosh:2009:BAS

[1938] Saugata Ghosh. Bulk asymptotics of skew-orthogonal polynomials for quartic double well potential and universality in the matrix model. Journal of Mathematical Physics, 50(6):063515, June 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i6/p063515_s1.

Cong:2009:QSC

[1939] Hongzi Cong, Jianjun Liu, and Xiaoping Yuan. Quasiperiodic solutions for the cubic complex Ginzburg–Landau equation. Journal of Mathematical Physics, 50(6):063516, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i6/p063516_s1.

Park:2009:EEA

[1940] Jong Yeoul Park and Tae Gab Ha. Erratum: “Existence and asymptotic stability for the semilinear wave equation with boundary damping and source term”, [J. Math. Phys. 49, 053511 (2008)]. Journal of Mathematical Physics, 50(6):069901, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i6/p069901_s1. See erratum [1483].

Zhang:2009:PNI

[1941] Zhongzhi Zhang, Yichao Zhang, Shuigeng Zhou, Ming Yin, and Jihong Guan. Publisher’s note: “Influences of degree inhomogeneity on average path length and random walks in disassortative scale-free networks”, [J. Math. Phys. 50, 033514 (2009)]. Journal of Mathematical Physics, 50(6):

069902, June 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i6/p069902_s1. See [1818].

Biswas:2009:QSM

[1942] Indranil Biswas. Quantization of a symplectic manifold associated to a manifold with projective structure. Journal of Mathematical Physics, 50 (7):072101, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i7/p072101_s1.

Buyukasik:2009:ESQ

[1943] Sirin A. Büyükasik, Oktay K. Pashaev, and Esra Tigrak-Ulas. Exactly solvable quantum Sturm–Liouville problems. Journal of Mathematical Physics, 50(7):072102, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i7/p072102_s1.

Miyao:2009:RVW

[1944] Tadahiro Miyao and Herbert Spohn. The retarded van der Waals potential: Revisited. Journal of Mathematical Physics, 50(7):072103, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i7/p072103_s1.

Mosonyi:2009:GRE

[1945] Milán Mosonyi and Nilanjana Datta. Generalized relative entropies and the capacity of classical-quantum channels. Journal of Mathematical Physics, 50(7):072104, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i7/p072104_s1.

Khuri:2009:LEP

[1946] N. N. Khuri, André Martin, J.-M. Richard, and Tai Tsun Wu. Lowenergy potential scattering in two and three dimensions. Journal of Mathematical Physics, 50(7):072105, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i7/p072105_s1.

Bar-on:2009:DWF

[1947] T. Bar-on. Discrete Wigner function by symmetric informationally complete positive operator valued measure. Journal of Mathematical Physics, 50(7):072106, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i7/p072106_s1.

Meng:2009:KP

[1948] Guowu Meng. The Sp(1)-Kepler problems. Journal of Mathematical Physics, 50(7):072107, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i7/p072107_s1.

Domenech:2009:MWM

[1949] G. Domenech, H. Freytes, and C. de Ronde. Many worlds and modality in the interpretation of quantum mechanics: An algebraic approach. Journal of Mathematical Physics, 50(7):072108, July 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i7/p072108_s1.

Jureit:2009:CIA

[1950] Jan-Hendrik Jureit and Christoph A. Stephan. On a classification of irreducible almost-commutative geometries V. Journal of Mathematical Physics, 50(7):072301, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i7/p072301_s1.

Wickramasekara:2009:RDQb

[1951] S. Wickramasekara. Relativistic dynamics of quasistable states. II. Differentiable representations of the causal Poincaré semigroup. Journal of Mathematical Physics, 50(7):072302, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i7/p072302_s1.

Chae:2009:SNL

[1952] Myeongju Chae and Hyungjin Huh. Semi-nonrelativistic limit of the Chern–Simons–Higgs system. Journal of Mathematical Physics, 50(7): 072303, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i7/p072303_s1.

Giambo:2009:FRP

[1953] Roberto Giambò, Fabio Giannoni, and Antonio Masiello. Functional regularity properties for light rays in general relativity. Journal of Mathematical Physics, 50(7):072501, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i7/p072501_s1.

Tanabe:2009:AFS

[1954] Kentaro Tanabe, Norihiro Tanahashi, and Tetsuya Shiromizu. Asymptotic flatness at spatial infinity in higher dimensions. Journal of Mathematical Physics, 50(7):072502, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i7/p072502_s1.

Petters:2009:MTSa

[1955] A. O. Petters, B. Rider, and A. M. Teguia. A mathematical theory of stochastic microlensing. I. Random time delay functions and lensing maps. Journal of Mathematical Physics, 50(7):072503, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i7/
p072503_s1.

Avan:2009:CFT

[1956] Jean Avan and Geneviève Rollet. On Calogero–Françoise-type Lax matrices and their dynamical r-matrices. Journal of Mathematical Physics, 50 (7):072701, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i7/p072701_s1.

Xie:2009:CEO

[1957] H. Y. Xie, P. T. Leung, and D. P. Tsai. Clarification and extension of the optical reciprocity theorem. Journal of Mathematical Physics, 50(7): 072901, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i7/p072901_s1.

Khorrami:2009:PSS

[1958] Mohammad Khorrami, Amir H. Fatollahi, and Ahmad Shariati. Poisson structure on a space with linear SU(2) fuzziness. Journal of Mathematical Physics, 50(7):072902, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i7/p072902_s1.

Boland:2009:LCI

[1959] G. Boland. Long cycles in the infinite-range-hopping Bose–Hubbard model. Journal of Mathematical Physics, 50(7):073301, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i7/
p073301_s1.

Feral:2009:LES

[1960] Delphine Féral and Sandrine Péché. The largest eigenvalues of sample covariance matrices for a spiked population: Diagonal case. Journal of Mathematical Physics, 50(7):073302, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i7/p073302_s1.

Adams:2009:HER

[1961] Fred C. Adams and Anthony M. Bloch. Hill’s equation with random forcing parameters: The limit of delta function barriers. Journal of Mathematical Physics, 50(7):073501, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i7/p073501_s1.

Szablikowski:2009:BHS

[1962] Blazej M. Szablikowski, Maciej Blaszak, and Burcu Silindir. Bi-Hamiltonian structures for integrable systems on regular time scales. Journal of Mathematical Physics, 50(7):073502, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i7/p073502_s1.

Vacaru:2009:SFN

[1963] Sergiu I. Vacaru. Spectral functionals, nonholonomic Dirac operators, and noncommutative Ricci flows. Journal of Mathematical Physics, 50 (7):073503, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i7/p073503_s1.

Batalin:2009:CSL

[1964] Igor A. Batalin and Klaus Bering. A comparative study of Laplacians and Schrödinger–Lichnerowicz–Weitzenböck identities in Riemannian and antisymplectic geometry. Journal of Mathematical Physics, 50(7):073504, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i7/p073504_s1.

Price:2009:ISS

[1965] Richard H. Price, Charalampos Markakis, and John L. Friedman. Iteration stability for simple Newtonian stellar systems. Journal of Mathematical Physics, 50(7):073505, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i7/p073505_s1.

Wang:2009:RME

[1966] Dong Wang. Random matrices with external source and KP τ functions. Journal of Mathematical Physics, 50(7):073506, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i7/p073506_s1.

Tian:2009:LRS

[1967] Lixin Tian, Yuexia Chen, Xiuping Jiang, and Limeng Xia. Low-regularity solutions of the periodic Fornberg–Whitham equation. Journal of Mathematical Physics, 50(7):073507, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i7/p073507_s1.

Grundland:2009:SEH

[1968] A. M. Grundland and A. J. Hariton. N = 2 supersymmetric extension of a hydrodynamic system in Riemann invariants. Journal of Mathematical Physics, 50(7):073508, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i7/p073508_s1.

Kalkanli:2009:NSI

[1969] A. Karasu (Kalkanli) and P. G. L. Leach. Nonlocal symmetries and integrable ordinary differential equations: ẍ + 3xẋ + x^{3} = 0 and its generalizations. Journal of Mathematical Physics, 50(7):073509, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i7/p073509_s1.

Carillo:2009:NKV

[1970] Sandra Carillo and Cornelia Schiebold. Noncommutative Korteweg–de Vries and modified Korteweg–de Vries hierarchies via recursion methods. Journal of Mathematical Physics, 50(7):073510, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i7/p073510_s1.

Izaurieta:2009:DFL

[1971] F. Izaurieta, A. Perez, E. Rodriguez, and P. Salgado. Dual formulation of the Lie algebra S-expansion procedure. Journal of Mathematical Physics, 50(7):073511, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i7/p073511_s1.

Hayn:2009:DSM

[1972] Michael Hayn and Matthias Holschneider. Directional spherical multipole wavelets. Journal of Mathematical Physics, 50(7):073512, July 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i7/p073512_s1.

Bohata:2009:MVB

[1973] Martin Bohata and Jan Hamhalter. Maximal violation of Bell’s inequalities and Pauli spin matrices. Journal of Mathematical Physics, 50(8): 082101, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i8/p082101_s1.

Friesecke:2009:ETQ

[1974] Gero Friesecke and Mario Koppen. On the Ehrenfest theorem of quantum mechanics. Journal of Mathematical Physics, 50(8):082102, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i8/p082102_s1.

Mkrtchyan:2009:EPD

[1975] A. R. Mkrtchyan, R. M. Avakyan, A. G. Hayrapetyan, B. V. Khachatryan, and R. G. Petrosyan. Equations for probability density and for the phase of wave function in quantum mechanics and superconductivity. Journal of Mathematical Physics, 50(8):082103, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/p082103_s1.

Karlsson:2009:TPC

[1976] Bengt R. Karlsson. Two-parameter complex Hadamard matrices for N = 6. Journal of Mathematical Physics, 50(8):082104, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/p082104_s1.

Duman:2009:FAE

[1977] Ali Nabi Duman. The fusion algebra of an extraspecial p-group. Journal of Mathematical Physics, 50(8):082301, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p082301_s1.

Aazami:2009:UMTb

[1978] A. B. Aazami and A. O. Petters. A universal magnification theorem. II. Generic caustics up to codimension five. Journal of Mathematical Physics, 50(8):082501, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i8/p082501_s1.

Kunduri:2009:CNH

[1979] Hari K. Kunduri and James Lucietti. A classification of near-horizon geometries of extremal vacuum black holes. Journal of Mathematical Physics, 50(8):082502, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i8/p082502_s1.

vandenHoogen:2009:CCS

[1980] R. J. van den Hoogen. A complete cosmological solution to the averaged Einstein field equations as found in macroscopic gravity. Journal of Mathematical Physics, 50(8):082503, August 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i8/p082503_s1.

Werner:2009:GUM

[1981] M. C. Werner. Geometry of universal magnification invariants. Journal of Mathematical Physics, 50(8):082504, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p082504_s1.

Karachalios:2009:RDA

[1982] Nikos I. Karachalios. A remark on the dimension of the attractor for the Dirichlet problem of the complex Ginzburg–Landau equation. Journal of Mathematical Physics, 50(8):082701, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p082701_s1.

Pandey:2009:GTIa

[1983] S. N. Pandey, P. S. Bindu, M. Senthilvelan, and M. Lakshmanan. A group theoretical identification of integrable cases of the Liénard-type equation ẍ + f(x)ẋ + g(x) = 0. I. Equations having nonmaximal number of Lie point symmetries. Journal of Mathematical Physics, 50(8): 082702, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i8/p082702_s1.

Koop:2009:CBC

[1984] M. J. Koop, R. B. Mann, and M. Rohanizadegan. Chaotic behavior in a charged three-body self-gravitating system. Journal of Mathematical Physics, 50(8):082703, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i8/p082703_s1.

Ren:2009:NNS

[1985] Yong Ren and Li Chen. A note on the neutral stochastic functional differential equation with infinite delay and Poisson jumps in an abstract space. Journal of Mathematical Physics, 50(8):082704, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i8/p082704_s1.

Gasull:2009:ODQ

[1986] Armengol Gasull, Jaume Llibre, and Xiang Zhang. One-dimensional quaternion homogeneous polynomial differential equations. Journal of Mathematical Physics, 50(8):082705, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p082705_s1.

Grillo:2009:HOC

[1987] Sergio D. Grillo. Higher order constrained Hamiltonian systems. Journal of Mathematical Physics, 50(8):082901, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p082901_s1.

Yeung:2009:ASN

[1988] Ling Hei Yeung and Yuen Manwai. Analytical solutions to the Navier– Stokes equations with density-dependent viscosity and with pressure. Journal of Mathematical Physics, 50(8):083101, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/p083101_s1.

Eyink:2009:SLM

[1989] Gregory L. Eyink. Stochastic line motion and stochastic flux conservation for nonideal hydromagnetic models. Journal of Mathematical Physics, 50(8):083102, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i8/p083102_s1.

Fa:2009:ADG

[1990] Kwok Sau Fa. Anomalous diffusion in a generalized Langevin equation. Journal of Mathematical Physics, 50(8):083301, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/p083301_s1.

Aliev:2009:GFS

[1991] M. A. Aliev. Generating function of spin correlation functions for kinetic Glauber–Ising model with time-dependent transition rates. Journal of Mathematical Physics, 50(8):083302, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p083302_s1.

Hubener:2009:CSS

[1992] R. Hübener, M. Van den Nest, W. Dür, and H. J. Briegel. Classical spin systems and the quantum stabilizer formalism: General mappings and applications. Journal of Mathematical Physics, 50(8):083303, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i8/p083303_s1.

Malbouisson:2009:BEM

[1993] A. P. C. Malbouisson, J. M. C. Malbouisson, and R. C. Pereira. Boundary effects on the mass and coupling constant in the compactified Ginzburg– Landau model: The boundary dependent critical temperature. Journal of Mathematical Physics, 50(8):083304, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p083304_s1.

Nikolayenko:2009:MTR

[1994] S. O. Nikolayenko and Yu. V. Slyusarenko. Microscopic theory of relaxation processes in systems of particles interacting with the hydrodynamic medium. Journal of Mathematical Physics, 50(8):083305, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i8/p083305_s1.

Ataguema:2009:GAN

[1995] H. Ataguema, A. Makhlouf, and S. Silvestrov. Generalization of n-ary Nambu algebras and beyond. Journal of Mathematical Physics, 50(8): 083501, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i8/p083501_s1.

Ye:2009:BVS

[1996] Deping Ye. On the Bures volume of separable quantum states. Journal of Mathematical Physics, 50(8):083502, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p083502_s1.

Hu:2009:BBR

[1997] Qiaoyi Hu and Zhaoyang Yin. Blowup and blowup rate of solutions to a weakly dissipative periodic rod equation. Journal of Mathematical Physics, 50(8):083503, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i8/p083503_s1.

Han:2009:LBS

[1998] Jianzhi Han, Junbo Li, and Yucai Su. Lie bialgebra structures on the Schrödinger–Virasoro Lie algebra. Journal of Mathematical Physics, 50(8):083504, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i8/p083504_s1.

Park:2009:GDQ

[1999] Jong Yeoul Park and Sun Hye Park. General decay for quasilinear viscoelastic equations with nonlinear weak damping. Journal of Mathematical Physics, 50(8):083505, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i8/p083505_s1.

Liu:2009:CTA

[2000] I-Shih Liu. Constitutive theory of anisotropic rigid heat conductors. Journal of Mathematical Physics, 50(8):083506, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/p083506_s1.

Fredsted:2009:SFL

[2001] John Fredsted. Spinor fields without Lorentz frames in curved space–time using complexified quaternions. Journal of Mathematical Physics, 50(8): 083507, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i8/p083507_s1.

Demenchuk:2009:EPF

[2002] Alexandr K. Demenchuk and Evgenii K. Makarov. Explicit polynomial formulas for solutions of the matrix equation AX − XA = C. Journal of Mathematical Physics, 50(8):083508, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p083508_s1.

Hu:2009:BPE

[2003] Taotao Hu, Kang Xue, and Chunfeng Wu. Berry phase and entanglement of three qubits in a new Yang–Baxter system. Journal of Mathematical Physics, 50(8):083509, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i8/p083509_s1.

Lyakhovich:2009:NFG

[2004] S. L. Lyakhovich and A. A. Sharapov. Normal forms and gauge symmetry of local dynamics. Journal of Mathematical Physics, 50(8):083510, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i8/p083510_s1.

Lopes:2009:GWP

[2005] C. E. F. Lopes, I. A. Pedrosa, C. Furtado, and A. M. de M. Carvalho. Gaussian wave packet states of scalar fields in a universe of de Sitter. Journal of Mathematical Physics, 50(8):083511, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/p083511_s1.

Penson:2009:LTD

[2006] K. A. Penson, P. Blasiak, A. Horzela, G. H. E. Duchamp, and A. I. Solomon. Laguerre-type derivatives: Dobiński relations and combinatorial identities. Journal of Mathematical Physics, 50(8):083512, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i8/p083512_s1.

Weihua:2009:SBQ

[2007] Liu Weihua and Wu Junde. On supremum of bounded quantum observable. Journal of Mathematical Physics, 50(8):083513, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/
p083513_s1.

Chen:2009:GES

[2008] Guowang Chen, Hongxia Guo, and Hongwei Zhang. Global existence of solutions of Cauchy problem for generalized system of nonlinear evolution equations arising from DNA. Journal of Mathematical Physics, 50(8): 083514, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i8/p083514_s1.

Vertman:2009:ZDR

[2009] Boris Vertman. Zeta determinants for regular-singular Laplace-type operators. Journal of Mathematical Physics, 50(8):083515, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/
p083515_s1.

He:2009:SSS

[2010] Chun-Lei He and De-Xing Kong. Spherical symmetric solutions for the motion of relativistic membranes in the Schwarzschild spacetime. Journal of Mathematical Physics, 50(8):083516, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i8/p083516_s1.

Prodan:2009:ESC

[2011] Emil Prodan. The edge spectrum of Chern insulators with rough boundaries. Journal of Mathematical Physics, 50(8):083517, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i8/
p083517_s1.

Yang:2009:PFE

[2012] Wen-Li Yang and Yao-Zhong Zhang. Partition function of the eightvertex model with domain wall boundary condition. Journal of Mathematical Physics, 50(8):083518, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i8/p083518_s1.

Sergeev:2009:SS

[2013] Sergey M. Sergeev. Supertetrahedra and superalgebras. Journal of Mathematical Physics, 50(8):083519, August 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i8/p083519_s1.

Its:2009:ISI

[2014] Alexander Its, Michio Jimbo, and Jean-Michel Maillet. Introduction to special issue: Integrable Quantum Systems and Solvable Statistical Mechanics Models. Journal of Mathematical Physics, 50(9):095101, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i9/p095101_s1.

Fischer:2009:NSG

[2015] André Fischer and Olaf Lechtenfeld. The noncommutative sine-Gordon breather. Journal of Mathematical Physics, 50(9):095201, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i9/p095201_s1.

Marquette:2009:STOb

[2016] Ian Marquette. Superintegrability with third order integrals of motion, cubic algebras, and supersymmetric quantum mechanics. II. Painlevé transcendent potentials. Journal of Mathematical Physics, 50(9):095202, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i9/p095202_s1.

Arguin:2009:RPF

[2017] Louis-Pierre Arguin and Yvan Saint-Aubin. Restricted partition functions of the two-dimensional Ising model on a half-infinite cylinder. Journal of Mathematical Physics, 50(9):095203, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i9/p095203_s1.

Tracy:2009:TCF

[2018] Craig A. Tracy and Harold Widom. Total current fluctuations in the asymmetric simple exclusion process. Journal of Mathematical Physics, 50(9):095204, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i9/p095204_s1.

Kozlowski:2009:FSA

[2019] K. K. Kozlowski. Fine structure of the asymptotic expansion of cyclic integrals. Journal of Mathematical Physics, 50(9):095205, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i9/p095205_s1.

Boos:2009:CFB

[2020] H. Boos, M. Jimbo, T. Miwa, and F. Smirnov. Completeness of a fermionic basis in the homogeneous XXZ model. Journal of Mathematical Physics, 50(9):095206, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i9/p095206_s1.

Hastings:2009:LCM

[2021] M. B. Hastings. Light-cone matrix product. Journal of Mathematical Physics, 50(9):095207, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i9/p095207_s1.

Starr:2009:TLM

[2022] Shannon Starr. Thermodynamic limit for the Mallows model on S_{n}. Journal of Mathematical Physics, 50(9):095208, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i9/p095208_s1.

Kitanine:2009:TLF

[2023] N. Kitanine, K. K. Kozlowski, J. M. Maillet, N. A. Slavnov, and V. Terras. On the thermodynamic limit of form factors in the massless XXZ Heisenberg chain. Journal of Mathematical Physics, 50(9):095209, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i9/p095209_s1.

Mathieu:2009:PPC

[2024] Pierre Mathieu. Paths and partitions: Combinatorial descriptions of the parafermionic states. Journal of Mathematical Physics, 50(9):095210, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i9/p095210_s1.

Santos:2009:TCO

[2025] Lea F. Santos. Transport and control in one-dimensional systems. Journal of Mathematical Physics, 50(9):095211, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v50/i9/p095211_s1.

Faribault:2009:BAA

[2026] Alexandre Faribault, Pasquale Calabrese, and Jean-Sébastien Caux. Bethe ansatz approach to quench dynamics in the Richardson model. Journal of Mathematical Physics, 50(9):095212, September 2009. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i9/
p095212_s1.

Hamza:2009:AGS

[2027] Eman Hamza, Spyridon Michalakis, Bruno Nachtergaele, and Robert Sims. Approximating the ground state of gapped quantum spin systems. Journal of Mathematical Physics, 50(9):095213, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i9/
p095213_s1.

Caux:2009:CFI

[2028] Jean-Sébastien Caux. Correlation functions of integrable models: A description of the ABACUS algorithm. Journal of Mathematical Physics, 50(9):095214, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i9/p095214_s1.

Feigin:2009:CAD

[2029] B. Feigin, K. Hashizume, A. Hoshino, J. Shiraishi, and S. Yanagida. A commutative algebra on degenerate CP^{1} and Macdonald polynomials. Journal of Mathematical Physics, 50(9):095215, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i9/
p095215_s1.

Elci:2009:TST

[2030] A. Elçi and H. P. Hjalmarson. The tunneling solutions of the timedependent Schrödinger equation for a square-potential barrier. Journal of Mathematical Physics, 50(10):102101, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p102101_s1.

Bach:2009:BHF

[2031] Volker Bach, Jürg Fröhlich, and Lars Jonsson. Bogolubov–Hartree–Fock mean field theory for neutron stars and other systems with attractive interactions. Journal of Mathematical Physics, 50(10):102102, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p102102_s1.

Valluri:2009:LFQ

[2032] S. R. Valluri, M. Gil, D. J. Jeffrey, and Shantanu Basu. The Lambert W function and quantum statistics. Journal of Mathematical Physics, 50(10):102103, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i10/p102103_s1.

Liu:2009:PST

[2033] Zuhan Liu. Phase separation of two-component Bose–Einstein condensates. Journal of Mathematical Physics, 50(10):102104, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p102104_s1.

Gour:2009:TRF

[2034] Gilad Gour, Barry C. Sanders, and Peter S. Turner. Time-reversal frameness and superselection. Journal of Mathematical Physics, 50(10): 102105, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p102105_s1.

Lidar:2009:AAE

[2035] Daniel A. Lidar, Ali T. Rezakhani, and Alioscia Hamma. Adiabatic approximation with exponential accuracy for many-body systems and quantum computation. Journal of Mathematical Physics, 50(10):102106, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p102106_s1.

Miyadera:2009:NCT

[2036] Takayuki Miyadera and Hideki Imai. No-cloning theorem on quantum logics. Journal of Mathematical Physics, 50(10):102107, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p102107_s1.

Fan:2009:EPS

[2037] Hong yi Fan and Cui hong Lv. Explicit pure-state density operator structure for quantum tomography. Journal of Mathematical Physics, 50(10): 102108, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p102108_s1. See [2204].

Fucci:2009:NOL

[2038] Guglielmo Fucci. Nonperturbative one-loop effective action for electrodynamics in curved space–time. Journal of Mathematical Physics, 50(10): 102301, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p102301_s1.

Avramidi:2009:LEE

[2039] Ivan G. Avramidi and Guglielmo Fucci. Low-energy effective action in nonperturbative electrodynamics in curved space–time. Journal of Mathematical Physics, 50(10):102302, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p102302_s1.

Adam:2009:CSS

[2040] C. Adam, P. Klimas, J. Sánchez-Guillén, and A. Wereszczyński. Compact shell solitons in K field theories. Journal of Mathematical Physics, 50(10):102303, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i10/p102303_s1.

Kim:2009:SFT

[2041] Hyeong-Chan Kim, Youngone Lee, Chaiho Rim, and Jae Hyung Yee. Scalar field theory in κ-Minkowski spacetime from twist. Journal of Mathematical Physics, 50(10):102304, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p102304_s1.

Burnett:2009:WLT

[2042] James Burnett and Dmitri Vassiliev. Weyl’s Lagrangian in teleparallel form. Journal of Mathematical Physics, 50(10):102501, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p102501_s1.

Choudhury:2009:DEF

[2043] A. Ghose Choudhury, Partha Guha, and Barun Khanra. Determination of elementary first integrals of a generalized Raychaudhuri equation by the Darboux integrability method. Journal of Mathematical Physics, 50(10):102502, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i10/p102502_s1.

Pandey:2009:GTIb

[2044] S. N. Pandey, P. S. Bindu, M. Senthilvelan, and M. Lakshmanan. A group theoretical identification of integrable equations in the Liénardtype equation ẍ + f(x)˙(x) + g(x) = 0. II. Equations having maximal Lie point symmetries. Journal of Mathematical Physics, 50(10):102701, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p102701_s1.

Kundu:2009:NLE

[2045] Anjan Kundu. Nonlinearizing linear equations to integrable systems including new hierarchies with nonholonomic deformations. Journal of Mathematical Physics, 50(10):102702, October 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i10/p102702_s1. See

publisher’s note [2205].

Cerkaski:2009:EM

[2046] M. Cerkaski. Ellipsoids of U(3) model. Journal of Mathematical Physics, 50(10):102703, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i10/p102703_s1.

Gershon:2009:KKT

[2047] Avi Gershon and Lawrence Horwitz. Kaluza–Klein theory as a dynamics in a dual geometry. Journal of Mathematical Physics, 50(10):102704, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p102704_s1.

Llibre:2009:DTI

[2048] J. Llibre and Ch. Pantazi. Darboux theory of integrability for a class of nonautonomous vector fields. Journal of Mathematical Physics, 50(10): 102705, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p102705_s1.

Gong:2009:ANT

[2049] Dongshan Gong, Haitao Song, and Chengkui Zhong. Attractors for nonautonomous two-dimensional space periodic Navier–Stokes equations. Journal of Mathematical Physics, 50(10):102706, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p102706_s1.

Zhu:2009:ASS

[2050] Peicheng Zhu. Asymptotic stability of the stationary solution to an initial boundary value problem for the Mullins equation of fourth order. Journal of Mathematical Physics, 50(10):102707, October 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i10/p102707_s1.

Li:2009:MRR

[2051] Xicheng Li and Mingyu Xu. A model for reversible reaction in a subdiffusive regime. Journal of Mathematical Physics, 50(10):102708, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p102708_s1.

Ren:2009:SON

[2052] Yong Ren and Dandan Sun. Second-order neutral impulsive stochastic evolution equations with delay. Journal of Mathematical Physics, 50(10): 102709, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p102709_s1.

Habibullin:2009:CLD

[2053] Ismagil Habibullin, Natalya Zheltukhina, and Asli Pekcan. Complete list of Darboux integrable chains of the form t_{1x} = t+x+d(t, t_{1}). Journal of Mathematical Physics, 50(10):102710, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p102710_s1.

Woodlin:2009:CCC

[2054] Mervin Woodlin and Zhifu Xie. Collinear central configurations in the n-body problem with general homogeneous potential. Journal of Mathematical Physics, 50(10):102901, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i10/p102901_s1.

Hone:2009:NPS

[2055] A. N. W. Hone and M. Senthilvelan. Note on the Poisson structure of the damped oscillator. Journal of Mathematical Physics, 50(10):102902, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p102902_s1.

Romano:2009:CD

[2056] Giovanni Romano, Raffaele Barretta, and Marina Diaco. On continuum dynamics. Journal of Mathematical Physics, 50(10):102903, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p102903_s1.

Frasca:2009:DLS

[2057] Marco Frasca. Dual Lindstedt series and Kolmogorov–Arnol’d–Moser theorem. Journal of Mathematical Physics, 50(10):102904, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p102904_s1.

Nunez:2009:BRE

[2058] Manuel Núñez. Basic results on the equations of magnetohydrodynamics of partially ionized inviscid plasmas. Journal of Mathematical Physics, 50(10):103101, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i10/p103101_s1.

Aulisa:2009:AGF

[2059] Eugenio Aulisa, Lidia Bloshanskaya, Luan Hoang, and Akif Ibragimov. Analysis of generalized Forchheimer flows of compressible fluids in porous media. Journal of Mathematical Physics, 50(10):103102, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103102_s1.

Oikonomou:2009:NDD

[2060] Thomas Oikonomou and G. Baris Bagci. A note on the definition of deformed exponential and logarithm functions. Journal of Mathematical Physics, 50(10):103301, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i10/p103301_s1.

Halim:2009:CCM

[2061] Zeghdoudi Halim and Boutabia Hacène. Computation for the canonical measures of a colored disordered lattice gas and spectral gap. Journal of Mathematical Physics, 50(10):103302, October 2009. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p103302_s1.

Tian:2009:NSB

[2062] Qianzhu Tian and Jie Sun. Nonlinear stability of boundary layer solution to the Boltzmann equation with diffusive effect at the boundary. Journal of Mathematical Physics, 50(10):103303, October 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i10/p103303_s1.

Hiai:2009:QHT

[2063] Fumio Hiai, Milán Mosonyi, and Masahito Hayashi. Quantum hypothesis testing with group symmetry. Journal of Mathematical Physics, 50(10): 103304, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p103304_s1.

Mano:2009:SMP

[2064] Toshiyuki Mano. Studies on monodromy preserving deformation of linear differential equations on elliptic curves. Journal of Mathematical Physics, 50(10):103501, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i10/p103501_s1.

Jorgensen:2009:EWI

[2065] Palle E. T. Jorgensen and Myung-Sin Song. An extension of Wiener integration with the use of operator theory. Journal of Mathematical Physics, 50(10):103502, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i10/p103502_s1.

Keller:2009:EEG

[2066] Kai J. Keller. Euclidean Epstein–Glaser renormalization. Journal of Mathematical Physics, 50(10):103503, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p103503_s1.

Bruzon:2009:STW

[2067] M. S. Bruzon, M. L. Gandarias, M. Torrisi, and R. Tracinà. Some traveling wave solutions for the dissipative Zabolotskaya–Khokhlov equation. Journal of Mathematical Physics, 50(10):103504, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103504_s1.

Zhang:2009:NSM

[2068] Xi Zhang. A note of Sasakian metrics with constant scalar curvature. Journal of Mathematical Physics, 50(10):103505, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103505_s1.

Wang:2009:GEC

[2069] Yu-Zhu Wang. Global existence of classical solutions to the minimal surface equation in two space dimensions with slow decay initial value. Journal of Mathematical Physics, 50(10):103506, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103506_s1.

Ding:2009:EBS

[2070] Qing Ding. Explicit blow-up solutions to the Schrödinger maps from R^{2} to the hyperbolic 2-space H^{2}. Journal of Mathematical Physics, 50(10): 103507, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p103507_s1.

Iohara:2009:NSM

[2071] Kenji Iohara and Yoshiyuki Koga. Note on spin modules associated to Z-graded Lie superalgebras. Journal of Mathematical Physics, 50(10): 103508, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p103508_s1.

Sasaki:2009:ESB

[2072] Ryu Sasaki. Exactly solvable birth and death processes. Journal of Mathematical Physics, 50(10):103509, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p103509_s1.

Ardavan:2009:ICT

[2073] Houshang Ardavan, Arzhang Ardavan, John Singleton, Joseph Fasel, and Andrea Schmidt. Inadequacies in the conventional treatment of the radiation field of moving sources. Journal of Mathematical Physics, 50(10):

103510, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p103510_s1.

Hayashi:2009:FSP

[2074] Nakao Hayashi and Pavel I. Naumkin. Final state problem for the cubic nonlinear Klein–Gordon equation. Journal of Mathematical Physics, 50(10):103511, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i10/p103511_s1.

Matache:2009:SNB

[2075] Mihaela T. Matache and Valentin Matache. On the sensitivity to noise of a Boolean function. Journal of Mathematical Physics, 50(10):103512, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p103512_s1.

Martin:2009:SGL

[2076] Antonio J. Calderón Martín. On the structure of graded Lie algebras. Journal of Mathematical Physics, 50(10):103513, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103513_s1.

Vignat:2009:SOP

[2077] C. Vignat and P. W. Lamberti. A study of the orthogonal polynomials associated with the quantum harmonic oscillator on constant curvature spaces. Journal of Mathematical Physics, 50(10):103514, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103514_s1.

Carinena:2009:QLS

[2078] José F. Cariñena, P. G. L. Leach, and Javier de Lucas. Quasi-Lie schemes and Emden–Fowler equations. Journal of Mathematical Physics, 50(10): 103515, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p103515_s1.

Hay:2009:CSD

[2079] Mike Hay. A completeness study on discrete, 2 × 2 Lax pairs. Journal of Mathematical Physics, 50(10):103516, October 2009. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p103516_s1.

DiCriscienzo:2009:FDH

[2080] Roberto Di Criscienzo and Sergio Zerbini. Functional determinants in higher derivative Lagrangian theories. Journal of Mathematical Physics, 50(10):103517, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i10/p103517_s1. See comment [2201] and er-

ratum [2203].

Melikyan:2009:QIL

[2081] A. Melikyan and A. Pinzul. On quantum integrability of the Landau– Lifshitz model. Journal of Mathematical Physics, 50(10):103518, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i10/p103518_s1.

Brezhnev:2009:UBC

[2082] Yu. V. Brezhnev. On uniformization of Burnside’s curve y^{2} = x^{5} − x. Journal of Mathematical Physics, 50(10):103519, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103519_s1.

Cardanobile:2009:TGT

[2083] Stefano Cardanobile and Delio Mugnolo. Toward a gauge theory for evolution equations on vector-valued spaces. Journal of Mathematical Physics, 50(10):103520, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i10/p103520_s1.

Egorova:2009:IST

[2084] Iryna Egorova, Johanna Michor, and Gerald Teschl. Inverse scattering transform for the Toda hierarchy with steplike finite-gap backgrounds. Journal of Mathematical Physics, 50(10):103521, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103521_s1.

Qu:2009:NEE

[2085] Changzheng Qu and Shoufeng Shen. Nonlinear evolution equations admitting multilinear variable separable solutions. Journal of Mathematical Physics, 50(10):103522, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i10/p103522_s1.

Skrypnyk:2009:ISG

[2086] T. Skrypnyk. Integrability and superintegrability of the generalized n-level many-mode Jaynes–Cummings and Dicke models. Journal of Mathematical Physics, 50(10):103523, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p103523_s1.

Hatzinikitas:2009:FSO

[2087] Agapitos Hatzinikitas. The fractional Schrödinger operator and Toeplitz matrices. Journal of Mathematical Physics, 50(10):103524, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/
p103524_s1.

Dong:2009:RCW

[2088] Bo-Qing Dong and Zhi-Min Chen. Regularity criteria of weak solutions to the three-dimensional micropolar flows. Journal of Mathematical Physics, 50(10):103525, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i10/p103525_s1.

Ardenghi:2009:NLC

[2089] Juan S. Ardenghi, M. Castagnino, and R. Campoamor-Stursberg. The nonrelativistic limit of (central-extended) Poincaré group and some consequences for quantum actualization. Journal of Mathematical Physics, 50(10):103526, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i10/p103526_s1.

Ciccariello:2009:CFP

[2090] Salvino Ciccariello. The correlation functions of plane polygons. Journal of Mathematical Physics, 50(10):103527, October 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i10/p103527_s1.

Amdeberhan:2009:CFE

[2091] Tewodros Amdeberhan, Victor H. Moll, and Armin Straub. Closedform evaluation of integrals appearing in positronium decay. Journal of Mathematical Physics, 50(10):103528, October 2009. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p103528_s1.

Barros:2009:CVA

[2092] Manuel Barros and Angel Ferrández. A conformal variational approach for helices in nature. Journal of Mathematical Physics, 50(10):103529, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p103529_s1.

Miyamoto:2009:HSC

[2093] Yasuhito Miyamoto. The “hot spots” conjecture for a certain class of planar convex domains. Journal of Mathematical Physics, 50(10): 103530, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i10/p103530_s1.

Weihua:2009:FPL

[2094] Liu Weihua and Wu Junde. On fixed points of Lüders operation. Journal of Mathematical Physics, 50(10):103531, October 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i10/p103531_s1.

Kamvissis:2009:CER

[2095] Spyridon Kamvissis. Comment on “Existence and regularity for an energy maximization problem in two dimensions”, [S. Kamvissis and E. A. Rakhmanov, J. Math. Phys. 46, 083505 (2005)]. Journal of Mathematical Physics, 50(10):104101, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i10/p104101_s1. See [330].

Gunara:2009:PNF

[2096] Bobby Eka Gunara and Freddy Permana Zen. Publisher’s note: “Flat Bogomolnyi–Prasad–Sommerfeld domain walls on two-dimensional Kähler–Ricci soliton”, [J. Math. Phys. 50, 063514 (2009)]. Journal of Mathematical Physics, 50(10):109901, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i10/p109901_s1. See [1937].

Tyson:2009:EME

[2097] Jon Tyson. Erratum: “Minimum-error quantum distinguishability bounds from matrix monotone functions: A comment on ‘Two-sided estimates of minimum-error distinguishability of mixed quantum states via generalized Holevo–Curlander bounds’ ”, [J. Math. Phys. 50, 062102 (2009)]. Journal of Mathematical Physics, 50(10):109902, October 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i10/

p109902_s1. See [1907, 1796].

Exner:2009:GPC

[2098] Pavel Exner and Martin Fraas. On geometric perturbations of critical Schrödinger operators with a surface interaction. Journal of Mathematical Physics, 50(11):112101, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i11/p112101_s1.

Benes:2009:DPL

[2099] Thomas Benes and Dietrich Burde. Degenerations of pre-Lie algebras. Journal of Mathematical Physics, 50(11):112102, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i11/
p112102_s1.

Aleixo:2009:MDT

[2100] A. N. F. Aleixo and A. B. Balantekin. Multiparameter deformation theory for quantum confined systems. Journal of Mathematical Physics, 50(11):112103, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i11/p112103_s1.

Robinson:2009:LSQ

[2101] Dean J. Robinson, Eliot Kapit, and André LeClair. Lorentz symmetric quantum field theory for symplectic fermions. Journal of Mathematical Physics, 50(11):112301, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i11/p112301_s1.

Gouesbet:2009:OTN

[2102] G. Gouesbet. On the optical theorem and non-plane-wave scattering in quantum mechanics. Journal of Mathematical Physics, 50(11):112302, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p112302_s1.

Sarisaman:2009:PBS

[2103] Mustafa Sarisaman. Pseudoduality between symmetric space sigma models. Journal of Mathematical Physics, 50(11):112303, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i11/
p112303_s1.

Hollands:2009:PQF

[2104] Stefan Hollands and Heiner Olbermann. Perturbative quantum field theory via vertex algebras. Journal of Mathematical Physics, 50(11):112304, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p112304_s1.

Vogeler:2009:LDS

[2105] Kirsten Vogeler and Michael Flohr. On a logarithmic deformation of the supersymmetric bc-system on curved manifolds. Journal of Mathematical Physics, 50(11):112305, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i11/p112305_s1.

Lubbe:2009:ETC

[2106] Christian Lübbe and Paul Tod. An extension theorem for conformal gauge singularities. Journal of Mathematical Physics, 50(11):112501, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p112501_s1.

Lubbe:2009:CET

[2107] Christian Lübbe. A conformal extension theorem based on null conformal geodesics. Journal of Mathematical Physics, 50(11):112502, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i11/p112502_s1.

Mota:2009:QSR

[2108] Herondy F. S. Mota and V. B. Bezerra. Quantum scattering of relativistic particles in Safko–Witten spacetime. Journal of Mathematical Physics, 50(11):112503, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i11/p112503_s1.

Barrett:2009:AAE

[2109] John W. Barrett, Richard J. Dowdall, Winston J. Fairbairn, Henrique Gomes, and Frank Hellmann. Asymptotic analysis of the Engle–Pereira– Rovelli–Livine four-simplex amplitude. Journal of Mathematical Physics, 50(11):112504, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i11/p112504_s1.

Park:2009:GES

[2110] Jong Yeoul Park and Jum Ran Kang. Global existence and stability for a von Karman equations with memory in noncylindrical domains. Journal of Mathematical Physics, 50(11):112701, November 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i11/p112701_s1.

Zhao:2009:NSH

[2111] Fukun Zhao, Leiga Zhao, and Yanheng Ding. A note on superlinear Hamiltonian elliptic systems. Journal of Mathematical Physics, 50(11):112702, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i11/p112702_s1.

Gurses:2009:DSP

[2112] Metin Gürses, Gusein Sh. Guseinov, and Kostyantyn Zheltukhin. Dynamical systems and Poisson structures. Journal of Mathematical Physics, 50(11):112703, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i11/p112703_s1. See comment [2209].

Castilho:2009:PUM

[2113] César Castilho and Marcelo Marchesin. A practical use of the Melnikov homoclinic method. Journal of Mathematical Physics, 50(11):112704, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p112704_s1.

Cuevas:2009:LUE

[2114] J. Cuevas, N. I. Karachalios, and F. Palmero. Lower and upper estimates on the excitation threshold for breathers in discrete nonlinear Schrödinger lattices. Journal of Mathematical Physics, 50(11):112705, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p112705_s1.

Yaremko:2009:IRT

[2115] Yurij Yaremko. Interference in the radiation of two pointlike charges. Journal of Mathematical Physics, 50(11):112901, November 2009. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i11/
p112901_s1.

Lupinski:2009:ANM

[2116] L. W. Lupinski and M. J. Madsen. Angular normal modes of a circular Coulomb cluster. Journal of Mathematical Physics, 50(11):112902, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p112902_s1.

Laugesen:2009:MNF

[2117] R. S. Laugesen and B. A. Siudeja. Maximizing Neumann fundamental tones of triangles. Journal of Mathematical Physics, 50(11):112903, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p112903_s1.

Yoshida:2009:CPB

[2118] Z. Yoshida. Clebsch parameterization: Basic properties and remarks on its applications. Journal of Mathematical Physics, 50(11):113101, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p113101_s1.

Wada:2009:GLL

[2119] Tatsuaki Wada. Generalized log-likelihood functions and Bregman divergences. Journal of Mathematical Physics, 50(11):113301, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i11/p113301_s1. See erratum [2199].

Jun:2009:SRI

[2120] Shen Jun and Wu Junde. Spectral representation of infimum of bounded quantum observables. Journal of Mathematical Physics, 50(11):113501, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p113501_s1.

Guo:2009:IHO

[2121] Jia-Feng Guo, Shi-Kun Wang, Ke Wu, Zhao-Wen Yan, and Wei-Zhong Zhao. Integrable higher order deformations of Heisenberg supermagnetic model. Journal of Mathematical Physics, 50(11):113502, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i11/
p113502_s1.

Jorgensen:2009:AUO

[2122] Palle E. T. Jorgensen. Analysis of unbounded operators and random motion. Journal of Mathematical Physics, 50(11):113503, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i11/
p113503_s1.

Garcia-Compean:2009:MDN

[2123] Hugo García-Compeán and Aldo A. Martínez-Merino. On the Moyal deformation of Nahm equations in seven dimensions. Journal of Mathematical Physics, 50(11):113504, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i11/p113504_s1.

Bilson-Thompson:2009:PTB

[2124] Sundance Bilson-Thompson, Jonathan Hackett, and Louis H. Kauffman. Particle topology, braids, and braided belts. Journal of Mathematical Physics, 50(11):113505, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i11/p113505_s1.

Liu:2009:GDR

[2125] Wenjun Liu. General decay rate estimate for a viscoelastic equation with weakly nonlinear time-dependent dissipation and source terms. Journal of Mathematical Physics, 50(11):113506, November 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i11/p113506_s1.

Su:2009:AGD

[2126] Ting Su, Hui-Hui Dai, and Xianguo Geng. On the application of a generalized dressing method to the integration of variable-coefficient coupled Hirota equations. Journal of Mathematical Physics, 50(11):113507, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p113507_s1.

Nucci:2009:MOQ

[2127] M. C. Nucci and P. G. L. Leach. The method of Ostrogradsky, quantization, and a move toward a ghost-free future. Journal of Mathematical Physics, 50(11):113508, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i11/p113508_s1.

Majewski:2009:CPP

[2128] Wladyslaw A. Majewski, Takashi Matsuoka, and Masanori Ohya. Characterization of partial positive transposition states and measures of entanglement. Journal of Mathematical Physics, 50(11):113509, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i11/p113509_s1.

Ji:2009:PSS

[2129] Shuguan Ji. Periodic solutions on the Sturm–Liouville boundary value problem for two-dimensional wave equation. Journal of Mathematical Physics, 50(11):113510, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i11/p113510_s1.

Ohkubo:2009:GSM

[2130] Atsushi Ohkubo. Ground states of the massless Dereziński–Gérard model. Journal of Mathematical Physics, 50(11):113511, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i11/
p113511_s1.

Avan:2009:BLP

[2131] Jean Avan and Anastasia Doikou. Boundary Lax pairs from non-ultra-local Poisson algebras. Journal of Mathematical Physics, 50(11):113512, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p113512_s1.

Laskin:2009:SAF

[2132] Nick Laskin. Some applications of the fractional Poisson probability distribution. Journal of Mathematical Physics, 50(11):113513, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i11/p113513_s1.

Figueroa-OFarrill:2009:DA

[2133] José Miguel Figueroa-O’Farrill. Deformations of 3-algebras. Journal of Mathematical Physics, 50(11):113514, November 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i11/p113514_s1.

Winfield:2009:WPE

[2134] Christopher J. Winfield. Well-posedness of Einstein’s equation with redshift data. Journal of Mathematical Physics, 50(11):113515, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i11/p113515_s1.

Collins:2009:SPO

[2135] Benoît Collins and Sho Matsumoto. On some properties of orthogonal Weingarten functions. Journal of Mathematical Physics, 50(11):113516, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p113516_s1.

Petz:2009:MPG

[2136] Dénes Petz and József Pitrik. Markov property of Gaussian states of canonical commutation relation algebras. Journal of Mathematical Physics, 50(11):113517, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i11/p113517_s1.

Salvarani:2009:ABL

[2137] F. Salvarani. Asymptotic behavior for logarithmic diffusion. Journal of Mathematical Physics, 50(11):113518, November 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i11/p113518_s1.

Chung:2009:HUS

[2138] Jae-Young Chung, Dohan Kim, and John Michael Rassias. Hyers–Ulam stability on a generalized quadratic functional equation in distributions and hyperfunctions. Journal of Mathematical Physics, 50(11):113519, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p113519_s1.

Kanna:2009:PSS

[2139] T. Kanna, K. Sakkaravarthi, C. Senthil Kumar, M. Lakshmanan, and M. Wadati. Painlevé singularity structure analysis of three component Gross–Pitaevskii type equations. Journal of Mathematical Physics, 50(11):113520, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i11/p113520_s1.

Bakke:2009:GPS

[2140] Knut Bakke, I. A. Pedrosa, and Claudio Furtado. Geometric phases and squeezed quantum states of relic gravitons. Journal of Mathematical Physics, 50(11):113521, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i11/p113521_s1.

Dey:2009:EGP

[2141] Rukmini Dey. Erratum: “Geometric prequantization of the moduli space of the vortex equations on a Riemann surface”, [J. Math. Phys. 47, 103501 (2006)]. Journal of Mathematical Physics, 50(11):119901, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i11/p119901_s1. See [784].

Nachtergaele:2009:AJM

[2142] Bruno L. Z. Nachtergaele. Announcement: Journal of Mathematical Physics introduces a new section — Quantum Information and Computation. Journal of Mathematical Physics, 50(12):120201, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p120201_s1.

Pellegrini:2009:DAS

[2143] C. Pellegrini and F. Petruccione. Diffusion approximation of stochastic master equations with jumps. Journal of Mathematical Physics, 50(12):122101, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p122101_s1.

Marquette:2009:SMO

[2144] Ian Marquette. Supersymmetry as a method of obtaining new superintegrable systems with higher order integrals of motion. Journal of Mathematical Physics, 50(12):122102, December 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i12/p122102_s1.

Accardi:2009:QEV

[2145] Luigi Accardi and Ameur Dhahri. Quadratic exponential vectors. Journal of Mathematical Physics, 50(12):122103, December 2009. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i12/
p122103_s1.

Hayashi:2009:GME

[2146] Masahito Hayashi, Damian Markham, Mio Murao, Masaki Owari, and Shashank Virmani. The geometric measure of entanglement for a symmetric pure state with non-negative amplitudes. Journal of Mathematical Physics, 50(12):122104, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i12/p122104_s1.

Znojil:2009:CSI

[2147] Miloslav Znojil. Complete set of inner products for a discrete P T - symmetric square-well Hamiltonian. Journal of Mathematical Physics, 50(12):122105, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p122105_s1.

Albertini:2009:AQW

[2148] Francesca Albertini and Domenico D’Alessandro. Analysis of quantum walks with time-varying coin on d-dimensional lattices. Journal of Mathematical Physics, 50(12):122106, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i12/p122106_s1.

Li:2009:NLB

[2149] Yuan Li and Xiu-Hong Sun. A note on the logic of bounded quantum observables. Journal of Mathematical Physics, 50(12):122107, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p122107_s1.

Ballesteros:2009:ABE

[2150] Miguel Ballesteros and Ricardo Weder. The Aharonov–Bohm effect and Tonomura et al. experiments: Rigorous results. Journal of Mathematical Physics, 50(12):122108, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i12/p122108_s1.

Thuillier:2009:DBC

[2151] F. Thuillier. Deligne–Beilinson cohomology and Abelian link invariants: Torsion case. Journal of Mathematical Physics, 50(12):122301, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p122301_s1.

Neznamov:2009:FWW

[2152] V. P. Neznamov and A. J. Silenko. Foldy–Wouthuysen wave functions and conditions of transformation between Dirac and Foldy–Wouthuysen representations. Journal of Mathematical Physics, 50(12):122302, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p122302_s1.

Petters:2009:MTSb

[2153] A. O. Petters, B. Rider, and A. M. Teguia. A mathematical theory of stochastic microlensing. II. Random images, shear, and the Kac–Rice formula. Journal of Mathematical Physics, 50(12):122501, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i12/
p122501_s1.

Miao:2009:TLG

[2154] S. P. Miao, N. C. Tsamis, and R. P. Woodard. Transforming to Lorentz gauge on de Sitter. Journal of Mathematical Physics, 50(12):122502, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p122502_s1.

Agnew:2009:MPS

[2155] Alfonso F. Agnew and Scot P. Childress. Matrix projective spaces and twistorlike incidence structures. Journal of Mathematical Physics, 50(12):122503, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p122503_s1.

Gomez-Lobo:2009:SCF

[2156] Alfonso García-Parrado Gómez-Lobo and José M. Martín-García. Spinor calculus on five-dimensional spacetimes. Journal of Mathematical Physics, 50(12):122504, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i12/p122504_s1.

Gravanis:2009:SSG

[2157] E. Gravanis and S. Willison. Singular sources in gravity and homotopy in the space of connections. Journal of Mathematical Physics, 50(12):122505, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p122505_s1.

Cook:2009:TGA

[2158] Samuel A. Cook and Tevian Dray. Tensor generalizations of affine symmetry vectors. Journal of Mathematical Physics, 50(12):122506, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p122506_s1.

Bruschi:2009:IAI

[2159] M. Bruschi, F. Calogero, and R. Droghei. Integrability, analyticity, isochrony, equilibria, small oscillations, and Diophantine relations: Results from the stationary Korteweg–de Vries hierarchy. Journal of Mathematical Physics, 50(12):122701, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i12/p122701_s1.

Lee:2009:ASS

[2160] M. Howard Lee. Analytical study of the superstable 3-cycle in the logistic map. Journal of Mathematical Physics, 50(12):122702, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p122702_s1.

Tarasov:2009:DMM

[2161] Vasily E. Tarasov. Discrete map with memory from fractional differential equation of arbitrary positive order. Journal of Mathematical Physics, 50(12):122703, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p122703_s1.

Ngome:2009:CMC

[2162] J.-P. Ngome. Curved manifolds with conserved Runge–Lenz vectors. Journal of Mathematical Physics, 50(12):122901, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i12/
p122901_s1.

Harland:2009:KCL

[2163] Derek Harland. Kinks, chains, and loop groups in the CP^{n} sigma models. Journal of Mathematical Physics, 50(12):122902, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i12/
p122902_s1.

Celerier:2009:CMC

[2164] Marie-Noëlle Célérier. Chaotic motion in classical fluids with scale relativistic methods. Journal of Mathematical Physics, 50(12):123101, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p123101_s1.

Bihlo:2009:LSE

[2165] Alexander Bihlo and Roman O. Popovych. Lie symmetries and exact solutions of the barotropic vorticity equation. Journal of Mathematical Physics, 50(12):123102, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i12/p123102_s1.

Ghaffari:2009:SGC

[2166] A. Ghaffari, A. Alinejad, and M. Eshaghi Gordji. On the stability of general cubic-quartic functional equations in Menger probabilistic normed spaces. Journal of Mathematical Physics, 50(12):123301, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i12/
p123301_s1.

Muslih:2009:SMA

[2167] Sami I. Muslih and Om P. Agrawal. A scaling method and its applications to problems in fractional dimensional space. Journal of Mathematical Physics, 50(12):123501, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i12/p123501_s1.

Zhou:2009:MHS

[2168] Ruguang Zhou. Mixed hierarchy of soliton equations. Journal of Mathematical Physics, 50(12):123502, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i12/p123502_s1.

Kytola:2009:SIV

[2169] Kalle Kytölä and David Ridout. On staggered indecomposable Virasoro modules. Journal of Mathematical Physics, 50(12):123503, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p123503_s1.

Chen:2009:NTI

[2170] Jinbing Chen. Neumann type integrable reduction for nonlinear evolution equations in 1+1 and 2+1 dimensions. Journal of Mathematical Physics, 50(12):123504, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p123504_s1.

Ren:2009:SGE

[2171] Bo Ren, Xue jun Xu, and Ji Lin. Symmetry group and exact solutions for the 2 + 1 dimensional Ablowitz–Kaup–Newell–Segur equation. Journal of Mathematical Physics, 50(12):123505, December 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i12/p123505_s1.

Liu:2009:RPS

[2172] Hongyu Liu, Hai Zhang, and Jun Zou. Recovery of polyhedral scatterers by a single electromagnetic far-field measurement. Journal of Mathematical Physics, 50(12):123506, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i12/p123506_s1.

Hovhannisyan:2009:NDE

[2173] Gro Hovhannisyan and Wen Liu. On nonautonomous Dirac equation. Journal of Mathematical Physics, 50(12):123507, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i12/
p123507_s1.

Yang:2009:CNE

[2174] Jianwei Yang and Shu Wang. Convergence of the nonisentropic Euler– Maxwell equations to compressible Euler–Poisson equations. Journal of Mathematical Physics, 50(12):123508, December 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i12/p123508_s1.

Gudder:2009:QMI

[2175] Stan Gudder. Quantum measure and integration theory. Journal of Mathematical Physics, 50(12):123509, December 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i12/p123509_s1.

Conrady:2009:QGP

[2176] Florian Conrady and Laurent Freidel. Quantum geometry from phase space reduction. Journal of Mathematical Physics, 50(12):123510, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p123510_s1.

El-Jaick:2009:LCH

[2177] Léa Jaccoud El-Jaick and Bartolomeu D. B. Figueiredo. A limit of the confluent Heun equation and the Schrödinger equation for an inverted potential and for an electric dipole. Journal of Mathematical Physics, 50(12):123511, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p123511_s1. See erratum [2207].

Yokomizo:2009:MCL

[2178] N. Yokomizo and J. C. A. Barata. Multiple classical limits in relativistic and nonrelativistic quantum mechanics. Journal of Mathematical Physics, 50(12):123512, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i12/p123512_s1.

Wang:2009:PSE

[2179] Deng-Shan Wang and S. Y. Lou. Prolongation structures and exact solutions of K(m, n) equations. Journal of Mathematical Physics, 50 (12):123513, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p123513_s1.

Zhou:2009:RCN

[2180] Yong Zhou and Milan Pokorný. On a regularity criterion for the Navier– Stokes equations involving gradient of one velocity component. Journal of Mathematical Physics, 50(12):123514, December 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i12/p123514_s1.

Becker:2009:IIW

[2181] Peter A. Becker. Infinite integrals of Whittaker and Bessel functions with respect to their indices. Journal of Mathematical Physics, 50(12):123515, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p123515_s1.

Demskoi:2009:NAT

[2182] Dmitry K. Demskoi and Jyh-Hao Lee. On non-Abelian Toda A_{2}^{(1)} model and related hierarchies. Journal of Mathematical Physics, 50(12):123516, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p123516_s1.

Ehrlich:2009:NSF

[2183] Paul E. Ehrlich, Jong Ryul Kim, and Seon-Bu Kim. Nonisolated spacelike focal and conjugate points in Lorentzian geometry. Journal of Mathematical Physics, 50(12):123517, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i12/p123517_s1.

Camargo:2009:ADF

[2184] R. Figueiredo Camargo, E. Capelas de Oliveira, and J. Vaz, Jr. On anomalous diffusion and the fractional generalized Langevin equation for a harmonic oscillator. Journal of Mathematical Physics, 50(12):123518, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p123518_s1.

Soloviev:2009:QFT

[2185] M. A. Soloviev. Quantum field theory with a fundamental length: A general mathematical framework. Journal of Mathematical Physics, 50(12):123519, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i12/p123519_s1.

Baleanu:2009:AIC

[2186] Dumitru Baleanu and Octavian G. Mustafa. On the asymptotic integration of a class of sublinear fractional differential equations. Journal of Mathematical Physics, 50(12):123520, December 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i12/p123520_s1.

Dunkl:2009:VSU

[2187] Charles Dunkl and Karol Zyczkowski. Volume of the set of unistochastic matrices of order 3 and the mean Jarlskog invariant. Journal of Mathematical Physics, 50(12):123521, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v50/i12/p123521_s1.

Lewandowski:2009:QGC

[2188] Jerzy Lewandowski and Andrzej Okolów. Quantum group connections. Journal of Mathematical Physics, 50(12):123522, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i12/
p123522_s1.

Toppan:2009:APC

[2189] Francesco Toppan and Piet W. Verbeek. On alphabetic presentations of Clifford algebras and their possible applications. Journal of Mathematical Physics, 50(12):123523, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v50/i12/p123523_s1.

Li:2009:FOD

[2190] Junjie Li. On a fourth order degenerate parabolic equation in higher space dimensions. Journal of Mathematical Physics, 50(12):123524, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i12/p123524_s1. See erratum [2211].

Ramm:2009:CDP

[2191] A. G. Ramm. Creating desired potentials by embedding small inhomogeneities. Journal of Mathematical Physics, 50(12):123525, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p123525_s1.

Suzuki:2009:SFI

[2192] Tatsuo Suzuki. Some formulas for invariant phases of unitary matrices by Jarlskog. Journal of Mathematical Physics, 50(12):123526, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p123526_s1.

Caroca:2009:GHO

[2193] R. Caroca, N. Merino, A. Perez, and P. Salgado. Generating higher-order Lie algebras by expanding Maurer–Cartan forms. Journal of Mathematical Physics, 50(12):123527, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i12/p123527_s1.

Lopez-Ruiz:2009:GSC

[2194] R. López-Ruiz, Á. Nagy, E. Romera, and J. Sañudo. A generalized statistical complexity measure: Applications to quantum systems. Journal of Mathematical Physics, 50(12):123528, December 2009. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v50/i12/p123528_s1.

Angelescu:2009:ELS

[2195] N. Angelescu, R. A. Minlos, J. Ruiz, and V. A. Zagrebnov. Erratum: “Lower spectral branches of a spin-boson model”, [J. Math. Phys. 49, 102105 (2008)]. Journal of Mathematical Physics, 50(12):129901, December 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v50/i12/p129901_s1. See [1615].

Castro:2010:CWF

[2196] L. B. Castro and A. S. de Castro. Comment on “Wave functions for a Duffin–Kemmer–Petiau particle in a time-dependent potential”, [J. Math. Phys. 48, 073515 (2007)]. Journal of Mathematical Physics, 51(3): 034101, March 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i3/p034101_s1. See [1108].

Chen:2010:CIR

[2197] Yong Chen. Comment on “On the imaginary-real ratio rule of power spectra”, [J. Math. Phys. 50, 063301 (2009)]. Journal of Mathematical Physics, 51(4):044101, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p044101_s1. See [1921].

Park:2010:CAT

[2198] Choonkil Park and M. Eshaghi Gordji. Comment on “Approximate ternary Jordan derivations on Banach ternary algebras”, [Bavand Savadkouhi et al. J. Math. Phys. 50, 042303 (2009)]. Journal of Mathematical Physics, 51(4):044102, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i4/p044102_s1. See [1830].

Wada:2010:EGL

[2199] Tatsuaki Wada. Erratum: “Generalized log-likelihood functions and Bregman divergences”, [J. Math. Phys. 80, 113301 (2009)]. Journal of Mathematical Physics, 51(4):049901, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v51/i4/p049901_s1. See [2119].

Yang:2010:ESS

[2200] Yisong Yang and Ruifeng Zhang. Erratum: “Steady state solutions for nonlinear Schrödinger equation arising in optics”, [J. Math. Phys. 50, 053501 (2009)]. Journal of Mathematical Physics, 51(4):049902, April 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i4/p049902_s1. See [1882].

Andrzejewski:2010:CFD

[2201] Krzysztof Andrzejewski, Joanna Gonera, and Paweł Maślanka. Comment on “Functional determinants in higher derivative Lagrangian theories”, [J. Math. Phys. 50, 103517 (2009)]. Journal of Mathematical Physics, 51 (5):054101, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p054101_s1. See [2080].

Wang:2010:CGN

[2202] Hai-Jun Wang. Comment on “General nonlocality in quantum fields”, [J. Math. Phys. 49, 033513 (2008)]. Journal of Mathematical Physics, 51(5): 054102, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v51/i5/p054102_s1. See [1407].

DiCriscienzo:2010:EFD

[2203] Roberto Di Criscienzo and Sergio Zerbini. Erratum: “Functional determinants in higher derivative Lagrangian theories”, [J. Math. Phys. 50, 103517 (2009)]. Journal of Mathematical Physics, 51(5):059901, May 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i5/p059901_s1. See [2080].

Lv:2010:EEP

[2204] Cui hong Lv and Hong yi Fan. Erratum: “Explicit pure-state density operator structure for quantum tomography”, [J. Math. Phys. 50, 102108 (2009)]. Journal of Mathematical Physics, 51(8):089901, August 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v51/i8/p089901_s1. See [2037].

Kundu:2010:PNN

[2205] Anjan Kundu. Publisher’s note: “Nonlinearizing linear equations to integrable systems including new hierarchies with nonholonomic deformations”, [J. Math. Phys. 50, 102702 (2009)]. Journal of Mathematical Physics, 51(10):109903, October 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v51/i10/p109903_s1. See [2045].

Bisi:2011:EFD

[2206] M. Bisi and G. Spiga. Erratum: “Fluid-dynamic equations for granular particles in a host medium”, [J. Math. Phys. 46, 113301 (2005)]. Journal of Mathematical Physics, 52(3):039901, March 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v52/i3/p039901_s1. See [424].

El-Jaick:2011:ELC

[2207] Léa Jaccoud El-Jaick and Bartolomeu D. B. Figueiredo. Erratum: “A limit of the confluent Heun equation and the Schrödinger equation for an inverted potential and for an electric dipole”, [J. Math. Phys. 50, 123511 (2009)]. Journal of Mathematical Physics, 52(4):049901, April 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v52/i4/p049901_s1. See [2177].

Rabeie:2011:CCS

[2208] A. Rabeie and Z. Jalilian. Comment on “Coherent states on spheres”, [J. Math. Phys. 43, 1211 (2002)]. Journal of Mathematical Physics, 52(8): 084101, August 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v52/i8/p084101_s1. See [1, 211].

Haas:2011:CDS

[2209] F. Haas. Comment on “Dynamical systems and Poisson structures” [J. Math. Phys. 50, 112703 (2009)]. Journal of Mathematical Physics, 52(12):124101, December 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v52/i12/p124101_s1. See [2112].

Dokovic:2012:EPI

[2210] Dragomir Ž. – Doković and Andreas Osterloh. Erratum: “On polynomial invariants of several qubits” [J. Math. Phys. 50, 033509 (2009)]. Journal of Mathematical Physics, 53(4):049902, April 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i4/p049902_s1. See [1813].

Li:2012:EFO

[2211] Junjie Li. Erratum: “On a fourth order degenerate parabolic equation in higher space dimension” [J. Math. Phys. 50, 123524 (2009)]. Journal of Mathematical Physics, 53(7):079901, July 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v53/i7/p079901_s1.

Garat:2014:ETG

[2212] Alcides Garat. Erratum: “Tetrads in geometrodynamics” [J. Math. Phys. 46, 102502 (2005)]. Journal of Mathematical Physics, 55(1):019902, January 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [377].

Schwartz:2014:ERQ

[2213] Charles Schwartz. Erratum: “Relativistic quaternionic wave equation” [J. Math. Phys. 47, 122301 (2006)]. Journal of Mathematical Physics, 55(5):059902, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [833].

Han:2015:EGN

[2214] Lei Han and Hai-Jun Wang. Erratum: “General nonlocality in quantum fields” [J. Math. Phys. 49, 033513 (2008)]. Journal of Mathematical Physics, 56(2):029901, February 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1407].

Fioresi:2015:EMC

[2215] R. Fioresi, M. A. Lledó, and V. S. Varadarajan. Erratum: “The Minkowski and conformal superspaces” [j. math. phys. 48, 113505 (2007)]. Journal of Mathematical Physics, 56(10):109902, October 2015. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1251].

Bender:2016:CLH

[2216] Carl M. Bender, Mariagiovanna Gianfreda, Nima Hassanpour, and Hugh F. Jones. Comment on “On the Lagrangian and Hamiltonian description of the damped linear harmonic oscillator” [j. math. phys. 48, 032701 (2007)]. Journal of Mathematical Physics, 57(8):084101, August 2016. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [946].

Kundu:2017:CCL

[2217] Prasun K. Kundu. Comment on “Comments on ‘The Euclidean gravitational action as black hole entropy, singularities and space-time voids’ ” [J. Math. Phys. 50, 042502 (2009)] — Schwarzschild black hole lives to fight another day. Journal of Mathematical Physics, 58(11):114101, November 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1834].